A Complete Bibliography of Publications in Limnology and Oceanography: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.05

Title word cross-reference

$1.50 [Fer70]. $1.75 [Fer70]. $10 [FW74]. $10.00 [Gol74a, Ril71a, Edm72b, Gro74b, Woo73]. $10.50 [Eme78]. $11.00 [Hyn71]. $11.50 [Hyn76, Liv72b]. $11.95 [Mil73a, Mil73b]. $12.00 [Eme76, Kil76, Pai70]. $12.50 [Ril73, Ste70]. $12.95 [Day79, Str75b]. $13.00 [Zar78]. $13.50 [Kra79, Man70, Smi76b, Bar71, Fag72, Ton72]. $14.50 [Gor79a, Gor79b, Koh71a, Par72, Sch76]. $14.60 [Man75b]. $14.95 [Koh77]. $15.00 [Gol74b, Koh73]. $16.00 [Gro71]. $16.50 [Liv76, Mil78]. $17.50 [Gro75b, Gro78, Ich71, Pom79, Stu71]. $17.95 [Sch77a]. $18.50 [Hed70, Pai75a]. $18.90 [Enr76]. $19.00 [Hor75]. $19.50 [Par72, Csa75, Gor78]. $19.80 [Boo75a]. $2.00 [Fer70]. $2.25 [Fer70]. $20.00 [Gro75a, Koz70, LaF77]. $21.50 [Lym73, Sha73a]. $21.75 [Wal76a]. $22.00 [How72, How78]. $22.50 [Hob74]. $24.00 [Eme76, Par75, Sta74b]. $24.50 [Pla78]. $24.75 [Phl73]. $24.80 [Hun77]. $24.95 [Liv72a, O’C71]. $25.00 [Cum71, Ebe72, Hun75, Ori73, Woo79]. $25.50 [Smi74a]. $26.95 [Smi76a]. $27.50 [Edm71a, Ill75, Kap78, Lee77b, Smi78a]. $28.90 [Car75].

$29.50 [Man75a, Woo77]. $29.70 [Cha74]. $29.80 [Man76]. $3.00 [Fer70, Cul73a, Pom74]. $3.95 [Sta72a]. $30.00 [Wet76]. $31.10 [Whe73]. $32.00 [Ste79a]. $32.50 [Ban73, Lun79]. $33.00 [Gro74a]. $34.00 [Ban73]. $34.50 [Tyl71]. $34.80 [Kel75]. $34.95 [Edm74a]. $35.00 [Bou75, Edm72a, Gro77, Val74b]. $36.10 [Smi78b]. $37.50 [DH72]. $39.50 [Rig76]. $39.60 [Hob79]. $39.95 [Gus79]. $4.75 [Fer70]. $40.00 [Dug71, LeC77]. $41.00 [Ori77]. $43.60 [Ril77]. $45.00 [Eme74]. $47.00 [Gro76b]. $48.00 [Fan77]. $49.00 [Pyt76b, Gol77a, Ste79b]. $49.20 [Koh74]. $49.50 [Ril78a]. $49.95 [Tea78]. $5.00 [d’A75]. $5.25 [Hor76]. $5.95 [Smi75]. $51.50 [Smi76c]. $51.95 [Liv77b]. $55.00 [Epp73]. $57.00 [Bow79]. $59.50 [Mac79]. $6.00 [Hut72]. $6.95 [Val71a]. $62.50 [Edw77]. $7.00 [And71, Fra72, Hob73]. $7.50 [Cle72]. $7.95 [Sch75b, Wal74b]. $8.00 [Hut72, Zar78, Ril71a]. $8.75 [Kin70, Ril74]. $8.90 [Par70]. $8.95 [Kra79, Col72, Fry71]. $9.00 [Rig72]. $9.25 [Gro76a]. $9.50 [Dee74, Par73, Sta75]. $9.80 [Liv78]. ^{13} [SCA77]. ^{133} [SRR79]. ^{137} [KB76]. ^{14} [Bro78a, Bro79, CE79, GV73, IBM76, LS77, LB79, Mcm73, Pug70, SD78, Taf76, The75, TP73, WWR70]. ^{15} [SCA77]. ^{2+} [Dav77, Lev71b, Wil78]. ^{2−} [Dav77, Whi71]. ^{210} [AW75, VBT79]. ^{35} [JDL78]. ^{55} [KLAS73, LS77]. ^{90} [KB76]. _{12} [BS78c, CB72, Nat70, OT72, Par77]. _{2} [BBG75, Bro79, CPOT79, Dav77, Eme75a, HW78, Mac79, MEW^{+}78, OT77, SR77, Taf76, TJ79]. _{2}^{0} [Bra73a]. _{3} [IBM76, Toe71]. _{4} [JDL78, OT77, Toe71]. _{4}^{2} [Lev71b]. _{4}^{2−} [SM76]. : [CAE78, DDS78, Rhe78]. μ [MCB79, Ogu70].

0.5- [Ogu70]. 0.95 [Cas71]. 000-year [Liv71].

1 [Enr77b, Fan77, Hut72, KOW79, SO78, Smi74a, VWB77a]. 1.90 [Aro73]. 10-year [GS72a]. 10.50 [LaF75]. 10s [Bro70]. 12.50 [But76]. 120 [Sio77]. 125 [Liv77b]. 13-1 [Sma74a]. 14 [Hyn78]. 14.00 [Mil77]. 162 [Edw77]. 18.50 [Pyt76b]. 182 [d’A75]. 19.50 [Smi76c]. 1967 [NW70]. 1969 [Bar71]. 1970 [HC75b, Fer70]. 1972 [MSR^{+}75]. 1973 [Val74b]. 1974 [Ano72c]. 1976 [Ano74d, Gus79]. 1st [Hob74].

2 [Ban73, EH77, OS78, VWB77b, WK79]. 2-month [RB71]. 2.95. [Koh72]. 20th [Lym73]. 22- [MCB79]. 22.40M [Sta72b]. 24-h [ES75a]. 25 [Hob79]. 25^{o} [PH74, Whi74b, Whi75]. 29 [Sio77]. 2nd [But76, Gol79, Gro76b, Mil77, Pyt76b, Smi76c].

3 [Ban73, Koh74, Pom74, Sch73]. 3.00 [Mof73]. 3.50 [Man73a]. 35 [Woo79]. 36th [Ano74b]. 3A [Gol74a].

4 [Mil78]. 4.00 [Hyn73]. 40th [Val74b]. 41st [Ano79d]. 45s [Par70].

5 [Mas70]. 5. [Man70]. 5.50 [Bri72]. 50 [Gro79a]. 5th [Hyn76].

60 [And71]. 66 [Fot70]. 69 [Sho74].

7.00 [Liv77a].

85 [Sch73]. 8th [Gus79].

90F [Mar78a]. 98F [Cus76a]. 98F. [Cus76a]. 9th [Bow79].

= [Pro71].

A$4.00 [Man72a]. Aagaard [LaF77]. AAPSE [Cul73a]. Aaronson [Sha78]. abiotic [LR74a]. Above [Smi71]. absolute [Bat73]. Absorption [WWR70, Car70, DG78, Hoj75, KB76, MB70a, TBD77]. Abundance [BSB71, RC72, DEWN76, Geh74, Gou71, Kar78]. abundans [Gid77]. abyssal [BHSJ78, Lau74]. Acad [Bar71, Ori73]. Academia [Sch73]. Academic [Aro73, Bri72, But76, Csa75, Gol79, Gor79a, Gor79b, Gor78, Gro76b, Gro78, Hed70, How78, LaF75, Man75b, Mas70, Mil77, Pai75a, Pom79, Pyt76b, Rig72, Rig76, Smi76c, Sta74b, Stu71]. Academy [Hyn76]. Acantharia [BSB71]. acantharians [BS70]. Acanthocyclops [LL79]. Acartia [BGMW74, DD78, KVW79, Kim79, Lan75a, MJH77, NN76, OSD76, RLS75]. Accommodation [KO76]. accompanying [SN71]. account [And71]. Accumulation [BA71b, Cam75, SCLM70, WL78]. accurate [Sal79]. Acer [LI76]. Acetate [SS77, CS73a]. Acetylene [CB76, RB70, VHMW74]. Acid [GCL79, AC76, CAE78, CE79, CW78b, Jef72, KH76, Koe76, MCHP73, NG78, Ros79a, SS74, TB75, WNS74, Wri75, WH78]. acid-bubbling [CE79, TB75]. Acidification [WSR79, Ram76]. acidity [GCL79]. Acids [BCB71, MQ71, CJN72, CC75b, GL75a, HWL70, Jef70, LB77, LP70, MQM74, Mey76, OW73, QM71, RHG73, Sch74, SN71, dH77, dlCP75]. Acorn [Hur73]. Acoustical [Gre79]. across [HC78, JPW79, Ver75]. acrylic [Ris73]. activated [Coo73a]. active [Boo76]. activities [CP79, LNS^{+}76]. Activity [Pyt75, YYP72, ATM74, AH77, Bro78a, DP78, FR77b, HP70, Har73c, HHHP^{+}72, KP75, Ley73, NCK79, Ore79, PHR71, RS79, SM79, SS74, SD78, Sta78, TA78, WBM^{+}77]. acuatica [Zar78]. adaptation [Ban79]. Adaptations [Jon78, DM79a, KW77]. adapted [FF79]. adaption [How78]. Adaptive [Dod74, Enr77b, EH77, Hai79a, Koh71b, Pea79]. added [RAH77]. addition [GHM78]. Additional [Mac78a]. adenosine [BPW76, Bul78, KA77]. adhesion [CNC73]. Adirondack [GL79]. adjacent [How72]. adjustments [Man73b]. Adriatic [SP73]. adsorbent [Moe71]. adsorption [CACB75, Ros79a, Ros79b]. Adsorptive [KV73]. adult [Mal70, NN76]. advantage [Enr77b]. advantages [KK79]. aerial [DL74]. Aerobic [Fit70, Har72, BT70, OH72a]. Affairs [Edm70a]. affecting [LL79, Sch75a]. affinis [GBB72, Rat74]. affinities [Wil72b]. Africa [Liv77a, Hul76, Jan75, LK74, NG78]. African [KH73, MK74, SMD76]. after [Bro78a, Dun70, KB76, McL75, SD78, WSR79]. Ag [And76]. agardhii [Sau72a]. age [KOW79, Thr79]. aged [NWEG75]. agents [MM79]. ages [TTC^{+}77]. aggregates [All79, RSS78]. aggregation [Pae74]. Agiosperm [Wil72a]. Agnes [Hoo74]. Aids [GT74]. air [BS74, DVCP79, HC79a, Ver75]. air-water [Ver75]. Airborne [GH76, NW70]. Akademiai [Hor75]. al. [SCA79]. Alabama [May73]. Alaska [Gol78, Gol74a, Gro75a, Ril71a, ICO^{+}74, LH72, Sch74]. alewife [Wel70]. Alewives [War72]. Aleyev [Ham79, Wu79]. Alga [YYP72, Rhe72, Rom78, SH73, SL78]. Algae [Sec71, Arn71, Bro78b, Bro74, CS78, CVV78, CK79, CB75b, EGTA70, FP79b, FC70, GG81, GA75, JC75, JSS^{+}71, MW77, MM79, MCSK79, MYY71b, MB71b, SL76, Smi78b, SL70, SP73, VVT76, Wal73, Sta74b]. Algal [HA72, Por72, Ban79, BLM74, BH78, CNC73, DB76a, Dun70, GD74, GIK77, GC74a, GLSD74, Gri78, KK76a, Lea70, Loh74, OH72a, OH72b, Rhe78, Sen78, SG76, WEMD72, WF72, YK73]. Aliphatic [WC76]. Alkaline [Ber70b, YYP72, AFK73, HK73, MK74, RS79]. Alkalinity [BG76, Kin78, Won79, BST70, MSH74]. Allen [Sha73a]. Allomorphotic [Ton71]. allow [GTF78]. along [BP74b, CL76, Sha74a, Sha77a]. alongshore [HPS75]. Alosa [War72]. Also [Par70]. alter [Dav74b]. altered [KO76]. alternate [SW73]. alterniflora [CW78a, GK74, Rei72a]. altiplano [HBLP76]. altitude [Mal73]. Aluminum [Moe71]. Alversos [How72]. Amazon [FP79b, Man79]. ambient [HP77]. Ambleside [Cas71]. Amer [And71, Cul73a, Gro71, Hut72]. Amerasian [KLG71]. America [Zar78, Dug71, Wil72b, WS70]. American [Ano70g, Ano71i, Ano73f, Ano74b, Ano79d, Col72, Gro74a, Lau70, Pai70, Sau72b, Ano71b, Ano71a, Ano72a, Ano73a, Ano74a, Ano75a, Ano75n, Ano76a, Ano76e, Ano77b, Ano77a, Ano77g, Ano78b, Ano78a, Ano78c, Ano78g, Ano79a, DD71]. americanus [FS72a]. amines [Nor75]. Amino [BCB71, HWL70, Ros79a, WNS74, dlCP75, CJN72, CC75b, GL75a, LB77, LP70, Sch74, SN71]. ammonia [Deg73, GHM78, GG75a, LTB75]. Ammonium [ERHC79, Ros79b, Bat76, BS77, Bie75, Big77, EG79, ECHH^{+}71, Gar78, GM78, MYY71a, MD78, Tho70a, TR76]. among [TGK^{+}78]. Ampeliscidae [Mil71]. amphipod [Har70, Har71]. Amphipoda [Mil71]. Amphistegina [Mul74]. Amphora [And76]. amplification [BK72]. AMSA [Man72a]. Amsterdam [Bow79, Car75, Edw77, Gus79, Liv77b, Smi76a, Ste79b, Tea78, Wal76a]. Anabaena [Hor79, HDFG72]. Anaerobic [CW78a, PB73, RB74, OH72b]. analys [SS78a]. Analyses [Buz71, EG79, HAG^{+}72, KB76]. Analysis [LP70, MP77b, ABGCP75, BGZS75, Bas75, Bie75, BM73, BH70, Cas71, CR76, DCC73, EK73, GH72, GMTA79, GS74, GB74, HPM72b, JM78, JS73, KV73, Lew77, Man72a, MQM74, MSH74, Mos76, RHW79, Rig76, SH78, SM76, Sha74b, Wil73, Ril77, Ste79a]. Analytical [Par73]. analyzed [KMP^{+}75]. analyzer [Taf76]. Analyzing [Kno77]. ancient [Sio77]. Andersen [Mac79]. Anderson [Bal76, Hyn78]. Angeles [Man70].

Angiospermae [Hut70]. Anikouchine [Mil73a, Mil73b]. Animal [Sno72, HCW70]. animals [Chi71, Hyn74, Jac70, Enr76, Rig72]. Ann [Leh75, Sch76, Woo77]. annexes [Edm70a]. Announcement [Ano72c, Ano72b, Ano73b, Ano75b, Ano75c, Ano75d, Ano76b, Ano77b, Ano77c, Ano77d, Ano78c, Ano74d]. Announcements [Ano70a, Ano70b, Ano72d, Ano78d, Ano79b]. Annual [Ano74b, Ano79d, BP74a, DB71, TS78, WC73, Dee71, GS73a, GS73b, Sma74a, SA74, Tim75]. Anodic [Bra73a]. Anomalies [Lev71b]. Anoxic [BY73, GP72a, BST70, Ber77, BS71, DH78b, MB77, MBR78, RAC71, RG76a, Ros79a, Ros79b]. Anstalt [Koz70]. Antarctic [WRO73, AVZ75, Hed70, Mas70, RSMC76]. Antarctica [HSW72, Mac70, RD76]. antenna [Ton71]. Anvers [WRO73]. apatites [AP77]. Aphanizomenon [HSC79, Hor79]. aphotic [And75, Fou70, Fou71, Fou73, TPG77]. Apparatus [BS74, DS75, BHD76, CE79, O’C74, SB75, Wal73, WB74]. apparent [Fou73, Glo78, MCHP73, Pyt71, PIM74, Zir75]. appendicularian [All76]. applicability [KK76a]. Application [MCA77, WQ73, RHW79, SO78, Wil73]. Applications [Ber72, BYR74, CA72]. Applied [Man73a, WWH70]. appliquee [Par78]. Applying [Cha79a]. approach [GC74a, HCW70]. aqua [Fra72]. aquaria [FR77a]. Aquasol [IBM76]. Aquatic [GP77, Kat71, Wil72a, Zar78, BHK78, CK79, Cha79b, Coo73b, CSHE78, DH75, GG81, Hob74, Hou74, Hyn78, JF72, LV76, LL77, MB76a, McL75, MS74b, Pae78, Par70, Pug73, Rhe72, Rig72, SSHV76, SOP73, Sch76, SA74, SS77, Swe76, Edm72a, O’C71]. Aqueous [Bra73a, DSC77, DH72]. aragonite [CB73]. Arbor [Leh75, Sch76, Woo77]. Archipelago [RB71]. Archon [Ori73]. Arctic [Gol78, Gor79a, Gor79b, Kel75, CL73, CO75, CK74, Daw78, Dod75, Dod79, HA72, KLG71, MG75, OS79, RC72, Wel74]. Area [Fee76, BGZS75, EP71, ES77, GG78, GC73, Har72, Owe74, TS78]. areas [AZ70, Jac70, Pai70]. arising [Mar77]. armatum [CL76]. Army [Cle72]. Arsenate [Pil74]. Arsenic [And79, Cre75, GB74]. Artemia [SM79]. Artemis [Kra79]. artificial [CK79, GG81, Hil74]. Arts [Ori73]. Aspects [WDW^{+}75, BBBS79, LaF75, RMG70, SF75, Koh71a, Wal76a]. asplanchna [Gil71]. Ass [Cas71, Cul73a, Gro71, Man72a]. assay [CAE78, GG75a]. assays [PG72]. assemblage [Til73]. assemblages [RK79, Sil75, WM72b]. Assembly [Ano74d, Hyn76]. assessing [BB77]. Assessment [HB79, Hob73, Ton72, Gol78]. Assimilation [LL73, NG78, Arn71, Bie75, ME72, Pet75, RLS75, TP74, TJ79]. Assoc [Gro79a]. Associated [Cul70, GP77, HC75a, Lee77b, McL74, MGC73]. Association [Ano73h, Gro74a, Whi74b]. associations [SAH72]. assumptions [LBL75]. Asterionella [TKK76]. Aswan [Sha77a]. asymmetry [Boo75b]. Atlantic [CP75, GN79, MC78, Pai70, WP76, AGM70, BGMW74, BSB71, BNV^{+}74, Fou71, Kel70, Mil71, SA71]. Atlas [Pai70, Dug71]. atmosphere [HIJ78, Whi74b, Kra79]. Atmospheric [GL79, Hea73, MCHP73, PH71, Whi75]. Atoll [QS70, GM74, SM73, WDW^{+}75]. atom [FHV79]. atomic [Car70, Rhe78]. ATP [Bal72, Cav76, CMDC78, CW78b, Jas75, SS78a, SOHH76]. attached [GP77, WM72b]. attack [FR77a]. attained [MCSK79]. Attention [Kro71]. Attenuation [MCSK79]. attenuator [KS70]. Attributes [HC79b]. Aufwuchs [CMDC78]. August [Hyn76]. Auke [ICO^{+}74]. Aust [Man72a]. Austin [Cul73a]. austral [Zar78]. Australia [Edm76a, Gro71, Tim72]. Author [Ano72e]. Authors [Ano73c, Ano75e, Ano76c, Ano77e, Ano78e, Ano79c, Val74a]. autoanalyzer [HAG^{+}72]. autolysis [OW74]. Automated [GB74, WCR71, And77, CC75b, DCC73, KHC74, Sta74a]. autoradiography [CA72, FC70, KK76a, KK76b, KK79, PS79, SD78, SS77]. Autumn [HDFG72, LI76]. autumn-shed [LI76]. availability [Ber70b, MTT77]. Available [LC71a, Edm70a, Par70]. averaged [TG77]. avoidance [Egg78, ZS76]. Awda [Bax77].

B [Boo75a, Bro70, Cha74, Cum71, Ebe72, Edm72a, Gar74, Gro71, Gro79a, Har78a, Hun77, Hyn71, Hyn73, Hyn78, Koh74, LaF77, LeC77, Lee77b, Liv76, Rig76, Woo79, SL76, BS78c, CB72, Høj78, KVW79, Nat70, OT72, Par77, SC76a]. Ba [SRR79]. Bacteria [Dal74, Wei73b, BLM74, BT70, BC72, BS74, BB70, Cas73, CP79, CACB75, DH75, FR76, FP79a, FR77b, GMCG78, HPM72a, Jon74, KB71, MH74, PHH78, RN78, Sor73, TeI70, TVB79, TJ72, Wal73, Wri75]. Bacterial [Bot75, BB73, DM79a, Jan73, Pae74, Dis73, GV78, Jas75, JP78, RHG73, SS77, CKS77b]. bacteriochlorophyll [MB76a]. bacteriochlorophyll-c [MB76a]. bacteriological [Sch70]. Bacterium [PA70, Rhe72]. bags [May76, TP73]. bahamense [Mac77]. Bahía [SCL^{+}71, CMB70]. Baja [Ber70a, Bla77, EB79, NG78, WKW^{+}74, WWK^{+}77]. balances [Woo77]. Balanus [Hur73]. Baltic [WL78]. Baltimore [Koh71a, Sta75]. Band [Wet75]. Bank [BBBS79]. Barbara [KLAS73, SG71]. barbipes [KA71]. Barcelona [Ori77, Par78]. Bardawil [GG78, Lev71b]. Barents [Kel70]. Barnacle [Hur73, GCL70]. Barne [HSW72]. Barnes [Man73a]. Barobiology [FW74]. barrier [Mey76]. Baruch [Har78a]. based [Gro79b]. basin [BBG75, Kir75b, MLCG79, PF73, RAC71, TC74, KLG71, KLAS73, SG71, Zo73]. baskets [PN79]. Bastian [TLR^{+}70]. Bathymetry [Kje78]. bathyoxymeter [JFKG72]. Bay [BS78c, DD78, Gro71, HPM72b, ICO^{+}74, KM73, May73, Mos76, RSS78, RB71, VHMW74, SCLM70, SK73b, GBK76, HS77b, HS77c, HC75b, LB77, Mar70a, MTT77, Pin73, Sma74a, TLT77, TS78, Won79]. Bayly [Edm76a]. bcnthic [JC77]. Bd [d’A75]. beach [GGW78, PH71]. Beaches [BA71b, Cub70, Mak72]. Bead [MB71b]. Beadle [Liv77a]. bear [Col72]. bearing [LB71]. bed [CH70]. Bedford [Zo73]. Beeton [Edm74a]. Behavior [Fro72, Kib71, Lev71a, Woo71, AF74, AP77, BR74, Cha79b, Egg78, Fro75, Fro77, HMA^{+}75, HW76, Lan78, Sta78]. Beige [Pyt76a]. Belle [Har78a]. Belowground [VTP76]. Benchmark [Woo79]. beneath [MG75, PDF^{+}78].

Benthic [HLH71, BHSJ78, Cas71, DC75, DB76a, ES77, FG75, Fou73, GT74, GIK77, Har70, JSS^{+}71, Kap78, KMR78, McL75, Mul74, PN79, Pec79, Ped74, Phl73, RHR71, SCE^{+}76, SWLH78, TKH75, TGK^{+}78, WL78, WW74]. benthonic [Sch71]. benthos [Hil74, HPM72b, TWPC75, Fag72, Har78a]. benzoate [dH77]. berg [Par75]. Bering [Gro75a, CT70, CO75, LaF77]. Berkeley [Dug71, LeC77, Lun79, Man70]. Berkshire [Lud76]. Berlin [Boo75a, Cha74, Hob79, Koh74, Liv78, Ste79a, Smi74a]. Bermuda [Dee71, DB71, GGW78, SA71]. Between [GBB72, Kro71, Kro72a, BGMW74, Beg76, Bla73a, BP70, CPOT79, CK79, CB75b, Fee79, GLSD74, GG81, Hai76, Hai79b, Hu78, JP76, Ker78, Kir75b, Lan75a, LJ73, MQ71, MS74b, MYY71b, Mor71, Pat80, Pou78, Rhe72, RHG73, SB75, SA71, WWHH77, YY79]. bias [Ano73g]. Bibliography [Ano71c, Edm70a, Gro74b]. Bicarbonate [PH74]. bicarbonates [Pyt75]. Big [KGP^{+}78]. Bight [SM78, MC79, WWB^{+}78, WP76]. Bikini [NWEG75]. billows [TH74]. binding [GC74b]. Binnengewasser [Edm76b, Wet75, Fot70]. bio [SB78b]. bio-optical [SB78b]. Bioassay [HS77b, Sma74a, CB72, GV78]. Bioenergetic [Swe78, KA71]. Bioenergetics [MMC79]. Bioenvironmental [How72]. Biogenic [Cul77]. biogeochemistry [Woo77, Liv78]. Biogeography [Koh71c]. Biol [Cas71, Hor75, Koz70, Rzo79, Sio77, Zar78]. Biological [Ano78h, Gor79a, Gor79b, Hyn76, JPW79, PKC77, Pec79, And79, Bla73a, BK72, HHGRW76, PME73, Smi76b, Bri79, Epp73, Ril74, Str75b]. biological-chemical [And79]. Biologie [Fot70]. biologique [Par78]. Biology [Edm76b, FS72a, Hor76, Sio77, FW74, Gro74b, HMA^{+}75, Koh71a, Koh74, LV76, Lun79, Man75b, Mul74, Hyn71, Gro74a, Koz70, Pai75a]. biology. [Ril73]. Bioluminescence [RD76, SCL^{+}71, EC72]. bioluminescent [SCLM70]. biomass [Cav76, ES75b, Jas75, JW70, JS79, Kal72, MCB79, PF73, TP78, VTP76]. biosphere [Man76]. biosynthesis [WBM^{+}77]. Biota [Zar78, Zar78]. biotin [CB72, OT72]. bipyridyl [HD77]. birth [Cas72b, Edm72c, Pal74, Thr79]. Biscay [Pin73]. Biscayne [LB77]. Bivalvia [Lev71a]. Biwa [iK75]. Black [Gro74a, Gro74b, Kar78]. Blackwell [Fag72, Fra72, Hob73, Hyn78, Par70, Ton72]. Bldg [Cul73a]. Block [BS78c]. Bloom [HDFG72, Dun70, Glo78, Har73c, HSK76, HS77b, HS77c, SK73b, TS78]. Blooms [CL76, HSC79, Hor79, ICO^{+}74]. Blue [RHG73, YYP72, Arn71, Bro78b, CC74, Gri78, Ham75, JC75, MB71b, Rom78, SH73, Sta74b, Wal73, WF72]. Blue-Green [YYP72, Arn71, Bro78b, CC74, Gri78, JC75, MB71b, Rom78, SH73, Sta74b, Wal73, WF72]. blue-water [Ham75]. BOD [CP73]. Body [Dod79, MB70b, Ton71]. bog [KH76, Koe76]. Bold [Kro71, Kro72a]. Boltovskoi [Rzo79]. Boney [Hor76]. Book [And71, Ano74b, Ano79d, Aro73, Ban73, Bar71, Boo75a, Bou75, Bow79, Bri79, Bri72, Bro70, But76, Car75, Cas71, Cha74, Cle72, Col72, Csa75, Cul73a, Cum71, Cus76a, Day79, Dee74, DH72, Dug71, Ebe72, Edm70a, Edm71a, Edm72a, Edm72b, Edm74a, Edm76a, Edm76b, Edw77, Eme74, Eme76, Eme78, Enr76, Epp73, Fag72, Fan77, Fer70, FW74, Fot70, Fra72, Fry71, Gol78, Gol74a, Gol74b, Gol79, Gol77a, Gor79a, Gor79b, Gor78, Gro71, Gro74a, Gro74b, Gro75a, Gro75b, Gro76a, Gro76b, Gro77, Gro78, Gro79a, Gus79, Ham79, Har78a, Hed70, Hob73, Hob74, Hob79, Hor75, Hor76, How72, How78, Hun75, Hun77, Hut72, Hyn71, Hyn73, Hyn76, Hyn78, Ich71, Ill75, Kap78, Kel75, Kil76, Kin70, Koh71a, Koh72, Koh73, Koh74, Koh77, Koh79, Koz70, Kra79, LaF75, LaF77]. Book [Lau74, LeC77, Lee77b, Leh75, Liv72a, Liv72b, Liv76, Liv77a, Liv77b, Liv78, Lun79, Lym73, Mac79, Man72a, Man75a, Man70, Man73a, Man75b, Man76, Man79, Mar78a, Mas70, Mil73a, Mil73b, Mil78, Mil77, Mof73, O’C71, Ori73, Ori77, Pai70, Pai75a, Pai75b, Par70, Par72, Par73, Par75, Par78, Phl73, Pla78, Pom74, Pom79, Pyt76a, Pyt76b, Rig72, Rig76, Ril71a, Ril73, Ril74, Ril77, Ril78a, Ril78b, Rzo79, Sch77a, Sch73, Sch75b, Sch76, Sha73a, Sio77, Smi74a, Smi75, Smi76b, Smi76c, Smi76a, Smi78b, Smi78a, Sta72a, Sta72b, Sta74b, Sta75, Ste79a, Ste70, Ste79b, Str75b, Stu71, Tea78, Ton72, Tyl71, Val71a, Val74b, Wal74b, Wal76a, Wet75, Wet76, Whe73, Woo73, Woo77, Woo79, Wu79, Zar78, d’A75]. Books [Ano70c, Ano70d, Ano71d, Ano71e, Ano72f, Ano73d, Ano75f, Ano75g, Ano76d, Ano77f, Ano78f, Ano79e, Edm72b, Ori73]. borderland [BBG75]. Bormann [Liv78]. Bosmina [Ste76, Ker78, OS79]. Boston [Cle72]. Bot [Lun79]. botany [Wet76]. bottle [Sch70]. bottles [MHK75]. bottom [Cli72, CP70, CP75, DV78, FN70, GB72c, Hak76, Mar78b, Rei72b, SS70, Sho70, dBSA74, Gus79]. bottom-water [Sho70]. Bougis [Cus76a, Par78]. boulderfield [Fot74]. bound [HBLP76]. boundaries [Ver75]. boundary [BP70, Kap78, Rei72b]. box [KC79]. Boyd [Bro70, Edm72a, Man70]. Brace [Sch75b]. brackish [EG79, Hob79]. Brauer [FW74]. Braunschweig [Ril74]. Breach [Par72]. Breeding [Zo73, Pro71]. Breitner [Dro78]. breve [MSR^{+}75]. brightwelli [AS77]. brightwellii [PG75]. brine [BYS77, BBBS79, CMB70, LJ73]. brines [SRR79, TL75]. Brinkhurst [Edm72a, Val71a]. British [Ano79d, CGPD79, FS72a, Hyn78, JPW79, Pug77]. Broecker [Sch75b]. bromocarbons [HH78]. Brook [HL73b]. Brown [Smi75, Sta72a, Col72]. bryophytes [Mey79]. Bubble [BS74, BS78b, Joh76, Mac78a]. bubbles [BC74, WD78]. bubbling [CE79, TB75]. Budapest [Hor75]. budget [BA71a, CL73, CO75, Dil75, FD72, GN79, Har71]. budgets [AL78, Tim75, Wak77]. Buffer [PA75, Whi74b]. Buffering [BY73]. buildup [GB72b]. bulk [KB76]. Bull [d’A75]. Bulnheim [Koz70]. buoyancy [KS78]. buoyant [HW76]. Bureau [Fer70]. burg [Woo79]. Burial [GE75]. Burrell [Gol78]. burrowing [Jac70, JF72]. bursting [BC74]. Burton [Gro78]. Butterworths [Koh71a].

C [Boo75a, Bow79, Bro78a, Bro79, CPOT79, Ebe72, Eme76, Gar74, Gol78, Gor79a, Gor79b, Gus79, Har78a, Hyn71, Ich71, Koh71a, Koh77, Lee77b, Liv77a, Rig76, Ril73, Smi75, Smi76b, Smi78a, Taf76, MB76a, Zar78, CE79, GV73, IBM76, LS77, LB79, Mcm73, Pug70, PH74, SD78, SCA77, The75, TP73, WWR70, Whi74b, Whi75]. C-diatom [Pug70]. C [?]NaHCOIverson:1976:LRD. C. [Pro71, SK73a]. c/o [Zar78]. Ca [Lev71b]. Cadmium [BKM78, SB79, ZY72, KM73, MB71a, MH74]. calanoid [BGMW74, Bay78]. Calanoida [Coo71]. Calanus [Bal72, Daw78, Fro72, Fro75, Fro77, MB70b, Paf76]. calcification [Kin78, Smi73b]. Calcite [MQ71, CB73, Rey78]. Calcium [CS70, ML70, OW73, Smi72, SE78]. calcul [Pyt76a]. calculated [KB76]. Calculating [BY70, GB79, LP77, Leh77, SW73]. Calculation [Pal74, FHV79, SO78]. Calculations [NN79]. calibrating [HB72a]. Calif [Dug71, Man70, Zar78]. California [LeC77, Lun79, RMG70, TPG77, BBG75, Ber70a, BP70, Bla77, BP74b, CJN72, ERHC79, ERH79, EB79, FN70, GMTA79, HSK76, HSC79, HJG71, HG72, HDFG72, Hor79, Jac77, JP75, KA74, Kim79, KM73, NG78, Nor75, PG72, PRLG75, RH76, RSEG78, RSS78, Sil75, Smi72, WKW^{+}74, WWK^{+}77]. California-Nevada [TPG77, GMTA79, PRLG75]. Caloric [PI73, Sno72, Ker72]. Cambridge [Wal74b]. Camera [RC71]. Cameron [Dil75]. Can [Che73, CE73, d’A75]. Canada [Hyn78, Val71a, Edm70a]. canadensis [FR77b]. Canadian [Eme75b]. Canal [HES75, MM73b, Mor75]. capacity [GV78, PME73, Whi74b]. Cape [HSW72, SM78, SA71]. Capillary [Bro70]. capitata [Ten77]. Capitella [Ten77]. capture [SF75]. carbohydrates [GH72]. Carbon [BW70, Bun71, CK74, Fie72, Kel70, MF71, SS74, Smi73b, SK75, SM74, TP74, WAA72a, Woo77, AGA78, AGF71, Ban74c, BL79, Bra73b, Cam75, CS78, CL76, DDS78, ES75a, GF75, GS74, GG75b, Hai76, HS75, HL73b, JM70, JPW79, KA77, Lam78a, LLS75, LI76, LH72, MWW77, MYY71a, MYY71b, MK79, NS78, PM77, Pae78, PCWR70, PS77, PSK78, Pet78b, SSHV76, Sal79, Sha73c, Sha74b, SCA77, SM73, SN71, SS78b, TCTW75, Ver75, Wan74, WM72a, Whe77, Whi74b, Whi75, WWHH77, YK73]. Carbon-13 [SM74]. Carbon-14 [TP74]. carbon-to-nitrogen [Ban74c, CL76]. Carbonate [BST70, WL78, CS70, Cha74, OW73, PH74, SB73, Smi72, SM74, SE78, Whi75, Gro71]. carbonates [OW72, Pyt75, Smi74a]. carbonic [MCHP73]. carboxylating [GM79]. Carefoot [Day79]. Cariaco [TJ79]. Caribbean [CP77, AGM70, KTTF76]. carinata [OV78]. Carnivorous [Lam78b, Mul79]. Carolina [FW74, Har78a, MG78, SM78, Wol71]. Carotenoid [Rie78]. carotenoids [Gri78]. Carrod [Bri72]. carterae [BL76]. case [Man79]. Castle [JP75]. cation [Sta74a]. Cattaneo [GG81]. Causes [Bar71]. Cayman [Mey76]. Cayuga [PBS74]. Cd [EA79]. CEBADEAU [Pyt76a]. Cebulski [Gro71]. Cell [RW71, And76, BH78, CAE78, HK74, Mal75, McA71, PT73, Rhe78, TK76, YY79]. cells [And75, OH72a, OH72b, Sha77b, Whi72]. cellular [AS77, TD72]. Cellulose [FM78]. census [WP76]. Center [Edm71a, Ich71]. Central [LK74, Ven74, BHSJ78, Daw78, ERVM73, ES75a, Hoo74, Jan75, Per76, SB72, WH71, Wal74a, Wal75]. Centre [Bri79, Pyt76a]. centrifugation [PRG78]. century [Cle72]. Ceratium [Wea79]. Ceratophyllum [Toe71]. Ceratophyllum-Periphyton [Toe71]. Certain [RMG70, BFMF72]. Cesium [BNV^{+}74, Wol71]. cesium-137 [Wol71]. CH [OT77]. Chaetoceros [MGHR75, CL76]. Chaetognath [Zo73]. Chaetognatha [FR77a]. chain [KM72b, ME74, WWB^{+}78]. chains [Man70]. chamber [CA72, GDHZ79, GTF78]. chambers [HDS79]. change [Dod74, HF70]. Changes [CS74, FR77b, GS72a, KA77, LhC73, SEA71, AS77, BG76, Cov76, DP78, DB79, GIK77, Hea73, KGP^{+}78, Kin78, LF76, PH71, PBK70, RB71, RN78, TKK76, WSR79, Car75]. changing [Lan79, Lym73]. Channel [JPW79, San77]. Chant [Val71a]. Chaoboridae [Fed75a]. Chaoborus [Fed75b, FS72a, Fed75a, LaR70, SF75]. Chapel [FW74]. Chapman [Tea78]. Chara [FR77b, Pro71]. Characteristic [BG71, Doy75, KMP^{+}75]. Characteristics [WRO73, AS77, Fed75a, Gar75, HHGRW76, Leh77, Sch76, SB72, Sma74b, TD72, Whi72]. characterization [PL73, Pra70]. charcoal [Coo73a]. charge [NL72, Pra70, WK77]. charts [Dug71, d’A75]. chelation [GV78]. chelator [JM78]. Chemical [BBBS79, Fan77, Gid77, HHGRW76, May76, Mor71, NK72, PBK70, RB71, RAC71, ZY72, All79, AR73, And79, BR72, Bro78a, CM79, CKGS77, CB75b, DLMS76, DG76a, Gar75, GMT^{+}72, Hut70, JS73, LK74, ME72, Sch76, SEA71, SG71, SD78, WSR79, WL73, But76, Gol79, Gro76b, Mil77, Pyt76b, Sch75b, Smi76c]. Chemically [Eme75a]. Chemie [Ril77]. Chemistry [HT70, WH78, And79, CP75, Cul73a, GS72a, GLSD74, Gro74a, Gro74b, IWC^{+}77, Lym73, MB77, MBR78, Par73, Pen73, Ser71, Gro78, O’C71, Stu71]. chemostat [FF79, Gid77, Jan74, MMD79, Per76, TD72, dO74]. chemostat-grown [TD72]. Cheng [Edw77, Koh71a]. Chesapeake [MTT77, TLT77, TS78, Won79]. Chester [Gol79, Mil77, Stu71]. Chief [Epp73]. Chilean [HBLP76]. chimie [Pyt76a]. chironomid [Lod74]. Chlamydomonas [Kro71, Kro72a]. chloride [Sta74a]. chlorinity [AGM70, AFK73, MSH74]. chloro [HH78]. chloro- [HH78]. Chlorococcum [Kro71, Kro72a]. Chlorophyll [Ber72, Ban74b, BP74a, DOP77, Den77, DR74, Fee76, KA77, Lor70, MB76a, Mar77, RSEG78, Rie78, Ril75, SL76, SL75, SH77, TKH75, TCTW75, TKH77, Tho70a, TP74, WD79]. chlorophyll-a [KA77, MB76a, Mar77, SL76, SH77, TKH75, TKH77, WD79]. CHN [Taf76]. Chromatographic [DBM77]. chromatography [MCA77, SM76]. chronology [VBT79]. Chroococcoid [JS79]. chydorid [Gou71]. Cie [Cus76a]. Circulation [Gor78, HES75, MM73b, Ben75, BY75, BG71, CGPD79, Csa75, Kje78, yLKR75, Pic76, PB77, Sha77a, TS78]. circulations [HL73a, Sta71]. City [Ano74b]. civil [Cle72]. Cladocera [HH73, Pre78, Fre71, Gou71, Zar72, DD71]. cladoceran [GMTA79, Thr79, WP78]. cladocerans [ES77]. Clair [Mon75]. clams [BG72, Wol71]. Clarifying [Mur71]. Clark [Hyn73]. class [Zar72]. classes [FH79, OSD76, Sha73c]. classification [SB78c]. classify [Swe70a]. clausi [BGMW74, DD78, Lan75a, NN76, OSD76]. clausii [KVW79]. clavipes [Geh74]. clay [Bas75]. clays [BGZS75]. Clear [HSC79, HJG71, HG72, HDFG72, Hor79, SH78]. clearance [Fro78]. Cleve [YY79]. Cliffs [Mil73a, Mil73b]. clocks [Str75b]. clone [Doy75, Sma74a]. close [Bla79, CRA72, Hes76]. close-interval [Bla79, CRA72]. closed [CR70, Fry71, LJ73]. closing [HBM73, KAG77]. Clouds [Ant77]. clutch [Mal73]. CO [Bro79, Mac79, Taf76, Bow79, Car75, Col72, Dee74, Edm70a, Edw77, Eme76, Eme78, Gus79, Hor76, Hyn71, Hyn78, Ill75, Kra79, Liv72b, Liv76, Liv77b, Par70, Phl73, Sch77a, Smi75, Smi76b, Smi76a, Smi78a, Ste79b, Wal76a, Eme75a, HW78, SR77, TJ79]. Co-operation [Edm70a]. Coachman [LaF77]. coast [BP74b, CL76, FN70, MSR^{+}75, PK74, RAC71, Sha74a, Sha77a, WRO73, WS70, BBPK71]. Coastal [Pai70, AZ70, BFFB75, CJN72, CACB75, Dug71, ERHC79, ERH79, FR76, GV73, Glo78, GC73, GG75b, HC79a, HIJ78, Har78b, Hea73, HH78, Høj78, Hol70, Jør77, KMP^{+}75, LNS^{+}76, Lee77b, MP77a, Man75b, Nor75, RE73, Rem71, Sha77a, Ste79a, Tea78, TP78, TLP78, Tyl71, Whe76, Whe77, Win71]. coasts [Eme78]. coated [HH73]. COBLAMED [Sho74]. coccolithophorid [BL76]. code [Swe70a]. Codium [BA71b, HC75a]. coefficient [Ley73, Swe70b]. coefficients [Fag73, Pyt75]. coelenterates [DM79b]. coexisting [Pou78]. coffeaeformis [And76]. Coffey [Sch77a]. cohort [GB79, LP77]. Cold [JB71, Ris73]. cold-cure [Ris73]. Cole [Kil76]. collapse [Whe75]. collect [HC78]. collected [Cli72]. collecting [GCB76, Jac70, PDF^{+}78]. Collection [PRG78, BS74, BS75, Fuh73, HB79]. collector [MMC79]. College [Ril71a]. colloidal [KH76, Koe76]. Colloq [Bow79, Gus79]. colonies [TKK76]. color [Hai79a, MP77b, PK74, STG73, WK79, Smi75]. Colorimetric [BS71]. Columbia [Har78a, How72, CGPD79, CR70, FS72a, JPW79, PCWR70, Ano79d]. column [BDW76, HHGRW76, ICO^{+}74, Kar78, Wor77]. columns [Cas72a]. com.bustion [MMW71]. Combat [Ker78]. combination [Zir75]. Combined [IJ79, CM75]. combustion [Sal79]. commemorative [Epp73]. Comment [Bro78a, GG81, Pat80, Bal76, Ban75, BS78a, Cus77, DH78a, Edm72d, Edm74b, Edm77, FHV79, LP78, Meg75, Nor76, PT74, PT75, Sha78, SCA79, Tho78, Vly77]. Comments [Dro78, Gar74, LR74b, Lew77, Pyt72, NC78, SK73a]. Committee [Hyn76]. communicating [Man72b]. communis [Wet75]. communities [Cas72a, CMDC78, GD74, GP77, GA75, HCW70, Jef70, Jon78, LL77, Mal71, Mon75, Pat71, RHW79, SWLH78, SS77, WNLS77]. Community [Toe71, Tur78, ATM74, BB77, CPOT79, CW78a, DV78, Dun70, Eag71, EB79, GM74, HD70, Hoo74, JG74, KHC74, Lyn77, Lyn79, Smi73b, SCE^{+}76]. Compact [Whi71]. Company [Val71a]. Comparative [FS72a, KS70, LC74, PM77]. compared [GL75b]. Comparison [CP79, Cli72, Fal75, Har77, KB76, LC71b, Pou78, SH72, Tho77, WN71, DC75, HAG^{+}72, JM78, ME72, MHK75, WC73, YY79]. comparisons [Gar74, HP74]. compartments [SSL78]. compensates [Lan79]. compensation [HW78]. Competition [Rhe72, Lyn77, Lyn79]. compilers [Fra72]. complanata [Str76]. Complementary [Dod70]. complete [PRLG75]. Completed [Ano71f]. Complex [Bra73a, Doy79, FH79, Pro71, MB71a]. Complexation [Eld75, PME73, SL78, SWM^{+}78]. complexes [ZH70]. complexing [LR74a, MM79]. complexity [Fot74]. component [GMCG78, Pin73]. Components [Edm71b, DG76a, HS75, HPM72b]. Composition [BDM77, DB71, DM79b, Por72, All78, All79, BR72, Bri79, BL72b, CM79, CB75b, DLMS76, Dee71, DB79, ECHH^{+}71, Gid77, Hai76, Har78a, Jef70, JW70, LLS75, LB71, Mos70, Rhe78, RN78, SEPS74, SEA71, Ste76, Wei73a, Gor79a, Gor79b, Gus79, Man75a]. compounds [FH79, GC74b, HC78, JKC79, MMW71, RAH77, WM72a]. comprehensive [Tyl71]. compressibility [Pla78]. computer [CSHE78, LBL75, Wie71]. computing [Ric78]. concentrating [MY72]. Concentration [Ber72, BC74, Fro72, MB70b, NWEG75, PE76, Par77, Ban74b, BC72, CS78, DS75, EGTA70, HP77, HDS79, HHPP70, KA74, LF76, LhC73, MSF75, RFFH76, SB73, Tho70a, VB75, Whi75, Win71]. concentrations [CS73a, FL75, GHM78, JM70, KB76, LT71, MP78, SCLM70, Sen78, VHMW74]. concept [BH78, Ids73]. Concepts [Woo79, MLKT75]. Concerning [Woo71, Kil71]. concretion [SC76b]. concretions [HT70]. Conditions [Kro71, CR72, FP79a, Geh74, GS73c, Hea73, KO76, Pec70, Ree74, SMD76, SH73, Pom74]. conductance [Sta74a]. confederata [HG76]. Conflict [HR71]. Congress [Ano72c, Ano73h, Ano70f]. Conifer [Rau78]. Conn [Fan77, Ori73]. Connecticut [War72]. consequences [Bar71, CMB70, Mil70, MB76b]. Conservation [MNM71, Bri72, Fra72]. considerations [WL73]. Constant [Bra73a, HD70, LM70]. constants [Fal75, MCHP73, PIM74]. constituents [BL72a, LB71]. Consumption [SB78a, Pam71, RH77, SWLH78]. contained [KLM74, dO74]. containers [LhC73]. containing [IBM76, Jer70, WD79]. Contamination [Dis73]. Content [Bal72, Sno72, WAA72a, AGF71, CM77, GCL70, Gro79b, Har72, Jan73, Lor70, PI73, PH71, SS74, Ste73b, Whi75]. contents [BSS78]. continental [Eme78, HPS75, Huy77, MC78]. Continents [Gro75b]. Continuous [DBH78, DM71, KHC74, MSJG74, BL72a, BFFB75, Gol77b, Hal71, MO72, MGHR75, MY72, dO74]. continuously [BDW76]. contrasting [Dev79]. Contribution [BA71a, FR76]. contributors [BGZS75]. Control [Lev71a, Wan72a, Gil71, Gou71, HK74, KS78, Mal75, PT73, Pyt72, Tim75, Hyn71]. controlling [LH73b, RFFH76]. convenient [HDS79]. conventional [MHK75]. conversion [Paf76]. cooling [Smi73a]. Cooper [Dee74]. Coordinator [Woo73]. Copepod [Bur70, Fro72, HC79b, MB70b, RC72, AF74, BGMW74, Enr77c, EC72, FP77, Ker77, LK74, LF76, Lam78b, Lan75a, Lan75b, Lan78, LWR73, LL79, LM70, Mal70, Mar70b, MJH77, Mul79, NN76, OS79, OSD76, Paf76, RSMC76, RLS75, Rom78, SB79]. Copepoda [Coo71, Hai79a, Hai79b, KVW79]. copepodite [NN76]. Copepods [Cun72, Enr77a, HR71, NN71, Wil72b, Bay78, Boy76, Dag77, ES77, Har73a, Kat73, Ker78, LH73a, Mal73, Pou78, RC74, Str77, Vly70]. Copper [AM78, ZY72, GV78, KV73, Kim73a, Kim73b, KM73, LWR73, MPN70, MM79, PME73, RMC79, SL78, SWM^{+}78]. copper-chelation [GV78]. copper-complexing [MM79]. Coprecipitation [OW72]. coral [Ano71h, BM74, BB77, CS73b, CB76, DM79b, GS73c, HC79b, JW70, Kin78, KD79b, LLS75, MQM74, NWEG75, PP77, RLS75, RMS77, WDW^{+}75, Koh79]. Corallinaceae [Lit73]. corals [BM74, D’E77, DM79a, JCK70, LT71, MD78, RLS75, Sor73]. Cord [Rei72a]. core [KM72a, SW72, VBT79]. Corer [KSGM71, BB79, Cub70, Mur76, Red75, SH72, TWPC75]. cores [BL74, PE76]. coronatus [SB79]. Corp [Bou75, Kap78, Mac79]. Corps [Cle72]. correction [Ker72]. correctives [Bar71]. Correlation [PSK78, Glo78]. correlations [SR77]. cosmopolitan [Pro71]. Costa [RAC71]. costatum [HS77b, YY79]. Costlow [Par72]. Coull [Har78a]. counterstaining [WW74]. Counting [WN71, IBM76, Loh74, Pug70, Pug73, She79, The75]. counts [DH75, Jon74]. County [WSR79]. cover [PI72]. crabs [Hai76]. Craig [Aro73]. Crane [Eme76, Eme78, Hor76, Kra79, Phl73, Smi76b]. Crater [Lam72, NND71, NND72, STG73]. Creatine [WD72]. Credibility [Wic71]. creeks [ES75b]. Cricosphaera [BL76]. crisis [Kat71]. Cristobal [Rog79]. criteria [FPO79]. critique [KK76a, Van77]. Crop [Por72, FR76, GLSD74, Kro72b, Sch78a]. crops [Cow70, OS78]. crustacea [JSS^{+}71, Zar72, DD71]. crustacean [GO78]. crustose [Lit73]. Cryolite [JF72]. cryptic [EA70]. Cs [KB76]. ctenophores [RWI78]. Cu [EA79, VM77, Wil78]. Cultivation [Gol77a, FF79, Koz70]. Culture [TLR^{+}70, DGEB70, DBH78, Fry71, Gol77b, Per76, YY79]. cultured [CB75b, Ten77]. cultures [Boo75c, CL76, DM71, ECHH^{+}71, Gid77, MGHR75, MMD79]. culturing [dO74]. Cumulative [MB70b]. cuneata [Wol71]. cure [Ris73]. Current [Sho74, CB70, HP76, Mor75, CB70, Edm70a, Ich71, Sil75]. Currents [CT70, FN70, GF70a, Lud74, RMS77, TG72]. curve [CSHE78, Sil79]. curves [Doy75, Gar75, Leh76b]. cuticles [DG78]. cyanobacteria [JS79]. cyanobacterial [JC77, KCS77]. Cyanophyta [Mel79]. Cyanophyte [RO70]. Cycle [DB71, RC72, Cre75, Dee71, GL75a, GS73a, GS73b, GMV72, Jør77, JC77, Sma74a, SS78b, WDW^{+}75]. cycles [CAE78, Mac78b, Wal75, Woo77]. Cyclic [PH71]. Cycling [CR70, ML70, MBN72, ERVM73, KH76, PRLG75, RH78]. cyclomorphosis [Dod74]. cyclonic [Ben75]. Cyclopoid [Bur70]. Cyclops [Ker78]. Cyclosalpa [HM79].

D [Bri72, Bro70, Cul73a, Ebe72, Edm70b, Edm71a, Edm74a, Edm76a, Eme74, Fag72, Fan77, Gol78, Gol74b, Gor79a, Gor79b, Gro71, Gro74a, Gro75a, Gro76a, Gro77, Gro78, Hor76, How72, Ill75, Kin70, Koh79, Leh75, Lun79, Lym73, Par72, Par73, Pom79, Ril71a, Ril78a, Rzo79, Sch77a, Smi74a, Sta74b, Str75b, Val71a, d’A75]. D. [Bal76, NC78, Tho78]. D.C [Bar71, Fer70, Hyn76]. Daget [Mar78a]. daily [MCSK79]. Dam [Sha77a]. Dams [Cle72]. Dana [Kim79]. danica [Aar71, BA76]. Danish [WAA72a, WAA72b]. dansylation [LP70]. Daphnia [Arn71, BR72, BW70, BR74, CB75b, Cro73, Cun72, EP71, HH73, HP77, Kib71, KO76, OV78, PHH78, PR78, Pre78, Sno72, Sta76, Sta78, Sta71, Ste75, WWR70, PR73]. dark [BS77, BL72b, CAE78, ES75a, MYY71a, MYY71b, TJ79]. Darwin [RB71]. Data [BL72b, Pai75b, Ban74c, Chu72, GB79, Mon78, NWEG75, RC74, SA74, Swe70a, Gol74a]. data-some [GB79]. Dating [RMG73, AW75]. Davies [Gro71, Ril78b]. Davis [Fag72, Fra72, Par70, Ton72, Dro78]. Davis/Breitner/Harrison [Dro78]. Davoll [Bro78a]. day [Cli72]. daylight [Che73]. Deacon [Woo79]. dead [GP77, LH73c, OH72a, OH72b, BAHA77, BYS77]. Deason [Kro71, Kro72a]. death [PR78]. debate [Gro75b]. decay [HM75, KTTF76]. decelerating [Leh77]. decomposing [KA77]. Decomposition [Fen70, GK74, Har73a, WM72a, BB73, CA79, GA75, Har72, Hob73, Hyn78, OH72a, OH72b, SK73b, dH77]. Deep [Mil71, Eme78, Far75, FW74, JFH73, KHHG^{+}72, LS77, MC78, ME74, PKC77, RMS77, Sch72, SB78a, SWLH78, SSM76, TJ76, Bou75, HSW72, Ebe72]. deep-living [KHHG^{+}72, SSM76]. Deep-sea [Mil71, Eme78, JFH73, MC78, Sch72, SWLH78, Bou75]. Deevey [Ori73]. defense [PS79]. deficiency [Cav76, Tho70b, TD72]. deficient [CR72]. deficits [Las75]. defined [Sil75]. Degens [Gro74a]. degradation [Rie78, SL75, WD79]. degree [Dil75]. Dekker [Par73, Pla78]. Delaware [GD74]. delayed [SR77]. demands [Smi74b]. demersal [PP77]. Denitrification [KVT79, BBG75, CR76, GC70, GM77, VB75]. Denmark [Jør77]. Densities [Buz71, Con71, CACB75, Ada76a, BSS78]. densitometry [EA70]. Density [Mal73, CM77, HSK76, Hil74, Huy77, KK76a, KK79, MLCG79, PS79, PRG78, SW73, SK72, SA71, Sta78, TGK^{+}78]. Dependence [Sen78, Pac73, Shu78, Smi79]. Dependent [ZY72, SP78, TG77, WP78]. depletion [May73]. deposit [Doy79, Har70, Har71, Nic74, TSJ78]. deposit-feeding [Har70, Har71, Nic74]. Deposition [MH74, GL79]. deposits [Hak76, Mur76, SL70]. Dept [Zar78]. Depth [ATM74, AFK73, Chi71, Lor70]. depths [Jer70]. derivatives [BDM77]. derive [RC74]. Derived [Fen70, LI76, Ten77]. describing [SW72, Wil78]. description [Ban79, Val71b]. Desert [Nai76]. Design [Whi71, Boo76, MM73a]. designed [Sch77a]. designs [WL79]. destruction [Dun70]. Desvaux [Pro71]. Detailed [Ste73b, Gag76]. Details [EH77]. detection [FR77a, PG72]. determinants [Pec79]. Determination [Bra73a, CB72, CS73a, CB73, DSC77, GHM78, JM70, SRR79, SHW^{+}70, WBM^{+}77, Whi71, Ada76a, BS71, Car70, Deg73, DL74, FR77a, HD77, JW70, JS75, KK76a, KK76b, LTB75, MPN70, MPN72, MY70, Rie78, Sal79, Sta74a, SWM^{+}78, WWH70, WD79]. Determinations [SH71, CW78b, SH77]. determine [BFM76]. determined [AW75, CACB75, KB76, MMD79, SS78a]. Determining [WWR70, BS74, Bro79, Bru71, GL75b, Hoi78, Lam74, Ree74, SB73]. detrital [Pae74]. Detritus [Fen70, HM75, CL79, FPO79, GM74, Har72, PS73, Ten77, THDW79, WBM^{+}77, dlCP75, Hob74]. d’etude [Pyt76a]. Development [Bur70, Edm70a, GC72, Las75, And71, Coo71, Eag71, Lan75a, Lan75b, Sch72, Sta76]. developmental [Swe78]. developments [Smi78b]. Deviations [BHK78]. Device [NN71, Smi71, BL76, Bru71, FF79, Fuh73, FF75, HC78, Lee77a, PN79, Ped74, Ree74]. DeW [Ril78b]. Dfl [Edw77, Liv77b, Sio77]. diagenesis [Cha74, MG78, Ros79a, VWB77a, VWB77b]. dialysis [May76, SRR79, YY79]. diaptomid [Mal73]. Diaptomus [Coo71, Geh74, GC72, Hai79a, Hai79b, Mal78]. diatom [DBH78, FP77, GM78, HS77b, HS77c, KK78, PE76, Pat71, Per76, Pug70, Sma74a, SK73b, TD72, YY79]. Diatoms [HK73, Bat73, Bra71, CAE78, CL76, DBH78, Dro78, HB72c, HC75b, Kil71, MMD79, PRJR73, Sma74b, WM72b, Lun79]. dictionary [Edm72b]. Diego [Smi74b, Zar78]. Diel [AS77, DG76b, MP77a, Mac78b, RS79, Sta78, Hor79]. diet [Hu78]. diets [Fed75b]. difference [SM76]. Differences [CB75b, HB72c]. Different [Kro71, Jer70, PP77, SKF74, Ten77, THDW79]. Differential [DB73, Mul79, Ber70a, BP72]. differentiation [SC76b]. Difficulties [Mon78, SOHH76]. diffraction [Str75a]. diffusion [BH75, IE78, JP75, KD79a, Mur72, ST78]. diffusivity [DL74, Swe70b]. digestion [DSC77, FM78]. dilute [DM71]. dimethyl [SL76]. dimorphism [Bay78]. Dinobryon [Leh76a]. dinoflagellate [EC72, Glo78, Har73c, HSK76, KS78, Mac77, RS79, TS78]. dinoflagellates [BS77, PRG78, SCLM70, SSM76]. Dioxide [WAA72a, Bra73b, CK74, GF75, JPW79, Kel70, MYY71a, MYY71b, PCWR70, SS74, Smi73b, SK75, Ver75, Whi74b, Whi75]. Diptera [Fed75a, Lod74]. Direct [DH75, HLH71, HMWW72, IBM76, Jon74]. Direction [Ano71f]. Directorate [Edm70a]. disadvantages [KK79]. disappearance [GMTA79]. disasters [Cle72]. Discarded [All76]. discaudatus [AF74, Mul79]. discharge [Boo75b]. discharges [Cow70]. discontinuity [GWN70]. disk [Hol70]. Dispersal [GC73]. dispersion [BHSJ78, RMG70]. dispersive [Bas75]. Dissertations [Ano71f]. dissociation [MCHP73]. Dissodinium [BS77]. dissolution [BP72, Joh76]. Dissolved [Bun71, CJN72, Fie72, LB77, Mey76, MF71, Nri78, Ogu70, Whe77, AGA78, AZ70, BLL70, CS70, CC75b, D’E77, Dux75, EK73, HS74, HC78, HL73b, HK79, KLM74, KLG71, Koe76, Lam78a, LP70, LI76, MMW71, Mar74, MSJG74, OH72a, OH72b, OW74, Pae74, PS77, QM71, RH75, SB78a, Sha73b, Sor73, SS78b, WM72a, MB77]. distance [FR77a]. distinct [FH79]. Distributed [Hyn78]. Distribution [DCP73, FP79a, GS73c, Kar78, KM72a, LH72, Nat70, ATM74, AH77, BL74, BH75, BFMF72, BS78c, CP77, CKS77b, CH70, Dal74, Daw78, DEWN76, DB76a, DE71, ERHC79, Fee76, Fou71, Fre71, Geh74, Gou71, HES75, HHHP^{+}72, HC75b, Hyn74, LK74, LH73a, MM75, MG75, MM73b, MC78, OV78, PHR71, RSEG78, Rem71, Sch71, SPS72, SC70, Sta71, Ste73a, SW72, Thr79, TG77, VB75, Ven74, Wie70, CKS77a, Phl73]. distributional [MB76b]. distributions [KMP^{+}75, MCB79, Pec79, RH77, Sta71, TS78, Van77]. disturbance [Bla79]. disturbances [ESH73]. disturbed [VWB77a, VWB77b]. Ditylum [AS77, PG75]. Diurnal [Enr77b, EH77, GF70b, OT77, Til73, Toe71, WCR71, Beg76, KZ77, PM77]. diurnally [Swe76]. divers [GT74, Ham75]. diverse [JS79]. Diversity [Koh71b, Cov76, SG70, WF72, Koh77]. division [CAE78, YY79]. DM [Fot70]. DM46.80 [Edm76b]. DM97.50 [Wet75]. DNA [SM79]. Do [Sha77b]. Doctoral [Ano71f, Koh71c]. Documentation [Pyt76a]. dodge [Che73, CE73]. Dolphins [Sch77a]. dominance [MMD79]. dominated [KC79, Mac78b]. double [The75]. Douglas [GF70a]. Dowden [Eme74, Par75, Woo79]. down [Enr79]. downwelling [DB79]. Dr [Ham79, Man79, Rzo79, Wu79]. drag [Leh77]. drainage [Boo75b, Kir75b, San77]. drift [All78, BB76, Con75, GB72c, Hil74, MHK75, WS70, Wun73a, Wun73b]. drift-bottles [MHK75]. drift-envelopes [MHK75]. Drifters [Fol71]. drifting [SRSW78]. driven [ICO^{+}74, LH78]. droebachiensis [Fie72]. drogue [MM73a]. drogues [CB70]. Droop [DH78a]. drop [BS78b, Mac78a]. drops [BS75]. dry [McL74, McL75]. Dual [Edm77]. Dubuque [Sta72a]. Dulce [RAC71]. durable [Lan79]. during [And76, Beg70, BB73, Cli72, FP79a, Gar75, Har73c, HSK76, HS77b, JP75, PBK70, Sho74, SK73b]. Dutch [GV73]. dye [Hal71]. Dyer [Mof73]. Dynamic [DVCP79]. Dynamics [SR77, BM75, Dev79, ERVM73, GMTA79, HCW70, KK78, LL77, LBL75, Mey79, Smi73b, VTP76, WWB^{+}78, Ste79b]. Dyrssen [Lym73].

Early [MG78, Omo79]. earth [AFK73, Gro75b, Val71a]. East [Ich71, MSR^{+}75, BBBS79, KH73, LK74, MK74]. East-West [Ich71]. eastern [Bla73a, CR72, CR76, CP75, GWN70, HP70, KD79a]. Eaton [Liv78]. eaux [Pyt76a]. ebullition [ST78]. echinulata [RO70]. eclipse [BFMF72]. Ecol [Hob79, Hyn78]. Ecología [Ori77]. Ecological [Koh74, Leh76a, Man76, Ori73, Pom79, RO70, Bay78, Bla73a, FM78, KH73, Man72b, MH72, RMG70, SG70, TK76, Wal76a, Woo77, Koh77]. Écologie [Cus76a, Mar78a]. Ecology [Ano72c, Enr76, Pet71, TLR^{+}70, Wic71, WF72, Ban73, Cum71, Fre71, Har73b, HK73, Hob79, Hut70, Jan74, Jef72, LK74, Lau74, LeC77, ME74, Phl73, Tyl71, Val74b, Har78a, Day79, Edm76a, Gol77a, Hed70, LeC77, Mas70, Rig76, Sta72a]. Economic [Edm70a]. Ecosystem [MBN72, KVT79, Mey79, SSL78, SP78, Ste79a, VTV^{+}78, WKW^{+}74, WWK^{+}77, WWHH77, WQ73, Liv78]. Ecosystems [Ano72c, HC79b, Hyn76, Pae78, Hob74, Tea78, Wal74b]. Ed [Ban73, Bou75, Bow79, Cha74, Dee74, DH72, Edw77, Eme74, Fan77, FW74, Gol77a, Gor79a, Gor79b, Gor78, Gus79, Har78a, Hed70, Hob79, How78, Hun75, Hut72, Ich71, Kap78, Kel75, Kin70, Koh71a, Koh74, LeC77, Lun79, Man75a, Man70, Mas70, Mil78, Ori73, Pai75a, Par70, Par75, Par78, Rig72, Rig76, Ril71a, Ril73, Ril77, Sio77, Smi78b, Smi78a, Tea78, Tyl71, Woo77, Woo79, Zar78, But76, Epp73, Gro76b, Mil77, Par72, Pyt76b, Smi76c]. Edaphic [MH72, GD74]. edax [Con71]. eddy [IE78, JP75, KD79a]. Ediciones [Ori77]. Edinburgh [Bro70, Edm72a, Fag72, Fra72, Man70, Ton72]. Edition [Gol79]. Editorial [Edm72d, Edm74b, Edm77]. Edmondson [Ton72]. EDs [Lym73, Bri72, But76, Edm74a, Fag72, Gol78, Gol74b, Gol79, Gro75a, Gro76b, Gro77, Gro78, Gro79a, Hob73, Hob74, Hor75, How72, Hyn73, Hyn78, Koh79, Koz70, LaF75, Lau74, Leh75, Mac79, Man73a, Man76, Mil77, Pom79, Pyt76b, Ril78a, Sch73, Sha73a, Smi76c, Smi76a, Ton72, Gro74a]. EDTA [MB71a]. Edwards [Bri72]. Eelgrass [MBN72, HM75, MB70a, PS77]. Effect [BL79, Bur70, CM77, EC72, GC72, Gie70, Kes72, Kib71, LaR70, MB71b, MB70b, Sha77a, Sno72, SL78, Tho70a, WWR70, dH77, Bra73b, CM79, CNC73, Coo71, CSHE78, GC74a, Har70, HP77, Hoo74, KA74, LM70, LJ79, MYY71a, OV78, Ram76, RH77, RMC79, Swe76, THDW79, Whi75]. effective [Owe74]. Effects [Bat76, BFMF72, CA79, DE71, Fro72, HC79c, KD79b, MCB79, NN76, PA70, Por72, Rhe78, SR79, TR76, War72, Wel70, WK77, AS77, Bel78, Car70, CS70, Dag77, Dav74a, ESH73, FG75, Gol77b, HM75, LL77, Lan75b, MPN70, MPN72, Pat71, RH78, Wie71, Gor79a, Gor79b]. efficiencies [DB76b, HM79, NN76]. Efficiency [HHPP70, Bla73a, HB79, NN79, Tyl75, ME74]. efficient [HW76, RG76a, Rup72, TTW75]. effluent [DM71, WF72]. egg [Thr79]. Eggs [Bur70, Coo71, JSS^{+}71, Kim79, Lan75b]. Egyptian [Sha74a, Sha77a]. Eh [GKA73, Whi71, Whi74a]. Ehrenb. [Leh76a]. Eises [Fot70]. electrical [Sta74a]. electrochemical [Whi72]. electrode [GHM78, GTF78, Whi74a]. electrodes [Man72a]. Electromagnetic [KL75]. electron [Bas75, CP79, HMWW72, KP75, PHR71, PS73]. electronic [Cul73b, Pen73]. electrophoresis [BL72a]. electrophoretic [AVZ75]. Electrostatic [BS75]. elegans [Zo73]. element [And79, LT71, TKC73]. elemental [GG75b]. elements [BG72, Cre75]. elevated [KD79b]. Elliott [Cas71]. ellipsoideum [Kro71, Kro72a]. elntriation [WB74]. Elodea [FR77b]. Elsevier [Bow79, Car75, Gus79, Liv77b, Smi76a, Ste79b, Tea78, Wal76a]. Eltringham [Koh72]. elutriation [Ped74]. embayment [TP78]. embayments [Cha79a, Hea73]. emergence [SA74]. Emission [NC76]. emphasis [Beg70, GMV72, Wal76a]. Enclosed [HDS79, LhC73]. enclosure [Lan79]. encyclopedia [Sch77a, Edm71a]. encyclopedic [Edm72b]. Enderby [Mac70]. Energetically [HW76]. Energy [BM74, Pai71, SP78, Vly70, Bas75, Doy79, Har71, MS74b, SSHV76]. energy-dispersive [Bas75]. energy-optimization [Doy79]. Enewetak [WDW^{+}75]. Eng [Cul73a]. engineering [Woo73]. Engineers [Cle72]. England [HT79]. Englewood [Mil73a, Mil73b]. English [Edm76b, Koh72, GLSD74]. enhanced [Eme75a]. Enhancement [Coo73b, GL79]. enigma [PR78]. Eniwetok [WDW^{+}75, GM74, Smi73b, SM73]. enriched [SM74, VTP76]. Enrichment [EA79, GL75b, Boo75c, FHV79, LL77, SG76, TR76]. enrichments [SRSS74]. Enright [Vly77]. ensembles [Owe74]. enthalpy [Con70]. entrainment [Bla73b]. entrance [TC74]. Entropy [Pla78]. envelopes [MHK75]. Environment [Pai70, BBPK71, HWL70, Lau74, TLP78, WRO73, Edm72b, Gol78, Hob79, Ste70]. environment. [Man73a]. Environmental [Ban73, Gol74a, Gou71, HK74, Mal75, PT73, SC76b, Tyl71, GCL70, Gro77, Kat71, Mil70, Edm74a, Gol74b, Woo77]. environments [All76, And75, BD72, Dag77, Fou70, Fou71, Fou73, Gor79a, Gor79b, Gro71, LC74, MB76a, MG78, RH77, SK75, Kra79, Par70, Sta75]. enzymatic [ME72]. enzyme [DP78]. enzymes [GM79]. Ephemeroptera [MBB73, MMC79]. Epibenthic [ES77, Lea70, VVT76]. epicontinental [Cha74]. epifaunal [DV78]. epifluorescence [DH75]. epilimnetic [LR74a]. epilimnia [Bla73b]. epiphyte [CC74]. epiphytes [CK79, GP72b, GG81, PS77]. epiphytic [Bro74, FR77b]. Epischura [Ker78]. equation [BHK78, SP78]. equations [Ban74b, War78]. equatorial [BP70, LB77, Mel79]. équilibres [Pyt76a]. equilibria [HK79, WK77]. equilibrium [PIM74, Wil78]. Erie [Nri78, NCK79, TTC^{+}77]. Erken [Pec70, Cus76b, GL71, Kno77, Lew77]. erotic [Koh73]. Errata [Ano73e, Ano75h, Ano75i]. Erratum [Ano74c, Ano75j, Ano75k, Ano75l]. Error [SD78, Mcm73, Wie71]. errors [LB79]. erythraea [YYP72]. especially [Bat73, Bro78b]. essay [Sta75]. essays [Ori73, Hyn73]. esters [LH73a]. Esthwaite [Gri78]. estimate [BYS77]. estimated [TP73]. Estimates [KD79a, CSHE78, Fee76, MCB79, Van77]. Estimating [Ada76b, Thr79, WN71, Ben79, CP73, Enr77c, GV73, SA74]. Estimation [BB76, Fag73, TKH77, AGF71, Bat73, CTB72, Gre79, MB76a, Mar77, Rig72]. estimations [Jas75]. estuaries [Bow79, Kje78, Lee77a, Lee77b, Man75b, Mof73]. Estuarine [HR71, Bat76, BG71, BBPK71, BH70, Con75, HWL70, HC79a, Hoi78, LC74, Man73a, MH74, OSD76, Sha77a, WBM^{+}77, Win71, WM72b, WQ73, Gor78, Gro78, How78]. Estuary [MG78, Wol71, CCK71, Dux79, Fer70, FMJC72, How72, LH72, NS78, Tur78, WEMD72, Won79, BG71, MEW^{+}78, Gor78, How78]. Ethiopia [BG70]. Ethiopian [BWP73]. Eubosmina [DD71]. eudiaptomus [Ton71]. Eunice [FS72a]. Euphausia [AVZ75]. euphausiid [FS72b]. euphausiids [BSS78, Hu78]. euphotic [BSB77, Lor70, WH71]. Europe [Edm70a]. eurycercus [Fre71]. eutrophic [CA79, Eme75a, GV73, RH78, SS78b]. eutrophication [CCK71, Edm70a, Fee79, HB72c, Imb74, Leh75, Pat80, Sch78b, Edm70a, Hut72, Bar71]. eutrophy [Dil75]. evacuated [JS73]. evaluation [Boo76, EK73, Jas75, Kir75a, LB79, Lit73, MCA77, Pug73]. Evelyn [Ano71c, Ano71f, Ori73]. events [WWB^{+}78]. Evidence [Kat73, NK72, RMS77, TH74, PRLG75, Bra71, Fou73]. Evolution [Eme78, Liv72b]. ex [KS78]. exalbescens [Wil72a]. examination [HMWW72, Kir75b]. exchange [Eme75a, Eme75b, JPW79, Sch75a, SR77, SWLH78, Sta74a, WS70]. Exchanges [WWHH77, Mor71, SEPS74, VTV^{+}78]. excreted [MWW77]. Excretion [Aar78, Fog77, PS77, Sha77b, SKF74, Big77, Pet75]. Exeter [Hyn73]. existence [TL75]. exoskeletons [Mar70b]. expansibility [Pla78]. expatriation [Ber70a]. expendable [JFKG72]. expenditure [Vly70]. Experimental [Har78b, Kro71, FR77a, FW74, Geh74, HCW70, Lit73, Lyn79, WM72a, Fee76, Pai75a]. experimentally [VTP76]. experiments [Cul73a, Fee79, GL75b, Koe76, Pam71, Pat80, PBK70, Sma74b]. explanation [Ben75, MSR^{+}75]. Exploitation [Par78]. Explorers [Col72]. export [Kir75b, MK79, Nai76]. exposed [TP78]. exposure [Gar75, PH71]. expression [LP77]. Extensive [May73]. Extracellular [Aar71, Kro72a, AP76, BLM74, SS78b]. extract [LWR73]. Extraction [CMDC78, LC71b, MLKT75, BPW76, Bul78, CS73a, Gar77, KV73, RG76a, SL76, dBSA74]. Extrusion [SN71]. exudates [SWM^{+}78].

F [Cul73a, Eme78, Fag72, Fra72, Gar74, Gro71, Liv72a, Liv72b, Liv78, Mil78, Par70, Par73, Pom79, Smi75, Smi76a, Ton72, Whe73]. F. [Bal76, BS78a]. f.120 [Ham79, Wu79]. f.140 [Rzo79]. f.20 [Man79]. face [Ebe72]. Factor [LC71a, MP78, Wal76b]. factorial [SRSS74]. Factors [Dal74, RFFH76, Sch75a, Sch78a, Ban73, BM73, Edm70a, GCL70, LL79, Mal78, MH72, TP78, Tyl71, WEMD72]. Failure [BGMW74]. Fairbanks [Gro75a, Gol74a]. Fairy [Kac71]. Falkenborg [Leh75]. fall [Smi73a]. falling [CWM75]. Fallout [Wol71]. family [Mil71]. fans [Gri72]. fast [GM78]. fast-growing [GM78]. fasting [SR79]. fate [Gor79a, Gor79b, LI76, Mac79, MC79]. Fatty [Jef72, MQ71, Jef70, MQM74, Mey76]. Fatty-acid [Jef72]. Faulkner [Ill75]. fauna [Chu72, Cli72, CH70, HF70, McL74, McL75]. Fay [Sta74b]. Fayetteville [Lud74]. Fe [Dav77, KLAS73, LS77]. feasibility [Hal71]. features [Beg70, BG71, Liu70]. fecal [BK72, FP77, FS72b]. Fecundity [Hur73, Sch72]. fed [Arn71]. Fee [Pat80]. feeder [Doy79, Leh76b]. feeders [TSJ78]. Feeding [Fed75a, Fro72, Fro77, Kib71, LL73, Lev71a, Paf76, AF74, BR74, Boy76, BCK72, Dod70, DG76b, Egg78, Fro75, Gro79b, HMA^{+}75, HG76, HM79, Har70, Har71, KO76, Lam78b, Lan78, Leh76b, Nic74, Omo79, Pet75, Pou78, Sor73, Sta78]. Ferguson [Cul73a]. Fern [Wil72a]. ferromanganese [HT70, SC76b]. Ferrous [DH78b, HD77]. Fertility [Par72]. Fiber [LC71b, She72]. fiddler [Hai76]. Fidelity [JC75]. Field [Ada76a, GB72b, MB79, BHD76, BH75, CL76, Cul73b, Fag73, GB79, HPS75, Pen73, RC74, Sha73b, Sta72a]. Fifty [HF70]. fight [Val71a]. Figs [Edm70a]. filamentous [Bro78b, Cas73, WP78]. filled [PN79]. filling [KC79]. filling-box [KC79]. film [BS75, DLMS76, Gal75]. films [CACB75, GH79, Whe75]. filter [Boy76, HM79, HMWW72, HHPP70, IR73, Leh76b, Mcm73, Pug70]. filter-feeder [Leh76b]. filter-feeding [Boy76, HM79]. filter-pump [IR73]. filterable [SS77]. filtered [PS73]. filterers [WP78]. Filtering [Cro73, EP71, HP77, LF76]. Filters [LC71b, EG79, GS74, QM71, She72, The75]. Filtration [GS74, BHD76, Har77, HDS79, JDL78, MPN70, MPN72, MCA77, MY72]. Financial [Ano74a, Ano75a, Ano76a, Ano77a, Ano78b, Ano79a, Ano71b, Ano72a, Ano73a]. Findley [Ric79]. fine [BL79, HL73b, LR72]. finite [Cas72b]. finmarchicus [Bal72]. Fire [Wea79]. first [MG75, Ano72c]. first-year [MG75]. Fischer [Sta72b]. Fish [d’A75, CB75a, JSS^{+}71, Mil73c, Ore79, Fry71]. Fisher [SCA79]. Fisheries [Fer70]. fishes [BM74]. fitting [Sil79]. five [KK78, Pou78]. Fixation [Ano78h, Bun71, HDFG72, Bro78a, BBPK71, CPOT79, CM75, CP77, CVV78, GP72b, GL71, HC75a, HSC79, HG72, Hor79, KB71, MGC73, Mil78, MYY71a, MYY71b, PM77, Pre78, RFFH76, RB70, SA70, Sha73b, SD78, TVB79, TL76, VHMW74, WWH75]. fixed [BS70, BFM76]. fjordlike [FMJC72]. fjords [Bow79]. Flaco [HBLP76]. flat [SM73, Smi77, Smi73b]. Flax [WWH75, WWHH77]. Flemming [Smi78a]. flexilis [HW78]. Flint [Liv72a]. floating [DL74]. flocs [CS73b]. flood [Sha77a]. Flooding [Hoo74]. floor [Eme78]. Florida [BBPK71, CB70, KB71, MSR^{+}75, SB72]. Flow [Pai71, RW71, Bru71, DE71, Hal71, HDS79, HHPP70, KL75, LJ79, Taf76, TGK^{+}78, Wor77]. Flow-Cell [RW71]. flow-through [Wor77]. Flower [BBBS79]. flowing [Edm70a, Ree74]. flowing-water [Ree74]. flowmeter [LV76]. flowmeters [HB72a]. flows [LH78]. fluctuating [Bot75, Jon78, Swe76]. fluctuation [DOP77, GGW78, MB71b]. fluctuations [BL79, CS74, DW72, Kim73a, Kim73b, OT77]. Flügel [Smi78b]. fluid [Kra79]. Fluorescence [Ber72, Bro78b, Gro79b, Jon74, MY70, SH77]. fluorescent [CA72]. Fluoride [BY75, KH73, Win71]. fluorometric [JS75, Tit75]. Flushing [Hea73, Dil75, Dux79, SG71]. flux [ES75b, HC78, Lee77a, SP78, SRSW78]. fluxes [GG75b, VTV^{+}78]. Fogg [Sta74b]. Folio [Pai70]. foliosus [Geh74]. following [FR77b, McL74]. Food [Fro72, MB70b, All76, Dag77, DEWN76, Fro77, Hil74, KM72b, LF76, Lam74, Man70, ME74, MSF75, Paf76, RWI78, RG76b, Sta78, WWB^{+}78]. forager [Leh76b]. foraminifer [Mul74]. Foraminifera [Ber70a, BP72, BHSJ78, ML70, Sch71, WL78]. Foraminiferal [Phl70]. foraminiferids [Phl73]. Forbes [Sno72]. forested [Liv78, MK79]. form [Gri72, MC78]. Formation [Whe75, Ada76b, And76, HM75, Joh76, LH73c, MB71a]. formosa [TKK76]. forms [KM72a, SKF74]. formulation [JP76]. Fosforescente [CMB70, SCL^{+}71]. Fossil [GS72b, BDM77, DBM77, Mac79, Pai75b, SG72, Smi78b]. Found [Zar78]. Four [GL75b, Bri76, BL72b, Gar74, HP74, Hu78, SN75]. fraction [Wor77]. Fractionation [Whe76, Kro75]. fractions [MCB79, MC79, SSL78]. fragile [BA71b, HC75a]. fragmentation [HM75]. Francisco [But76, Dee74, Gor79a, Gor79b]. Free [BCB71, CJN72, CC75b, Sch74, Sch77b, Sho70, SCE^{+}76, SRSW78, TVB79]. free-drifting [SRSW78]. free-living [TVB79]. freely [CWM75]. Freeman [Dee74]. freezing [SW72]. French [Edm70a]. frequency [MCB79]. fresh [Hob73, MPN72, Ton72, Gro79a]. Freshwater [Cas71, HJG71, RE73, BR72, BA76, CB75b, EG79, Far75, Fro78, HCW70, HS74, HT70, Jon74, JS75, JDL78, Kil71, KK78, LC74, LR74b, LH73c, MPN70, Mos70, OL74, OW73, Pae74, Ram76, RH77, RHW79, Sma74b, Sta74a, Tit75, TK76, Wri75, Pom79]. freshwaters [GC73, JS73, LP70, Sch78a, Wil78]. fringing [Lit73, Mey76]. fringing-reef [Lit73]. frits [EG79]. frosted [Pre78]. frustules [FP77, PE76]. Fucus [TR76]. fuel [Mac79]. Fugitive [Hur73]. Fulvic [CW78b, dH77]. Function [Ano72c, CS78, Har72, Ley73, MSF75]. functional [HPM72b]. Functionally [FH79]. fundamentals [Edm70a, Ste79b]. funerralis [Pai71]. fungi [CACB75]. Further [Hyn74, WWK^{+}77]. fusiformis [BS77, Sil75, SSM76]. future [Bax77].

G [And71, But76, Dug71, Edm72a, Gro71, Gro76b, Ham79, Hob79, Hut72, Kil76, Kra79, LaF75, Lau74, Liv76, Liv78, Ori73, Pai70, Pyt76b, Rig72, Ril78b, Smi76c, Sta72b, Sta74b, Ton72, Wet76, Wu79, d’A75]. G. [Ano71c, Ano71f, Ban75, Koh71c]. Gabe [Bro70]. Galapagos [RB71]. galeata [HP77]. Galilee [Ber74, DB76b]. Gambusia [GBB72]. Gap [Wic71]. Garden [BBBS79]. Garrels [Liv72b]. Gas [Eme75b, Coo73a, Eme75a, GHM78, SA70, Sch75a, Wal73]. gas-sensing [GHM78]. gases [Hun75, Mar74, MB77, RH75]. Gastropod [Pai71]. gauge [Tem73]. Geiger [WN71]. Geiger-Mueller [WN71]. gel [MCA77]. gelatinous [Big77, HMA^{+}75]. General [Buz71, Hyn76, Ban74a, CGPD79, Eme75a, Man72b]. generated [BG76]. Generation [SS78a]. gently [GCB76]. genus [MJH77]. geochemical [Cre75, HK73, IWC^{+}77, KH73, SN75]. Geochronological [KLAS73]. Geographical [Pai70, LH73a]. Geol [Gro71, Woo79]. geological [Ano70f, Eme78]. Geologists [Gro74a]. Geology [Gro74a, Eck70, Eme76, Gro74b, Kel75, Liv72a]. geomagnetic [San77]. geomorphology [Eme76]. geophysics [Boo75a]. George [Bur70, Lau74, Sch77a]. Georgia [Whe76, Win71]. Geothermal [Pug77]. German [Edm70a]. gerrid [DG78]. Gerridae [LC74]. Gerris [LC74]. Gewässer [Sta72b]. Geyser [MB71b]. giant [Jac77, KW77, TP73]. gibberum [Ste73a]. Gibbs [Man75a]. Giesbr [Lan75a]. Giesbrecht [BGMW74, KVW79]. gigas [RSMC76]. Gjessing [Sch76]. glacial [LH72, Liv72a]. Glass [LC71b, EG79, She72]. Glass-Fiber [LC71b, She72]. Gleotrichia [RO70]. Global [BY75]. globosa [Kro71, Kro72a]. globularis [Pro71]. glucose [JS75, Ram76]. glycine [WNLS77]. glycolic [Wri75]. glyptotendipes [KA71]. Gnotobiotic [TLR^{+}70]. Goldberg [Fan77, Gro76a, Ril78a]. Golfo [RAC71]. Golterman [Liv77b]. Gonyaulax [AM78, Mac78b]. good [Val71a, Pom79]. Gordon [Par72, Sta72a]. governing [BM73]. grab [SH72]. gracilis [Rom78]. gradient [BL76, MH72]. gradients [CWM75, KZ77, PRG78, Pug77]. grain [KK76a, KK79, PS79]. grains [DB73, Dav73]. Grand [Mey76]. Grane [WAA72a, WAA72b]. Grant [FW74]. Grass [Fen70, Rei72a, GK74]. grasses [GP72b]. Grasshoff [Ril77]. gravel [O’C74]. grazer [FG75]. grazers [MSF75]. Grazing [Cus76b, Han71, Por72, Pou73, Coo73b, GMV72, Kno77, Lam78a, Lew77, Mar70a, MB79, NN76, PHH78]. Great [Lan70, BH75, Mor71, Mos76, SG76, SE78, Wei73a]. Green [Lud74, Man73a, YYP72, And75, Arn71, Bro78b, CC74, Gri78, JC75, MB71b, OH72a, OH72b, Rom78, SH73, Sta74b, Wal73, WF72, HC75b, VHMW74]. Greenwich [Fan77]. Greifensee [IE78]. Greville [YY79]. Griffith [Edm74a]. Grindavik [DW72]. groove [RMS77]. Gross [Ril78b]. groundwater [LG79, Pet78a, VTV^{+}78]. Group [Liv77a]. grow [BT70]. growing [CK79, GM78, GG81]. grown [TD72, YY79]. Growth [Boo75c, BB70, Kro71, Kro72a, MJH77, MB70b, Omo79, PA70, Ten77, Zo73, Ban74a, Ban79, Bot75, BG76, CSHE78, DBH78, DDS78, ECHH^{+}71, ERHC79, GC74a, Gol77b, GM78, Gri72, JM78, KOW79, KD79b, MGHR75, MS74a, Mel79, Paf76, PRJR73, Rhe78, SM79, She79, Sma74a, Str75a, Str76, SSM76, TeI70, TR76, WEMD72, Ori73]. Guatemala [Cov76]. guide [Day79, Pom74, SG72]. Guillén [Dug71]. Guinea [Mac77]. Gulf [BBPK71, Cha74, GS73c, d’A75, BBBS79, KA74, MHU72, SSA72]. gun [WMM74]. Gustav [Sta72b]. gut [Gro79b]. Gymnodinium [MSR^{+}75]. Gypsum [HBLP76]. Gypsum-bound [HBLP76]. Gyre [Ven74, ERVM73].

H [CPOT79, Cha74, Cum71, Edm70b, Fra72, Fry71, Hob73, Koz70, LR74b, Leh75, Liv77b, Liv78, Man70, Man76, Mil78, Ril78a, Sha73a, SK73a, Wal74b, Zar78, d’A75, ES75a, SS78a]. H.-P [Koz70]. habitat [Sil75]. habitats [All76, JC75]. habits [BG78]. Hague [Ham79, Man79, Rzo79, Sio77, Wu79]. half [Fal75]. half-saturation [Fal75]. Halifax [WSR79]. Hall [Mil73a, Mil73b]. Halobates [Che73, CE73, LC74]. Haloragaceae [Hut70]. Halsted [Hyn78]. Hamburg [Koz70]. Hamden [Ori73]. Hamilton [CSHE78, PP76]. Hampshire [VBT79]. Handbook [Fag72, Fra72, Hob73, Man72a, Par70, Ton72, Edm74a]. Hansen [Omo79]. Hanson [Par75]. Harbor [DW72, PP76, BCB71]. Harcourt [Sch75b]. hard [CA79, LI76, ST78, WM72a]. hard-water [CA79, LI76, ST78, WM72a]. Hardcover [Ril71a, Koh72]. Hardy [Ano73g]. harpactacoid [Rom78]. Harrison [Dro78]. Harvard [Wal74b]. Harvey [Hyn73, Kra79]. Hatcher [Gar74]. Hawaiian [GMT^{+}72, Lit73]. Hayq [BG70]. headwater [Hoo74, Mey79]. health [Gro76a]. healthy [Sha77b]. heat [AL78, Ste73b, Tim75, WS70]. heavy [BK72, Wor77]. Hecky [PT74]. Heezen [Ebe72]. Heidelberg [Boo75a, Cha74, Hob79, Koh74, Liv78, Smi74a, Ste79a]. Helgoland [Koz70]. Helgoländer [Koz70]. helgolandicus [MB70b, Paf76]. helium [TTC^{+}77]. Hellawell [Bri79]. Helmholtz [TH74]. herbicide [FR77b]. herbivore [Coo73b, SL75]. Herbivorous [Pai71, BG78, NN76]. Herbivory [Wal76b]. Herman [Kel75]. Herrick [GC72]. heterocysts [HG72]. heterogeneity [DV78, MB76b, TP78]. heterogeneous [Wil73]. heteropsis [Bel78]. Heteroptera [LC74]. Heterotrophic [PG72, SSL78, HP70, RHG73]. Heterotrophy [Sau72a, HA72, Mon78]. Hickling [Smi75]. High [Sha77a, Wei73b, Enr77a, GF75, Mal73, Sal79, SCLM70, Taf76, Til73, RC72]. high-speed [Enr77a]. Hill [FW74, Gro75b]. Hissel [Pyt76a]. Histioteuthis [Bel78]. historic [SN75]. History [Col72, And71, DBM77, SG72, Sta78, Woo79]. Hitchcock [NC78]. Holberg [DO75]. hold [PN79]. Holdgate [Hed70, Mas70]. Hole [Gro74b]. Holland [Edw77]. Hollister [Ebe72]. Holme [Fag72]. Holocene [Cha74]. holography [BKS70]. holomictic [Lud76]. Holopedium [Ste73a]. homing [Enr76]. Hong [FT73]. Honolulu [Ich71]. honor [Ori73]. Hood [Gol78, Gol74b, Gro75a, Ril71a]. Hopton [Hob74]. Horizontal [Geh74, HC75b, IE78, CR76, Sta71]. horizontal-vertical [Sta71]. Horse [DH72]. Hot [Cas72a, Cas73, JC75]. Hoult [Kin70]. houses [All76]. Hrbácek [Sch73]. HRIR [SSA72]. Hubbard [HL73b]. Hudson [MC79]. Humic [NK72, PRJR73, WK77, Wil78]. humid [Bri76]. humus [Sch76]. Hung [Hor75]. Hurlbert [Zar78, Zar78]. Huron [TTC^{+}77, KMR78]. Hurricane [Hoo74]. hurry [Enr77a]. Hutchinson [Eme74, Ori73, Par75, Woo79, And71, Ano71c, Ano71f, Edm71b, Koh71c, Val71b, Wet76]. Hutchinsonia [Koh71c]. hyalina [BR72, ML70]. Hydra [SR79]. Hydrobiological [Sch73]. hydrobiology [Hyn73]. Hydrocarbon [Wak77]. hydrocarbons [WC76]. Hydrochemistry [Dep76]. hydrocoral [Str76]. Hydrodynamics [Gus79, Hun77, Bow79]. hydrogen [HIJ78, Ley73]. hydrographic [WEMD72]. hydrologic [BL79]. hydrology [GG78]. Hydrolysis [VM77, RAH77]. Hydrostatic [PA70, BT70]. hydroxylarnine [BWP73]. Hynes [Ben79, Cum71, CSHE78]. Hynes/Hamilton [CSHE78]. hyperboreus [Daw78]. hypereutrophic [ST78]. hypolimnetic [Dav77, DH78b]. hypothesis [Dod74, Doy79, Kil71, LR74a, MB79, Sta71].

I40 [Mof73]. Ianula [And71]. IBP [Fag72, Fra72, Hob73, Hob74, Par70, Ton72]. IBP-UNESCO [Hob74]. Ice [Cun72, Ada76b, Ada76a, AL78, BL72b, HA72, MS70, MG75, PI72, PDF^{+}78]. Iceland [DW72]. ideal [SR77]. Ideas [Ril78a]. Idemitsu [Epp73]. identification [Bas75]. identify [HPM72b, PN79]. Idrobiol [Hob74]. II [VM77, CE73, DB71, Fou71, HDFG72, MB71a, OH72b, TD72]. III [Fou73]. illumination [GTF78]. imagery [Lan70]. imagines [MBB73]. Immediately [Smi71]. impact [Fed75a]. Implications [EG79, Fee76, HC79c, Ker77, FM78, HK73, Lew73, TSJ78, TK76]. Importance [CGPD79, BK72, Dil75, HS77c, Koz70, Mal71, Mar70a, MP78, Thr79, Tur78, dBSA74]. important [GMCG78]. Improved [Fuh73, GH72, JDL78, SA70, Sha74b, SL76, BYR74, Bul78, CE79, FF75, KLM74, Pre78, Swe70a]. Inc. [Ano70g, Ano71b, Ano71a, Ano71i, Ano72a, Ano73a, Ano73f, Ano74a, Ano75a, Ano75n, Ano76a, Ano76e, Ano77b, Ano77a, Ano77g, Ano78b, Ano78a, Ano78c, Ano78g, Ano79a, Lau70, Sau72b]. incident [GTF78]. including [Ban79, RB71]. Incomplete [Gar77, CM79]. Incorporation [BW70, AC76, CAE78, SB79, TA78, TP73]. increased [BT70]. increases [Tho70a, Whi75]. increasing [Pat71]. incubated [KS70]. incubation [SA70, SMD76, SH73]. incubations [ES75a]. Independent [BYS77, BHD76, MW77]. Inderbitzen [Bou75]. Index [Ano70e, Ano71g, Ano72e, Bra76, Car77, Lor70]. India [OV78]. Indian [Koh74]. indicated [BSS78]. Indicator [Cav76, Kje78, Sil75]. indices [CTB72, MH72]. individual [FC70]. induced [FT73]. induction [San77]. inexpensive [Bak70, BL76, JFH73, LV76, Tem73, Web74]. infauna [KWK74]. infaunal [BB77, DC75]. infertile [Hoo74]. infestation [JSS^{+}71]. Influence [DLMS76, HP76, Koh71c, KMR78, PRJR73, WAA72b, And79, BM75, Cow70, ES75b, GCL70, KH76, Mal70, Ste76, WM78, Wil78]. Influences [Sta76, LR74a, Sta78]. influencing [Mal78]. Information [Edm71a, Man72b]. infrared [CB73, Lan70]. Ingestion [Arn71, Doy79, GM74, MB70b, MSF75, Rom78, EC72, RWI78]. Inhibition [Cro73, Rey78]. inhibitions [WP78]. inhibitors [AS77]. Initiation [FPO79, HS77c]. injection [Hal71]. inland [Fra72, GG78, Liv77a, Edm76a]. Inlet [Wea79]. Inlets [DO75]. Inorganic [SMD76, ZH70, AGA78, CS78, HPM72a, HS75, Jer70, JM70, Mac78b, SAH72]. input [RE73]. inputs [VTV^{+}78]. insect [CSHE78]. Insecta [MMC79]. insects [SA74, Edw77]. Inshore [MF71, Kim79, SOHH76]. Inst [Gol78, Gol74a, Gol74b, Gro75a, Ril71a]. Instantaneous [Edm72c, Cas72b, CSHE78, Pal74]. instantaneous-growth [CSHE78]. Instar [Fed75b]. Institution [Gro74b]. instruments [Cul73b, Pen73]. Int [Bow79, Gus79, Koz70]. integral [MCSK79]. integrated [Tyl71]. intensity [PA75]. Inter [Edm74a, Lau74]. Interaction [CM72, MQ71, OW73, CGPD79]. Interactions [Kro71, Kro72a, And79, BCK72, CK79, DH78b, GG81, JM78, LL79, MCA77, Pec79, SB75]. interbreeding [BGMW74]. intercalibration [War78]. interchangeable [GDHZ79]. interdisciplinary [Woo73]. Interface [KSGM71, SH71, Smi71, DVCP79, HC79a]. interference [Rie78]. interferences [CW78b]. interferes [GG75a]. Internal [Sch77b, iK75]. International [Ano70f, Ano71h, Ano72c, Ano73h, Ano75m, Hyn76]. interpretation [Ban74c, Mon78]. interpreting [Fee79, Pat80]. Interscience [Ban73, DH72, Fry71, Lym73, Mof73, O’C71, Sha73a, Str75b, Tyl71]. Interspecies [HR71]. Interspecific [BCK72]. Interstitial [MB77, MBR78, WAA72a, And79, Mak72, PH71, STND74, SM74, VB75, Wei73a]. intertidal [Dal74, Day79, Gar75, GGW78, Lea70]. interval [Bla79, CRA72, Hes76]. intestines [Ore79]. intracellular [Sen78]. intracellularis [Ven74]. Intrazooplankton [Con71]. Introduction [Ril71b, Hun77, Liv77a, Mil73a, Mil73b, Ste70, Boo75a, Eme76, Smi76b, Stu71]. introduction. [Mof73]. intrusion [Ahr75]. intussusception. [Ori73]. Invariance [PIM74]. invertebrate [CH70, Fry71, OV78]. invertebrates [DV78, Koh71b, Pai75b, Pec79, RHR71, SSHV76, WM78, Cas71]. investigating [Bro70]. investigation [HA72, SAH72]. Investigations [TG72, LS72]. invisible [Zar72]. involving [Eld75, PBK70]. Ion [Kes72, Wan72a, Wan72b, AS77, Ley73, MYY71a, Pyt72, PH74, SB73, Whi74b, Whi75, WK77, Man72a]. ion-association [Whi74b]. ion-pairs [PH74]. ion-transport [AS77]. ionic [KS78, Kim73a, Kim73b, Ley73]. ions [PC76]. Iron [Ant77, Glo78, LC71a, BL74, BH70, DH78b, GC74b, HD77, KH76, Koe76, LhC73]. iron-binding [GC74b]. iron-phosphorus [KH76].

Iron-rich [Ant77]. iron-sulfide [DH78b]. irradiance [Boo76, DB79, Jon78]. Island [BS78c, GB72c, MB77, MBR78, Ril78b, SC70, WRO73, Wea79, WWHH77, Hsu75, AW75, DO75, FP79a, RH76, WWH75]. Islands [GS73c, Koh71b, GMT^{+}72]. isobutyl [CS73a]. isolated [HPM72a]. Isothrix [Sau72a]. Isotope [GG78, Hai76, Kro75, PSK78]. isotopes [FHV79, SCA77]. Isotopic [NK72, LLS75, ME72]. Israel [BGZS75, DB79]. Ital [Hob74]. Italian [AGA78]. Italy [And71]. IV [Mar70a, Pam71]. ix [Bou75, Cus76a, Eme76, Gro74a, Kel75, Man75b, Smi76b, Smi76a, Wet75].

J [Bow79, Bri79, Bri72, Bro78a, But76, Cas71, Col72, Cul73a, Dee74, Dug71, Edm74a, Eme74, Fra72, Gol74b, Gol79, Gro71, Gro75a, Gro77, Gro78, Gro79a, Gus79, Hob74, Hor75, Hyn73, Hyn78, Ich71, Koh73, Kra79, Lau74, Leh75, Liv78, Man75a, Man70, Man73a, Man75b, Man79, Mar78a, Mil77, O’C71, Par72, Par75, Pat80, Phl73, Pla78, Pyt72, Pyt76a, Pyt76b, Ril78a, Sch77a, Sha73a, Sio77, Smi74a, Smi76c, Smi76a, Sta75, Ste79a, Str75b, Stu71, Tea78, Vly77, Wal74b, Whe73, Woo73, Woo77, d’A75, Sch73]. J. [Edm70b, Gro76b, Meg75]. Jagner [Lym73]. Jamaica [Rei72b]. James [Won79]. Jamieson [Edm72a]. Japan [Ich71, SS74, SK73b]. Jena [Sta72b]. Jenkins [Cul73a]. Jerlov [LaF75]. Jersey [Mil73a, Mil73b]. jet [BS75, BS78b, Mac78a]. Johannes [Koh79, Smi76a]. John [Gol77a, Gro77, Koh77, Lee77b, Ril78a, Ste70, Wet76]. Johnson [Liv78]. Joint [Ano74d]. Jordan [Ser74]. Jovanovich [Sch75b]. July [HC75b]. June [Ano74b, Ano79d]. Junk [Ham79, Man79, Rzo79, Sio77, Wu79]. juvenile [Egg78].

Kadota [Hob73]. Kalff [GG81]. Kamenkovich [Ste79b]. Kamloops [CGPD79]. Kaplan [Hun75]. Kara [Kel70]. Kariba [Beg76, Beg70, War79]. Kavaratti [QS70]. Kcs [Sch73]. Kelley [Gol74b, Gro75a]. kelp [Jac77, KW77, TP73]. Kelvin [TH74]. Kenya [Mel79]. Kiado [Hor75]. Kilham [PT74]. Kinetic [VWB77a, VWB77b, GC74a, Wil73]. Kinetics [BH78, BHK78, GF75, CCK71, LR74a, YK73]. King [Eme76, Smi76b]. Kinne [Ban73, Ban73, Gol77a, Koz70, Tyl71]. Kinneret [BGZS75, Ber70b, BP74a, Ber74, Cav76, DB76b, DB79, HS74, Ser71, SEPS74, Ser74, SM74]. Kirchner [SG72]. Kivu [Jan75]. Klamath [HPM72a]. Knights [Kim73a, Kim73b]. Knoechel [Cus77]. known [OS78]. Koehn [Fan77]. Koehn-Schneider [Fan77]. Koenig [Enr76]. Kol [Fot70]. Kolisko [Edm76b]. Kong [FT73]. Koppelman [Ril78b]. Kramer [Sha73a]. Kremer [Ste79a]. Kryobiologie. [Fot70]. Kryovegetation [Fot70]. Kuroshio [Ich71]. Kutz [And76].

L [Bou75, Cul73a, Edw77, Gro79a, How72, LaF77, Liv77a, Liv77b, Pom79, Ril78b, Wet75]. L. [CA79, HM75, MD71, MBN72]. Lab [Cul73a]. Labidocera [Lan78]. Laboratory [Gar74, HP74, Kac71, Mon78, RWI78, WM72b, CL76, Coo73b, CB75b, Cul73a, Pam71, RSS78, Sta72a, SG76, Eme74]. Labrador [Pin73]. Laccadives [QS70]. lacustrine [KB71, Ric75]. lacustris [Fro78]. Lago [And71, Cov76, HBLP76]. Lagoon [Lev71b, GM74, Har73b, KA71, NWEG75]. lagoonal [Hsu75]. Lake [Ahr75, Ano74b, BDM77, CKGS77, CKS77a, CKS77b, DBM77, Dil75, Ecc74, FS72a, GF70a, GF70b, GKA73, HF70, HSW72, HSC79, HJG71, HG72, HDFG72, Hor79, JP75, JC77, KGP^{+}78, KCS77, Lam72, Lud74, Man73b, MS74a, NND71, NND72, PI72, RHG73, RC72, SG76, Val71b, VHMW74, VBT79, Wei73a, WAA72a, WAA72b, Ada76a, AL78, BGZS75, BWP73, CBN78, CGPD79, CA79, CL79, CS74, Cov76, Cul77, Dav73, Dav74a, Dav74b, Eme75a, Fed75a, Fee79, Fit70, FH79, GL79, GS72b, Ids73, Imb74, JG74, JM70, JKC79, Kal72, KH76, Koe76, LR74a, LB71, LL75, MMW71, McL74, McL75, Mel79, OW72, Pat80, Pet78a, Pug77, Rau78, RAH77, RHC74, RH75, RH78, SG70, STND74, Sha73b, SHW^{+}70, SL70, SS78b, ST78, TG72, Tho77, TBOO75, Til73, Tim72, TA78]. lake [TL75, TG77, Wet70, ATM74, Ant77, BGZS75, BG70, Beg70, Beg76, Ber70b, BP74a, BR74, Ber74, BG78, BB73, Bra71, Bur70, CB72, Cav76, CM77, Cre75, Cro73, Cus76b, DP78, DB76b, DB79, Ecc74, EA70, Egg78, ESH73, ES77, FP77, FG75, GL75a, GMV72, GMTA79, GL71, GE75, Gro79b, HS74, HL73a, HPM72a, Hit78, HB72c, HHGRW76, IWC^{+}77, Jan75, iK75, KB71, KM72a, KHHG^{+}72, Kno77, KMR78, LR72, Lew73, Lew77, yLKR75, LH78, Liu70, LG79, Mar78b, Meg72, Mel79, Mon75, NW70, NC76, OT72, PG72, PRLG75, PB73, PE76, Par77, Pec70, PBS74, Pic76, PB77, Rey78, RMG70, Ric79, SH78, SRSS74, Sch77b, Ser71, SEPS74, Ser74, SEA71, Smi73a, STG73, SM74, TPG77, TL76, TP74, WC76, Wak77, War79, War72, WL73, Wel70]. lake-averaged [TG77]. lake-river [CGPD79]. Lakes [NCK79, SC76b, AGA78, And77, BA71a, Bax77, Ben75, Bla73b, BSB77, BL74, BM73, BM75, Car77, DB73, DG76a, Dev79, DR74, Dil75, Edm70a, Eme75b, Far75, Fee76, FP79b, GH76, HK73, HNT75, HBLP76, KC79, Kos70, LK74, Lan70, Las75, LR74b, Lee77a, LJ73, Lew73, Lud76, May76, MK74, OS78, OL74, RFFH76, RB70, SN75, SS74, SO78, SB72, Smi79, Ste73b, TeI70, TG77, Ver75, WWH70, Wal74a, Wal75, WSR79, Wel74, Wil72b, WH78, Wun73a, Wun73b, BH75, EA79, GLSD74, Mor71, Nri78, SE78, TTC^{+}77, Wei73a]. Laking [Gro74b]. Lamarck [Str76]. Laminated [Lud76, Ant77]. laminations [EA70]. Lanao [Lew73]. Lancaster [Smi75, CO75]. Land [Mac70]. Langmuir [HL73a, Sta71]. Langsø [WAA72a, WAA72b]. Large [GBB72, Ben75, BS77, CM79, Cov76, FR77a, MGHR75, PN79, STND74, SH72, TWPC75, Wun73a, Wun73b]. large-volume [STND74]. larger [WP78]. Larvae [LaR70, Fed75a, Lod74, SF75]. larval [Mil73c, Omo79]. laser [Str75a]. Latin [Zar78]. Latium [And71]. Lawrence [Ada76b, d’A75]. layer [Ban74a, GWN70, JB71, KD79a, Lor70, Tab79, WH71, Kap78]. layers [DM79a, Jer70, RSEG78]. leachates [LH73c, Tur78]. leaching [HM75]. Lead [ZY72, KM73]. leaves [KA77, LI76, LH73c, WM72a]. LeBlond [LR74b]. Leckie [Cul73a]. leei [ML70]. Legend [Edm71b]. Leinebø [Gro79a]. Length [DD78, Ton71]. Leslie [Hyn73]. Leucothrix [JSS^{+}71]. level [DW72, Hil74, QPB70, Car75]. Levels [Bun71, CS74, Lan79, Ten77]. Leydig [BR72]. Leyendekkers [Pla78]. Library [Har78a]. Liege [Bow79, Gus79, Pyt76a]. Lieth [Man76]. Life [GC72, RC72, Lan75a, Tyl71, Koh72, Rzo79]. ligands [SL78]. Light [BS77, DB76b, PBHC70, RK79, Ste75, VVT76, Bat76, CAE78, GIK77, GB72a, HS77c, Hoj75, JP76, KS70, KA74, Kos79, MG75, MCSK79, MYY71b, Mur76, Pac73, Pat71, Sen78, Sta76, Str76]. light-saturated [Sen78]. Light-scattering [PBHC70]. lightweight [BHD76]. like [TJ72]. Likens [Hut72, Liv78]. Lillj [Coo71]. Limfjorden [Jør77]. limit [Ban74b]. Limitation [Hul76, DBH78, GL75b, PG74, Rhe78, TKK76, VVT76]. limitations [Whi74a]. limited [DBH78, Gid77, PG75, Per76, SRSS74]. Limiting [LC71a]. limits [TL75]. Limnetic [Egg78, PL73]. Limnocalanus [RC72]. Limnol [Hut72]. limnologic [IE78]. Limnological [Beg70, Ber72, SB72, Wet76, BYR74, Mac70]. Limnologie [Fot70]. Limnology [Ano70g, Ano71b, Ano71a, Ano71i, Ano72a, Ano73a, Ano73f, Ano73h, Ano74a, Ano74b, Ano75a, Ano75n, Ano76a, Ano76e, Ano77b, Ano77a, Ano77g, Ano78b, Ano78a, Ano78c, Ano78g, Ano79a, Ano79d, Lau70, Sau72b, BG70, CKGS77, Ecc74, Eck70, Mor73, TTC^{+}77, WL73, Wet76, Hor75, Liv76, Kil76, Liv77a, Liv77b, Ano70e, Ano71g, Ano72e, Ano73c, Ano75e, Ano76c, Ano77e, Ano78e, Ano79c, Val74a]. Limnothrissa [Beg76]. Linear [Buz71, Str75a]. lipid [AS77, WBM^{+}77]. lipids [BA76, LC74]. Lippincott [Hyn78]. Liquid [Pug70, WWR70, WN71, BC74, IBM76, Pug73]. Lisitzin [Car75]. Liss [Gro78]. litterfall [GH76]. Little [BDM77, DBM77]. littoral [FG75, Fot74]. Live [SL70, IR73]. Living [NN71, DM79a, Ill75, KHHG^{+}72, Phl73, SSM76, TVB79]. loading [Cha79a, SO78, Tho77]. loadings [OS78, SO78]. lobate [RWI78]. Lobelia [WAA72a, WAA72b]. location [Swe70a]. Loch [Mor73, SW73, SK72, SK73a, TH74]. Logan [Gro71]. Logistics [Ham75]. Loisel [GK74, Rei72a]. London [Aro73, Bri72, But76, Ebe72, Edm74a, Gol79, Gor79a, Gor79b, Gor78, Gro76b, Gro78, Hed70, Hob73, How78, Hun75, Kin70, Koh71a, Koh72, LaF75, LaF77, Lau74, Liv76, Liv77a, Man75a, Man73a, Man75b, Mas70, Mil77, Mof73, Pyt76b, Sch77a, Smi76c, Sta74b, Stu71, Wet76, Woo73]. Long [Ril78b, Wea79, CS74, iK75, AW75, FP79a, GB72c, MB77, MBR78, WWH75, WWHH77]. long-period [iK75]. long-term [CS74]. Longhurst [Ano73g]. Longman [Edm76a, Liv77a]. Longshore [Sha74a]. Lookout [SM78, SA71]. Loring [d’A75]. Loss [Bro78a, IBM76, JG74, SD78, ST78, ES75a]. losses [Bri76]. lotic [CR70]. Lough [GS72a, GS73a, GS73b]. Louis [Kil76]. Low [Wei73b, CS73a, JM70, MYY71a]. low-nutrient [MYY71a]. lower [HS77b, Sma74a]. Ltd [Man73a, Sch77a]. luminous [RN78]. lunula [BS77]. luteum [TBD77]. Luther [Fra72]. lutheri [SL78]. Lux [Tyl73].

M [Bal76, Bri79, Bro78a, Cas71, Csa75, Dee74, Dug71, Edm72a, Edm74a, Eme76, Gor78, Hed70, How78, Hyn78, Liv72b, Liv78, Man72a, Mas70, Ril74, Ril78b, Sch73, Smi76b, Sta75, Ste79b, Woo79, Ber77, Dee71, DB71, MCB79]. M. [DH78a]. Macartney [FT73]. Macfadyen [Hyn78]. MacIntyre [BS78a]. Mackenzie [Liv72b]. Mackereth [BB79]. Macmillan [Sch77a, Smi75, Smi78a, Val71a]. macroalga [HC75a]. macrobenthic [Hoo74, KWK74]. Macrocystis [Jac77, TP73]. macrofauna [Ped74]. macroinvertebrates [KMR78, WB74]. Macromolecules [Aar71]. macrophyte [AGA78, HD70, OW74]. macrophytes [McL75]. macrophytic [Tur78]. macroscopic [All79, RSS78]. Macrosetella [Rom78]. macrozooplankton [Mil73c]. macrurus [RC72]. madagascariensis [Mul74]. Madison [TL76, Ste73b]. magna [HP77]. Magnetic [TBOO75]. Maine [FMJC72, Glo78, SN75]. mainland [Smi72]. Major [DG76a, GC72, LB71, PC76]. making [CA72]. Malahoff [Mac79]. Malawi [Ecc74]. Malina [Cul73a]. Malins [Gor79a, Gor79b]. Malone [PT75]. Maloney [Leh75]. mammals [Sch77a]. Management [Ano72c, Edm70a, Fry71, HNT75, Pom79, Gro77]. mandibular [Sta78]. Manganese [GBK76, BL74, BS71, GCL70, KM73]. Manitoba [RHG73]. manual [HAG^{+}72, Par70, Ton72, Cul73a]. maple [LI76]. mapping [KWD75]. maps [Cha74]. Mar [Gol78, Gol74a, Gol74b, Gro75a, Har78a, Man72a, Pla78, Ril71a, Woo73]. Marcel [Par73, Pla78]. Margalef [Ori77]. Margaret [HPM72b]. margins [Eme78]. mariculture [Par75]. marin [Cus76a]. marina [HM75, MBN72, PS77, TLR^{+}70]. Marine [BY73, Bou75, Bun71, Fro72, LC71a, Man75a, MB70b, NK72, Pai70, PA70, Ril78a, ST79, TLR^{+}70, YYP72, AZ70, And75, AC76, AH77, BGMW74, BPW76, Bas75, BLM74, BST70, BC72, Bul78, CS78, CTB72, CB72, CRW72, CAE78, DGEB70, DBH78, DCP73, Eag71, EGTA70, Eme76, Fag72, Fal75, Fou70, Fou71, Fou73, GM79, Gol78, GM78, GC74b, Gor79a, Gor79b, GG75b, GM77, HIJ78, Har73a, Har73b, Har78a, HC75a, HK79, Hun75, Jør77, KS78, KOW79, KP75, KM72b, Koh71b, Koz70, Lam78b, Lan75a, Lan75b, Lan78, LH73a, LC74, LWR73, LP70, LM70, MP77a, MO72, ME72, McC72, MH74, Mil70, MJH77, MYY71b, Mul79, PHR71, Pac73, Pae74, Paf76, Pai75a, PBHC70, Phl70, PF73, PH71, Pom74, Pou78, PRJR73, PRG78, RH77, Ril73, Sch74, Sha77b, She72]. marine [SL75, SB79, SMD76, SCA77, SK75, Smi76a, Sta75, Ste79a, SN71, Ste70, Str75b, Stu71, Swe70a, TCTW75, VB75, WD78, Wal74b, WNS74, Wil72b, WK79, YY79, Aro73, Ban73, Dug71, Edw77, Gol77a, Gro77, Kel75, Man70, Par73, Smi74a, Tyl71, Wal76a, Par78]. marinus [CM72, PA70]. Mariscal [Pai75a]. markers [DL74]. marl [OW72, Wet70]. Marr [Ich71]. Marsh [Rei72a, SG72, AW75, Boo75b, Cam75, CVV78, ES75b, GD74, Gal75, GP77, Gar75, GK74, Hai76, HT79, Jef72, KVT79, LH73b, Phl70, San77, TVB79, Tur78, VTP76, VTV^{+}78, VVT76, WWH75, WWHH77, dlCP75]. Martin [Par73]. Mary [WL73]. mass [BP74b, Woo77, VWB77a]. Massachusetts [Gro74b, Wal74b, CVV78, Kim73a, Kim73b]. masses [AGM70, PG72]. Masson [Cus76a, Mar78a, Par78]. mat [WF72]. material [BK72, Gor78, Moe71, Pug70]. materials [BH75, WK77, Wil78]. Mathematical [JP76, GB72b, YY79]. mathématiques [Mar78a]. mats [DB76a, HMS^{+}77, JC77, KCS77]. Matter [Ogu70, Pou73, Aar78, All76, AZ70, Bri76, CM79, CS70, Fog77, Fuh73, GH76, HHPP70, HK79, Jan73, KLG71, KH76, MHU72, MSJG74, MP78, Nai76, NL72, OH72a, OH72b, OW74, Pae74, PSK78, Pou78, PBK70, QS70, Rei72b, Ril71a, Sha77b, Sor73, SK73b, SRSW78, TP74, WD78]. Matthews [Gro79a]. maturity [MH72]. Maximum [PA70, CM77, RSEG78]. mayfly [Swe78, WC73]. McCave [Ril78a, Kap78]. McGraw [Gro75b]. McGraw-Hill [Gro75b]. McIntyre [Fag72]. McMurdo [RD76]. mean [BH75, Ste73b, Wun73a, Wun73b]. measure [CS74, FC70, Gar78, Ids72, JP78, MW77, SC76a, TA78]. measured [And77, BL76, KO76, KB76, MYY71a, SS77]. Measurement [GCL79, Han71, MCHP73, RHC74, She79, SRSW78, Zir75, BPW76, BGZS75, Boo76, Bra76, Dav77, GTF78, Har77, HC78, Høj78, IJ79, KHC74, LB79, Pet75, Pom74, PME73, SCA77, Sta74a, TKH75, ZH71]. Measurements [CL79, Cul70, San77, Bla73a, CB70, DBM77, ES75a, Fro78, HP76, Har78b, Hoj75, MO72, Nor76, Ram76, RSS78, SD78, SWLH78, Str75a, SOHH76, TB75, The75, Whi74a]. measures [MWW77]. Measuring [HLH71, CE79, GV78, Lee77a, Mar74, ME72, Mcm73, Par70, QPB70, SCE^{+}76, Tit75, Wal73]. Mechanical [KW77, Bou75]. mechanism [Ore79, ZS76]. Mechanisms [HIJ78, LH73b, MS74a, SCLM70, Mor71]. Mechums [TGK^{+}78]. media [DGEB70, JM78]. Mediterranean [Eme74, Sha74a, Sha77a, Sho74]. medium [JF72]. Meeresuntersuch [Koz70]. Meeting [Ano74b, Ano79d]. Méhauté [Hun77]. meiobenthos [Rup72]. meiofauna [TTW75, dBSA74]. Melchiorri [Hob74]. Melchiorri-Santolini [Hob74]. Mem [Gro71, Hob74]. members [Hut72]. Membrane [LC71b, Mcm73, RW71, CACB75, HMWW72, SRR79, She72]. membrane-adsorption [CACB75]. Membranes [Aar71]. Menai [MF71]. Mendota [TL76, EA79, GL75a]. mendotae [HP77]. Menten [BHK78]. Menzies [Lau74]. Mercury [FL75, KM72b, Jer70]. meromictic [Eck70, GS72b, KGP^{+}78, Tim72, Wal74a, WL73]. meromixis [Cul77]. Mersey [BG71]. mesh [dBSA74]. Mesocyclops [Con71]. mesopelagic [Bel78]. messenger [HBM73]. messenger-operated [HBM73]. Metabolic [Wel74, Enr77b, Kar78, SS74]. metabolism [BB70, CW78a, Gal75, GP77, Har70, PB73, PBK70, RH78, SK75, SCE^{+}76, WH71, Wri75]. metal [GL79, JM78, WK77]. metal-chelator [JM78]. metallic [MCA77]. Metals [Woo77, BK72, DGEB70, DS75, DCP73, Eld75, Mar70b]. meter [Hoj75, SC76a]. Methane [Jan75, RH78, SM78, RH77, RHC74, RFFH76, SB78a, ST78]. methanogenesis [Ore79]. Methanogenic [Ore79]. methanol [Mar77, TKH77]. Method [BY70, Ert71, Han71, HLH71, JW70, KSGM71, Pom74, Por72, QPB70, AGF71, Bat73, Ben79, BT70, BFM76, BB77, Bul78, CE79, DLMS76, EK73, Gar78, GHM78, GV73, GCB76, HB72b, Hoi78, KK76b, Lam74, LB79, Loh74, Mar74, MWW77, Mcc70, Mcm73, MQM74, MY70, MY72, PDF^{+}78, Rup72, Sal79, SW73, SB73, Sta74a, SW72, TKH75, TB75, TTC^{+}77, WWH70, WD79, YK73]. Methodol [Koh79, Mil78]. methodology [BB70, Kir75a, Lit73]. Methods [GB79, WN71, Bro70, Cas71, CP79, GL75b, MO72, ME72, Par70, Ril73, Ton72, WC73, Fag72, Rig72, Ril77, Koh79, Par73]. methyl [Jer70]. Mexico [Bra71, BBBS79, MHU72]. Mich [Woo77]. Michaelis [BHK78]. Michigan [EA79, Sch76, BG78, ES77, FP77, GF70a, Gro79b, HB72c, Liu70, Mar78b, NW70, PE76, SRSS74, VHMW74, Wel70]. micro [Bro70]. micro-organisms [Bro70]. Microalgae [Bun71, BL72b, PRG78, TKH75]. microalgal [TGK^{+}78]. Microbial [JKC79, Pae78, RH77, TJ76, TJ79, Cas72a, Cav76, CW78a, ES75b, GMCG78, GA75, Hob73, Mon78, OH72a, OH72b, Ram76, RHC74, Sta72a, TA78, WBM^{+}77, Wil73, dH77, Hob79, Sta75]. microbiological [GMT^{+}72]. microbiology [Web74]. microcosm [CR70]. Microcosms [GBB72, Coo73b, RB74]. microcrustaceans [ES77]. microelectrode [Zir75]. microflora [Har70]. Microfluorometric [Gar78]. microfossil [Bat73]. microheterotrophs [AH77]. microlayer [EA79, Gar74, HB72b, HP74]. microlayers [BC74]. microorganisms [BHK78, CKS77a, Fou70, Fou71, Fou73, GC74b, HHHP^{+}72, Kar78, SS74, SB75, Shu78, Sta75]. microplankton [DP78, Mon75, SS78a]. microsampler [Bak70]. Microscale [ST79]. microscope [HMWW72, Jon74]. microscopy [Bas75, Bro78b, CA72, PS73]. Microstructure [NND71, CBN78]. mid [HC75b]. mid-July [HC75b]. middle [Høj78, SC76a, WP76]. middle-ultraviolet [SC76a]. Middlebrooks [Leh75]. midge [KA71]. Midultraviolet [KVW79]. midwater [Chi71]. migrating [Bel78, PP77]. Migration [LaR70, LB71, BFM76, Enr77a, Enr77b, EH77, GF70b, HC79c, Hu78, KZ77, LL75, MBB73, MM75, Mil70, Pea79, ZS76, Enr76]. Migrations [Cun72]. migrators [Kos79]. Mikrobiologie [Sta72b]. Milan [Par78]. miliaris [Kim79]. Millepora [Str76]. Miller [EK73, WWH70]. Milliman [Smi74a]. Mimeographed [Edm70a]. Mineral [ML70, Bas75, DG76a]. Mineralization [ML70, BSB77, CL79]. minimal [Bla79]. Minnesota [Ant77, DG76a, Meg72, MS74a, SG72]. Minnetonka [Meg72]. minutus [MMC79, Pou73]. miodon [Beg76]. Mirror [VBT79]. Missouri [Cow70]. Mitchell [Edm74a]. mixed [Ban74a, KD79a, Kje78, Tab79]. mixing [DO75, HNT75, IWC^{+}77, ICO^{+}74, KMR78, PRLG75, Ser74, TK76]. mixture [Doy79]. mixtures [Fro77]. Mize [KB71]. Mn [Dav77, Wil78]. Mobile [May73]. mode [Sch72]. Model [Buz71, LF76, Mur71, ZY72, ABGCP75, BH75, DBH78, Dro78, Eme75a, Enr77b, EH77, GIK77, GB72a, GB72b, Imb74, KC79, LL77, LBL75, yLKR75, PH74, Ril76, TSJ78, TG77, Van77, Whi74b, Wie71, Wil78, WK79, WQ73, YY79]. modèles [Mar78a]. modeling [DH78a, Leh75, Ril78a]. Models [Mor72, Ste72, Ber77, Cha79a, OS78, Sch78b, Tho77, VWB77a, VWB77b, Dee74]. modern [LT71]. modes [HNT75]. modification [Ben79, OSD76]. Modifications [BB79, Ids72]. modified [Bro79, CB72, Cas72a, DH75, GTF78, PRG78, TB75, WW73]. modifiers [AS77]. Mohave [Nai76]. Moina [Rat74]. molal [PC76]. molecular [Whe76]. mollusc [CB73]. Mollusca [Lev71a]. molluscan [CS74, Cov76, HF70]. monitored [BDW76]. monitoring [BFFB75, MSJG74, RH75]. Monochrysis [SL78]. Monogr [Koh79, Lun79, Mil78, Pom74, Rzo79, Sio77]. monster [Mor73, SK73a]. monsters [SW73, SK72]. Monterey [KM73, RSS78]. Monterosi [And71]. month [RB71]. moonlight [CE73]. Mordukhai [Rzo79]. Mordukhai-Boltovskoi [Rzo79]. Morel [Bal76]. Morgan [O’C71, Cle72]. morphology [Dod74, Kir75b, OS79, SC76b, d’A75]. Morphometric [TKK76, Tim75]. Morphometry [Ric75, BM75, Fee79, Kos70, Pat80]. Mortality [Van77, Fag73, Omo79, Thr79]. Mortimer [SK73a]. Mosby [Kil76]. Mosquitofish [GBB72]. motile [KZ77]. motion [CNC73, FPO79, Gro75b, TC74]. Motoda [Epp73]. moulted [Mar70b]. mountain [GMCG78, Til73, Wil72b]. Movement [Con75, Rei72a, TL75]. movements [Beg76, PI72]. mucor [JSS^{+}71]. mucus [BM74, CS73b, DM79a, DM79b, RLS75]. Mud [KSGM71, SH71, Koh72, McL74, SEPS74, Web74]. Mud-Water [KSGM71, SH71]. muddy [dBSA74]. mudflat [Lea70]. muds [Fit70]. Mueller [WN71]. Multiple [Tay78, WCR71]. multispecies [MMD79]. Multivariate [Buz71, HPM72b]. multivoltine [Ben79, RC74]. Murphy [Ste71]. Murray [KS78, Phl73]. Museum [Col72]. mussels [BG72]. Myriophyllum [ATM74, CA79, Hut70, MD71, Wil72a]. Mysis [BG78, FM78, Gro79b, LL73].

N [Cum71, Fag72, Gro74b, Kap78, LaF75, Liv78, Mac79, Pai75a, Ril78a, Smi78a, Ste79a, BBG75, CPOT79, MEW^{+}78, OT77, Rhe78, SCA77]. N.Y [DH72]. N0 [Toe71]. Nägele [Fot70]. NaHCO [IBM76]. Najas [HW78]. Nanaimo [NS78]. nannoplankton [Mal71, PM77]. nanoplankton [Kal72]. Napeague [BS78c]. Narragansett [DD78, GBK76, HS77b, HS77c, Mar70a, Sma74a]. NAS [Hyn76]. NAS-NRC [Hyn76]. Nat [Bar71]. National [MB71b, Hyn76, Fer70]. Natural [Col72, Con71, IE78, Pai71, Sau72a, AC76, BL72a, Bri72, BLL70, CK79, Cha79b, DM71, Eld75, Eme74, Gar78, GHM78, GG81, GA75, HD77, Jon78, LL79, LG70, Mac78b, Man72a, McC72, MCA77, Moe71, PHH78, RK79, Rie78, Ril71a, Ris73, Sch75a, SRR79, Sha73a, SC76a, SB78c, SL78, Tho70a, Tyl75, WNLS77, Whi72, YY79, Hun75]. Naturally [Pou73, MP78, Pou78]. Nature [Gor79a, Gor79b, Fan77, Man79, Zar72]. nauplii [SM79]. Neagh [GS72a, GS73a, GS73b]. Near [CP75, Bru71, BS78c, CP70, GB72c, SS70, SA76, TP78]. Near-bottom [CP75, CP70, GB72c, SS70]. near-surface [TP78]. Nearshore [Fol71, ES77, Hsu75, Ros79a, Ros79b, RN78, WRO73]. Needham [Chu72]. needle [Rau78]. Negast [Bax77]. negatively [HW76]. negligens [RLS75]. nekton [LH73a, Ham79, Wu79]. Nematode [TLR^{+}70]. neretic [All79]. neritic [FP79a, Mal71]. Ness [Mor73, SW73, SK72, SK73a, TH74]. Net [CM72, Kal72, BFM76, BSS78, Coo73b, Ids72, Kje78, Mil73c, PM77, Pet78b, WF72]. Netherlands [dH77]. netplankton [Mal71]. nets [Che73, CE73]. Neuse [Wol71]. neuston [Dis73]. Nevada [GMTA79, PRLG75, RMG70, TPG77, KGP^{+}78]. NH [Toe71]. niches [Dod70]. Nielsen [LaF75, Wal76a]. night [Cli72]. Nihoul [Bow79, Gus79]. Nile [Man79, Sio77, Sha77a]. Nimbus [SSA72]. Nitrate [ERH79, Fal75, Har73c, MGHR75, Bat76, BS77, Bie75, CR72, ECHH^{+}71, HAG^{+}72, KD79a, Pac73, PRLG75, Rhe78, RB71, Tho70a, TR76, VB75]. nitrate-ammonium [Bie75]. Nitrification [GM77, BBG75, VB75]. Nitrite [WH71]. Nitrogen [Ano78h, BBPK71, CM75, CP77, CVV78, GWN70, GP72b, GKA73, GL71, HC75a, HSC79, HG72, HDFG72, Hor79, KB71, MGC73, TVB79, TL76, VHMW74, WWH75, Ban74c, CL76, DSC77, DDS78, Edm70a, ECHH^{+}71, GL75a, Gid77, GS74, Har78b, KM72a, Ker72, KD79b, KA77, Mac78b, MMW71, ME72, MYY71a, PSK78, PBK70, RFFH76, RB70, SA70, SRSS74, Sha74b, Shu78, SMD76, SCA77, THDW79, Tho70b, TD72, TR76, WDW^{+}75, WM72a, Woo77, VWB77b]. Nitrogenous [MTT77, CM75]. Nixon [Ste79a]. NKr [Gro79a]. No [Fag72, Fra72, Har78a, Koh79, Man72a, Par70, Ton72]. Nobel [Lym73]. Noctiluca [FT73, Kim79, BS77, KS78, RS79, SSM76]. Nocturnal [Gro79b]. Nonliving [Joh76]. nonmarine [Bay78]. nonmembers [Hut72]. Nonmetallic [Ped74]. nonmotile [KS78]. nonresonant [TC74]. nontidal [Con75]. North [Edw77, Epp73, FW74, MG78, SM78, SA71, Wol71, CR72, CT70, ERVM73, Kal72, Kel70, Mil71, SB72, WH71, BHSJ78, CR76, CP75, ES75a, Fou71, GN79, MC78, Per76, Ven74, Wil72b]. North-Holland [Edw77]. northeast [AZ70, BKM78]. northeastern [Ant77, ZH71]. Northern [Lev71b, BP74b, Epp73, FN70, MHU72]. Northwest [Cul70, RE73, BA71a, NG78, SMD76]. northwestern [BBBS79, Fee76]. Norton [Liv72b]. Norwegian [Gro79a, WH78]. Nota [d’A75]. Note [Woo71, BG70, Bax77, GC70, Smi77]. Notes [LG70, Gil71]. Nova [HPM72b, BSS78]. Novel [Whi71]. November [MSR^{+}75, Pic76]. NRC [Hyn76]. Nriagu [Woo77]. nuclides [KTTF76]. numbers [Bat73]. Nuphar [TBD77]. Nutrient [CO75, MMD79, SEPS74, SWLH78, Smi79, SA71, VTV^{+}78, ABCP75, ABGCP75, Ban74a, Ban79, BA71a, Ber77, BM75, Doy75, ERVM73, FP79b, GL75b, LL77, LG79, MYY71a, PG74, PG75, Ser71, ST79, TK76, VVT76, VHMW74, Wal76b, MBR78, MTT77]. Nutrient-density [SA71]. Nutrient-determined [MMD79]. nutrient-saturated [Ban79]. Nutrients [Jac77, PCWR70, CM75, CA79, HB72c, MPN72, NS78, Hut72, Sha73a]. Nutrition [Wil72a, JCK70, MTT77]. nutritional [Leh76a, Sta78]. Nyasa [Ecc74].

O [Ban73, Cul73a, Dug71, Edm72a, Gol77a, Koz70, Tyl71, Val71a, Woo77, Zar78, Dav77, MEW^{+}78, OT77, SR77]. Oak [MG78]. Obermiller [Fot70]. obligately [BT70]. O’Brien [Ril78a]. Observation [Str77]. Observations [BH70, CS73b, Fie72, HP70, HL73a, HS77a, Mac77, MBB73, RH76, RO70, Beg70, GMT^{+}72, Ham75, HMA^{+}75, HSK76, JM78, Jon74, JS75, LLS75, Mac70, RAC71, Ril78a, Sho74, SE78]. Observed [PB77, Ben75]. obtained [Jon74]. obtaining [MWW77]. Occas [Gol78, Gol74a, Gol74b, Gro75a, Ril71a]. Occurrence [TJ72, BWP73, Chi71, FT73]. Occurring [Pou73, MP78, Pou78]. Ocean [Bow79, Gus79, Big77, FL75, HHHP^{+}72, How72, Hul76, Kra79, Mil73a, Mil73b, MP77b, MY70, Mur72, SB78a, SPS72, SSP73, SB78b, Ste79b, Whi75, WK79, ABCP75, AGM70, AZ70, BSB71, BNV^{+}74, CR72, CL73, CO75, Csa75, Daw78, ERVM73, GN79, GWN70, GMT^{+}72, Kel70, Koh74, LS77, MM75, MG75, Mil71, MC78, Nat70, WH71, ZH71, Koh73]. Ocean-ogr [Ste79b]. Ocean. [Epp73]. oceanic [BS77, Boo75c, BFMF72, CM79, Koh71b, Kro75, Mal71, Omo79, Ore79, Per76, Rem71, RS79, Sma74a, SSM76, TL75, Wie70, Tho70b, TD72]. Oceanogr [Car75, Hut72, Koh79, Mil78, Pom74]. Oceanographers [Gro79a]. Oceanographic [Ano74d, Ano75m, Gro74b, BYR74, Bla73a, Ril74]. Océanographie [Par78]. Oceanography [Ano70g, Ano71b, Ano71a, Ano71i, Ano72a, Ano73a, Ano73f, Ano74a, Ano74b, Ano75a, Ano75n, Ano76a, Ano76e, Ano77b, Ano77a, Ano77g, Ano78b, Ano78a, Ano78c, Ano78g, Ano79a, Ano79d, Lau70, Sau72b, Smi76a, Ste70, AR73, Eme78, Epp73, Kel75, Lee77b, Gro75a, Woo79, But76, Gol79, Gro76b, LaF75, LaF77, Mil73a, Mil73b, Mil77, Pyt76b, Sch75b, Smi76b, Smi76c, Ano70e, Ano71g, Ano72e, Ano73c, Ano75e, Ano76c, Ano77e, Ano78e, Ano79c, Val74a]. oceans [LNS^{+}76, Pyt71, Smi75, Tyl71, CP75, KLG71, Gro76a, Lau74, Lym73, Mac79]. Ochromonas [Aar71, AP76, BA76]. October [NW70]. off [Ber70a, BP70, BSS78, Dee71, DB71, FN70, GMT^{+}72, Hul76, HPS75, Huy77, Jac77, PK74, RE73, RH76, RSEG78, WRO73, WS70]. Officer [Boo75a, Lee77b]. offshore [AZ70, Jan73]. ogr [Ste79b]. OH [Bra73a]. oil [Con75, Mur72, Kin70, Woo73]. Okla [Gro71]. Oklahoma [Gro74a]. Old [Fan77]. oligactis [SR79]. Oligochaeta [Edm72a]. oligochaetes [BCK72, Lod74]. oligotrophic [CL79, Per76]. olive [And75]. olive-green [And75]. Oliver [Bro70, Edm72a, Man70]. Olympic [CL76]. Omega [Ori77]. Omnivorous [CB75a]. Oneida [HF70]. Ontario [EA79, Nri78, BA71a, BDM77, CM77, DBM77, Dil75, Fee76, GF70b, GMV72, HL73a, Hit78, KM72a, Las75, LR72, yLKR75, NC76, NCK79, Pic76, PB77, Sch77b, SC76b, TTC^{+}77, Val71b, Wei73a]. open [Big77, FL75, MY70, Nor76, Par75]. open-ocean [FL75]. operated [HBM73, WMM74]. operation [Edm70a]. ophiuroids [Sch72]. Optical [STG73, SB78c, SB78b, KOW79, LaF75]. Optically [Owe74]. optima [Bot75]. optimal [Leh76b]. optimization [Doy79]. Orca [MLCG79]. Oregon [BP74b, HPS75, Huy77, Lam72, NND71, NND72, PK74]. oregonensis [Coo71]. Organic [BW70, Bri76, Fen70, Fie72, FPO79, MF71, MK79, Ogu70, SM73, Aar78, All76, AZ70, BL79, Cam75, CS70, Fog77, GH76, GH79, Har72, HS75, HL73b, HK79, Joh76, KLG71, KH76, Lam78a, LI76, LH72, MMW71, MSJG74, MWW77, Nai76, OH72a, OH72b, OW73, OW74, Pae74, Pae78, PS77, PSK78, QS70, QM71, Rei72b, RHG73, RAH77, SSHV76, SAH72, Sha73c, Sha74b, Sha77b, Smi73b, SHW^{+}70, Sor73, SK73b, SRSW78, SS78b, SL78, THDW79, TP74, WD78, Wan74, WM72a, Whe76, Whe77, Ril71a]. organic-inorganic [SAH72]. Organism [RC71, MMC79, TS78]. organisms [BDW76, Gor79a, Gor79b, Hyn78, JF72, Koz70, SOP73, Swe76, Bro70, Str75b]. Organization [Edm70a]. organo [MCA77]. organo-metallic [MCA77]. Orientation [Gri72]. Origin [NK72, And75, HT70]. Orthophosphate [Dux75]. Osborn [LR74b]. oscillations [Sch77b]. Oscillatoria [CM75, CP77, FF79, RMC79, Sau72a, YYP72]. oscillaxanthin [GE75]. Oslo [Gro79a]. osmotic [Man73b]. ostracods [ES77]. other [Cle72, Cre75, HG76, NWEG75, PRG78, SD78, WSR79]. Our [Kra79, Val71a]. outburst [Pec70]. outdoor [MGHR75]. outflow [CO75]. outline [Ecc74]. Output [HL73b]. ovei [HPS75]. override [MH72]. overturn [Smi73a]. Oxford [Car75, Ebe72, Fag72, Fra72, Hob73, Hyn78, Liv77b, Par70, Ril74, Ton72]. Oxidation [SH71, Jan75, RHC74, RFFH76, SM78]. Oxidation-Reduction [SH71]. oxide [MW77, Moe71, SC76b]. Oxygen [ABCP75, Bun71, CR72, LaR70, Pam71, Smi74b, ABGCP75, Ano75m, Coo73a, CS74, CP73, Dux75, EK73, GTF78, HP77, IJ79, JPW79, KLM74, Kro75, Las75, May73, Pyt71, RFFH76, SWLH78, WWH70]. Oxygen-nutrient [ABCP75, ABGCP75].

P [Bro78a, But76, Cus76a, Edm72b, Eme78, Gol79, Gro74b, Gro76b, Gro78, Kin70, Koz70, LR74b, MM75, Mil77, Pai70, Pai75b, Par78, Pyt72, Pyt76b, Ril78b, Smi75, Smi76c, Sta74b, Ste70, Stu71, Whe73, Woo73, And71, Aro73, Ban73, Bar71, Boo75a, Bou75, Bow79, Bri79, Bri72, Bro70, Car75, Cas71, Cha74, Cle72, Col72, Csa75, Cum71, Day79, Dee74, DH72, Dug71, Ebe72, Edm70a, Edm71a, Edm72a, Edm74a, Edm76b, Edw77, Eme74, Eme76, Enr76, Epp73, Fag72, Fan77, FW74, Fot70, Fra72, Fry71, Gol74a, Gol74b, Gol77a, Gor79a, Gor79b, Gor78, Gro71, Gro74a, Gro75a, Gro75b, Gro76a, Gro77, Gro79a, Gus79, Ham79, Har78a, Hed70, Hob73, Hob74, Hob79, Hor75, Hor76, How72, How78, Hun75, Hun77, Hut72, Hyn71, Hyn73, Hyn76, Hyn78, Ich71, Ill75, Kap78]. p [Kel75, Kil76, Koh71a, Koh72, Koh73, Koh74, Koh77, Kra79, LaF75, LaF77, LeC77, Lee77b, Liv72a, Liv72b, Liv76, Liv77a, Liv77b, Liv78, Lun79, Lym73, Mac79, Man72a, Man75a, Man70, Man73a, Man75b, Man76, Man79, Mar78a, Mas70, Mil73a, Mil73b, Mil78, Mof73, O’C71, Ori73, Pai75a, Par70, Par72, Par73, Par75, Phl73, Pla78, Pom74, Pom79, Rig72, Rig76, Ril71a, Ril73, Ril74, Ril77, Ril78a, Rzo79, Sch77a, Sch73, Sch75b, Sch76, Sha73a, Sio77, Smi74a, Smi76b, Smi76a, Smi78b, Smi78a, Sta72a, Sta72b, Ste79a, Ste79b, Str75b, Tea78, Ton72, Tyl71, Val71a, Val74b, Wal74b, Wal76a, Wet75, Wet76, Woo77, Woo79, Wu79, Zar78, d’A75, OS78, Rhe78, Tab79]. p.

[Edm76a, Gol78, Koh79, Lau74, Leh75, Pai75b, Par78, Pyt76a, Ril78b]. Pa [Eme74]. Pacific [Dug71, Epp73, ABCP75, AZ70, BGMW74, BHSJ78, Bla73a, BP70, BKM78, CR72, CR76, CP75, Day79, ERVM73, ES75a, GWN70, GMT^{+}72, HP70, KD79a, LS77, LB77, Nat70, Per76, RE73, RAC71, RG76b, Tho70a, Tho70b, TD72, Ven74, WH71, ZH71]. pacificus [Fro72, Fro75, Fro77, Hur73]. paged [Cul73a, Edm70a]. Pairing [Kes72, Wan72a, Wan72b, Pyt72]. pairs [PH74]. Palaemon [SKF74]. Paleobiology [Pai75b, Dee74]. Paleolimnology [Bra71]. pallidus [GC72]. Palmer [Str75b]. Panama [GS73c]. paperback [Eme76, Smi76b, Kra79]. Papers [Woo79, Pug70]. Papua [Mac77]. paralimnology [Mor73]. parameters [MP77a, Tho70b]. Parana [Dep76]. paraquat [FR77b]. Paris [Cus76a, Edm70a, Gro76a, Koh79, Mar78a, Mil78, Par78, Pom74]. Park [Koh71a, Sta75, MB71b]. Parker [Gar74]. Parsons [Ril74]. Part [And71, Gol77a, Pla78, Ban73, Edm76b, HES75, MM73b, Smi74a, Tyl71, Enr77b, EH77, KOW79, OS78, SO78, VWB77a, VWB77b, WK79]. Partial [PC76, CK74, PI72]. Particle [KMP^{+}75, NN76, OSD76, BL72a, Boy76, HM79, Har72, Joh76, NN79, Owe74, She79, TSJ78]. Particle-size [OSD76]. Particles [Fro72, CM79, Doy79, FPO79, Fro77, Gar77, HDS79, LH73c, MCB79, PBHC70, SPS72, SSP73]. particular [Ben79, BH70, Edm70a, GMV72, SSP73]. Particulate [KLG71, LS77, Pou73, Wan74, All76, BL79, CM79, GS74, HS75, HL73b, HHPP70, Jan73, KLAS73, LhC73, MP78, NL72, Pet78b, Pou78, PBK70, QS70, Rei72b, Sha74b, SRSW78, TP74, VTV^{+}78, WD78]. particulates [Cha79b, GB72a]. Pass [SA76]. patches [ST79]. patchiness [Ril76, Wie71]. patchy [Dag77]. Patten [Rig76]. Pattern [Eag71, RHW79, SH78, Sha77a]. Patterns [MM75, Ric79, SH73, CGPD79, Dod79, Gro79b, HS77a, JP75, yLKR75, LJ79, MA76, PM77, Pro71, RS79, Ser74, TS78, Wal76b]. Pb [AW75, EA79, VBT79]. Pearl [GS73c]. peat [Red75]. Pegea [HG76]. Pelagic [Cul70, RG76b, SSL78, CC74, Fou73, GM74, HG76, KM72b, SP78]. Pelamis [DE71]. pellets [FP77, FS72b]. penetration [SC76a, The75]. Peninsula [CL76, WRO73]. Penn [Par70, Par75]. Pennant [SKF74]. Pennsylvania [Hoo74]. percent [FHV79]. Perfil’ev [Bro70]. performances [Str77]. Pergamon [Ril74]. period [GS72a, iK75, RB71]. Periodicities [PP76]. periodicity [Leh76a, Til73]. periphytic [BB70, WEMD72]. Periphyton [Ert71, Toe71, AGF71, CR70, LJ79]. Perkins [Man75b]. permanent [BKS70, CA72]. Persian [Cha74, SSA72]. Perspectives [Par75]. persulfate [DSC77]. Peru [KWD75]. Peten [Cov76]. Petrol [Gro71]. Petroleum [Gro74a, Gor79a, Gor79b]. Pettaquamscutt [GP72a]. Pettijohn [Whe73]. Ph [Rzo79, Pyt72, BY70, BY73, BYS77, BYR74, Dav74b, GCL79, GF75, Gie70, GGW78, Kes72, KO76, Kro71, PIM74, Wan72a, Wan72b, Wei73b, Whi71, WCR71, Won79, ZY72, Zir75]. pH-Dependent [ZY72]. phase [McL74, McL75]. phased [CAE78]. phases [EB79]. phenolhypochlorite [GG75a]. pheo [GMV72]. pheo-pigments [GMV72]. pheophytin [Mar77, TKH75, TKH77]. pheophytin-a [Mar77, TKH77]. pheromones [Kat73]. Phil [And71]. Philadelphia [Fag72, Fra72, Hyn71, Liv76, Par70, Ton72]. Philippines [Lew73]. philosophy [DH78a]. phorbin [BDM77]. phorbins [DBM77]. Phosphatase [YYP72, RS79]. phosphatases [AP76, Ber70b]. Phosphate [MB70a, Per76, TLT77, Wil72a, BH78, BHK78, CS73a, FC70, GB74, HAG^{+}72, HS74, LR74a, Leh76a, MBR78, OW72, Pyt71, RK79, Rhe72, Sha73b, TKK76]. phosphates [HPM72a]. phosphomonoesters [TLT77]. Phosphorus [BB75, BL74, Cul70, FC70, Imb74, MBN72, OS78, PR73, Pet75, Rei72a, SO78, TCTW75, TBD77, AP76, Ber70b, Cav76, Cha79a, D’E77, DR74, Dil75, Edm70a, Edm74a, Fit70, FH79, HL73b, IE78, KD79b, Kir75b, KH76, LR74a, LJ79, Meg72, Mey79, MC78, Per76, PL73, PBS74, Ric79, RAH77, SRSS74, Sen78, Shu78, SHW^{+}70, Woo77]. Phosphorus-33 [FC70]. phosphorus-chlorophyll [DR74]. phosphorus-limited [Per76]. photochemical [GG75a]. photography [DL74]. photometry [MSJG74]. Photorespiration [Hou74, HW78]. Photosynthesis [MD71, Mel79, AGA78, And76, BP74a, GIK77, JP76, KHC74, KS70, Kro75, Lor79, MW77, Meg72, MCSK79, MS74b, SR77, Sen78, Wal76a]. Photosynthetic [CS78, ES75a, GM79, MK74, TeI70, ATM74, Cas73, CKS77a, MP77a, Tho70b, TD72]. photosynthetically [Boo76]. phototrophic [Cas73, JS79]. Phragmites [Wet75]. phycoerythrin [MY70]. Phylogeny [Koh71c]. Physalia [Woo71]. Physical [Bou75, Fan77, yLKR75, Ecc74, LaF77, Mof73, RAC71, SG71, Smi76b, TLP78, TK76, TTC^{+}77, CKGS77, Lee77b, Sch76]. physical-chemical [SG71]. Physicochemical [Eck70, PP76]. physics [Aro73, Csa75]. Physiological [TLR^{+}70, MYY71a, PBS74, Liv77b]. physiology [TR76]. Phytoflagellate [Aar71, BA76]. Phytoneuston [Har73b]. phytoplancton [Cus76a]. phytoplankters [CTB72, CRW72, SN71, WNS74]. Phytoplankton [CCK71, ECHH^{+}71, KHHG^{+}72, LC71a, Mar70a, McA71, Meg72, SG76, SS78b, Aar78, AZ70, AC76, Ban74a, Ban74c, Bat76, BL72a, BP74a, Bie75, Boo75c, BG76, Bro78a, Bro78b, BS78c, CP79, CL79, CA72, DDS78, DB76b, DM71, ERHC79, ERH79, EB79, Fal75, Fog77, GF75, GM79, GWN70, Gol77b, HS77a, HK74, Hul76, ICO^{+}74, JM78, JG74, Jas75, JP76, Jon78, KZ77, KA74, KOW79, KD79a, KM73, KK76b, Lam72, LBL75, Lor79, MP77a, Mac78b, Mal71, Mal75, MCB79, MC79, ME72, McC72, MTT77, MB79, MW77, MS74a, MK74, MY70, MYY71a, MY72, MB76b, NG78, Nor75, OS78, Pac73, PT73, PG74, Pec70, PBS74, PI73, PF73, PRJR73, RSEG78, RK79, RHW79, Rie78, SH78, SRSS74, Sch78a, SCL^{+}71, Sha77b, She79, SD78, SMD76, SCA77, SH77]. phytoplankton [SWM^{+}78, TLT77, TCTW75, TP78, TLP78, Tho77, Tho70a, Tho70b, TD72, Til73, TPG77, Tit75, TK76, Tyl75, VHMW74, Wal75, WNLS77, WK79, Hor76]. Phytoplankton-zooplankton [Mar70a]. pictorial [Sta75]. Piece [MNM71]. Pielou [Koh77]. Pierce [Liv78]. Pigment [LC71b, TGK^{+}78]. Pigmentation [Sec71, Hai79b]. pigmented [Fou70, Fou71, Fou73]. pigments [GMV72, GS72b, MY70, SG70, SG72]. Pilsbry [Hur73]. piston [Cub70, Red75]. plancton [Cus76a]. planctonicis [Hut74]. planktic [Rod74]. planktivorous [CB75a]. Plankton [Rod74, Ano73g, BS70, BKS70, BFM76, Cow70, Dev79, Dod74, ERVM73, FR76, Gal75, Ham75, HB72a, Jef70, Kal72, KTTF76, LL77, Lyn77, MTT77, ME74, Ore79, PHR71, PM77, PS73, PP77, Ril76, SSL78, Tur78, Edm76b]. Planktonic [Ber70a, BP72, Fro72, NN71, RO70, All76, AF74, BB76, BDW76, FF79, GCB76, HC75b, Kat73, Kil71, KK78, LG70, MCSK79, ME74, MSF75, Paf76, Rod74, Str77, Wal73, WEMD72]. plants [CK79, GP77, GG81, Hai76, Hou74]. plastic [Ris73]. Plate [Mac77]. plateau [Liv71]. platensis [Mel79]. Plates [Fot70, Pai70, Ris73, Wet75]. platinum [Whi74a]. plea [Boy76]. Plecoptera [MBB73]. Plenum [Bou75, Hun75, Kap78, Kin70, Mac79, Man75a]. plots [Tho77]. plume [MC79]. pneumatically [WMM74]. Pocillopora [Pil74]. point [HW78]. polarographic [Dav77, KLM74]. Polarography [MB71a]. Policy [Edm74b]. pollen [DB73, Dav73, Liv71, RMG70]. pollutant [GB72b]. pollution [CMB70, Hyn71, Smi76a]. polychaete [Nic74]. polyedra [Mac78b]. polyethylene [TP73]. polymerase [SM79]. polymeric [WK77]. Polymorphism [OV78, Hai79a, Zar72]. Polyphenol [Rey78]. polythionates [NCK79]. Pond [Kim73a, Kim73b, WWH75, WWHH77, Dun70, Eck70, Geh74]. ponds [Dod75, OV78]. Ponyi [Hor75]. pool [BBBS79]. pools [Mos70]. poor [Tho70b]. Population [Aar71, BSS78, Pai71, Sau72a, GMCG78, GMTA79, HSK76, KK78, LBL75, Mac78b, Mul74, RC74, SW73, SK72, WC73]. populations [AC76, BGMW74, BL72a, Ben79, CS74, DM79a, DM71, Gre79, HS77a, HA72, KHHG^{+}72, LG70, McC72, Phl70, Rie78, RHG73, Ste76, Tho70a, Tyl75, Wel70, Wil73, YY79]. Pore [BY73, BST70, DCP73, Hes76, Hoi78, LH73b, Mar74, Nri78, RG76a, SEPS74]. porpoises [Sch77a]. Port [Edm71a, Gol74a, Gol74b]. Portable [BD72, RW71, Wal73]. Porter [Cle72]. Positive [KS78]. possible [And75, BK72, Cas73, Mar70b, MH72, Sch72]. Postdepositional [Man73b]. postembryonic [Sta76]. postglacial [DBM77, SG72, SL70, Wet70]. postlarval [Omo79]. Potamogeton [Geh74]. Potential [Far75, Sau72a, CS73b, Dav74b, Pom79, Sma74a]. potentials [MO72]. Potentiometric [SWM^{+}78, Whi71]. Potomac [MEW^{+}78]. Potter [Whe73]. Powell [Rey78]. power [BHD76]. power-independent [BHD76]. Practical [Sil79]. practique [Pyt76a]. Praeger [Ril78b]. Prague [Sch73]. prawn [SKF74]. precipitation [GCL79, Rey78, VTV^{+}78, WH78]. precipitations [SE78]. precise [Hoi78, MWW77]. Predation [Con71, Dod75, Kim79, Lev71a, Lod74, Lyn79, All78, Dod70, Dod74, Fed75a, Ker77, Mal70, MB79, Mul79, OV78, OS79, Ste73a, Wel70]. predator [Egg78, LL79, Ste76, ZS76]. predator-prey [LL79]. Predators [Zar72]. Predatory [Lan78, AF74, Ker78]. predicted [Con75, Leh76b, YY79]. Predicting [Bax77, TSJ78, DDS78, Lan75b]. predictions [MA76].

Predictive [Sch78b]. preferences [MTT77]. preformed [Pyt71]. Preliminary [Ano74d, LLS75, SWLH78, SAH72]. Prentice [Mil73a, Mil73b]. Prentice-Hall [Mil73a, Mil73b]. prepaid [Zar78]. Preparation [HMWW72, PS73]. Presence [Ogu70, GB74, Jas75, MB76a]. Present [SN75]. presented [Hyn73]. Preservation [BSB71, BS70, HH73, Mil78]. Press [Aro73, Bri72, But76, Csa75, Cum71, Day79, Dug71, Ebe72, Gol79, Gor79a, Gor79b, Gor78, Gro76b, Gro78, Har78a, Hed70, How72, How78, Hun75, Hyn78, Ich71, Kin70, Koh71a, Koh72, LaF75, LaF77, Lun79, Man75a, Man70, Man75b, Mas70, Mil78, Mil77, Ori73, Pai75a, Pom79, Pyt76b, Rig72, Rig76, Ril73, Ril74, Smi76c, Sta74b, Sta75, Stu71, Wal74b, Woo73]. Press-Butterworths [Koh71a]. Pressure [PA70, Sec71, Bel78, BT70, MCHP73, Wal73, Whi74b]. pressures [CK74]. Prey [Con71, FR77a, Ker78, LL79, SF75, Zar72]. Primary [BSB77, CK79, GD74, GBB72, GG81, HLH71, MF71, Nai76, Nor75, WN71, CE79, CKS77a, Coo73b, CP73, Fee76, Fee79, FG75, GV73, LB79, Lor70, Mal71, Mcm73, Par70, Pat80, Pom74, PBK70, Pug73, Smi79, TB75, The75, WF72, Man76]. probable [BWP73, Mor71]. probe [CWM75, CP73, FR77a, KLM74]. problem [Leh76a]. Problems [Enr77c, MO72, BS78b, HMA^{+}75, Mac78a, Mar77, Par75]. Proc [Bow79, Gus79, Hob74, Hyn76]. Procedure [WWR70, Bro79, CS73a, CACB75, FL75]. procedures [Ham75, JDL78]. Proceedings [Ano75n, Ano76e, Ano77g, Ano78g, Bar71, FW74, Ano70g, Ano71i, Ano73f, Lau70, Sau72b, Ano74b, Ano79d]. process [PP76, Leh75]. Processes [Pom79, BB73, Gor78, GC73, How78, Phl70, TLP78, TK76, DH72, Hyn78, Ril74]. processing [Rau78]. produced [Bro79, Sch77a]. producers [Mal71]. Production [Ban74b, CC74, GC74b, GBB72, HLH71, KA71, MEW^{+}78, MF71, Mur71, OH72a, OH72b, QS70, Smi72, SK73b, TP73, VTP76, VVT76, Ben79, Ber70a, BSB77, CP77, CK79, CE79, CKS77a, CKS77b, CSHE78, ERH79, Fee76, GD74, GB79, GG81, HCW70, Hob73, Jac77, JPW79, JP78, Kal72, KK76a, LP77, Lea70, LB79, Lit73, MC79, Mcm73, Mul74, Nai76, Par70, Pet78b, Pom74, PBK70, Pug73, RB74, RLS75, Ric78, Rig72, Sch78a, SSP73, Smi73b, SM73, SA74, SG76, ST78, WC73, WL78, Wet70, WF72]. Productivity [GBB72, WN71, Bri72, BM73, BM75, CPOT79, Coo73b, CP73, DP78, Fee79, FG75, Gid77, GV73, HNT75, Hou74, KK76b, KK78, Lam72, Lit73, Lor70, Man76, Pat80, PF73, Pug70, Ric75, Rog79, SD78, Smi79, TB75, The75, Ton72, Hyn76]. Products [Kro72a, BLM74, Rie78, WD79]. Professors [Cul73a]. Profile [RC71, WAA72b, Gag76]. Profiles [Ber72, Dav74b, Ids72]. profundal [Dav74a, Dav74b, DG76a, KMR78]. Program [FW74, Hyn76]. progress [Ril78a]. Project [Fra72]. Properties [DVCP79, Bla79, KOW79, STG73, Tab79, TeI70, VWB77a, Bou75]. proposed [Fra72]. Prospects [AR73, Par75]. Protein [CM72]. proton [WK77]. proton-metal [WK77]. Protonation [HK79]. protozoa [GCB76]. Pruter [How72]. pS [Whi71]. Pseudocalanus [Pou73]. Pseudodiaptomus [SB79]. pseudoharengus [War72]. pseudonana [Sma74a]. Psychrophilic [PA70]. pteropod [TKC73]. pteropods [PT70]. Publ [Bar71, Bou75, Bow79, Car75, Cas71, Col72, Edw77, Fag72, Fra72, Gol78, Gol74a, Gol74b, Gro75a, Gus79, Hob73, Hyn78, Kap78, Koh73, Liv77b, Mac79, Par70, Ril71a, Sch77a, Smi76a, Smi78a, Ste79b, Ton72, Wal76a, Woo77]. public [Man72b]. Publication [Ano77b, Ano78c, Edm77, Tay78]. Publishers [Man73a, Ril78b]. Puerto [CMB70, Rog79, SCL^{+}71, AFK73]. Puget [Woo73, Dux75, FD72, GM77]. pulex [Arn71, BW70, BR74, Cro73, KO76, Sta76, Sta78, Ste75, WWR70]. pulicaria [Sno72]. pump [IR73]. pumping [JFH73]. Purser [Cha74]. push [Mil73c]. push-net [Mil73c]. pyrifera [Jac77]. Pyrocystis [BS77, KS78, RS79, SSM76]. Pyrodinium [Mac77].

Quadrangle [Edm72b, Ill75]. Quadrangle/The [Ill75]. Quanta [CC75a, MS74b, Tyl73]. quantifying [Bro78b, YK73]. Quantitative [Ban79, Ert71, IR73, LC71b, MB76a, PP77, SL75, CA72, KK76a, LG70, Mil73c, Rup72]. quantitatively [BB77, TWPC75]. Quantity [DB71, Dee71]. quantum [Ban74b, Boo76, Tyl75]. Quaternary [Liv72a]. quota [BH78, TCTW75]. quotas [Shu78].

R [Aro73, Bri72, Bro70, DH78a, DH72, Dug71, Edm70a, Edm72a, FW74, Gar74, Gol79, Gro71, Gro79a, Hun75, Hyn73, Koh79, LaF77, Lau74, LR74b, Liv72a, Liv72b, Liv76, Liv78, Mac79, Man75a, Man76, Mil73a, Mil73b, Mil77, Mof73, NC78, Ori77, Pai75a, Par70, Pom79, Ril74, Sha73a, Smi76a, Sta72a, Ste79b, Stu71, Val71a, Whe73]. R. [Cus77]. R.I. [DD78]. radiation [Ids72, Koh71b, Lor79, SC76a]. radioactive [Ram76]. radioactivity [SS77]. Radiocarbon [Pet78b, IBM76, SH73]. Radioecological [LS72]. radiographs [EA70]. radiometers [Ids72]. radionuclides [LS77, NWEG75]. radium [EGTA70]. Radok [Ste79b]. radon [IE78]. Rafter [McA71]. rain [GH79]. Rainbird [Sch77a]. rainfall [Pet78a]. Rangia [Wol71]. Rapid [SSA72, AGF71, GIK77, Hoi78, Loh74, Sal79, YK73]. rapidly [Gar77]. Rate [Cro73, GC72, MB70b, Ver75, BS74, Cam75, Chi71, Dil75, Doy79, KOW79, KO76, Nor76, Pal74, Par77, RLS75, SR79, YY79]. Rates [Han71, AW75, Bot75, CS78, Cas72b, Dod75, DDS78, Edm72c, Eme75b, EC72, FS72b, Fro78, HG76, HM79, Har77, HP77, JG74, KHC74, KD79a, Lan75b, MS74a, MCSK79, MK74, PG75, PR78, SM79, Sen78, SCA77, SSM76, Taf76, Thr79, Tit75, Wel74, Wet70, WBM^{+}77]. ratio [AGM70, Ban74c, CL76, PF73, TCTW75]. rationales [LBL75]. ratios [AFK73, BNV^{+}74, CB73, DDS78, Mal70, PSK78, Rhe78]. ray [Bas75]. Reaction [Wic71, GG75a, GTF78]. reactions [Eld75, RH77]. Reactive [Cul70, CL73, CO75]. reactivity [BLL70]. Read [Gro71]. reappraisal [Chu72]. Rearing [Kac71]. reason [Boy76]. rebus [Hut74]. rebuttal [PT74]. Received [Ano71d, Ano70c, Ano70d, Ano71e, Ano72f, Ano73d, Ano75f, Ano75g, Ano76d, Ano77f, Ano78f, Ano79e]. recession [Gar75]. record [BKS70, Liv71, Smi73b, Pai75b]. recorded [SSA72]. recorder [Ano73g]. Recording [EA70, WCR71, Tem73]. Recovery [Dun70, McL74, Taf76, CM79, McL75, Pae78]. recreated [Cas72a]. Red [Bla77, HJG71, FT73, MA76, MSR^{+}75, TS78, Col72, Eck70, JB71]. Redeposition [Dav73, BP72]. Redfield [ABGCP75]. Redox [WAA72b, Dav74b, MO72, Whi72]. reductase [Har73c, Pac73]. Reduction [SH71, CR72, CB76, HT79, LWR73, Pil74, RB71, RB70, VHMW74]. Reef [Rog79, Smi73b, BM74, D’E77, DM79b, HC79b, Kin78, KD79b, LLS75, Lit73, Mey76, MD78, PP77, QS70, RLS75, RMS77, SM73, WDW^{+}75, Ill75]. reefs [Ano71h, GS73c, Koh79]. referees [Edm72d]. reference [Bat73, BH70, Edm70a, SSP73, WH78]. Reflectance [KOW79, WK79]. refraction [CTB72]. regeneration [Ber77, MBR78, Sch74]. regime [Bot75, CR76, CSHE78, Lew73, Mor75, Sta76]. region [Bla77, KMP^{+}75, LS72, SMD76]. regional [LaF77]. regression [ABGCP75, OS78]. Regressions [Bla73a]. regulate [MS74a]. regulating [Sch78a]. regulatory [Mor71]. Reinheimer [Hob79, Sta72b]. Related [MD71, Dal74, KOW79, TP78]. Relation [GKA73, Hai76, MS74b, MF71, CO75, Cre75, Dev79, Fee79, Geh74, GS73c, Hil74, Pae74, Pat80, Pet78b, SSHV76, SS78b, Str76, TLP78, VHMW74, WEMD72]. Relations [RC71, EP71, Ril75]. Relationship [AGA78, CPOT79, MYY71b, Pet78a, AGM70, Beg76, DR74, Hai79b, JP76, Kir75b, Lan75a, RHG73, SSL78, Ton71]. Relationships [GBB72, Hu78, TGK^{+}78, ABCP75, ABGCP75, DD78, GB79, GLSD74, Mar70a, SN75, SA71, Whe77]. relative [HM75, Mal71, Wal73, WBM^{+}77]. Release [AZ70, Fie72, Jer70, Lam78a, MM79, OW74, PR73, BB75, D’E77, HIJ78, MEW^{+}78, MD78, SS78b]. released [DM79b, PL73]. releases [Hit78]. Reley [Stu71]. reliability [Cul73b, Pen73]. relicta [BG78, Gro79b, LL73]. remineralization [Har78b]. Remote [PK74, SB78b]. Removal [DGEB70, BC74, Coo73a, CSHE78, Fit70, GH79, MBR78]. removal-summation [CSHE78]. Rep [Fan77]. Repetitive [HSC79]. replicate [KAG77]. replicate-sample [KAG77]. replicated [Ris73]. Reply [Bal76, Ban75, BS78a, Cus77, LP78, Meg75, NC78, Sha78, SK73a, SCA79, Tho78, Vly77, Wan72b, DH78a]. reported [JM78]. Reports [Val74b]. represented [BG71]. reproduction [Arn71, Hai79b]. requirement [AP76]. Res [Bri79, Fan77, d’A75]. Research [Koh79, Edm70a, Mac78a, Ril73]. researches [BS78b]. reservoir [Cow70, HSK76, RMG73]. reservoirs [BM73]. Residual [GB72c]. Resistance [GA75]. Resonant [TC74]. resources [Koh71b]. respect [MYY71a, PBS74]. Respiration [And76, Bel78, Big77, KP75, Kro75, MD71, RSMC76, Bro79, Dev79, DG76b, HP77, KHC74, SR77, Smi73b, Tur78]. Respiratory [Chi71, CP79, AS77, DP78, KP75, PHR71]. Response [FR77a, And77, Dod74, GIK77, LF76, MP78, PT75, SG76, Swe78]. Responses [SRSS74, BL76, OS78, SO78, Ste75]. resting [And76, Coo71]. result [Hoo74]. resultant [PB77]. Results [KWD75, SMD76, Smi78b, TTC^{+}77]. resuspension [SS70]. Retention [QM71, HM79, Mcm73, MD78, NN76, NN79]. retrieval [Pai75b]. retrieved [NWEG75]. reusable [Lan79]. reverse [HDS79, HHPP70]. reverse-flow [HHPP70]. Review [And71, Ano74b, Ano79d, Aro73, Bar71, Boo75a, Bou75, Bow79, Bri79, Bri72, Bro70, But76, Car75, Cas71, Cha74, Cle72, Col72, Csa75, Cul73a, Cum71, Cus76a, Day79, Dee74, DH72, Dug71, Ebe72, Edm71a, Edm72a, Edm72b, Edm74a, Edm76a, Edm76b, Edw77, Eme74, Eme76, Eme78, Enr76, Epp73, Fag72, Fan77, Fer70, FW74, Fot70, Fra72, Fry71, Gol78, Gol74a, Gol74b, Gol79, Gol77a, Gor79a, Gor79b, Gor78, Gro71, Gro74a, Gro74b, Gro75a, Gro75b, Gro76a, Gro76b, Gro77, Gro78, Gro79a, Gus79, Ham79, Har78a, Hed70, Hob73, Hob74, Hob79, Hor75, Hor76, How72, How78, Hun75, Hun77, Hut72, Hyn71, Hyn73, Hyn76, Hyn78, Ich71, Ill75, Kap78, Kel75, Kil76, Kin70, Koh71a, Koh72, Koh73, Koh74, Koh77, Koh79, Koz70, Kra79, LaF75, LaF77, Lau74, LeC77]. Review [Lee77b, Leh75, Liv72a, Liv72b, Liv76, Liv77a, Liv77b, Liv78, Lun79, Lym73, Mac79, Man72a, Man75a, Man70, Man73a, Man75b, Man76, Man79, Mar78a, Mas70, Mil73a, Mil73b, Mil78, Mil77, Mof73, O’C71, Ori73, Ori77, Pai70, Pai75a, Pai75b, Par70, Par72, Par73, Par75, Par78, Phl73, Pla78, Pom74, Pom79, Pyt76a, Pyt76b, Rig72, Rig76, Ril71a, Ril73, Ril74, Ril77, Ril78a, Ril78b, Rzo79, Sch77a, Sch73, Sch75b, Sch76, Sha73a, Sio77, Smi74a, Smi75, Smi76b, Smi76c, Smi76a, Smi78b, Smi78a, Sta72a, Sta72b, Sta74b, Sta75, Ste79a, Ste70, Ste79b, Str75b, Stu71, Tea78, Ton72, Tyl71, Val71a, Val74b, Wal74b, Wal76a, Wet75, Wet76, Whe73, Woo73, Woo77, Woo79, Wu79, Zar78, d’A75]. Reviews [Ban73, Edm70a]. revision [MLKT75]. Rhabditis [TLR^{+}70]. Rheumatobates [LC74]. Rhincalanus [RSMC76]. Rhizosolenia [HMS^{+}77]. rhizosphere [OT77, TVB79]. rhodanese [NCK79]. Rhode [SC70]. Rhodesia [Beg76]. rhythmically [Ant77]. rhythms [DG76b]. Rica [RAC71]. rich [Ant77, Høj78, Tho70b]. Richelia [Ven74]. richness [MB79]. Ricker [LP78]. Rico [AFK73, CMB70, Rog79, SCL^{+}71]. Riley [Gro76b, Ban75, But76, Gol79, Mil77, Pyt76b, Smi76c]. rising [BS74]. River [How72, MC79, MG78, Rzo79, TGK^{+}78, Wol71, CGPD79, GIK77, GB72b, HD70, KHC74, KC79, LJ79, Web74, Ada76b, Cow70, CR70, Dep76, GP72a, GC73, LeC77, NS78, PCWR70, Ser74, Won79, Sio77]. river-dominated [KC79]. rivers [Bax77, GT74, Man79, PN79, SR77, WL79, Bri79]. Robertson [SC76a]. rock [BB77, PN79]. rock-filled [PN79]. rocks [Liv72b]. Rodewald [Wet75]. Rodewald-Rudescu [Wet75]. Role [Kro71, Kro72a, AP76, AL78, DO75, ERHC79, GL75a, GMV72, HC79b, Hob74, HG72, Hyn78, JCK70, KA71, Lud74, McL75, Mey79, RFFH76]. Rosalina [ML70]. rosea [Kib71]. Ross [Gro74a, Woo79, Eme74, Par75]. rotary [HB72a]. rotifer [Gil71]. rotifers [Edm76b, GF70b, LG70]. Rough [Ert71]. Round [BDM77, DBM77]. Rowe [Lau74]. Rudescu [Wet75]. Rudloe [Koh73]. rules [MJH77]. running [Cum71]. Runoff [Gar75]. Rupert [DO75]. Russak [Eme76, Eme78, Hor76, Kra79, Phl73, Smi76b].

Ruttner [Edm76b]. Ruttner-Kolisko [Edm76b]. ry [KB71]. Rzóska [Fra72, Man79, Sio77].

S [Fry71, Gro78, Gro79a, JDL78, Koh72, Liv78, Man73a, Ori73, Ril78b, Sch75b, Ste79a, Zar78, Dav77]. S. [Sha78]. S.A [Ori77]. sabkhas [GG78]. saccharum [LI76]. safety [Ham75]. Sagami [OT72]. Sagitta [Zo73]. Saint [Kil76, Ada76b]. Sakshaug [Gro79a]. Salanki [Hor75]. salina [SM79]. saline [Ahr75, HK73, MK74, SB73, WWH70, Wal75]. salinities [LH73b]. Salinity [Car70, CM72, MD71, PA70, DE71, HES75, Huy77, MM73b, Mos76, SC70, WP76]. Salinometer [RW71]. salmon [Col72, Egg78]. Salpa [Sil75]. salps [HG76, HC79c]. Salt [Ano74b, Rei72a, SG76, TL75, Boo75b, Cam75, CVV78, CM77, ES75b, GD74, Gal75, GP77, GK74, Hai76, HT79, KVT79, LJ73, LH73b, MPN72, San77, TVB79, Tur78, VTP76, VTV^{+}78, VVT76, WWH75, WWHH77, dlCP75]. saltwater [MPN70]. Samalga [SA76]. sample [KAG77]. sampled [Chu72]. sampler [Bla79, Chu72, CRA72, CP70, GDHZ79, GO78, Hes76, JS73, Kro72b, MS70, Mak72, Sho70, SH72, Web74]. samplers [CM79, DC75, Gar77, Gar74, HP74, RH75]. samples [BS70, BKS70, BSS78, Bro79, Cas71, Cli72, CL76, DSC77, Deg73, Fag73, HMWW72, HB72b, IJ79, KS70, Ore79, TTW75, WD79, WW74]. Sampling [Ano73g, Ert71, FL75, HMA^{+}75, KSGM71, Sch71, Smi71, Wan74, And77, Cub70, CSHE78, DLMS76, GT74, HC79a, HB72a, HSC79, IR73, KWK74, LG70, May76, McA71, Mur76, O’C74, PN79, PP77, RHR71, Rup72, Sch70, SA70, SS70, TWPC75, Wie71, dO74]. San [But76, Dee74, Gor79a, Gor79b, Zar78, Rog79, Smi74b]. sand [Cub70, DV78, Sha74a, WM78, Whe73, Koh72]. sandstone. [Whe73]. sandy [Mak72]. Sanit [Cul73a]. Santa [KLAS73, SG71]. Santolini [Hob74]. Santoro [Dug71]. sardine [Beg76]. Sargasso [CC74, CM75, CP77, Cul70, Dee71, DB71, Gag76, GB72a, LB77, SSP73, SS78a]. Sargassum [CC74, Cul70]. Sargent [Cle72]. Sarnoff [Edm72b]. Sars [RC72]. Saskatchewan [Kos70]. Satellite [SE78]. saturated [Ban74a, Ban79, Sen78]. saturation [CS70, Fal75]. Saunders [Hyn71, Liv76]. saycia [Fre71]. scalar [Boo76]. scale [CWM75, HS77a, Wie70]. scanning [Bas75, PS73]. scattering [GB72a, PBHC70]. SCDP [DBM77]. Scenedesmus [Gid77]. Schacht [Geh74]. Schilfrohr [Wet75]. Schindler [Tho78]. Schlieper [Ril73]. schmeil [Ton71]. Schmidt [Enr76, Ven74]. Schmidt-Koenig [Enr76]. Schnees [Fot70]. Schneider [Fan77]. Schopf [Dee74]. Schuett [KS78]. Schweizerbart’sche [Fot70, Wet75]. Sci [Bar71, Bow79, Car75, Cas71, Fag72, Fan77, Fra72, Gol78, Gol74a, Gol74b, Gro75a, Gus79, Har78a, Hyn78, Liv77b, Man72a, Ori73, Par70, Pla78, Ril71a, Smi76a, Ste79b, Ton72, Wal76a, Woo77]. Science [Bou75, Leh75, Man75a, Man73a, Sch76, Edm74a, Kat71, Lau74]. Sciences [Hyn76]. Scientific [Edm70a, Hob73]. scintillans [FT73]. Scintillation [WWR70, WN71, IBM76, Loh74, Pug70, Pug73]. Scirpus [OW74]. scleractinian [JCK70, Sor73]. Scotia [BSS78, HPM72b]. screens [MCB79]. SCUBA [GT74]. SDSU [Zar78]. Sea [Eme74, Fie72, Gro74a, Gro74b, Woo71, BL72b, DOP77, DE71, Eme78, FW74, GP72b, Gri72, HAG^{+}72, HC79a, Ham75, Hoj75, HA72, JFH73, JPW79, JS79, KWD75, MG75, Mil71, MC78, Owe74, Par75, PK74, PKC77, Pin73, Pra70, QPB70, Ril78a, Ril78b, Sch77a, Sch72, Smi75, SWLH78, SSA72, TJ72, TJ76, Wal76b, AGM70, BAHA77, BYS77, Ber74, CC74, CM75, Car75, Cul70, Dee71, DB71, DB76b, Eck70, Gag76, GB72a, Gro75a, JB71, Kar78, LB77, SSP73, SS78a, SK73b, SP73, Bou75, Cha74, Gro79a, Kin70, Par72]. sea-ice [BL72b]. Sea-level [Car75]. Seabed [Fol71, Pam71]. seagrass [Smi77]. seagrasses [MGC73]. seas [All79, Ril78a, CP77, Kel70, Kel75]. seascapes [Sta75]. seashores [Day79]. Season [Sno72]. Seasonal [BP74b, BS78c, DP78, DB79, HPS75, Huy77, Jef70, Kim73a, Kim73b, KM73, Lan75b, Leh76a, MF71, PRLG75, RC72, RN78, WS70, BP74a, CGPD79, Fee76, GMV72, Glo78, Mar70a, SC70, SS78b, Wal75, HG72]. Seattle [Day79, How72, Hyn76, LaF77, Ril73, Woo73]. Seawater [BY70, Gie70, Kes72, MQ71, Ogu70, Wan72a, ZY72, And79, Ano75m, AP77, AFK73, BC74, Bra73b, Bul78, BH70, Car70, CS73a, CS70, Con70, CC75b, DS75, Deg73, DCC73, DVCP79, GDHZ79, GV78, GC70, GS74, Hoi78, JM78, KV73, KS70, LhC73, Ley73, LTB75, MB71a, MLKT75, Mcc70, MCHP73, Mey76, MSH74, NL72, Pla78, PC76, PME73, Pyt72, PH74, Pyt75, PA75, QM71, Ril77, RN78, SM76, SRR79, Sha73c, Sha74b, Whi74b, WD72, ZH70, Zir75, Fan77]. secchi [Hol70]. secondary [Ben79, Ton72]. secretion [BA76, TBD77]. sections [NWEG75]. Sedgwick [McA71]. Sediment [GKA73, Mul74, Nic74, RC71, Smi71, WAA72a, WAA72b, BL74, Bro79, Bul78, Cam75, Dav73, Dav74b, EA70, GLSD74, GE75, Gro77, Hak76, Har72, HC78, HB79, HPM72a, Hoi78, Jør77, KM72a, KAG77, Kir75a, LWR73, MC78, PE76, RG76a, STND74, SS70, SRSW78, TTW75, WL79]. Sediment-Water [Smi71]. sedimentary [Hak76, Liv72b, MG78, PSK78, Smi74a]. Sedimentation [AW75, AS77, And77, BSB77, Cha74, DB73, Eme74, Gro71, Lud74, WW73]. sedimenting [Fuh73]. Sediments [BY73, BCB71, FMJC72, Ant77, BPW76, BGZS75, BBG75, BST70, Ber77, Bou75, BB73, CW78b, Dal74, Dav74a, DG76a, DCP73, FP77, GL79, GH72, GS72b, Gor78, GG75b, Gri78, GM77, HIJ78, Hit78, Jer70, Kar78, KB71, KH73, KMR78, LB71, LJ73, LL75, Lud76, MO72, Man73b, MB77, MBR78, May76, MH74, Mey79, Mor71, NC76, Nri78, NCK79, PB73, Pra70, Pug77, RMG73, RAH77, Ros79a, Ros79b, RH75, SG70, STND74, Ser71, SEPS74, SEA71, Smi74b, SHW^{+}70, SM74, SW72, ST78, TVB79, TBOO75, TA78, VB75, VWB77a, VWB77b, WC76, Wei73a, d’A75, Hun75]. sediments-measurement [BGZS75]. see [Kos79]. seepage [Lee77a, Pet78a]. Seiche [DW72]. Seiches [War79, ESH73, GF70a]. Selection [Boy76, Doy75, Lam74, Ree74, TSJ78, WM78]. Selective [BLM74, BCK72, Dod70, Dod74, Doy79, HDS79, Man72a, Enr77b]. selenium [SOP73, Woo77]. Self [WWR70, KLM74, dO74]. Self-Absorption [WWR70]. self-contained [KLM74, dO74]. seligo [DD71]. Seneca [Ahr75]. sensing [GHM78, PK74, SB78b]. sensitive [MY70]. sensitivity [AM78, KVW79]. Sensor [Whi71, BYR74]. sensors [KL75]. separately [Edm70a]. Separating [BL72a]. Separation [HS75, She72]. September [Ano72c, Hyn76]. Ser [Car75, Hor76, Man75a, Pla78, Ste79b]. Sergestes [Omo79]. Serial [Pai70]. Series [Smi76a, KTTF76]. serratus [SKF74]. Sessile [GMCG78]. Seston [Mos70, AGF71, She72]. set [War79]. set-up [War79]. Settling [Cha79b, Ris73, BP72, CA72, Gar77, KLAS73]. Seven [BS78b, Arn71, Fal75]. several [ESH73]. sewage [DM71]. sex [Kat73, Mal70]. sexual [Bay78]. sexuality [Gil71]. Shagawa [MS74a]. shallow [Cha74, CB70, DV78, DL74, DCP73, HP76, Hor75, KWK74, Koh71b, Rei72b, Smi77, TWPC75]. shallow-bottom [Rei72b]. shallow-water [KWK74, Koh71b]. shape [Doy75]. shaped [TC74]. shapes [Leh76b]. Shark [Gro71]. Shebandowan [SC76b]. shed [LI76]. sheets [Ada76a]. shelf [HPS75, Huy77, Smi72]. shellfish [Pai70]. shells [CB73, GCL70, PT70, Str75a]. Shepard [Eme78]. shield [Kos70, RFFH76, RH78, Eme75b]. shielded [Ids72]. Shiels [Gol74b]. shipboard [ECHH^{+}71, Pam71]. ships [QPB70]. Shoestring [Ori73]. shoots [GP77]. shore [Eck70]. Short [And77, Den77]. shoshone [Mal78]. Shoten [Epp73]. Shrimp [Kac71, BG72, Omo79]. Si [DBH78]. Si-limited [DBH78]. side [Eld75, GG75a, The75]. sieboldi [Gil71]. Sieburth [Sta75]. Siever [Whe73]. sieving [Ped74]. Significance [HB72c, Ste76, Bay78, Ben79, Bla73a, CS73b, Enr77b, EH77, Gal75, Hai79a, KH73, Kno77, Nat70, Pea79, SG70, Ven74]. SIL [Ano73h]. silica [Kil71, Nri78, PRG78, VWB77a]. silicate [CL73, CO75, HAG^{+}72, TKK76]. Silicic [AC76, CAE78, NG78]. Silicon [LNS^{+}76, BLL70, DBH78, Dro78]. Silicon-32 [LNS^{+}76]. Silver [Bro78a, GS74]. Simbi [Mel79]. similis [Omo79]. Simple [KSGM71, NN71, BL76, BH75, BHK78, Cub70, GCB76, HB72b, Jac70, Loh74, Wil73, WK79]. simplified [LP77]. Simpson [Pom79]. simulated [SMD76]. Simulation [Ste79a, CSHE78, Rig76]. simultaneous [SCA77]. Sinai [CKGS77, CKS77a, CKS77b, JC77, KCS77, GG78, JKC79, Lev71b]. single [AFK73]. Sinking [FS72b, TK76, SR79, Sma74b, Tit75]. siphon [MS70]. Situ [Han71, BYR74, DS75, Fro78, Hes76, KK78, Koe76, KB76, Ore79, RSS78, SMD76, Smi74b, SWLH78, ST78, SSM76, TP73, Tyl75, dO74, NK72, Por72, RC71, Whi71, BY70, MO72]. six [BL74]. Size [AH77, EP71, Fro72, Sha73c, She72, Shu78, Sno72, WWR70, All78, BR72, Dod70, Dod74, Dod79, HK74, KMP^{+}75, LF76, LM70, Mal75, MCB79, MC79, Mal73, Mal78, OSD76, PT73, SPS72, SSL78, Van77, WP78, WM78, dBSA74]. size-dependent [WP78]. size-frequency [MCB79]. size-selective [Dod70, Dod74]. sizes [Boy76]. Skeleton [Str76]. Skeletonema [HS77b, YY79]. skeletons [CB76]. Skirrow [But76, Gro76b, Pyt76b, Smi76c]. Skreslet [Gro79a]. Skuja [Sau72a]. slack [GB72b]. slack-water [GB72b]. slides [CA72]. sliding [CA72]. sliding-chamber [CA72]. slope [CBN78, MC78]. slurp [WMM74]. Small [Wie70, BA71a, BL79, CWM75, Eme75b, Fed75a, GMCG78, HS77a, HB72b, Lam78b, MH72, PF73, QPB70, RG76a, ST78, WL79, WH78]. Small-scale [Wie70, CWM75, HS77a]. Smith [Van77]. snake [DE71]. Snoeyink [Cul73a]. snow [AL78, BA71a, ST79, Kro71, Kro72a]. Soap [Man73b]. Soc [And71, Hut72, Hyn78]. Society [Ano70g, Ano71b, Ano71a, Ano71i, Ano72a, Ano73a, Ano73f, Ano74a, Ano74b, Ano75a, Ano75n, Ano76a, Ano76e, Ano77b, Ano77a, Ano77g, Ano78b, Ano78a, Ano78c, Ano78g, Ano79a, Ano79d, Lau70, Pai70, Sau72b]. sockeye [Egg78]. Soda [KGP^{+}78]. sodium [GCL70]. soft [Cul77, Mur76]. soft-water [Cul77]. Softcover [Koh72, Man79, Ril71a]. soil [CW78a]. soils [Hai76]. sol [PRG78]. Solar [Høj78, BFMF72, Ids72, SC76a, CKGS77, CKS77a, CKS77b, JC77, JKC79, KCS77]. solids [Man75a]. Solubility [AP77, Ano75m]. Solubilization [HPM72a]. solubilizing [Loh74]. soluble [BH75, LhC73, Moe71, PL73]. solutions [DH72, IBM76]. Some [Ber72, CB70, CMB70, Dag77, Edm71b, Fie72, Gil71, GLSD74, GMT^{+}72, HSK76, Jon74, JS75, Liu70, Mac70, Mal78, Mar77, MPN70, MPN72, Mil70, MB76b, Sma74b, SF75, TK76, WP78, WL73, AGA78, And75, BA71a, BR72, Beg70, Beg76, BLL70, CRW72, CB75b, GB79, JCK70, KB71, LLS75, LT71, MBB73, OS78, PBHC70, PC76, Pom74, SS74, SO78, Sch72, Sor73, Ste73b, SSM76, Cas71]. Somes [FMJC72]. sonar [BSS78]. Sons [Gol77a, Gro77, Koh77, Lee77b, Pai75b, Ril78a, Wet76]. Sorokin [Hob73]. Sorption [EG79, RAH77]. Sorting [NN71, FF75, TTW75, WW74, Wor77]. Sound [RD76, Woo73, MLCG79, BS78c, CO75, Dux75, FMJC72, FD72, GB72c, GM77, MB77, MBR78, Ril78b, SC70, WWHH77]. source [ECHH^{+}71, Fra72, GH76, Mcm73, SD78]. sources [All76, Ten77, THDW79]. Souriray [Kim79]. South [Dug71, Har78a, Mos76, AFK73, FP79a, WRO73, OV78, WS70]. Southeast [Sch74]. southeastern [ES77, KMR78]. southern [BBG75, BP70, CJN72, ERHC79, ERH79, HES75, Jac77, JB71, Kim79, Las75, MM73b, RSEG78, Smi72, Won79]. sp [Bur70, MGHR75]. Spartina [CW78a, GK74, Rei72a]. Spatial [BHSJ78, DV78, PF73, RSEG78, TP78, TLP78, HS77a, HB72c, Sch71, Wie70]. Spec [Hut72]. Special [Hyn76, Kro71, Val74b, Beg70, Pom74, Wal76a, WH78]. Speciation [ZY72, And79, Wil78]. Species [Buz71, Hur73, LL79, ML70, Por72, Arn71, Bat76, Cov76, CB75b, DD71, DB76a, EP71, Fal75, Fed75b, GF70b, Hai79a, Hai79b, HM79, Hu78, Hul76, Hut70, KK76a, KK76b, KK78, Mal70, MB79, Pou78, RSEG78, RN78, WF72]. Species-specific [LL79, Fed75b]. Specific [DEWN76, Gri78, BGZS75, Bra73b, Fed75b, LNS^{+}76, LL79]. spectra [CB73]. Spectral [Whe77, DB79, Hoj75, MG75]. Spectrophotometric [War78, TKH75]. spectrophotometry [Car70, Rie78]. spectroradiometer [BD72]. spectroscopy [KOW79, WK79]. spectrum [DOP77, Rei72b]. speculations [Mor71]. speed [Enr77a, MLCG79]. Spencer [Edm74a]. spicatum [ATM74, CA79, MD71]. spilled [Con75]. Spin [WKW^{+}74]. Spin-up [WKW^{+}74]. spindle [TC74]. spindle-shaped [TC74]. Spinner [Pai70]. spiralis [TR76]. Spiroloculina [ML70]. Spirulina [Mel79]. sponge [Fro78]. Spongilla [Fro78]. spore [CNC73]. Sport [Fer70]. Spotte [Fry71]. spp [CP77, Ste76]. spring [Cas72a, HS77b, HS77c, HSC79, JC75, LS72, Pec70]. Springer [Boo75a, Cha74, Enr76, Hob79, Hun77, Kel75, Koh74, Liv78, Man76, Smi74a, Smi78b, Ste79a, Whe73]. Springer-Verlag [Boo75a, Cha74, Enr76, Hob79, Hun77, Kel75, Koh74, Liv78, Man76, Smi74a, Smi78b, Ste79a, Whe73]. springs [Cas73]. squeezer [RG76a, STND74]. squid [Bel78]. Sr [KB76]. St [Mon75, Val74b, d’A75]. St. [HPM72b]. Stability [Bra73a, CM77, Cul77, Ids73, LR74b, OL74, Wal74a]. stabilization [KA71]. stable [FHV79, Hai76, PSK78]. staeger [KA71]. stage [Mor75]. Stages [GC72, Lan75a, NN76, Omo79]. stagnating [SAH72]. standard [GHM78]. Standing [Por72, Cow70, FR76, GLSD74, Kro72b, Nai76, OS78, Sch78a]. Stanley [Eme74, Gro77]. State [Zar78, Ban74a, Ban79, Bie75, Car77, Dev79, GM78, Jan74, Koe76, LJ73, MYY71a, PBS74, SB78b, TP74]. Statement [Ano74a, Ano75a, Ano76a, Ano77a, Ano78b, Ano79a, Ano71b, Ano72a, Ano73a]. States [Edm70a, Kar78, WWK^{+}77, And77]. Static [LR74b, OL74, CM77]. Station [MM75, AFK73, Tab79]. statistical [Cas71]. statistics [RC74, Ven71]. Status [GKA73, Edm70a, KGP^{+}78, TK76]. Steady [Bie75, GM78, Jan74, Ban74a, Ban79, LJ73, LH78, TP74]. steady-state [LJ73]. Steedman [Mil78]. Steele [Ril78a, Man70, Wal74b]. Steemann [Wal76a]. Steemaxn [LaF75]. stenolepis [FM78]. Stern [Csa75]. Sternberg [Mil73a, Mil73b]. sterols [Gag76, GN79]. Stevanage [Bri79]. Stevenson [Dug71]. Stewart [Sta74b]. stimulation [BLM74, RK79]. stirring [MY72]. stock [Nai76]. Stoddart [Koh79]. Stoichiometric [Ber77]. stomach [BSS78]. Stones [BA71b]. stony [Pec79]. storage [Deg73, FP79b]. stories [Man79]. strains [Sta76]. Strait [LaF77, CT70, CO75, MF71]. Straskraba [Sch73]. stratification [BDW76, JP75]. Stratified [TL75, Bla73b, FP79a, JKC79]. Stratigraphic [Dav74a]. Stratigraphy [SG72, BDM77]. stream [All78, BL79, Bot75, Chu72, Cli72, CH70, Hil74, Hoo74, Hyn74, LI76, MMC79, Mey79, MH72, Nai76, RHR71, WC73, WM72a, WM78]. streambed [SW72]. Streams [HLH71, GMCG78, O’C74, Pec79, WL79]. strength [Wal73]. stress [Hsu75]. stresses [How78]. Strickland [Edm70b]. striders [LC74]. Stripping [Bra73a, FP79b]. Stromatolitic [KCS77]. strong [MM79, RMS77]. Strongylocentrotus [Fie72]. Strontium [ML70, PT70, AGM70, BNV^{+}74, Car70]. Strouds [Par75, Woo79]. Strouds-berg [Par75]. Strouds-burg [Woo79]. Stroudsburg [Eme74]. Structure [Ano72c, DH72, BAHA77, Den77, HCW70, LR72, Lyn77, Lyn79, NND72, Pat71, SSL78, TG72, Wal74b]. Stuart [JPW79]. Stud [Hob79].

Students [Koh71c]. Studies [Fen70, Fou70, Fou71, Fou73, HDFG72, Hor76, Koh74, LeC77, Man76, ML70, Wri75, DVCP79, Gol74b, Hes76, Hor79, How72, Hyn74, JDL78, KLAS73, Lan70, Leh76a, Mil73c, Mon78, Pug70, Pug73, RWI78, SA70, SS70, Sch73, WM72b]. Study [Por72, RC71, AFK73, AVZ75, BDW76, Bro74, CK79, EA70, ERVM73, Fag72, Fer70, FD72, GG81, GB72b, HS74, HHHP^{+}72, IWC^{+}77, JF72, KS70, KK78, LC74, yLKR75, Lyn79, PM77, PS73, Ril78a, SG71, Smi74b, SG76, The75, Wie71, Woo73, WWHH77]. studying [SB75, Swe76, Web74]. Stumm [O’C71]. Stuttgart [Fot70, Wet75]. subalpine [Rau78]. subaqueous [Mur76]. subarctic [BP70, Gor79a, Gor79b, Nat70]. subject [MA76]. submarine [KA74]. submerged [Jac70]. submersed [Hou74, OW74]. submicromolar [GHM78]. subregions [FD72]. subsampling [Ven71]. subsistence [Shu78]. substance [Høj78]. Substances [NK72, PRJR73, Whe76]. Substrate [WM78, Bru71, CK79, GG81, McL74, O’C74, Pec79, TGK^{+}78]. Substrates [Ert71, PP77]. substratum [Hyn74, Ree74]. Subsurface [RHR71, BP70, HP76, TS78]. subterminalis [OW74]. subtropical [CCK71, GWN70, MYY71a]. Succession [HD70]. successional [WF72]. successions [SCL^{+}71]. Sudamerica [Zar78]. Suez [HES75, MM73b, Mor75]. Sugar [HH73, Pre78]. Sugar-coated [HH73]. Sugar-frosted [Pre78]. Sulfate [HT79, JP78, Mon75, Hoi78, MSH74, Mon78, SRR79, Sta74a]. Sulfide [Hit78, RB74, Cas73, DH78b, HIJ78, IJ79]. sulfoxide [SL76]. sulfur [DB76a, Jør77, JC77, JKC79, NC76, TeI70, Woo77]. Sullivan [Gro75b]. summation [CSHE78]. Summer [ICO^{+}74, OS78]. Sunfish [GF70b]. sunlight [SR77]. superba [AVZ75]. Superior [LH78, Nri78, Smi73a]. supernatant [HC78]. Supervisor [Woo73]. suppl [Hob74]. supplement [THDW79]. supplied [ECHH^{+}71]. supply [Ric79]. surber [Chu72, Kro72b]. surf [CL76, Sha74a]. surf-zone [CL76]. Surface [BC72, DB71, GF70a, Lor70, Pra70, AP76, All76, BGZS75, BFFB75, BC74, CK74, CACB75, DLMS76, DG76a, DVCP79, DE71, EA79, Gal75, Gar74, GS72b, GH79, HP76, Har72, HP74, Jan73, JPW79, Kel70, KWD75, Kro75, MHU72, NL72, Ore79, PK74, RHR71, SSP73, Sma74a, SSA72, Tab79, TG72, TP78, Whe75, Whi75, WS70]. surfaces [Dis73, HC78, Ris73]. surveillance [Bri79]. surveys [NW70]. Survival [Wei73b, Arn71, BL72b, Val71a]. susceptibility [GA75, TBOO75]. Suspended [Bas75, MHU72, FP79a, MCB79, SS70, Man75a]. suspension [GTF78]. suspensions [WP78]. Sustained [Enr77a, TPG77, Dod70]. Swarming [Rat74]. swarms [HC79b]. Sweden [Pec70]. Swift [Gro77]. swimming [HC79c, HW76, Str77, Vly70]. Swiss [BSB77]. Sydney [Edm74a, Man72a, Ril74, Wet76]. symbiosis [Koh71a]. symbiotic [LLS75]. Symp [Hob74, Hor75, Hut72, Hyn76, Hyn78, Koz70]. Symposium [Ano78h, Bar71, Lym73, Ano71h, Ich71, FW74]. synoptic [HSC79, HDFG72]. Synthesis [Aar71, CM72]. syntrophic [SB75]. syringe [CRA72]. System [WCR71, BFFB75, Boo75b, Boo76, HB72a, IR73, JFH73, KWK74, KWD75, KP75, Lan79, Mil73c, PHR71, SS70, SAH72, Swe76, Whi74b, dO74].

Systems [GKA73, Coo73a, Eld75, LL77, NG78, Nor76, Pae74, TP74, WF72, Woo73, Rig76, Fry71].

T [Col72, Day79, Dee74, Edm74a, Gro74a, How72, Koh71a, Lau74, LR74b, Leh75, Liv72b, Ril74, Sch76, Ton72, Vly77]. Tables [Ano75m]. Tahoe [GMTA79, PG72, PRLG75, RMG70, CB72, FG75, HHGRW76, IWC^{+}77, KHHG^{+}72, LG79, STG73, TPG77, TP74]. Takahashi [Ril74]. Takenouti [Epp73]. tamarensis [AM78]. tank [ME74]. Tasch [Pai75b]. taxonomy [Edm76b]. technique [Bro74, CTB72, CC75b, CA72, DH75, GH72, HH73, HHPP70, Pre78, Sha73b, Tit75, War78, WW74]. Techniques [HC79a, CB72, Ham75, Har77, Pug73, SA70, Hob73]. Tegula [Pai71]. Tellinacean [Lev71a]. temperate [Dod79, Har73b, Jef70, Kal72, LI76]. Temperature [BL76, Bur70, CM72, GC72, Gie70, Gol77b, Kib71, Lam72, LR72, NND71, Wei73b, Bot75, Bra76, CWM75, CA79, CM77, Coo71, DE71, GS73a, GS73b, GC74a, HD70, HC79c, Huy77, KZ77, Lan75a, Lan75b, MA76, MB71b, NND72, NW70, Pat71, PH71, Sal79, SR79, SC70, Smi77, Ste73b, Swe76, SSA72, TG77, WS70, WP76]. temperature/salinity [WP76]. temperatures [Ada76b, Far75, LM70, PK74]. Temporal [AFK73, DV78]. Tension [LaR70, DVCP79]. tentaculate [RWI78]. term [And77, CS74, Den77]. terms [Ban74b]. terrestrial [Hyn78, Swe70a]. Test [ABGCP75, LR74a, EH77]. tested [TB75]. Testing [Doy79]. tests [MB79, TKC73]. testudinum [CPOT79, Fen70, KA77, OT77]. Texas [Cul73a, ESH73]. Texcoco [Bra71]. Textbook [Kil76]. Thalassia [CPOT79, Fen70, KA77, OT77]. Thalassiosira [Sma74a]. Thaliacea [HC79c, HM79]. theibautii [RMC79]. Their [BSB71, Bla73a, CK79, CS73b, Dal74, Edm76a, GG81, Ker78, OS78, RH77, SO78, SSL78, Sta75]. Theoretical [JM78, Boo75a, GB72a, Lew73]. theory [Ban74a]. Thermal [CBN78, BAHA77, BL76, Bot75, GS73c, Lan70, Lew73, LS72, Nai76, Swe76, Swe78, WF72]. thermal-gradient [BL76]. thermistor [LV76]. thermocline [KD79a, Swe70b]. Thermocyclops [Bur70]. Thermodynamic [Whi74a]. thermodynamics [DH72, Pla78]. Thermometric [MSH74]. thermophilic [JC75, SH73]. thiamine [CB72, Nat70, OT72]. thiebautii [CM75]. thiobacillus [TJ72]. thiobacillus-like [TJ72]. Thiosulfate [NCK79, TJ76]. Thorium [SP73, EGTA70, KTTF76]. those [KB76]. three [GF70b, HM79, Hut70]. threshold [Fro75]. Thuillier [Pro71]. thymidine [TA78]. Tidal [Boo75b, VTV^{+}78, DO75, ES75b, Gar75, GB72b, Jef72]. Tide [HJG71, Bla77, FT73, MA76, MSR^{+}75, Tem73, TS78]. Tides [FN70, War79]. Time [Bur70, Ada76b, Cam75, MP78, SR77, SP78, Ste73b, TL75, TG77]. time-delayed [SR77]. time-dependent [SP78, TG77]. Times [Edm72b, Ill75, SS78a]. timing [Enr77b, EH77]. tissue [JW70, Loh74]. tissues [LLS75]. title [Ano72e]. titles [Ano73c, Ano75e, Ano76c, Ano77e, Ano78e, Ano79c, Val74a]. titration [MSH74]. Tjeukemeer [dH77]. Todd [Edm71a]. Tokyo [Epp73]. tolerance [Bra76]. tonsa [Kim79]. tool [Jac70]. Torlanus [AF74]. Toronto [Cum71, Ebe72, Edm72a, Edm74a, Hyn78, Liv76, Ril74, Val71a, Wet76, BCB71]. Tortanus [Mul79]. total [BFMF72, CC75b, DSC77, HS75, MS74b, OS78, SB73, SHW^{+}70, Whi74b]. toxic [Mac77, RMC79]. toxicity [Leh76a, LWR73, SL78]. Trace [BG72, JM78, LT71, TKC73, And79, BK72, DGEB70, DS75, DCP73, Eld75, Mar70b]. Tracer [ML70, KLAS73]. tracers [IE78, LL75]. Track [KK76b, KK79]. Tracking [Fol71]. Trans [And71, Ori73]. transect [Mil73c]. transfer [BM74, Gor78, VWB77a, Woo77]. Transformations [GKA73, JKC79]. Transient [LJ73, BH78]. transition [BP70, WWK^{+}77]. transitional [MG78, Mor75]. transl [Ste79b]. translation [Edm76b]. translocation [TBD77]. transmission [KA74]. transmissometry [BB76]. Transparency [Ril75, Lam72, OS78]. Transparency-chlorophyll [Ril75]. Transport [FP77, NS78, PG74, PG75, WD78, AS77, BL79, CP79, DH72, GN79, Gro77, KP75, LS77, LJ73, LG79, Mar70b, PHR71, PKC77, San77, Sha74a, TS78, Ver75]. transports [CO75]. Transuranics [NWEG75]. trap [Hak76, HB79, KAG77, Kir75a, SRSW78, WL79, WW73]. trapping [TKC73]. Travis [ESH73]. trawl [HBM73]. treated [DM71]. treatise [Tyl71, Wet76]. treatment [FR77b, TTW75]. Trench [TJ79]. Trends [MM73a]. Trichodesmium [CM75, RMC79, CP77, Rom78, YYP72]. Tricorythodes [MMC79]. Trinius [Wet75]. triphosphate [BPW76, Bul78, KA77]. Tripp [LaF77]. trispinosa [Lan78]. Tritiated [TA78]. Tritium [MW77, IWC^{+}77, TTC^{+}77]. tritium-helium-3 [TTC^{+}77]. trivittatus [FS72a]. Trophic [Fot74, Car77, CS73b, Dev79]. Tropical [Bur70, YYP72, AGM70, Bla73a, CR72, CR76, CS74, Cov76, Dun70, EGTA70, HP70, Hsu75, JSS^{+}71, KD79a, Koh71b, LH73a, Lew73, Liv77a, Mal71, McL74, McL75, RG76b, Tho70a, Tho70b, TD72, ZH71, Smi76a]. tropics [Bri76]. Trough [Smi74b]. Trout [All78]. tube [BS74, GO78]. tubes [JS73]. tubificid [BCK72]. Tubificids [Dav74b, Dav74a]. Tulsa [Gro71, Gro74a]. tunicata [Fou73, HC79c, HM79]. tunicate [HG76]. tunicates [MLKT75]. turbid [Hol70]. turbidity [BFFB75, Lud74]. turbulence [Gus79, RMS77]. Turbulent [Mur72, DL74]. turgor [Wal73]. Turnover [HS74, Cam75, Nic74, SCA77, TLP78]. Turtle [Fen70]. Two [EB79, ML70, RH75, WL79, All79, Bat76, BSB77, CB75b, Dev79, DC75, EP71, Fed75b, Hai79a, Hai79b, HG76, KS70, Mal70, Mal73, Mos70, OV78, OSD76, RHR71, Sma74b, Sta76, Wal75]. types [KS70, TJ72]. Typewriter [Bri79, Gor79a, Gor79b, Gus79, Har78a, Man75a].

U [Hob74]. U.K [Cas71]. U.S.A [Hyn78]. ubiquitous [JS79]. Uganda [Bur70]. Ultrasonic [TTW75]. Ultraviolet [Lor79, MMW71, Høj78, SC76a]. Undercurrent [RH76]. Underestimation [Kro72b]. undersea [Smi78a]. Underwater [Ham75, HMA^{+}75, BD72, Hal71]. undisturbed [KHC74]. UNESCO [Gro76a, Hob74, Koh79, Mil78, Pom74]. UNESCO. [Pom74]. unicellular [Shu78, SL78]. unique [KLAS73]. United [Edm70a, And77]. Univ [Cul73a, Cum71, Day79, Dug71, Ebe72, Edm72a, Gol74a, Gro75a, Har78a, How72, Koh71a, LaF77, Man70, Ril71a, Woo73, Zar78]. Universities [Koh72]. University [FW74, Hyn73, LeC77, Lun79, Ril73, Sta75, Wal74b]. Unpaged [Sta75]. Unreliability [GV73]. upland [MK79]. upper [Ban74b, Dee71, QPB70, HPM72a]. upper-level [QPB70]. upstream [Cow70, MBB73]. Uptake [LL75, SOP73, Sch74, Toe71, WNLS77, AC76, Bat76, BS77, BH78, BHK78, CM75, D’E77, Doy75, Dro78, Fal75, FC70, GF75, GWN70, GM78, HS74, JDL78, JP78, KD79a, LJ79, Mac78b, MGHR75, McC72, MW77, Mon75, Mon78, MD78, Pae74, Par77, PG74, PG75, Pet78b, Pil74, Ram76, RK79, Rhe78, SMD76, SH73, SS77, SN71, TLT77, WNS74, YK73]. upward [Enr77a]. Upwelling [Doy75, Dux79, SA76, Bla77, EB79, KWD75, KMP^{+}75, NG78, SMD76, WKW^{+}74, WWK^{+}77]. Uranium [KTTF76, EGTA70]. urban [Ril78b]. Urchin [Fie72]. Urea [Ber74, CRW72, DCC73, ECHH^{+}71, McC72, Mcc70, Rem71]. urease [Mcc70]. US$10.25 [Man79]. US$40.50 [Gol79]. US$70 [Rzo79]. USA [Par77]. Use [Bro78b, CAE78, CP73, HPM72b, SA74, BHD76, Cas73, CE79, Coo73a, Fee79, GTF78, IBM76, MS70, ME74, Pat80, Pug70, Pug73, RC74, SCA77, SC76a, WL79, Whi74a, WW74]. used [FC70, HB72a]. Uses [How78]. Using [Bra73a, ABGCP75, BKS70, CB72, DSC77, GHM78, GS74, Hal71, HS77b, JS73, Lan70, SL76, Sma74a, Sta74a, Str75a, The75]. Usinger [Chu72]. Utah [SG76, Ano74b]. Utilization [CRW72, Nor75, Dro78, DB76b, Fou73, Koh71b, Per76, Pyt71, RWI78, Ric79, RHG73, THDW79, TJ76, Wal76b]. UV [DG78, Høj78, KVW79, MSJG74, SC76a]. UV-B [Høj78, KVW79, SC76a]. UV-visible [MSJG74].

V [Bro70, Cul73a, Gol77a, Gor79a, Gor79b, Lun79, MLKT75, Pla78, Ril78a, Ste79b, Tea78, Woo77, Ban73, Bou75, But76, Edm76b, Enr76, Fer70, Fot70, Gol79, Gor78, Gro76b, Hed70, Hor75, How78, LeC77, Man76, Mas70, Mil77, Par73, Pyt76b, Smi76c, Smi76a, Tyl71, Wet76]. vacuolation [Wal73]. vacuoles [Wal73]. Vagners [Woo73]. Valdez [Gol74b, Gol74a]. validity [Wil73]. Valley [LG79]. values [Ker72]. vanadium [PT70, MLKT75]. Vancouver [DO75, RH76]. Var. [Sau72a]. variability [Chu72, Den77, PF73, Tab79, TLP78, Wan74]. Variation [Bay78, MF71, BP74b, Glo78, HG72, HPS75, Huy77, Sta78, Swe78, Tim75]. variations [BP74a, KM73, MP77a, MP77b, Mos76, Smi77, Ste73b, SSA72, WK77, WS70, CL76]. varieties [OV78]. various [QM71, Ten77]. variously [Cul73a]. varying [KA74, LM70, MP78, SH73]. vascular [Hou74]. vectors [KMP^{+}75, PBHC70]. Vegetation [WAA72b, VTP76]. vehicle [Sho70, SCE^{+}76]. velocity [Bru71, Enr77c, HPS75, TG72]. Verde [HBLP76]. Verduin [Meg75]. Verification [Bra76]. Verlag [Boo75a, Cha74, Enr76, Hob79, Hun77, Kel75, Koh74, Liv78, Man76, Ril77, Smi74a, Smi78b, Ste79a, Whe73]. vernalis [LL79]. verrucosa [Pil74]. versatile [GDHZ79, Sal79]. versions [Edm70a]. Vertical [Bla73b, Cun72, Daw78, DB76a, JP75, Kos79, LaR70, NND72, PHR71, Swe70b, VB75, ZS76, BFM76, BFMF72, CH70, Den77, Enr77b, EH77, Fee76, Gag76, HC79c, Hu78, IE78, KZ77, KD79a, MM75, Mil70, MHK75, PRLG75, Pea79, RS79, SRSW78, Sta71, SW72, ST78, TG77]. vertically [Bel78]. vi [Car75, Cul73a, Cus76a, Eme78, Ham79, Koh72, Man76, Sta74b, Wu79]. via [LG79, Mar70b]. Viability [CM72, TPG77]. vibrating [FR77a]. Vibrio [CM72, PA70]. Victoria [Ano79d]. vie [Par78]. view [Pet71]. vii [Aro73, Hor76, Hyn76, Kin70, Lun79, Sch76, Sta72a]. viii [Cha74, Fot70, Gro71, Hun77, Koh77, Liv77a, Man75a, Man70, Ril78b]. Virginia [TGK^{+}78]. visible [MSJG74]. visualization [Hal71]. Vitamin [OT72, Par77, BS78c, CB72, Nat70]. vitro [Aar78]. vivo [SH77, Ber72]. Vol [Par72, Sch73, Hob74]. Volatile [HH78]. Volga [Rzo79]. Vollenweider [Edm70a, Par70]. Voltammetric [ZH71]. Voltammetry [Bra73a]. Volume [Ano71g, Ano77e, Ano78e, Ano79c, Val74a, Ano70e, Ano72e, Ano73c, Bra73b, Epp73, Gol74a, Pla78, Rig76, San77, STND74, SH72]. volumes [PC76, RG76a]. volumetric [WP76]. vragilis [Pro71]. vs [CC75a, Tyl73]. vulgaris [Ton71, FR77b].

W [Bri72, Bro78a, Edm76a, FW74, Gol78, Gol74b, Gro71, Gro75a, Gro75b, Ham79, Har78a, Hed70, Hob74, Hyn71, Liv72b, Liv76, Man79, Mas70, Mil73a, Mil73b, O’C71, Phl73, Ril71a, Rzo79, Sch75b, Sta72a, Sta74b, Ste79a, Tho78, Ton72, Wu79]. W. [LP78]. Waccasassa [BBPK71]. Walker [Col72]. Walsby [Sta74b]. Wangersky [Pyt72]. Ward [LG79]. Warm [Mar78b]. Warren [Hyn71]. Wash [How72]. Washington [Bar71, Day79, Edm71a, Fer70, Hyn76, LaF77, Par77, Ril73, Woo73, CL76, Cre75, DP78, EA70, Egg78, GE75, Hyn76, Man73b, PB73, RH76, Ric79, SEA71, WC76, Wak77, Wal74a, Wal75]. waste [KA71]. Water [BY73, Bri79, FP79b, Fry71, GKA73, Gri78, KSGM71, LG79, Pen73, SH71, Smi71, TS78, WAA72a, Ada76b, AGM70, BWP73, Bla79, BS74, BP70, BP74b, Bro79, CM79, CA79, CNC73, CJN72, Cro73, CP70, Cul73a, Cul77, DLMS76, DV78, Dav77, DW72, DCP73, DE71, Edm71a, Eld75, FH79, FD72, GP72a, Gar77, GS72a, GLSD74, GB72b, Gro79a, HP76, Ham75, HC78, Hes76, HHHP^{+}72, Hob79, HHGRW76, Hun77, Hyn71, Ids72, IJ79, ICO^{+}74, JM70, JB71, KWK74, Kar78, KM73, Koh71b, Kro75, KB76, Lan79, LR74a, LC74, LB71, Ley73, LH73b, LI76, MS70, Mak72, Mar78b, MB77, MBR78, MPN72, May76, Moe71, MYY71a, Mor71, PG72, PDF^{+}78, PK74, Pin73, PH71, PBK70, Ree74, RG76a, RHC74, RH75, STND74, SEPS74, Sha73b, Sho70, Sma74a]. water [ST78, SP73, Tab79, TWPC75, Tho70b, VB75, Ver75, Wei73a, WM72a, Whe76, Whi75, WWHH77, dO74, DH72, Edm70a, Man75a]. water-soluble [Moe71]. water-striders [LC74]. Waters [MF71, And79, BFFB75, Ber74, BST70, BS71, BKM78, BH70, BLL70, CB72, CO75, CK74, DH78b, DL74, DCP73, Dug71, Edm70a, Edm76a, ERHC79, ERH79, FP79a, Fit70, FL75, FMJC72, Fra72, Gar78, GHM78, GV73, Glo78, GB72c, HC79a, Har78b, HD77, HH78, HS77b, Hoi78, Høj78, Hol70, Hsu75, Jan73, Kel70, KH73, Kim79, LNS^{+}76, LS77, Lee77b, Liv77a, MHU72, MMW71, Mar74, MCA77, MLCG79, Nor75, Nri78, Ore79, Per76, RE73, Rei72b, Rem71, Sch74, Sch75a, SSP73, SB73, SC76a, SB78b, SB78c, SM74, SOHH76, TP78, WD78, Whe77, Win71, WS70, Bri72, Cum71, Hor75, How72, Man72a, Man75b, Ril71a, Ton72, Tyl71]. waters. [Hob73, Sha73a]. watersheds [Bri76, HL73b, MK79]. watts [CC75a]. waves [HP76, iK75, Liu70, Hun77]. Wax [BM74, LH73a]. weak [MM79]. weather [ESH73]. webs [RG76b]. Weight [BR72, MB70b, DD78, Mur76, Whe76]. Weinheim [Ril77]. well [Kje78]. Werner [Lun79]. West [Ich71, Val74b, WS70, RHG73]. Western [Gro71, And77, CC74, CM75, CP77, GN79, Hul76, Mil71, Mon75, Sho74, WL78, Wil72b, SA71, Wei73a]. Westmorland [Cas71]. Wet [Tea78]. wetland [MK79]. wetlands [Pai70, Pom79]. Wetzel [Liv76]. Weyl [Ril78b, Ste70]. whale [BSS78]. whales [Sch77a]. Whigham [Pom79]. White [MG78]. Whitfield [Man72a]. Whittaker [Man76]. Whitton [LeC77]. who [Val74b, Val74b]. whole [Fee79, Pat80, RH78]. whole-lake [Fee79, Pat80]. wide [MB71b]. Wildlife [Fer70, Pai70]. Wiley [Ban73, DH72, Edm74a, Fry71, Gol77a, Gro77, Koh77, Lau74, Lee77b, Liv72a, Lym73, Mof73, O’C71, Pai75b, Ril78a, Sha73a, Ste70, Str75b, Tyl71, Wet76, Gor78, How78]. Wiley-Inter-science [Edm74a, Lau74]. Wiley-Interscience [Ban73, DH72, Fry71, Lym73, Mof73, O’C71, Sha73a, Str75b, Tyl71]. Williams [Edm76a]. Winberg [Ton72, Rig72]. Wind [Hsu75, LH78, WWB^{+}78, BB76, ICO^{+}74, Liu70, War79]. Wind-driven [LH78]. winds [QPB70]. windward [SM73]. Wingra [ATM74, BB73]. Winnebago [BR74]. Winogradski [Cas72a]. winter [BB73, CT70, HS77b, HS77c, OS78, Smi73a, Smi77, SK73b]. winter-spring [HS77b, HS77c]. Wintergreen [Cro73]. Wisconsin [BR74, BL74, RB70, TL76]. Wiss [Koz70]. within [GTF78, Hyn74, Mos76, Pec79, Pro71, WWB^{+}78]. Wm [Sta72a]. Wononskopomuc [War72]. Wood [Smi76a]. Woods [Gro74b]. Wootton [Hyn73]. works [Cle72]. World [Col72, Koh73, HNT75, Hyn76, LNS^{+}76, Lau74, Mil73a, Mil73b, Pet71, Sch78a, Edm72a]. Worldwide [Fre71].

x [Boo75a, Csa75, Edm71a, Ich71, Kap78, Koh71a, LeC77, Mas70, Pai75a, Ril74, Sha73a, Smi75, Ste79b, Str75b, Val71a, Wal74b, Wal76a, Bas75, EA70]. X-radiographs [EA70]. X-ray [Bas75]. xi [Ebe72, Edm72a, Gol78, Hyn78, Mil73a, Mil73b, Par72, Par73, Par75, Pla78, Ste79a]. xii [Bri79, Edw77, Enr76, Gro78, Hob79, Hyn73, Lee77b, Liv76, Liv78, Mac79, Man73a, Par78, Rig72, Tea78, Wet76]. xiii [Cle72, Epp73, Gol77a, How72, Koh74, Liv77b, Phl73, Ril73, Rzo79, Smi78b]. xiv [Bow79, Col72, FW74, Fra72, Fry71, Gro74b, Gro75b, Gus79, Liv72a, Mil77, Ril71a, Ril78a, Stu71, Zar78]. XIX [Ano73h]. xv [Bri72, Gro77, Koh79, LaF75, Mof73, O’C71, Rig76, Ril77, Smi74a, Tyl71]. xvi [Edm74a, Eme74, Gol79, Hob73, Hyn71, Liv72b, Whe73]. xvii [Gor79a, Gor79b, Har78a, Pom79, Ste70, Woo73, Woo79]. xviii [But76, Gor78, Gro76b, How78]. xx [Gor79a, Gor79b, Hed70, Pyt76b, Sio77, Smi76c]. xxi [Gro75a]. xxiii [Lau74]. xxiv [Cum71, Ton72].

Ycaza [Dug71]. year [GS72a, Liv71, MG75]. years [HF70, WSR79]. yellow [Høj78, Moe71, OW73]. Yellowstone [MB71b, WF72]. yield [Ban74b]. York [Aro73, Ban73, Boo75a, Bou75, Bow79, Bri72, But76, Car75, Cha74, Col72, Csa75, Ebe72, Edm71a, Edm72b, Edm74a, Edw77, Eme76, Eme78, Enr76, Fry71, Gol77a, Gor79a, Gor79b, Gor78, Gro75b, Gro76b, Gro77, Gro78, Gus79, Hed70, Hob79, Hor76, How78, Hun75, Hun77, Hyn78, Ill75, Kap78, Kel75, Kin70, Koh73, Koh74, Koh77, Kra79, LaF75, Lau74, Lee77b, Liv72b, Liv77b, Liv78, Lun79, Lym73, Mac79, MC79, Man75a, Man75b, Man76, Mof73, O’C71, Pai70, Pai75a, Pai75b, Par72, Par73, Par78, Phl73, Pla78, Pom79, Pyt76b, Rig76, Ril74, Ril77, Ril78a, Ril78b, Sch77a, Sch75b, Sha73a, Smi74a, Smi75, Smi76b, Smi76c, Smi76a, Smi78b, Smi78a, Sta74b, Ste79a, Ste70, Ste79b, Str75b, Stu71, Tea78, Tyl71, Val74b, Wal76a, WWB^{+}78, Wet76, Whe73, Ahr75, BS78c]. York [HF70, Lud74, Mos76, OS78, SO78, Wea79]. York-Heidelberg-Berlin [Koh74]. Yu [Ham79, Wu79].

Zaddach [Ste73a]. Zaire [Man79]. Zambia [Liv71]. Zeitzschel [Koh74]. Zinc [ZY72, CR70, KM73, MH74, PKC77, PT70, SKF74, ZH70, ZH71]. zinc-65 [CR70, PKC77]. Zn [Bra73a, EA79, Wil78]. zone [BP70, BSB77, CL76, FG75, Kal72, Sha74a, Pai70]. Zoophysiology [Enr76]. zooplancton [Cus76a]. zooplankter [AVZ75]. zooplankters [BR72, BFMF72]. Zooplankton [DB71, Dev79, DEWN76, Han71, Lyn77, MB79, Por72, War72, Beg76, Big77, BB75, CB75a, Dee71, Dod75, Dod79, Dun70, DG76b, Edm72c, Fag73, FF75, GO78, GMV72, Gre79, HMA^{+}75, Har77, HW76, IR73, JCK70, KP75, Lam74, Lam78a, LH73a, Leh77, Lyn79, MM75, Mar70a, MP78, Mil70, MB76b, PL73, Pet75, Wel70, Wie70, Wie71, Wor77, ZS76, Mil78]. zooxanthellae [LLS75]. Zostera [HM75, MBN72, PS77].

References

Aaronson:1971:SEM

[Aar71] S. Aaronson. The synthesis of extracellular macromolecules and membranes by a population of the phytoflagellate Ochromonas danica. Limnology and Oceanography, 16(1):1–9, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Aaronson:1978:EOM

[Aar78] S. Aaronson. Excretion of organic matter by phytoplankton in vitro. Limnology and Oceanography, 23(4):838, July 1978. CO-DEN LIOCAH. ISSN 0024-3590. See reply [Sha78].

Alvarez-Borrego:1975:ONR

[ABCP75] Saul Alvarez-Borrego, Charles H. Culberson, and P. Kilho Park. Oxygen-nutrient relationships in the Pacific Ocean. Limnology and Oceanography, 20(4):806–814, September 1975. CODEN LI-OCAH. ISSN 0024-3590.

Alvarez-Borrego:1975:TRM

[ABGCP75] Saul Alvarez-Borrego, Donald Guthrie, Charles H. Culberson, and P. Kilho Park. Test of Redfield’s model for oxygen-nutrient relationships using regression analysis. Limnology and Oceanography, 20(4):795–805, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Azam:1976:SAU

[AC76] F. Azam and S. W. Chisholm. Silicic acid uptake and incorporation by natural marine phytoplankton populations. Limnology and Oceanography, 21(3):427–435, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Adams:1976:FDD

[Ada76a] W. P. Adams. Field determination of the densities of lake ice sheets. Limnology and Oceanography, 21(4):602–608, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Adams:1976:EWT

[Ada76b] Charles E. Adams, Jr. Estimating water temperatures and time of ice formation on the Saint Lawrence River. Limnology and Oceanography, 21(1):128–137, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Ambler:1974:FBP

[AF74] Julie W. Ambler and Bruce W. Frost. The feeding behavior of a predatory planktonic copepod, Torlanus discaudatus. Limnology and Oceanography, 19(3):446–451, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Atwood:1973:TDS

[AFK73] Donald K. Atwood, Phillip N. Froelich, and W. Frank Kinard. Temporal and depth study of alkaline earth chlorinity ratios in seawater at a single station south of Puerto Rico. Limnology and Oceanography, 18(5):771–783, September 1973. CODEN LI-OCAH. ISSN 0024-3590.

Adams:1978:RDI

[AGA78] Michael S. Adams, Piero Guilizzoni, and Susan Adams. Relationship of dissolved inorganic carbon to macrophyte photosynthesis in some Italian lakes. Limnology and Oceanography, 23(5):912– 919, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Armstrong:1971:RME

[AGF71] Richard Armstrong, Charles R. Goldman, and Dennis K. Fujita. A rapid method for the estimation of the carbon content of seston and periphyton. Limnology and Oceanography, 16(1):137–139, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Andersen:1970:RSC

[AGM70] N. R. Andersen, J. D. Gassaway, and W. E. Maloney. The relationship of the strontium: chlorinity ratio to water masses in the tropical Atlantic Ocean and Caribbean Sea. Limnology and Oceanography, 15(3):467–472, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Azam:1977:SDA

[AH77] F. Azam and R. E. Hodson. Size distribution and activity of marine microheterotrophs. Limnology and Oceanography, 22(3): 492–501, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Ahrnsbrak:1975:SIS

[Ahr75] William F. Ahrnsbrak. A saline intrusion into Seneca Lake, New York. Limnology and Oceanography, 19(4):275–278, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Adams:1978:RIS

[AL78] W. P. Adams and D. C. Lasenby. The role of ice and snow in lake heat budgets. Limnology and Oceanography, 23(5):1025–1028, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Alldredge:1976:DAH

[All76] Alice L. Alldredge. Discarded appendicularian houses as sources of food, surface habitats, and particulate organic matter in planktonic environments. Limnology and Oceanography, 21(1):14–24, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Allan:1978:TPS

[All78] J. David Allan. Trout predation and the size composition of stream drift. Limnology and Oceanography, 23(6):1231–1237, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Alldredge:1979:CCM

[All79] Alice L. Alldredge. The chemical composition of macroscopic aggregates in two neretic seas. Limnology and Oceanography, 24(5): 855–866, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1978:CSG

[AM78] D. M. Anderson and F. M. M. Morel. Copper sensitivity of Gonyaulax tamarensis. Limnology and Oceanography, 23(2):283– 295, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1971:BRH

[And71] Roger Y. Anderson. Book review: Hutchinson, G. E., and others. 1970. Ianula: an account of the history and development of the Lago di Monterosi, Latium, Italy. Trans. Amer. Phil. Soc. 60 (Part 4) 178 p. $7.00. Limnology and Oceanography, 16(3):594–595, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1975:POS

[And75] O. Roger Anderson. A possible origin of some olive-green cells in aphotic marine environments. Limnology and Oceanography, 20 (3):1000–1004, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1976:RPD

[And76] O. Roger Anderson. Respiration and photosynthesis during resting cell formation in Amphora coffeaeformis (Ag.) Kutz. Limnology and Oceanography, 21(3):452–456, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1977:STS

[And77] Roger Y. Anderson. Short term sedimentation response in lakes in western United States as measured by automated sampling. Limnology and Oceanography, 22(3):423–433, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Andreae:1979:ASS

[And79] Meinrat O. Andreae. Arsenic speciation in seawater and interstitial waters: The influence of biological-chemical interactions on the chemistry of a trace element. Limnology and Oceanography, 24(3):440–452, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1970:Aa

[Ano70a] Anonymous. Announcements. Limnology and Oceanography, 15 (2):332, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1970:Ab

[Ano70b] Anonymous. Announcements. Limnology and Oceanography, 15 (6):977, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1970:BRa

[Ano70c] Anonymous. Books received. Limnology and Oceanography, 15 (3):489, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1970:BRb

[Ano70d] Anonymous. Books received. Limnology and Oceanography, 15(6): 975–976, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1970:ILO

[Ano70e] Anonymous. Index to Limnology and Oceanography volume 15, 1970. Limnology and Oceanography, 15(6):978–983, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1970:IGC

[Ano70f] Anonymous. International geological congress. Limnology and Oceanography, 15(1):170, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1970:PAS

[Ano70g] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 15(1):171– 172, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1971:ASLb

[Ano71a] Anonymous. American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 18(5):840, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1971:ASLa

[Ano71b] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 16(1):155, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1971:BGE

[Ano71c] Anonymous. Bibliography of G. Evelyn Hutchinson. Limnology and Oceanography, 16(2):164–168, March 1971. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1971:BRa

[Ano71d] Anonymous. Books received. Limnology and Oceanography, 18(5): 843–844, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1971:BRb

[Ano71e] Anonymous. Books received. Limnology and Oceanography, 16(6): 1013–1014, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1971:DDC

[Ano71f] Anonymous. Doctoral dissertations completed under the direction of G. Evelyn Hutchinson. Limnology and Oceanography, 16(2): 169–172, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1971:ILO

[Ano71g] Anonymous. Index to Limnology and Oceanography volume 16, 1971. Limnology and Oceanography, 16(6):1008–1013, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1971:ISC

[Ano71h] Anonymous. International symposium on coral reefs. Limnology and Oceanography, 18(5):844, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1971:PAS

[Ano71i] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 16(3):590– 592, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1972:ASL

[Ano72a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 17(1):163– 164, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1972:Ab

[Ano72b] Anonymous. Announcement. Limnology and Oceanography, 17 (2):324, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1972:AFI

[Ano72c] Anonymous. Announcement: First International Congress of Ecology Structure, Function and Management of Ecosystems 8– 14 September 1974. Limnology and Oceanography, 17(4):650, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1972:Aa

[Ano72d] Anonymous. Announcements. Limnology and Oceanography, 17 (1):168, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1972:ATI

[Ano72e] Anonymous. Author and title index to Limnology and Oceanography volume 17, 1972. Limnology and Oceanography, 17(6):973– 979, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1972:BR

[Ano72f] Anonymous. Books received. Limnology and Oceanography, 17(6): 971–972, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1973:ASL

[Ano73a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 18(1):186– 187, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1973:A

[Ano73b] Anonymous. Announcement. Limnology and Oceanography, 18 (1):188, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1973:ATL

[Ano73c] Anonymous. Authors and titles in Limnology and Oceanography volume 18, 1973. Limnology and Oceanography, 18(6):1015–1020, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1973:BR

[Ano73d] Anonymous. Books received. Limnology and Oceanography, 18(6): 1013–1014, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1973:E

[Ano73e] Anonymous. Errata. Limnology and Oceanography, 18(5):809, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1973:PAS

[Ano73f] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 18(1):184– 185, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1973:SBL

[Ano73g] Anonymous. Sampling bias of a Longhurst–Hardy plankton recorder. Limnology and Oceanography, 18(3):500, May 1973. CO-DEN LIOCAH. ISSN 0024-3590.

Anonymous:1973:XCI

[Ano73h] Anonymous. XIX Congress of the International Association of Limnology (SIL). Limnology and Oceanography, 18(2):351, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1974:ASL

[Ano74a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 19(1):173– 174, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1974:BRP

[Ano74b] Anonymous. Book review: Proceedings of the American Society of Limnology and Oceanography, Inc. 36th Annual Meeting 11 June 1973 Salt Lake City, Utah. Limnology and Oceanography, 19(1): 171–172, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1974:E

[Ano74c] Anonymous. Erratum. Limnology and Oceanography, 19(6):1017, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1974:JOA

[Ano74d] Anonymous. Joint Oceanographic Assembly 1976 preliminary announcement. Limnology and Oceanography, 19(5):874, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:ASL

[Ano75a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 20(1):156, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:Aa

[Ano75b] Anonymous. Announcement. Limnology and Oceanography, 20 (6):451, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:Ab

[Ano75c] Anonymous. Announcement. Limnology and Oceanography, 20 (5):680, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:Ac

[Ano75d] Anonymous. Announcement. Limnology and Oceanography, 20 (3):1055, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:ATL

[Ano75e] Anonymous. Authors and titles in Limnology and Oceanography. Limnology and Oceanography, 20(3):1064–1070, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:BRa

[Ano75f] Anonymous. Books received. Limnology and Oceanography, 19 (4):304, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:BRb

[Ano75g] Anonymous. Books received. Limnology and Oceanography, 20 (3):1062, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:Ea

[Ano75h] Anonymous. Errata. Limnology and Oceanography, 20(2):210, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:Ed

[Ano75i] Anonymous. Errata. Limnology and Oceanography, 20(6):961, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:Eb

[Ano75j] Anonymous. Erratum. Limnology and Oceanography, 20(5):680, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:Ec

[Ano75k] Anonymous. Erratum. Limnology and Oceanography, 20(5):698, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:Ee

[Ano75l] Anonymous. Erratum. Limnology and Oceanography, 20(4):999, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:IOT

[Ano75m] Anonymous. International oceanographic tables for the oxygen solubility in seawater. Limnology and Oceanography, 19(4):298, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1975:PAS

[Ano75n] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 20(1):154– 155, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1976:ASL

[Ano76a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 21(1):190– 191, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1976:A

[Ano76b] Anonymous. Announcement. Limnology and Oceanography, 21 (2):348, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1976:ATL

[Ano76c] Anonymous. Authors and titles in Limnology and Oceanography. Limnology and Oceanography, 21(6):934–939, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1976:BR

[Ano76d] Anonymous. Books received. Limnology and Oceanography, 21 (4):636, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1976:PAS

[Ano76e] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 21(1):188– 191, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1977:ASLb

[Ano77a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 22(1):179, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1977:ASLa

[Ano77b] Anonymous. American Society of Limnology and Oceanography, Inc. publication announcement. Limnology and Oceanography, 22 (1):175, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1977:Aa

[Ano77c] Anonymous. Announcement. Limnology and Oceanography, 22 (2):369, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1977:Ab

[Ano77d] Anonymous. Announcement. Limnology and Oceanography, 22 (3):552, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1977:ATL

[Ano77e] Anonymous. Authors and titles in Limnology and Oceanography volume 22, 1977. Limnology and Oceanography, 22(6):1106–1111, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1977:BR

[Ano77f] Anonymous. Books received. Limnology and Oceanography, 22(5): 977–978, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1977:PAS

[Ano77g] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 22(1):176, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1978:ASLb

[Ano78a] Anonymous. American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 23(1):194, January 1978. CO-DEN LIOCAH. ISSN 0024-3590.

Anonymous:1978:ASLa

[Ano78b] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 23(1):193– 194, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1978:ASLc

[Ano78c] Anonymous. American Society of Limnology and Oceanography, Inc. publication announcement. Limnology and Oceanography, 23 (2):387, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1978:A

[Ano78d] Anonymous. Announcements. Limnology and Oceanography, 23 (5):1084, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1978:ATL

[Ano78e] Anonymous. Authors and titles in Limnology and Oceanography volume 23, 1978. Limnology and Oceanography, 23(6):1268–1274, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1978:BR

[Ano78f] Anonymous. Books received. Limnology and Oceanography, 23 (3):584, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1978:PAS

[Ano78g] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 23(1):190– 192, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1978:SBN

[Ano78h] Anonymous. Symposium on biological nitrogen fixation. Limnology and Oceanography, 23(1):194, January 1978. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1979:ASL

[Ano79a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 24(1):196– 197, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1979:A

[Ano79b] Anonymous. Announcements. Limnology and Oceanography, 24 (3):592, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1979:ATL

[Ano79c] Anonymous. Authors and titles in Limnology and Oceanography volume 24, 1979. Limnology and Oceanography, 24(6):1179–1184, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1979:BRP

[Ano79d] Anonymous. Book review: Proceedings of the American Society of Limnology and Oceanography, Inc. 41st Annual Meeting 20 June 1978 Victoria, British Columbia. Limnology and Oceanography, 24(1):194–195, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1979:BR

[Ano79e] Anonymous. Books received. Limnology and Oceanography, 24 (4):797–798, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Anthony:1977:IRR

[Ant77] R. S. Anthony. Iron-rich rhythmically laminated sediments in Lake of the Clouds, northeastern Minnesota. Limnology and Oceanography, 22(1):45–54, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Aaronson:1976:RSE

[AP76] S. Aaronson and N. J. Patni. The role of surface and extracellular phosphatases in the phosphorus requirement of Ochromonas. Limnology and Oceanography, 21(6):838–845, November 1976. CO-DEN LIOCAH. ISSN 0024-3590.

Atlas:1977:SBA

[AP77] E. Atlas and R. M. Pytkowicz. Solubility behavior of apatites in seawater. Limnology and Oceanography, 22(2):290–300, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Andersen:1973:PCO

[AR73] Neil R. Andersen and Francis A. Richards. Prospects in chemical oceanography. Limnology and Oceanography, 16(5):815–817, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Arnold:1971:IAS

[Arn71] Dean E. Arnold. Ingestion, assimilation, survival, and reproduction by Daphnia pulex fed seven species of blue-green algae. Limnology and Oceanography, 16(6):906–920, November 1971. CO-DEN LIOCAH. ISSN 0024-3590.

Arons:1973:BRC

[Aro73] Arnold Arons. Book review: Craig, R. E. 1973. Marine physics. Academic Press, London and New York. vii + 83 p. £1.90. Limnology and Oceanography, 18(4):695, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1977:DCS

[AS77] Lars W. J. Anderson and Beatrice M. Sweeney. Diel changes in sedimentation characteristics of Ditylum brightwelli: Changes in cellular lipid and effects of respiratory inhibitors and ion-transport modifiers. Limnology and Oceanography, 22(3):539–552, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Adams:1974:DDP

[ATM74] Michael S. Adams, John Titus, and Michael McCracken. Depth distribution of photosynthetic activity in a Myriophyllum spicatum community in Lake Wingra. Limnology and Oceanography, 19(3): 377–389, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Ayala:1975:ESA

[AVZ75] F. J. Ayala, James W. Valentine, and G. S. Zumwalt. An electrophoretic study of the Antarctic zooplankter Euphausia superba. Limnology and Oceanography, 20(6):635–640, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Armentano:1975:SRL

[AW75] T. V. Armentano and G. M. Woodwell. Sedimentation rates in a Long Island marsh determined by ^{210}Pb dating. Limnology and Oceanography, 20(6):452–456, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1970:RDO

[AZ70] G. C. Anderson and R. P. Zeutschel. Release of dissolved organic matter by marine phytoplankton in coastal and offshore areas of the northeast Pacific Ocean. Limnology and Oceanography, 15(3): 402–407, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Barica:1971:CSN

[BA71a] Jan Barica and F. A. J. Armstrong. Contribution by snow to the nutrient budget of some small northwest Ontario lakes. Limnology and Oceanography, 16(6):891–899, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Ben-Avkaham:1971:ASB

[BA71b] Zvi Ben-Avkaham. Accumulation of stones on beaches by Codium fragile. Limnology and Oceanography, 16(3):553–554, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Billmire:1976:SLF

[BA76] E. Billmire and S. Aaronson. The secretion of lipids by the freshwater phytoflagellate Ochromonas danica. Limnology and Oceanography, 21(1):138–140, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Ben-Avraham:1977:TSD

[BAHA77] Zvi Ben-Avraham, Ralph Hënel, and Gad Assaf. The thermal structure of the Dead Sea. Limnology and Oceanography, 22(6): 1076–1078, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Baker:1970:IM

[Bak70] A. L. Baker. An inexpensive microsampler. Limnology and Oceanography, 15(1):158–160, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Balch:1972:ACC

[Bal72] Norval Balch. ATP content of Calanus finmarchicus. Limnology and Oceanography, 17(6):906–908, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Baldridge:1976:RCF

[Bal76] H. David Baldridge. Reply to comment by F. M. Morel and D. M. Anderson. Limnology and Oceanography, 21(4):627–632, July 1976. CODEN LIOCAH. ISSN 0024-3590. See [MA76].

Banse:1973:BRK

[Ban73] Karl Banse. Book reviews: Kinne, O. [Ed.]. 1972. Marine ecology, v. 1. Environmental factors, part 2. Wiley-Interscience, New York. p. 683–1244. $32.50. Kinne, O. [Ed.]. 1972. Marine ecology, v. 1. Environmental factors, part 3. Wiley-Interscience, New York. p. 1245–1774. $34.00. Limnology and Oceanography, 18(2):349–350, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Bannister:1974:GTS

[Ban74a] T. T. Bannister. A general theory of steady state phytoplankton growth in a nutrient saturated mixed layer. Limnology and Oceanography, 19(1):13–30, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Bannister:1974:PET

[Ban74b] T. T. Bannister. Production equations in terms of chlorophyll concentration, quantum yield, and upper limit to production. Limnology and Oceanography, 19(1):1–12, January 1974. CODEN LI-OCAH. ISSN 0024-3590.

Banse:1974:IDC

[Ban74c] Karl Banse. On the interpretation of data for the carbon-to-nitrogen ratio of phytoplankton. Limnology and Oceanography, 20(2):695–699, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Bannister:1975:RCG

[Ban75] T. T. Bannister. Reply to comment by G. A. Riley. Limnology and Oceanography, 20(1):152–153, January 1975. CODEN LIOCAH. ISSN 0024-3590. See [Ril75].

Bannister:1979:QDS

[Ban79] T. T. Bannister. Quantitative description of steady state, nutrientsaturated algal growth, including adaptation. Limnology and Oceanography, 24(1):76–96, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Bartsch:1971:BRE

[Bar71] A. F. Bartsch. Book review: Eutrophication: Causes, consequences, correctives. Proceedings of a Symposium. 1969. Nat. Acad. Sci. Publ. 1700. Washington, D.C. 661 p. $13.50. Limnology and Oceanography, 16(1):153–154, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Bassin:1975:SMC

[Bas75] N. Jay Bassin. Suspended marine clay mineral identification by scanning electron microscopy and energy-dispersive X-ray analysis. Limnology and Oceanography, 20(1):133–137, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Battarbee:1973:NME

[Bat73] Richard W. Battarbee. A new method for the estimation of absolute microfossil numbers, with reference especially to diatoms. Limnology and Oceanography, 18(4):647–653, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Bates:1976:ELA

[Bat76] Stephen S. Bates. Effects of light and ammonium on nitrate uptake by two species of estuarine phytoplankton. Limnology and Oceanography, 21(2):212–218, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Baxter:1977:PFL

[Bax77] R. M. Baxter. Predicting the future with lakes and rivers: a note on the Awda Negast. Limnology and Oceanography, 22(4):777–780, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Bayly:1978:VSD

[Bay78] I. A. E. Bayly. Variation in sexual dimorphism in nonmarine Calanoid copepods and its ecological significance. Limnology and Oceanography, 23(6):1224–1228, November 1978. CODEN LI-OCAH. ISSN 0024-3590.

Bott:1970:GMP

[BB70] Thomas L. Bott and Thomas D. Brock. Growth and metabolism of periphytic bacteria: methodology. Limnology and Oceanography, 15(3):333–342, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Boylen:1973:BDP

[BB73] Charles Boylen and Thomas D. Brock. Bacterial decomposition processes in Lake Wingra sediments during winter. Limnology and Oceanography, 18(4):628–634, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Bishop:1975:PRZ

[BB75] John W. Bishop and John P. Barloiv. Phosphorus release by zooplankton. Limnology and Oceanography, 20(1):148–149, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Baker:1976:EPW

[BB76] Alan L. Baker and Kathleen Kromer Baker. Estimation of planktonic wind drift by transmissometry. Limnology and Oceanography, 21(3):447–452, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Brock:1977:MQA

[BB77] Richard E. Brock and Julie H. Brock. A method for quantitatively assessing the infaunal community in coral rock. Limnology and Oceanography, 22(5):948–951, September 1977. CODEN LI-OCAH. ISSN 0024-3590.

Barton:1979:MMC

[BB79] C. E. Barton and F. R. Burden. Modifications to the Mackereth corer. Limnology and Oceanography, 24(5):977–983, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Brooks:1979:CAB

[BBBS79] James M. Brooks, Thomas J. Bright, Bernie B. Bernard, and Claude R. Schwab. Chemical aspects of a brine pool at the East Flower Garden Bank, northwestern Gulf of Mexico. Limnology and Oceanography, 24(4):735–745, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Barnes:1975:AND

[BBG75] R. O. Barnes, K. K. Bertine, and E. D. Goldberg. N_{2}: Ar, nitrification and denitrification in southern California borderland basin sediments. Limnology and Oceanography, 20(4):962–970, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Brooks:1971:NFE

[BBPK71] Ralph H. Brooks, Jr., Patrick L. Brezonik, Hugh D. Putnam, and Michael A. Keirn. Nitrogen fixation in an estuarine environment: the Waccasassa on the Florida Gulf Coast. Limnology and Oceanography, 16(5):701–710, September 1971. CODEN LI-OCAH. ISSN 0024-3590.

Bezdek:1972:SCM

[BC72] H. F. Bezdek and A. F. Carlucci. Surface concentration of marine bacteria. Limnology and Oceanography, 17(4):566–569, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Bezdek:1974:CRL

[BC74] H. F. Bezdek and A. F. Carlucci. Concentration and removal of liquid microlayers from a seawater surface by bursting bubbles. Limnology and Oceanography, 19(1):126–132, January 1974. CO-DEN LIOCAH. ISSN 0024-3590.

Brinkhurst:1971:FAA

[BCB71] R. O. Brinkhurst, K. E. Chua, and E. Batoosingh. The free amino acids in the sediments of Toronto Harbor. Limnology and Oceanography, 16(3):555–559, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Brinkhurst:1972:IIS

[BCK72] Ralph O. Brinkhurst, Kian E. Chua, and Narinder K. Kaushik. Interspecific interactions and selective feeding by tubificid oligochaetes. Limnology and Oceanography, 17(1):122–133, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Burr:1972:PSU

[BD72] A. H. Burr and M. J. Duncan. Portable spectroradiometer for underwater environments. Limnology and Oceanography, 17(3): 466–475, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Brown:1977:CSF

[BDM77] S. R. Brown, R. J. Daley, and R. N. McNeely. Composition and stratigraphy of the fossil phorbin derivatives of Little Round Lake, Ontario. Limnology and Oceanography, 22(2):336–348, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Booker:1976:CMC

[BDW76] M. J. Booker, M. T. Dinsdale, and A. E. Walsby. A continuously monitored column for the study of stratification by planktonic organisms. Limnology and Oceanography, 21(6):915–919, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Begg:1970:LOL

[Beg70] G. W. Begg. Limnological observations on Lake Kariba during 1967 with emphasis on some special features. Limnology and Oceanography, 15(5):776–788, September 1970. CODEN LI-OCAH. ISSN 0024-3590.

Begg:1976:RBD

[Beg76] G. W. Begg. The relationship between the diurnal movements of some of the zooplankton and the sardine Limnothrissa miodon in Lake Kariba, Rhodesia. Limnology and Oceanography, 21(4): 529–539, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Belman:1978:REP

[Bel78] Bruce W. Belman. Respiration and the effects of pressure on the mesopelagic vertically migrating squid Histioteuthis heteropsis. Limnology and Oceanography, 23(4):735–739, July 1978. CO-DEN LIOCAH. ISSN 0024-3590.

Bennett:1975:AEO

[Ben75] John R. Bennett. Another explanation of the observed cyclonic circulation of large lakes. Limnology and Oceanography, 20(1): 108–110, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Benke:1979:MHM

[Ben79] Arthur C. Benke. A modification of the Hynes method for estimating secondary production with particular significance for multivoltine populations. Limnology and Oceanography, 24(1):168–171, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Berger:1970:PFD

[Ber70a] Wolfgang H. Berger. Planktonic Foraminifera: differential production and expatriation off Baja California. Limnology and Oceanography, 15(2):183–204, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Berman:1970:APP

[Ber70b] T. Berman. Alkaline phosphatases and phosphorus availability in Lake Kinneret. Limnology and Oceanography, 15(5):663–674, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Berman:1972:PCC

[Ber72] T. Berman. Profiles of chlorophyll concentration by In Vivo fluorescence: Some limnological applications. Limnology and Oceanography, 17(4):616–618, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Berman:1974:UWL

[Ber74] T. Berman. Urea in the waters of Lake Kinneret (Sea of Galilee). Limnology and Oceanography, 19(6):977–980, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Berner:1977:SMN

[Ber77] Robert A. Berner. Stoichiometric models for nutrient regeneration m anoxic sediments. Limnology and Oceanography, 22(5):781–786, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Berg:1975:CTM

[BFFB75] Walter W. Berg, Jr., Peter Fleischer, Gary R. Freitag, and Emmett L. Bryant. A continuous turbidity monitoring system for coastal surface waters. Limnology and Oceanography, 20(1):137– 141, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Bovbjerg:1976:FNM

[BFM76] Richard V. Bovbjerg, Judy Freitag, and Dana McHaney. A fixed net method to determine the vertical migration of plankton. Limnology and Oceanography, 21(6):920–922, November 1976. CO-DEN LIOCAH. ISSN 0024-3590.

Bright:1972:ETS

[BFMF72] Thomas Bright, Frank Ferrari, Douglas Martin, and Guy A. Franceschini. Effects of a total solar eclipse on the vertical distribution of certain oceanic zooplankters. Limnology and Oceanography, 17(2):296–301, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Baxter:1970:NLL

[BG70] R. M. Baxter and D. L. Golobitsh. A note on the limnology of Lake Hayq, Ethiopia. Limnology and Oceanography, 15(1):144– 149, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Bowden:1971:CFE

[BG71] K. F. Bowden and R. M. Gilligan. Characteristic features of estuarine circulation as represented in the Mersey Estuary. Limnology and Oceanography, 16(3):490–502, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Bertine:1972:TEC

[BG72] K. K. Bertine and E. D. Goldberg. Trace elements in clams, mussels, and shrimp. Limnology and Oceanography, 17(6):877–884, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Brewer:1976:ACG

[BG76] Peter G. Brewer and Joel C. Goldman. Alkalinity changes generated by phytoplankton growth. Limnology and Oceanography, 21 (1):108–117, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Bowers:1978:HHM

[BG78] James A. Bowers and Nevin E. Grossnickle. The herbivorous habits of Mysis relicta in Lake Michigan. Limnology and Oceanography, 23(4):767–776, July 1978. CODEN LIOCAH. ISSN 0024-3590.

B-G:1974:FIB

[BGMW74] Enrique Carrillo B.-G, Charles B. Miller, and Peter H. Wiebe. Failure of interbreeding between Atlantic and Pacific populations of the marine calanoid copepod Acartia clausi Giesbrecht. Limnology and Oceanography, 19(3):452–458, May 1974. CODEN LI-OCAH. ISSN 0024-3590.

Banin:1975:SSA

[BGZS75] A. Banin, M. Gal, Y. Zohar, and A. Singer. The specific surface area of clays in lake sediments-measurement and analysis of contributors in Lake Kinneret, Israel. Limnology and Oceanography, 19(4):278–282, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Burton:1970:OAI

[BH70] J. D. Burton and P. C. Head. Observations on the analysis of iron in seawater with particular reference to estuarine waters. Limnology and Oceanography, 15(1):164–167, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Boyce:1975:SDM

[BH75] F. M. Boyce and P. F. Hamblin. A simple diffusion model of the mean field distribution of soluble materials in the Great Lakes.

Limnology and Oceanography, 20(4):511–517, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Brown:1978:KAT

[BH78] Edward J. Brown and Robin F. Harris. Kinetics of algal transient phosphate uptake and the cell quota concept. Limnology and Oceanography, 23(1):35–40, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Blakar:1976:LPI

[BHD76] Inggard Arne Blakar, Dag Hongve, and Ivan Digernes. A lightweight, power-independent filtration apparatus for field use. Limnology and Oceanography, 21(4):617–619, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Brown:1978:KPU

[BHK78] Edward J. Brown, Robin F. Harris, and Joseph F. Koonce. Kinetics of phosphate uptake by aquatic microorganisms: Deviations from a simple Michaelis–Menten equation. Limnology and Oceanography, 23(1):26–34, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Bernstein:1978:SDB

[BHSJ78] Brock B. Bernstein, Robert R. Hessler, Robert Smith, and Peter A. Jumars. Spatial dispersion of benthic Foraminifera in the abyssal central North Pacific. Limnology and Oceanography, 23 (3):401–416, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Bienfang:1975:SSA

[Bie75] Paul K. Bienfang. Steady state analysis of nitrate-ammonium assimilation by phytoplankton. Limnology and Oceanography, 20 (4):402–411, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Biggs:1977:RAE

[Big77] Douglas C. Biggs. Respiration and ammonium excretion by open ocean gelatinous zooplankton. Limnology and Oceanography, 22 (1):108–117, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Boothe:1972:PIF

[BK72] Paul N. Boothe and George A. Knauer. The possible importance of fecal material in the biological amplification of trace and heavy metals. Limnology and Oceanography, 17(2):270–274, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Bruland:1978:CNP

[BKM78] Kenneth W. Bruland, George A. Knauer, and John H. Martin. Cadmium in northeast Pacific waters. Limnology and Oceanography, 23(4):618–625, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Beers:1970:PRP

[BKS70] John R. Beers, Cameron Knox, and John D. H. Strickland. A permanent record of plankton samples using holography. Limnology and Oceanography, 15(6):967–970, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Bayne:1972:SCN

[BL72a] David R. Bayne and John M. Lawrence. Separating constituents of natural phytoplankton populations by continuous particle electrophoresis. Limnology and Oceanography, 17(3):481–489, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Bunt:1972:DCD

[BL72b] J. S. Bunt and C. C. Lee. Data on the composition an dark survival of four sea-ice microalgae. Limnology and Oceanography, 17(3):458–461, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Bortleson:1974:PIM

[BL74] Gilbert C. Bortleson and G. Fred Lee. Phosphorus, iron, and manganese distribution in sediment cores of six Wisconsin lakes. Limnology and Oceanography, 19(5):794–801, September 1974. CO-DEN LIOCAH. ISSN 0024-3590.

Blankley:1976:TRC

[BL76] William F. Blankley and Ralph A. Lewin. Temperature responses of a coccolithophorid, Cricosphaera carterae, measured in a simple and inexpensive thermal-gradient device. Limnology and Oceanography, 21(3):457–462, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Bilby:1979:EHF

[BL79] Robert E. Bilby and Gene E. Likens. Effect of hydrologic fluctuations on the transport of fine particulate organic carbon in a small stream. Limnology and Oceanography, 24(1):69–75, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Blackburn:1973:RBB

[Bla73a] Maurice Blackburn. Regressions between biological oceanographic measurements in the eastern tropical Pacific and their significance to ecological efficiency. Limnology and Oceanography, 18(4):552– 563, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Blanton:1973:VEE

[Bla73b] Jackson O. Blanton. Vertical entrainment into the epilimnia of stratified lakes. Limnology and Oceanography, 18(5):697–704, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Blasco:1977:RTU

[Bla77] Dolors Blasco. Red tide in the upwelling region of Baja California. Limnology and Oceanography, 22(2):255–263, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Blakar:1979:CIW

[Bla79] Inggard Arne Blakar. A close-interval water sampler with minimal disturbance properties. Limnology and Oceanography, 24(5):983– 988, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Burton:1970:RDS

[BLL70] J. D. Burton, T. M. Leatherland, and P. S. Liss. The reactivity of dissolved silicon in some natural waters. Limnology and Oceanography, 15(3):473–476, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Bell:1974:SSM

[BLM74] Wayne H. Bell, Jeanne M. Lang, and Ralph Mitchell. Selective stimulation of marine bacteria by algal extracellular products. Limnology and Oceanography, 19(5):833–839, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Brylinsky:1973:AFG

[BM73] M. Brylinsky and K. H. Mann. An analysis of factors governing productivity in lakes and reservoirs. Limnology and Oceanography, 18(1):1–14, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Benson:1974:WCM

[BM74] A. A. Benson and L. Muscatine. Wax in coral mucus: Energy transfer from corals to reef fishes. Limnology and Oceanography, 19(5):810–814, September 1974. CODEN LIOCAH. ISSN 0024-

Brylinsky:1975:IMN

[BM75] M. Brylinsky and K. H. Mann. The influence of morphometry and of nutrient dynamics on the productivity of lakes. Limnology and Oceanography, 20(5):666–667, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Bowen:1974:CSR

[BNV^{+}74] Vaughan T. Bowen, Victor E. Noshkin, H. L. Volchok, H. D. Livingston, and K. M. Wong. Cesium 137 to strontium 90 ratios in the Atlantic Ocean 1966 through 1972. Limnology and Oceanography, 20(2):670–681, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Booker:1975:BRO

[Boo75a] John R. Booker. Book review: Officer, C. B. 1974. Introduction to theoretical geophysics. Springer-Verlag, New York, Heidelberg, Berlin, x + 385 p. $19.80. Limnology and Oceanography, 19(4): 299, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Boon:1975:TDA

[Boo75b] John D. Boon III. Tidal discharge asymmetry in a salt marsh drainage system. Limnology and Oceanography, 20(1):71–80, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Booth:1975:GOP

[Boo75c] Beatrice C. Booth. Growth of oceanic phytoplankton in enrichment cultures. Limnology and Oceanography, 20(4):865–869, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Booth:1976:DEM

[Boo76] Charles R. Booth. The design and evaluation of a measurement system for photosynthetically active quantum scalar irradiance. Limnology and Oceanography, 21(2):326–336, March 1976. CO-DEN LIOCAH. ISSN 0024-3590.

Bott:1975:BGR

[Bot75] Thomas L. Bott. Bacterial growth rates and temperature optima in a stream with a fluctuating thermal regime. Limnology and Oceanography, 20(2):191–197, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Bouma:1975:BRI

[Bou75] Arnold H. Bouma. Book review: Inderbitzen, A. L. [Ed.]. 1974. Deep-sea sediments: Physical and mechanical properties. Marine Science, v. 2, Plenum Publ. Corp., New York, ix + 497 p. $35.00. Limnology and Oceanography, 20(3):682, July 1975. CODEN LI-OCAH. ISSN 0024-3590.

Bowden:1979:BRN

[Bow79] K. F. Bowden. Book review: Nihoul, J. C. J. [Ed.] 1978. Hydrodynamics of estuaries and fjords. Proc. 9th Int. Liege Colloq. Ocean Hydrodynamics. Elsevier Sci. Publ. Co., Amsterdam and New York, xiv + 546 p. $57.00. Limnology and Oceanography, 24 (3):600, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Boyd:1976:SPS

[Boy76] Carl M. Boyd. Selection of particle sizes by filter-feeding copepods: a plea for reason. Limnology and Oceanography, 21(1):175–180, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Blanton:1970:SBB

[BP70] Jackson Blanton and June G. Pattullo. The subsurface boundary between subarctic Pacific water and Pacific equatorial, water in the transition zone off southern California. Limnology and Oceanography, 15(4):606–614, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Berger:1972:PFD

[BP72] Wolfgang H. Berger and David J. W. Piper. Planktonic Foraminifera: differential settling, dissolution, and redeposition. Limnology and Oceanography, 17(2):275–287, March 1972. CO-DEN LIOCAH. ISSN 0024-3590.

Berman:1974:ASV

[BP74a] Thomas Berman and Utsa Pollingher. Annual and seasonal variations of phytoplankton, chlorophyll, and photosynthesis in Lake Kinneret. Limnology and Oceanography, 19(1):31–55, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Bourke:1974:SVW

[BP74b] Robert H. Bourke and June G. Pattullo. Seasonal variation of the water mass along the Oregon–northern California coast. Limnology and Oceanography, 19(2):190–198, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Bancroft:1976:EMA

[BPW76] K. Bancroft, E. A. Paul, and W. J. Wiehe. The extraction and measurement of adenosine triphosphate from marine sediments.

Limnology and Oceanography, 21(3):473–480, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Baudouin:1972:WSC

[BR72] M. F. Baudouin and O. Ravera. Weight, size, and chemical composition of some freshwater zooplankters: Daphnia hyalina (Leydig). Limnology and Oceanography, 17(4):645–649, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Berman:1974:FBD

[BR74] Mark S. Berman and Sumner Richman. The feeding behavior of Daphnia pulex from Lake Winnebago, Wisconsin. Limnology and Oceanography, 19(1):105–109, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Bradbury:1971:PLT

[Bra71] John P. Bradbury. Paleolimnology of Lake Texcoco, Mexico. Evidence from diatoms. Limnology and Oceanography, 16(2):180–200, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Bradford:1973:DSC

[Bra73a] Wesley L. Bradford. The determination of a stability constant for the aqueous complex Zn(OH)_{2}^{0} using anodic stripping voltammetry. Limnology and Oceanography, 18(5):757–762, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Bradshaw:1973:ECD

[Bra73b] Alvin Bradshaw. The effect of carbon dioxide on the specific volume of seawater. Limnology and Oceanography, 18(1):95–105, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Bradley:1976:MTT

[Bra76] Brian P. Bradley. The measurement of temperature tolerance: Verification of an index. Limnology and Oceanography, 21(4):596– 599, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Briskhurst:1972:BRE

[Bri72] R. O. Briskhurst. Book review: Edwards, R. W., and D. J. Carrod [Eds.]. 1972. Conservation and productivity of natural waters. Academic Press, London and New York. xv + 318 p. £5.50. Limnology and Oceanography, 17(5):802–803, September 1972. CODEN LI-OCAH. ISSN 0024-3590.

Brinson:1976:OML

[Bri76] Mark M. Brinson. Organic matter losses from four watersheds in the humid tropics. Limnology and Oceanography, 21(4):572–582, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Brinkhurst:1979:BRH

[Bri79] R. Brinkhurst. Book review: Hellawell, J. M. 1978. Biological surveillance of rivers. Water Res. Centre, Stevanage. xii + 332 p. (Typewriter composition). Limnology and Oceanography, 24(4): 793–794, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Brock:1970:BRP

[Bro70] Thomas D. Brock. Book review: Perfil’ev, B. V., and D. R. Gabe. 1969. Capillary methods of investigating micro-organisms. Oliver and Boyd, Edinburgh. 627 p. £9 10s. Limnology and Oceanography, 15(1):168–169, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Brown:1974:NTS

[Bro74] Helen Davis Brown. A new technique for the study of epiphytic algae. Limnology and Oceanography, 19(1):151–152, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Brock:1978:CLA

[Bro78a] Thomas D. Brock. Comment on “Loss of ^{14}C activity after chemical fixation of phytoplankton” (M. W. Silver and P. J. Davoll). Limnology and Oceanography, 23(5):1083–1084, September 1978. CODEN LIOCAH. ISSN 0024-3590. See [SD78].

Brock:1978:UFM

[Bro78b] Thomas D. Brock. Use of fluorescence microscopy for quantifying phytoplankton, especially filamentous blue-green algae. Limnology and Oceanography, 23(1):158–160, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Brown:1979:MPD

[Bro79] Keith A. Brown. A modified procedure for determining [^{14}C]CO_{2} produced by respiration in water or sediment samples. Limnology and Oceanography, 24(6):1141–1145, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Brundritt:1971:DDV

[Bru71] J. K. Brundritt. A device for determining velocity of flow near the substrate. Limnology and Oceanography, 16(1):120–123, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Beers:1970:PAF

[BS70] John R. Beers and Gene L. Stewart. The preservation of acantharians in fixed plankton samples. Limnology and Oceanography, 15(5):825–827, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Brewer:1971:CDM

[BS71] P. G. Brewer and D. W. Spencer. Colorimetric determination of manganese in anoxic waters. Limnology and Oceanography, 16(1): 107–110, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Blanchard:1974:BTA

[BS74] Duncan C. Blanchard and Lawrence D. Syzdek. Bubble tube: Apparatus for determining rate of collection of bacteria by an air bubble rising in water. Limnology and Oceanography, 19(1):133– 138, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Blanchard:1975:ECJ

[BS75] Duncan C. Blanchard and Lawrence D. Syzdek. Electrostatic collection of jet and film drops. Limnology and Oceanography, 20(4): 762–774, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Bhovichitra:1977:LDU

[BS77] Mahn Bhovichitra and Elijah Swift. Light and dark uptake of nitrate and ammonium by large oceanic dinoflagellates: Pyrocystis noctiluca, Pyrocystis fusiformis, and Dissodinium lunula. Limnology and Oceanography, 22(1):73–83, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Blanchard:1978:RCF

[BS78a] Duncan C. Blanchard and Lawrence D. Syzdek. Reply to comment by F. MacIntyre. Limnology and Oceanography, 23(3):573, May 1978. CODEN LIOCAH. ISSN 0024-3590. See [Mac78a].

Blanchard:1978:SPB

[BS78b] Duncan C. Blanchard and Lawrence D. Syzdek. Seven problems in bubble and jet drop researches. Limnology and Oceanography, 23(3):389–400, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Bruno:1978:SVB

[BS78c] Stephen F. Bruno and Robert D. Staker. Seasonal vitamin B_{12} and phytoplankton distribution near Napeague Bay, New York (Block Island Sound). Limnology and Oceanography, 23(5):1045– 1051, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Bottazzi:1971:AAO

[BSB71] Elsa Massera Bottazzi, Bruno Schreiber, and Vaughan T. Bowen. Acantharia in the Atlantic Ocean, their abundance and preservation. Limnology and Oceanography, 16(4):677–684, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Bloesch:1977:PPM

[BSB77] J. Bloesch, P. Stadelmann, and H. Bührer. Primary production, mineralization, and sedimentation in the euphotic zone of two Swiss lakes. Limnology and Oceanography, 22(3):511–526, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Brodie:1978:PDE

[BSS78] P. F. Brodie, D. D. Sameoto, and R. W. Sheldon. Population densities of euphausiids off Nova Scotia as indicated by net samples, whale stomach contents, and sonar. Limnology and Oceanography, 23(6):1264–1267, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Berner:1970:CAP

[BST70] Robert A. Berner, Martha R. Scott, and Catherine Thomlinson. Carbonate alkalinity in the pore waters of anoxic marine sediments. Limnology and Oceanography, 15(4):544–549, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Berger:1970:MGO

[BT70] Leslie R. Berger and Leslie Q. Tam. A method to grow obligately aerobic bacteria at increased hydrostatic pressure. Limnology and Oceanography, 15(3):483–485, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Bulleid:1978:IME

[Bul78] N. C. Bulleid. An improved method for the extraction of adenosine triphosphate from marine sediment and seawater. Limnology and Oceanography, 23(1):174–178, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Bunt:1971:LDO

[Bun71] J. S. Bunt. Levels of dissolved oxygen and carbon fixation by marine microalgae. Limnology and Oceanography, 16(3):564–566, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Burgis:1970:ETD

[Bur70] Mary J. Burgis. The effect of temperature on the development time of eggs of Thermocyclops sp., a tropical cyclopoid copepod from Lake George, Uganda. Limnology and Oceanography, 15(5): 742–747, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Butler:1976:BRR

[But76] James N. Butler. Book review: Riley, J. P., and G. Skirrow [Eds.]. 1975. Chemical oceanography, v. 4. 2nd ed. Academic Press, New York, London, San Francisco, xviii + 363 p. £12.50. Limnology and Oceanography, 21(5):766, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Buzas:1971:ASD

[Buz71] Martin A. Buzas. Analyses of species densities by the multivariate general linear model. Limnology and Oceanography, 16(4):667– 670, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Bell:1970:IOC

[BW70] R. K. Bell and F. J. Ward. Incorporation of organic carbon by Daphnia pulex. Limnology and Oceanography, 15(5):713–726, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Baxter:1973:POH

[BWP73] R. M. Baxter, R. B. Wood, and M. V. Prosser. The probable occurrence of hydroxylarnine in the water of an Ethiopian lake. Limnology and Oceanography, 18(3):470–472, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Ben-Yaakov:1970:MCL

[BY70] S. Ben-Yaakov. A method for calculating the In situ pH of seawater. Limnology and Oceanography, 15(2):326–328, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Ben-Yaakov:1973:PBP

[BY73] Sam Ben-Yaakov. pH buffering of pore water of recent anoxic marine sediments. Limnology and Oceanography, 18(1):86–94, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Bewers:1975:FGC

[BY75] J. M. Bewers and P. A. Yeats. Fluoride: Global circulation. Limnology and Oceanography, 20(1):149–150, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Ben-Yacikov:1974:ISP

[BYR74] S. Ben-Yacikov and E. Ruth. An improved in situ pH sensor for oceanographic and limnological applications. Limnology and Oceanography, 19(1):144–151, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Ben-Yaakov:1977:IEP

[BYS77] Sam Ben-Yaakov and Eytan Sass. Independent estimate of the pH of Dead Sea brine. Limnology and Oceanography, 22(2):374–376, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Coulon:1972:SCP

[CA72] Christopher Coulon and Vera Alexander. A sliding-chamber phytoplankton settling technique for making permanent quantitative slides with applications in fluorescent microscopy and autoradiography. Limnology and Oceanography, 17(1):149–152, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1979:ENT

[CA79] Stephen R. Carpenter and Michael S. Adams. Effects of nutrients and temperature on decomposition of Myriophyllum spicatum L. in a hard-water eutrophic lake. Limnology and Oceanography, 24 (3):520–528, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Crow:1975:DBF

[CACB75] S. A. Crow, D. G. Ahearn, W. L. Cook, and A. W. Bourquin. Densities of bacteria and fungi in coastal surface films as determined by a membrane-adsorption procedure. Limnology and Oceanography, 20(6):644–646, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Chisholm:1978:SAI

[CAE78] S. W. Chisholm, F. Azam, and R. W. Eppley. Silicic acid incorporation in marine diatoms on light : dark cycles: Use as an assay for phased cell division. Limnology and Oceanography, 23(3):518–529, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Cammen:1975:ART

[Cam75] Leon M. Cammen. Accumulation rate and turnover time of organic carbon in a salt marsh sediment. Limnology and Oceanography, 20(3):1012–1015, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Carr:1970:SED

[Car70] Roderick A. Carr. Salinity effects on the determination of strontium in seawater by atomic absorption spectrophotometry. Limnology and Oceanography, 15(2):318–320, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Cartwright:1975:BRL

[Car75] D. E. Cartwright. Book review: Lisitzin, E. 1974. Sea-level changes. Elsevier Oceanogr. Ser. 8. Elsevier Sci. Publ. Co., Amsterdam, Oxford, New York, vi + 286 p. $28.90. Limnology and Oceanography, 20(3):681–682, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Carlson:1977:TSI

[Car77] Robert E. Carlson. A trophic state index for lakes. Limnology and Oceanography, 22(2):361–369, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Cassie:1971:BRE

[Cas71] R. Morrison Cassie. Book review: Elliott, J. M. 1971. Some methods for the statistical analysis of samples of benthic invertebrates. Sci. Publ. 25. Freshwater Biol. Ass., Ambleside, Westmorland, U.K. 144 p. £0.95. Limnology and Oceanography, 16(3):596–597, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Castenholz:1972:HSM

[Cas72a] Richard W. Castenholz. Hot spring microbial communities recreated in modified “Winogradski columns”. Limnology and Oceanography, 17(5):767–772, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Caswell:1972:IFB

[Cas72b] Hal Caswell. On instantaneous and finite birth rates. Limnology and Oceanography, 17(5):787–791, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Castenholz:1973:PPU

[Cas73] Richard W. Castenholz. The possible photosynthetic use of sulfide by the filamentous phototrophic bacteria of hot springs. Limnology and Oceanography, 18(6):863–876, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Cavari:1976:ALK

[Cav76] Benzion Cavari. ATP in Lake Kinneret: Indicator of microbial biomass or of phosphorus deficiency? Limnology and Oceanography, 21(2):231–236, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Chew:1970:SMC

[CB70] Frank Chew and George A. Berberian. Some measurements of current by shallow drogues in the Florida Current. Limnology and Oceanography, 15(1):88–99, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Carlucci:1972:DVB

[CB72] A. F. Carlucci and Peggy M. Bowes. Determination of vitamin B_{12}, thiamine, and biotin in Lake Tahoe waters using modified marine bioassay techniques. Limnology and Oceanography, 17(5): 774–777, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Compere:1973:DCA

[CB73] Edward L. Compere, Jr. and John M. Bates. Determination of calcite: aragonite ratios in mollusc shells by infrared spectra. Limnology and Oceanography, 18(2):326–331, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Confer:1975:OZP

[CB75a] John L. Confer and Pamela I. Blades. Omnivorous zooplankton and planktivorous fish. Limnology and Oceanography, 20(4):571– 579, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Cowgill:1975:DCC

[CB75b] U. M. Cowgill and C. W. Burns. Differences in chemical composition between two species of Daphnia and some freshwater algae cultured in the laboratory. Limnology and Oceanography, 20(3): 1005–1011, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Crossland:1976:ARC

[CB76] C. J. Crossland and D. J. Barnes. Acetylene reduction by coral skeletons. Limnology and Oceanography, 21(1):153–156, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Caldwell:1978:TML

[CBN78] Douglas R. Caldwell, John M. Brubaker, and Victor T. Neal. Thermal microstructure on a lake slope. Limnology and Oceanography, 23(2):372–374, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1974:PPS

[CC74] Edward J. Carpenter and James L. Cox. Production of pelagic Sargassum and a blue-green epiphyte in the western Sargasso Sea. Limnology and Oceanography, 19(3):429–436, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Chapman:1975:QVW

[CC75a] A. R. O. Chapman and C. C. M. Campbell. Quanta vs. watts. Limnology and Oceanography, 20(6):496, May 1975. CODEN LI-OCAH. ISSN 0024-3590.

Coughenower:1975:ATT

[CC75b] D. Douglas Coughenower and Herbert C. Curl, Jr. An automated technique for total dissolved free amino acids in seawater. Limnology and Oceanography, 20(1):128–131, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Caperon:1971:PKS

[CCK71] John Caperon, S. Allen Cattell, and George Krasnick. Phytoplankton kinetics in a subtropical estuary: eutrophication. Limnology and Oceanography, 16(4):599–607, July 1971. CODEN LI-OCAH. ISSN 0024-3590.

Cheng:1973:CHDb

[CE73] Lanna Cheng and J. T. Enright. Can Halobates dodge nets? II: By moonlight? Limnology and Oceanography, 18(4):666–669, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Chris:1979:IAU

[CE79] Wessels Chris and Birnbaum Eliezer. An improved apparatus for use with the ^{14}C acid-bubbling method of measuring primary production. Limnology and Oceanography, 24(1):187–188, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Carmack:1979:ILR

[CGPD79] Eddy C. Carmack, Colin B. J. Gray, Christopher H. Pharo, and Ralph J. Daley. Importance of lake-river interaction on seasonal patterns in the general circulation of Kamloops Lake, British Columbia. Limnology and Oceanography, 24(4):634–644, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Coleman:1970:VDI

[CH70] Mary J. Coleman and H. B. N. Hynes. The vertical distribution of the invertebrate fauna in the bed of a stream. Limnology and Oceanography, 15(1):31–40, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Chave:1974:BRP

[Cha74] Keith E. Chave. Book review: Purser, B. H. [Ed.]. 1973. The Persian Gulf. Holocene carbonate sedimentation and diagenesis in a shallow epicontinental sea. Springer-Verlag, New York, Heidelberg, and Berlin, viii + 471 p. + 3 maps. $29.70. Limnology and Oceanography, 19(2):376, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Chapra:1979:APL

[Cha79a] Steven C. Chapra. Applying phosphorus loading models to embayments. Limnology and Oceanography, 24(1):163–168, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Chase:1979:SBN

[Cha79b] Robert R. P. Chase. Settling behavior of natural aquatic particulates. Limnology and Oceanography, 24(3):417–426, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Cheng:1973:CHDa

[Che73] Lanna Cheng. Can Halobates dodge nets? I: By daylight? Limnology and Oceanography, 18(4):663–665, July 1973. CODEN LI-OCAH. ISSN 0024-3590.

Childress:1971:RRD

[Chi71] James J. Childress. Respiratory rate and depth of occurrence of midwater animals. Limnology and Oceanography, 16(1):104–106, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Chutter:1972:RNU

[Chu72] F. M. Chutter. A reappraisal of Needham and Usinger’s data on the variability of a stream fauna when sampled with a surber sampler. Limnology and Oceanography, 17(1):139–141, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Clark:1972:DFA

[CJN72] M. E. Clark, G. A. Jackson, and W. J. North. Dissolved free amino acids in southern California coastal water. Limnology and Oceanography, 17(5):749–758, September 1972. CODEN LI-OCAH. ISSN 0024-3590.

Coyne:1974:CDP

[CK74] Patrick I. Coyne and John J. Kelley. Carbon dioxide partial pressures in Arctic surface waters. Limnology and Oceanography, 19 (6):928–938, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Cattaneo:1979:PPA

[CK79] A. Cattaneo and J. Kalff. Primary production of algae growing on natural and artificial aquatic plants: a study of interactions between epiphytes and their substrate. Limnology and Oceanography, 24(6):1031–1037, November 1979. CODEN LIOCAH. ISSN 0024-3590. See comment [GG81].

Cohen:1977:SLSa

[CKGS77] Yehuda Cohen, Wolfgang E. Krumbein, Moshe Goldberg, and Moshe Shilo. Solar Lake (Sinai). 1. Physical and chemical limnology. Limnology and Oceanography, 22(4):597–608, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Cohen:1977:SLSb

[CKS77a] Yehuda Cohen, Wolfgang E. Krumbein, and Moshe Shilo. Solar Lake (Sinai). 2. Distribution of photosynthetic microorganisms and primary production. Limnology and Oceanography, 22 (4):609–620, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Cohen:1977:SLSc

[CKS77b] Yehuda Cohen, Wolfgang E. Krumbein, and Moshe Shilo. Solar Lake (Sinai). 3. Bacterial distribution and production. Limnology and Oceanography, 22(4):621–634, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Codispoti:1973:RSB

[CL73] L. A. Codispoti and Dorothy Lowman. A reactive silicate budget for the Arctic Ocean. Limnology and Oceanography, 18(3):448– 456, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Collos:1976:BSZ

[CL76] Yves Collos and Joyce Lewin. Blooms of surf-zone diatoms along the coast of the Olympic Peninsula, Washington. 7. Variations of the carbon-to-nitrogen ratio in field samples and laboratory cultures of Chaetoceros armatum. Limnology and Oceanography, 21(2):219–225, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Cole:1979:MMP

[CL79] Jonathan J. Cole and Gene E. Likens. Measurements of mineralization of phytoplankton detritus in an oligotrophic lake. Limnology and Oceanography, 24(3):541–547, May 1979. CODEN LI-OCAH. ISSN 0024-3590.

Cleaver:1972:BRM

[Cle72] Fred Cleaver. Book review: Morgan, A. E. 1971. Dams and other disasters. A century of the Army Corps of Engineers in civil works. Porter Sargent, Boston. xiii + 422 p. $7.50. Limnology and Oceanography, 17(1):166–167, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Clifford:1972:CSS

[Cli72] Hugh F. Clifford. Comparison of samples of stream bottom fauna collected during the day and at night. Limnology and Oceanography, 17(3):479–481, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Cooper:1972:IST

[CM72] Mary F. Cooper and Richard Y. Morita. Interaction of salinity and temperature on net protein synthesis and viability of Vibrio marinus. Limnology and Oceanography, 17(4):556–565, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1975:NFU

[CM75] Edward J. Carpenter and James J. McCarthy. Nitrogen fixation and uptake of combined nitrogenous nutrients by Oscillatoria (Trichodesmium) thiebautii in the western Sargasso Sea. Limnology and Oceanography, 20(4):389–401, May 1975. CODEN LI-OCAH. ISSN 0024-3590.

Chen:1977:ESC

[CM77] Chen-Tung Chen and Frank J. Millero. Effect of salt content on the temperature of maximum density and on static stability in Lake Ontario. Limnology and Oceanography, 22(1):158–159, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Calvert:1979:EIR

[CM79] S. E. Calvert and M. J. McCartney. The effect of incomplete recovery of large particles from water samplers on the chemical composition of oceanic particulate matter. Limnology and Oceanography, 24(3):532–536, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Cintron:1970:SCB

[CMB70] Gilberto Cintrón, William S. Maddux, and Paul R. Burkholder. Some consequences of brine pollution in the Bahía Fosforescente, Puerto Rico. Limnology and Oceanography, 15(2):246–249, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Clark:1978:EAA

[CMDC78] James R. Clark, David I. Messenger, Kenneth L. Dickson, and John Cairns, Jr. Extraction of ATP from Aufwuchs communities. Limnology and Oceanography, 23(5):1055–1059, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Charters:1973:EWM

[CNC73] A. C. Charters, M. Neushul, and D. Coon. The effect of water motion on algal spore adhesion. Limnology and Oceanography, 18 (6):884–896, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Codispoti:1975:NTT

[CO75] L. A. Codispoti and T. G. Owens. Nutrient transports through Lancaster Sound in relation to the Arctic Ocean’s reactive silicate budget and the outflow of Bering Strait waters. Limnology and Oceanography, 20(1):115–119, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Cole:1972:BRW

[Col72] Raelyn Cole. Book review: Walker, T. J. 1971. Red salmon, brown bear. the American Museum of Natural History’s The New Explorers. World Publ. Co., New York. xiv + 226 p. $8.95. Limnology and Oceanography, 17(3):502–503, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Connors:1970:ES

[Con70] Donald N. Connors. On the enthalpy of seawater. Limnology and Oceanography, 15(4):587–594, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Confer:1971:IPM

[Con71] John L. Confer. Intrazooplankton predation by Mesocyclops edax at natural prey densities. Limnology and Oceanography, 16(4): 663–666, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Conomos:1975:MSO

[Con75] T. J. Conomos. Movement of spilled oil as predicted by estuarine nontidal drift. Limnology and Oceanography, 20(2):159–173, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Cooley:1971:ETD

[Coo71] John M. Cooley. The effect of temperature on the development of resting eggs of Diaptomus oregonensis Lillj (Copepoda: Calanoida). Limnology and Oceanography, 16(6):921–926, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Cooke:1973:UAC

[Coo73a] Robert C. Cooke. The use of activated charcoal for the removal of oxygen from gas systems. Limnology and Oceanography, 18(1): 150–152, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Cooper:1973:ENP

[Coo73b] David C. Cooper. Enhancement of net primary productivity by herbivore grazing in aquatic laboratory microcosms. Limnology and Oceanography, 18(1):31–37, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Covich:1976:RCM

[Cov76] Alan Covich. Recent changes in molluscan species diversity of a large tropical lake (Lago de Peten, Guatemala). Limnology and Oceanography, 21(1):51–59, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Cowell:1970:IPD

[Cow70] Bruce C. Cowell. The influence of plankton discharges from an upstream reservoir on standing crops in a Missouri River reservoir. Limnology and Oceanography, 15(3):427–441, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Culberson:1970:NBW

[CP70] C. Culberson and R. M. Pytkowicz. A near-bottom water sampler. Limnology and Oceanography, 15(1):160–162, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Czaplewski:1973:UBO

[CP73] Raymond L. Czaplewski and Michael Parker. Use of a BOD oxygen probe for estimating primary productivity. Limnology and Oceanography, 18(1):152–154, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Culberson:1975:NBC

[CP75] C. H. Culberson and R. M. Pytkoivicz. Near-bottom chemistry in the eastern Pacific and North Atlantic Oceans. Limnology and Oceanography, 20(6):463–467, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1977:NFD

[CP77] Edward J. Carpenter and Charles C. Price IV. Nitrogen fixation, distribution, and production of Oscillatoria (Trichodesmium) spp. in the western Sargasso and Caribbean Seas. Limnology and Oceanography, 22(1):60–72, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Christensen:1979:RET

[CP79] J. P. Christensen and T. T. Packard. Respiratory electron transport activities in phytoplankton and bacteria: Comparison of methods. Limnology and Oceanography, 24(3):576–583, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Capone:1979:RBP

[CPOT79] Douglas G. Capone, Polly A. Penhale, Ronald S. Oremland, and Barrie F. Taylor. Relationship between productivity and N_{2} (C_{2} H_{2}) fixation in a Thalassia testudinum community. Limnology and Oceanography, 24(1):117–125, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Cushing:1970:CZC

[CR70] Colbert E. Cushing and Fred L. Rose. Cycling of zinc-65 by Columbia River periphyton in a closed lotic microcosm. Limnology and Oceanography, 15(5):762–767, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Cline:1972:ODC

[CR72] Joel D. Cline and Francis A. Richards. Oxygen deficient conditions and nitrate reduction in the eastern tropical north Pacific Ocean. Limnology and Oceanography, 17(6):885–900, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Codispoti:1976:AHR

[CR76] L. A. Codispoti and F. A. Richards. An analysis of the horizontal regime of denitrification in the eastern tropical North Pacific. Limnology and Oceanography, 21(3):379–388, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Clasby:1972:CIS

[CRA72] R. C. Clasby, W. S. Reeburgh, and Vera Alexander. A closeinterval syringe sampler. Limnology and Oceanography, 17(4):632– 633, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Crecelius:1975:GCA

[Cre75] Eric A. Crecelius. The geochemical cycle of arsenic in Lake Washington and its relation to other elements. Limnology and Oceanography, 20(6):441–451, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Crowley:1973:FRI

[Cro73] Philip H. Crowley. Filtering rate inhibition of Daphnia pulex in Wintergreen Lake water. Limnology and Oceanography, 18(3): 394–402, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1972:UUS

[CRW72] Edward J. Carpenter, Charles C. Remsen, and Stanley W. Watson. Utilization of urea by some marine phytoplankters. Limnology and Oceanography, 17(2):265–269, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Chave:1970:CCS

[CS70] Keith E. Chave and Erwin Suess. Calcium carbonate saturation in seawater: effects of dissolved organic matter. Limnology and Oceanography, 15(4):633–637, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Cescon:1973:DLP

[CS73a] B. S. Cescon and P. G. Scarazzato. Determination of low phosphate concentrations in seawater by an isobutyl acetate extraction procedure. Limnology and Oceanography, 18(3):499–500, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Coles:1973:OCM

[CS73b] Stephen L. Coles and Richard Strathmann. Observations on coral mucus “flocs” and their potential trophic significance. Limnology and Oceanography, 18(4):673–678, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Covich:1974:COM

[CS74] Alan Covich and Minze Stuiver. Changes in oxygen 18 as a measure of long-term fluctuations in tropical lake levels and molluscan populations. Limnology and Oceanography, 20(2):682–691, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Caperon:1978:PRM

[CS78] J. Caperon and D. F. Smith. Photosynthetic rates of marine algae as a function of inorganic carbon concentration. Limnology and Oceanography, 23(4):704–708, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Csanady:1975:BRS

[Csa75] G. T. Csanady. Book review: Stern, M. E. 1975. Ocean circulation physics. Academic Press, New York, x + 246 p. $19.50. Limnology and Oceanography, 20(3):1058–1059, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Cushman:1978:EGC

[CSHE78] R. M. Cushman, H. H. Shugart, Jr., S. G. Hildebrand, and J. W. Elwood. The effect of growth curve and sampling regime on instantaneous-growth, removal-summation, and Hynes/Hamilton estimates of aquatic insect production: a computer simulation. Limnology and Oceanography, 23(1):184–189, January 1978. CO-DEN LIOCAH. ISSN 0024-3590.

Coachman:1970:CNB

[CT70] L. K. Coachman and R. B. Tripp. Currents north of Bering Strait in winter. Limnology and Oceanography, 15(4):625–632, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Carder:1972:TEI

[CTB72] Kendall L. Carder, Richad D. Tomlinson, and George F. Beardsley, Jr. A technique for the estimation of indices of refraction of marine phytoplankters. Limnology and Oceanography, 17(6): 833–839, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Cubit:1970:SPC

[Cub70] John Cubit. A simple piston corer for sampling sand beaches. Limnology and Oceanography, 15(1):155–156, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Culliney:1970:MRP

[Cul70] John L. Culliney. Measurements of reactive phosphorus associated with pelagic Sargassum in the northwest Sargasso Sea. Limnology and Oceanography, 15(2):304–305, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Culver:1973:BRJ

[Cul73a] David A. Culver. Book review: Jenkins, D., V. L. Snoeyink, J. F. Ferguson, and J. O. Leckie. 1972. AAPSE water chemistry laboratory manual. Amer. Ass. Professors Sanit. Eng. (J. F. Malina, Eng. Lab. Bldg., Univ. Texas, Austin). vi + 9 variously paged experiments. $3.00. Limnology and Oceanography, 18(1):182, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Culver:1973:REF

[Cul73b] David A. Culver. The reliability of electronic field instruments. Limnology and Oceanography, 16(5):812–814, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Culver:1977:BMS

[Cul77] David A. Culver. Biogenic meromixis and stability in a soft-water lake. Limnology and Oceanography, 22(4):667–686, July 1977. CO-DEN LIOCAH. ISSN 0024-3590.

Cummins:1971:BRH

[Cum71] Kenneth W. Cummins. Book review: Hynes, H. B. N. 1970. The ecology of running waters. Univ. Toronto Press, xxiv + 555 p. $25.00. Limnology and Oceanography, 16(3):593–594, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Cunningham:1972:VMD

[Cun72] Linda Cunningham. Vertical migrations of Daphnia and copepods under the ice. Limnology and Oceanography, 17(2):301–303, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Cushing:1976:BRB

[Cus76a] D. H. Cushing. Book review: Bougis, P. 1974. Écologie du plancton marin. 1. Le phytoplancton. 2. Le zooplancton. Masson & Cie, Paris. 1. ix + 196 p. 98F; 2. vi + 200 p. 98F. Limnology and Oceanography, 21(1):183, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Cushing:1976:GLE

[Cus76b] D. H. Cushing. Grazing in Lake Erken. Limnology and Oceanography, 21(3):349–356, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Cushing:1977:RCR

[Cus77] D. H. Cushing. Reply to comment by R. Knoechel. Limnology and Oceanography, 22(5):970, September 1977. CODEN LIOCAH. ISSN 0024-3590. See [Kno77].

Carpenter:1978:NFA

[CVV78] Edward J. Carpenter, Charlene D. Van Raalte, and Ivan Valiela. Nitrogen fixation by algae in a Massachusetts salt marsh. Limnology and Oceanography, 23(2):318–327, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Christian:1978:AMC

[CW78a] R. R. Christian and W. J. Wiebe. Anaerobic microbial community metabolism in Spartina alterniflora soil. Limnology and Oceanography, 23(2):328–336, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Cunningham:1978:FAI

[CW78b] H. Wilson Cunningham and Robert G. Wetzel. Fulvic acid interferences on ATP determinations in sediments. Limnology and Oceanography, 23(1):166–173, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Caldwell:1975:FFP

[CWM75] Douglas R. Caldwell, Steve D. Wilcox, and Mark Matsler. A freely falling probe for small-scale temperature gradients. Limnology and Oceanography, 20(3):1034–1042, November 1975. CODEN LIOCAH. ISSN 0024-3590.

dAnglejan:1975:BRL

[d’A75] Bruno F. d’Anglejan. Book review: Loring, D. H., and D. J. G. Nota. 1973. Morphology and sediments of the Gulf of St. Lawrence. Bull. Fish. Res. Bd. Can. 182. 147 p. + 7 charts. $5.00. Limnology and Oceanography, 19(4):300–301, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Dagg:1977:SEP

[Dag77] Michael Dagg. Some effects of patchy food environments on copepods. Limnology and Oceanography, 22(1):99–107, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Dale:1974:BIS

[Dal74] Norman G. Dale. Bacteria in intertidal sediments: Factors related to their distribution. Limnology and Oceanography, 19(3):509–518, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Davis:1973:RPG

[Dav73] Margaret Bryan Davis. Redeposition of pollen grains in lake sediment. Limnology and Oceanography, 18(1):44–52, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Davis:1974:SET

[Dav74a] Ronald B. Davis. Stratigraphic effects of tubificids in profundal lake sediments. Limnology and Oceanography, 19(3):466–488, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Davis:1974:TAP

[Dav74b] Ronald B. Davis. Tubificids alter profiles of redox potential and pH in profundal lake sediment. Limnology and Oceanography, 19 (2):342–346, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Davison:1977:PMF

[Dav77] W. Davison. The polarographic measurement of O_{2}, Fe^{2+}, Mn^{2+}, and S^{2−} in hypolimnetic water. Limnology and Oceanography, 22 (4):746–753, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Dawson:1978:VDC

[Daw78] John K. Dawson. Vertical distribution of Calanus hyperboreus in the central Arctic Ocean. Limnology and Oceanography, 23(5): 950–957, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Dayton:1979:BRC

[Day79] Paul K. Dayton. Book review: Carefoot, T. 1978. Pacific seashores: a guide to intertidal ecology. Univ. Washington Press, Seattle. 208 p. $12.95. Limnology and Oceanography, 24(4):792, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Deevey:1971:ACQb

[DB71] Georgiana B. Deevey and Albert L. Brooks. The annual cycle in quantity and composition of the zooplankton of the Sargasso Sea off Bermuda. II. The surface to 2,000 m. Limnology and Oceanography, 16(6):927–943, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Davis:1973:DSP

[DB73] Margaret Bryan Davis and Linda B. Brubaker. Differential sedimentation of pollen grains in lakes. Limnology and Oceanography, 18(4):635–646, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Doemel:1976:VDS

[DB76a] William N. Doemel and Thomas D. Brock. Vertical distribution of sulfur species in benthic algal mats. Limnology and Oceanography, 21(2):237–244, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Dubinsky:1976:LUE

[DB76b] Z. Dubinsky and Thomas Berman. Light utilization efficiencies of phytoplankton in Lake Kinneret (Sea of Galilee). Limnology and Oceanography, 21(2):226–230, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Dubinsky:1979:SCS

[DB79] Zvy Dubinsky and Thomas Berman. Seasonal changes in the spectral composition of downwelling irradiance in Lake Kinneret (Israel). Limnology and Oceanography, 24(4):652–663, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Davis:1978:CCM

[DBH78] Curtiss O. Davis, Norman F. Breitner, and Paul J. Harrison. Continuous culture of marine diatoms under silicon limitation. 3. A model of Si-limited diatom growth. Limnology and Oceanography, 23(1):41–52, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Daley:1977:CSM

[DBM77] R. J. Daley, S. R. Brown, and R. N. McNeely. Chromatographic and SCDP measurements of fossil phorbins and the postglacial history of Little Round Lake, Ontario. Limnology and Oceanography, 22(2):349–360, March 1977. CODEN LIOCAH. ISSN 0024-3590.

deBovee:1974:IMS

[dBSA74] F. de Bovée, J. Soyer, and Ph. Albert. The importance of the mesh size for the extraction of the muddy bottom meiofauna. Limnology and Oceanography, 19(2):350–354, March 1974. CODEN LI-OCAH. ISSN 0024-3590.

Dickinson:1975:CTB

[DC75] John J. Dickinson and Andrew G. Carey, Jr. A comparison of two benthic infaunal samplers. Limnology and Oceanography, 20(4): 900–902, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Demanche:1973:AAU

[DCC73] J. Michael Demanche, Herbert Curl, Jr., and D. Douglas Coughenower. An automated analysis for urea in seawater. Limnology and Oceanography, 18(4):686–689, July 1973. CODEN LI-OCAH. ISSN 0024-3590.

Duchart:1973:DTM

[DCP73] Patricia Duchart, S. E. Calvert, and N. B. Price. Distribution of trace metals in the pore waters of shallow water marine sediments. Limnology and Oceanography, 18(4):605–610, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Deevey:1971:ASE

[DD71] Edward S. Deevey, Jr. and Georgiana B. Deevey. The American species of Eubosmina seligo (Crustacea, Cladocera). Limnology and Oceanography, 16(2):201–218, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Durbin:1978:LWR

[DD78] Edward G. Durbin and Ann Gall Durbin. Length and weight relationships of Acartia clausi from Narragansett Bay, R.I. Limnology and Oceanography, 23(5):958–969, September 1978. CODEN LI-OCAH. ISSN 0024-3590.

Donaghay:1978:PPG

[DDS78] Percy L. Donaghay, J. Michael DeManche, and Lawrence F. Small. On predicting phytoplankton growth rates from carbon : nitrogen ratios. Limnology and Oceanography, 23(2):359–362, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Dunson:1971:ETS

[DE71] William A. Dunson and Gary W. Ehlert. Effects of temperature, salinity, and surface water flow on distribution of the sea snake Pelamis. Limnology and Oceanography, 16(6):845–853, November 1971. CODEN LIOCAH. ISSN 0024-3590.

DElia:1977:URD

[D’E77] Christopher F. D’Elia. The uptake and release of dissolved phosphorus by reef corals. Limnology and Oceanography, 22(2):301– 315, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Deevey:1971:ACQa

[Dee71] Georgiana B. Deevey. The annual cycle in quantity and composition of the zooplankton of the Sargasso Sea off Bermuda. I. The upper 500 m. Limnology and Oceanography, 16(2):219–240, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Deevey:1974:BRS

[Dee74] E. S. Deevey. Book review: Schopf, T. J. M. [Ed.]. 1972. Models in paleobiology. Freeman, Cooper & Co., San Francisco. 250 p. $9.50. Limnology and Oceanography, 19(2):375–376, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Degobbis:1973:SSS

[Deg73] Danilo Degobbis. On the storage of seawater samples for ammonia determination. Limnology and Oceanography, 18(1):146–150, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Denman:1977:STV

[Den77] Kenneth L. Denman. Short term variability in vertical chlorophyll structure. Limnology and Oceanography, 22(3):434–441, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Depetris:1976:HPR

[Dep76] Pedro J. Depetris. Hydrochemistry of the Parana River. Limnology and Oceanography, 21(5):736–739, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Devol:1979:ZRR

[Dev79] Allan H. Devol. Zooplankton respiration and its relation to plankton dynamics in two lakes of contrasting trophic state. Limnology and Oceanography, 24(5):893–905, September 1979. CODEN LI-OCAH. ISSN 0024-3590.

Dodson:1976:ZSD

[DEWN76] Stanley I. Dodson, Candace Edwards, Fred Wiman, and James C. Normandin. Zooplankton: Specific distribution and food abundance. Limnology and Oceanography, 21(2):309–313, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Dean:1976:MCM

[DG76a] Walter E. Dean and Eville Gorham. Major chemical and mineral components of profundal surface sediments in Minnesota lakes. Limnology and Oceanography, 21(2):259–284, March 1976. CO-DEN LIOCAH. ISSN 0024-3590.

Duval:1976:DFR

[DG76b] W. S. Duval and G. H. Geen. Diel feeding and respiration rhythms in zooplankton. Limnology and Oceanography, 21(6):823–829, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Douek:1978:UAG

[DG78] Maurice Douek and David A. I. Goring. UV absorption by gerrid cuticles. Limnology and Oceanography, 23(3):554–556, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Davey:1970:RTM

[DGEB70] Earl W. Davey, John H. Gentile, Stanton J. Erickson, and Peter Betzer. Removal of trace metals from marine culture media. Limnology and Oceanography, 15(3):486–488, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Drost-Hasses:1972:BRH

[DH72] W. Drost-Hasses. Book review: Horse, R. A. [Ed.]. 1972. Water and aqueous solutions: Structure, thermodynamics and transport processes. Wiley-Interscience, N.Y. 837 p. $37.50. Limnology and Oceanography, 17(5):799, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Daley:1975:DCA

[DH75] Ralph J. Daley and John E. Hobbie. Direct counts of aquatic bacteria by a modified epifluorescence technique. Limnology and Oceanography, 20(4):875–882, September 1975. CODEN LI-OCAH. ISSN 0024-3590.

deHaan:1977:EBM

[dH77] H. de Haan. Effect of benzoate on microbial decomposition of fulvic acids in Tjeukemeer (the Netherlands). Limnology and Oceanography, 22(1):38–44, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Davis:1978:PMR

[DH78a] Curtiss O. Davis and Paul J. Harrison. On the philosophy of modeling — reply to a comment by M. R. Droop. Limnology and Oceanography, 23(2):385–387, March 1978. CODEN LIOCAH. ISSN 0024-3590. See [Dro78].

Davison:1978:FIS

[DH78b] W. Davison and S. I. Heaney. Ferrous iron-sulfide interactions in anoxic hypolimnetic waters. Limnology and Oceanography, 23(6): 1194–1200, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Dillon:1975:PBC

[Dil75] P. J. Dillon. The phosphorus budget of Cameron Lake, Ontario: The importance of flushing rate to the degree of eutrophy of lakes. Limnology and Oceanography, 20(1):28–39, January 1975. CO-DEN LIOCAH. ISSN 0024-3590.

Disalvo:1973:CSB

[Dis73] Louis H. Disalvo. Contamination of surfaces by bacterial neuston. Limnology and Oceanography, 18(1):165–168, January 1973. CO-DEN LIOCAH. ISSN 0024-3590.

Djuric:1974:DTD

[DL74] Dušan Djurić and Henri R. Leribaux. On the determination of turbulent diffusivity in shallow waters by aerial photography of floating markers. Limnology and Oceanography, 19(1):138–144, January 1974. CODEN LIOCAH. ISSN 0024-3590.

delaCruz:1975:AAS

[dlCP75] Armando A. de la Cruz and William E. Poe. Amino acids in salt marsh detritus. Limnology and Oceanography, 20(1):124–127, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Daumas:1976:ISM

[DLMS76] Raoul A. Daumas, Tierre L. Laborde, Jean Claude Marty, and Alain Saliot. Influence of sampling method on the chemical composition of water surface film. Limnology and Oceanography, 21 (2):319–326, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Dunstan:1971:CCN

[DM71] William M. Dunstan and David W. Menzel. Continuous cultures of natural populations of phytoplankton in dilute, treated sewage effluent. Limnology and Oceanography, 16(4):623–632, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Ducklow:1979:BPA

[DM79a] Hugh W. Ducklow and Ralph Mitchell. Bacterial populations and adaptations in the mucus layers on living corals. Limnology and Oceanography, 24(4):715–725, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Ducklow:1979:CMR

[DM79b] Hugh W. Ducklow and Ralph Mitchell. Composition of mucus released by coral reef coelenterates. Limnology and Oceanography, 24(4):706–714, July 1979. CODEN LIOCAH. ISSN 0024-3590.

deNoyelles:1974:SCS

[dO74] Frank deNoyelles, Jr. and W. John O’Brien. The in situ chemostat — a self-contained continuous culturing and water sampling system. Limnology and Oceanography, 19(2):326–331, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Drinkwater:1975:RTM

[DO75] K. F. Drinkwater and T. R. Osborn. The role of tidal mixing in Rupert and Holberg Inlets, Vancouver Island. Limnology and Oceanography, 20(4):518–529, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Dodson:1970:CFN

[Dod70] Stanley I. Dodson. Complementary feeding niches sustained by size-selective predation. Limnology and Oceanography, 15(1):131– 137, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Dodson:1974:ACP

[Dod74] Stanley I. Dodson. Adaptive change in plankton morphology in response to size-selective predation: a new hypothesis of cyclomorphosis. Limnology and Oceanography, 19(5):721–729, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Dodson:1975:PRZ

[Dod75] Stanley I. Dodson. Predation rates of zooplankton in Arctic ponds. Limnology and Oceanography, 20(6):426–433, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Dodson:1979:BSP

[Dod79] Stanley I. Dodson. Body size patterns in Arctic and temperate zooplankton. Limnology and Oceanography, 24(5):940–949, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Denman:1977:CFS

[DOP77] Kenneth Denman, Akira Okubo, and Trevor Platt. The chlorophyll fluctuation spectrum in the sea. Limnology and Oceanography, 22(6):1033–1038, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Doyle:1975:UCS

[Doy75] Roger W. Doyle. Upwelling, clone selection, and the characteristic shape of nutrient uptake curves. Limnology and Oceanography, 20 (6):487–489, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Doyle:1979:IRS

[Doy79] Roger W. Doyle. Ingestion rate of a selective deposit feeder in a complex mixture of particles: Testing the energy-optimization hypothesis. Limnology and Oceanography, 24(5):867–874, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Devol:1978:SCR

[DP78] Allan H. Devol and T. T. Packard. Seasonal changes in respiratory enzyme activity and productivity in Lake Washington microplankton. Limnology and Oceanography, 23(1):104–111, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Dillon:1974:PCR

[DR74] P. J. Dillon and F. H. Rigler. The phosphorus-chlorophyll relationship in lakes. Limnology and Oceanography, 19(5):767–773, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Droop:1978:CDB

[Dro78] M. R. Droop. Comments on the Davis/Breitner/Harrison model for silicon uptake and utilization by diatoms. Limnology and Oceanography, 23(2):383–385, March 1978. CODEN LIOCAH. ISSN 0024-3590. See reply [DH78a].

Davey:1975:ASC

[DS75] Earl W. Davey and Albert E. Soper. Apparatus for the in situ concentration of trace metals from seawater. Limnology and Oceanography, 20(3):1019–1023, November 1975. CODEN LIOCAH. ISSN 0024-3590.

DElia:1977:DTN

[DSC77] Christopher F. D’Elia, Paul A. Steudler, and Nathaniel Corwin. Determination of total nitrogen in aqueous samples using persulfate digestion. Limnology and Oceanography, 22(4):760–764, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1971:BRS

[Dug71] R. C. Dugdale. Book review: Stevenson, M. R., O. Guillén G., and J. Santoro de Ycaza. 1970. Marine atlas of the Pacific coastal waters of South America. Univ. Calif. Press, Berkeley. 20 p. + 99 charts. $40.00. Limnology and Oceanography, 16(1):152–153, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Dunn:1970:RTP

[Dun70] I. G. Dunn. Recovery of a tropical pond zooplankton community after destruction by algal bloom. Limnology and Oceanography, 15(3):373–379, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Duxbury:1975:ODO

[Dux75] Alyn C. Duxbury. Orthophosphate and dissolved oxygen in Puget Sound. Limnology and Oceanography, 19(4):270–274, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Duxbury:1979:UEF

[Dux79] Alyn C. Duxbury. Upwelling and estuary flushing. Limnology and Oceanography, 24(4):627–633, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Davis:1978:STH

[DV78] Noël Davis and Glenn R. VanBlaricom. Spatial and temporal heterogeneity in a sand bottom epifaunal community of invertebrates in shallow water. Limnology and Oceanography, 23(3):417–427, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Dragcevic:1979:PSA

[DVCP79] Dj. Dragcevic, M. Vukovic, D. Cukman, and V. Pravdic. Properties of the seawater–air interface. Dynamic surface tension studies. Limnology and Oceanography, 24(6):1022–1030, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Donn:1972:SWL

[DW72] William L. Donn and David M. Wolf. Seiche and water level fluctuations in Grindavik Harbor, Iceland. Limnology and Oceanography, 17(4):639–643, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1970:RDX

[EA70] W. T. Edmondson and David E. Allison. Recording densitometry of X-radiographs for the study of cryptic laminations in the sediment of Lake Washington. Limnology and Oceanography, 15(1): 138–144, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Elzerman:1979:EZC

[EA79] Alan W. Elzerman and David E. Armstrong. Enrichment of Zn, Cd, Pb, and Cu in the surface microlayer of Lakes Michigan, Ontario, and Mendota. Limnology and Oceanography, 24(1):133–144, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Eager:1971:PDM

[Eag71] Edward W. Eager. Pattern in the development of a marine community. Limnology and Oceanography, 16(2):241–253, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Estrada:1979:TPP

[EB79] Marta Estrada and Dolors Blasco. Two phases of the phytoplankton community in the Baja California upwelling. Limnology and Oceanography, 24(6):1065–1080, November 1979. CODEN LI-OCAH. ISSN 0024-3590.

Eberhart:1972:BRH

[Ebe72] Richard Eberhart. Book review: Heezen, B. C., and C. D. Hollister. 1971. The face of the deep. Oxford Univ. Press, New York, London, and Toronto. xi + 659 p. $25.00. Limnology and Oceanography, 17(5):801–802, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Esaias:1972:EDB

[EC72] Wayne E. Esaias and Herbert C. Curl, Jr. Effect of dinoflagellate bioluminescence on copepod ingestion rates. Limnology and Oceanography, 17(6):901–906, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Eccles:1974:OPL

[Ecc74] David H. Eccles. An outline of the physical limnology of Lake Malawi (Lake Nyasa). Limnology and Oceanography, 19(5):730– 742, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Eppley:1971:PGC

[ECHH^{+}71] R. W. Eppley, A. F. Carlucci, O. Holm-Hansen, D. Kiefer, J. J. McCarthy, Elizabeth Venrick, and P. M. Williams. Phytoplankton growth and composition in shipboard cultures supplied with nitrate, ammonium, or urea as the nitrogen source. Limnology and Oceanography, 18(5):741–751, September 1971. CODEN LI-OCAH. ISSN 0024-3590.

Eckstein:1970:PLG

[Eck70] Yoram Eckstein. Physicochemical limnology and geology of a meromictic pond on the Red Sea shore. Limnology and Oceanography, 15(3):363–372, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1970:BRV

[Edm70a] W. T. Edmondson. Book reviews: Vollenweider, R. A. 1968. Water management research. Scientific fundamentals of the eutrophication of lakes and flowing waters with particular reference to nitrogen and phosphorus as factors in eutrophication. Organization for Economic Co-operation and Development. Directorate for Scientific Affairs. Paris. Mimeographed. 159 p. + 34 Figs. + 2 separately paged annexes: Bibliography, 61 p; Current status of research on eutrophication in Europe, the United States and Canada, 20 p. (French and German versions available). Limnology and Oceanography, 15(1):169–170, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1970:JDH

[Edm70b] Yvette H. Edmondson. J. D. H. Strickland — 1920–1970. Limnology and Oceanography, 15(6):971–972, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1971:BRT

[Edm71a] W. T. Edmondson. Book review: Todd, D. K. 1970. The water encyclopedia. Water Information Center, Port Washington, New York, x + 559 p. $27.50. Limnology and Oceanography, 16(3):597, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1971:SCH

[Edm71b] Yvette H. Edmondson. Some components of the Hutchinson legend. Limnology and Oceanography, 16(2):157–163, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1972:BRB

[Edm72a] W. T. Edmondson. Book review: Brinkhurst, R. O., and B. G. M. Jamieson. 1971. Aquatic Oligochaeta of the world. Oliver and Boyd, Edinburgh. xi + 860 p. £12. 1972. Univ. Toronto. $35.00. Limnology and Oceanography, 17(1):166, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1972:BRS

[Edm72b] W. T. Edmondson. Book review: Sarnoff, P. 1972. The New York Times encyclopedic dictionary of the environment. Quadrangle Books, Inc., New York. 352 p. $10.00. Limnology and Oceanography, 17(3):502, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1972:IBR

[Edm72c] W. T. Edmondson. Instantaneous birth rates of zooplankton. Limnology and Oceanography, 17(5):792–795, September 1972. CO-DEN LIOCAH. ISSN 0024-3590.

Edmondson:1972:ECR

[Edm72d] Yvette H. Edmondson. Editorial comment: referees. Limnology and Oceanography, 17(6):963–964, November 1972. CODEN LI-OCAH. ISSN 0024-3590.

Edmondson:1974:BRG

[Edm74a] W. T. Edmondson. Book review: Griffith, E. J., A. Beeton, J. M. Spencer, and D. T. Mitchell [Eds.] 1973. Environmental phosphorus handbook. Wiley-Inter-science. New York, London, Sydney, Toronto, xvi + 718 p. $34.95. Limnology and Oceanography, 19 (2):369–375, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1974:ECP

[Edm74b] Yvette H. Edmondson. Editorial comment: Policy. Limnology and Oceanography, 19(1):167–168, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1976:BRB

[Edm76a] W. T. Edmondson. Book review: Bayly, I. A. E., and W. D. Williams. 1973. Inland waters and their ecology. Longman, Australia. 316 p. Limnology and Oceanography, 21(3):483–484, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1976:BRR

[Edm76b] W. T. Edmondson. Book review: Ruttner-Kolisko, A. 1974. Plankton rotifers. Biology and taxonomy. English translation of Die Binnengewasser v. 26, part 1. 146 p. DM46.80. Limnology and Oceanography, 21(1):183–184, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1977:ECD

[Edm77] Yvette H. Edmondson. Editorial comment: Dual publication. Limnology and Oceanography, 22(4):773, July 1977. CODEN LI-OCAH. ISSN 0024-3590.

Edwards:1977:BRC

[Edw77] John S. Edwards. Book review: Cheng, L. [Ed.] 1976. Marine insects. North-Holland Publ. Co., Amsterdam and New York, xii + 581 p. $62.50; Dfl. 162. Limnology and Oceanography, 22(6): 1104–1105, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Eaton:1979:SAG

[EG79] Andrew D. Eaton and Virginia Grant. Sorption of ammonium by glass frits and filters: Implications for analyses of brackish and freshwater. Limnology and Oceanography, 24(2):397–399, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Eggers:1978:LFB

[Egg78] Douglas M. Eggers. Limnetic feeding behavior of juvenile sockeye salmon in Lake Washington and predator avoidance. Limnology and Oceanography, 23(6):1114–1125, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Edgington:1970:CRT

[EGTA70] David N. Edgington, Solon A. Gordon, Michael M. Thommes, and Luis R. Almodovar. The concentration of radium, thorium, and uranium by tropical marine algae. Limnology and Oceanography, 15(6):945–955, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Enright:1977:DVMb

[EH77] J. T. Enright and H.-W. Honegger. Diurnal vertical migration: Adaptive significance and timing. Part 2. Test of the model: Details of timing. Limnology and Oceanography, 22(5):873–886, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Ellis:1973:EMM

[EK73] J. Ellis and S. Kanamori. An evaluation of the Miller method for dissolved oxygen analysis. Limnology and Oceanography, 18(6): 1002–1005, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Elder:1975:CSR

[Eld75] John F. Elder. Complexation side reactions involving trace metals in natural water systems. Limnology and Oceanography, 20(1):96– 102, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Emery:1974:BRS

[Eme74] K. O. Emery. Book review: Stanley, D. J. [Ed.]. 1972. The Mediterranean Sea: a natural sedimentation laboratory. Dowden, Hutchinson & Ross, Inc., Stroudsburg, Pa. xvi + 765 p. $45.00. Limnology and Oceanography, 19(3):547, May 1974. CODEN LI-OCAH. ISSN 0024-3590.

Emerson:1975:CEC

[Eme75a] Steve Emerson. Chemically enhanced CO_{2} gas exchange in a eutrophic lake: a general model. Limnology and Oceanography, 20 (4):743–753, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Emerson:1975:GER

[Eme75b] Steve Emerson. Gas exchange rates in small Canadian Shield lakes. Limnology and Oceanography, 20(4):754–761, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Emery:1976:BRK

[Eme76] K. O. Emery. Book review: King, C. A. M. 1975. Introduction to marine geology and geomorphology. Crane, Russak & Co., Inc., New York, ix + 309 p. $24.00 (paperback, $12.00). Limnology and Oceanography, 21(1):186, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Emery:1978:BRS

[Eme78] K. O. Emery. Book review: Shepard, F. P. 1977. Geological oceanography: Evolution of coasts, continental margins, and the deep-sea floor. Crane, Russak & Co., Inc., New York, vi + 214 p. $10.50. Limnology and Oceanography, 23(5):1085, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Enright:1976:BRS

[Enr76] J. T. Enright. Book review: Schmidt-Koenig, K. 1975. Migration and homing in animals. Zoophysiology and Ecology, v. 6. Springer-Verlag, New York, xii + 99 p. $18.90. Limnology and Oceanography, 21(6):931–932, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Enright:1977:CHS

[Enr77a] J. T. Enright. Copepods in a hurry: Sustained high-speed upward migration. Limnology and Oceanography, 22(1):118–125, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Enright:1977:DVMa

[Enr77b] J. T. Enright. Diurnal vertical migration: Adaptive significance and timing. Part 1. Selective advantage: a metabolic model. Limnology and Oceanography, 22(5):856–872, September 1977. CO-DEN LIOCAH. ISSN 0024-3590.

Enright:1977:PEC

[Enr77c] J. T. Enright. Problems in estimating copepod velocity. Limnology and Oceanography, 22(1):160–162, January 1977. CODEN LIOCAH. ISSN 0024-3590. See reply [Vly77].

Enright:1979:WW

[Enr79] J. T. Enright. The why and when of up and down. Limnology and Oceanography, 24(4):788–791, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Egloff:1971:SRF

[EP71] David A. Egloff and David S. Palmer. Size relations of the filtering area of two Daphnia species. Limnology and Oceanography, 16(6): 900–905, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Eppley:1973:BRT

[Epp73] Richard W. Eppley. Book review: Takenouti, A. Y. [Chief ed.]. 1972. Biological oceanography of the northern North Pacific Ocean. (Motoda commemorative volume.) Idemitsu Shoten, Tokyo. xiii + 626 p. $55.00. Limnology and Oceanography, 16(5):818, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Eppley:1979:NPP

[ERH79] R. W. Eppley, E. H. Renger, and W. G. Harrison. Nitrate and phytoplankton production in southern California coastal waters.

Limnology and Oceanography, 24(3):483–494, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Eppley:1979:ADS

[ERHC79] Richard W. Eppley, E. H. Render, W. G. Harrison, and J. J. Cullen. Ammonium distribution in southern California coastal waters and its role in the growth of phytoplankton. Limnology and Oceanography, 24(3):495–509, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Ertl:1971:QMS

[Ert71] Milan Ertl. A quantitative method of sampling periphyton from rough substrates. Limnology and Oceanography, 16(3):576–577, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Eppley:1973:SPD

[ERVM73] Richard W. Eppley, Edward H. Renger, Elizabeth L. Venrick, and Michael M. Mullin. A study of plankton dynamics and nutrient cycling in the central gyre of the north Pacific Ocean. Limnology and Oceanography, 18(4):534–551, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Eppley:1975:PMC

[ES75a] Richard W. Eppley and Jonathan H. Sharp. Photosynthetic measurements in the central North Pacific: The dark loss of carbon in 24-h incubations. Limnology and Oceanography, 20(4):981–988, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Erkenbrecher:1975:ITF

[ES75b] Carl W. Erkenbrecher and L. Harold Stevenson. The influence of tidal flux on microbial biomass in salt marsh creeks. Limnology and Oceanography, 20(6):618–625, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Evans:1977:EBM

[ES77] Marlene S. Evans and John A. Stewart. Epibenthic and benthic microcrustaceans (copepods, cladocerans, ostracods) from a nearshore area in southeastern Lake Michigan. Limnology and Oceanography, 22(6):1059–1066, November 1977. CODEN LI-OCAH. ISSN 0024-3590.

Ellis:1973:ESW

[ESH73] G. E. Ellis, J. A. Shooter, and C. W. Horton, Sr. The effects of several weather disturbances on the seiches in Lake Travis, Texas.

Limnology and Oceanography, 18(3):476–482, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Fager:1972:BRH

[Fag72] E. W. Fager. Book review: Holme, N. A., and A. D. McIntyre [Eds.]. 1971. Methods of the study of marine benthos. IBP Handbook No, 16. Blackwell Sci. Publ., Oxford and Edinburgh. F. A. Davis, Philadelphia. 346 p. $13.50. Limnology and Oceanography, 17(6):968–969, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Fager:1973:EMC

[Fag73] E. W. Fager. Estimation of mortality coefficients from field samples of zooplankton. Limnology and Oceanography, 18(2):297–300, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Falkowski:1975:NUM

[Fal75] Paul G. Falkowski. Nitrate uptake in marine phytoplankton: Comparison of half-saturation constants from seven species. Limnology and Oceanography, 20(4):412–417, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Fanning:1977:BRG

[Fan77] Kent A. Fanning. Book review: Goldberg, E. D. [Ed.] 1975. The nature of seawater. Physical and Chemical Sci. Res. Rep.

1. Koehn-Schneider, Old Greenwich, Conn. 719 p. $48.00. Limnology and Oceanography, 22(5):972, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Farmer:1975:PTD

[Far75] D. M. Farmer. Potential temperatures in deep freshwater lakes. Limnology and Oceanography, 20(6):634–635, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Fuhs:1970:PAU

[FC70] G. Wolfgang Fuhs and Edmondo Canelli. Phosphorus-33 autoradiography used to measure phosphate uptake by individual algae. Limnology and Oceanography, 15(6):962–967, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Friebertshauser:1972:WBS

[FD72] Mark A. Friebertshauser and Alyn C. Duxbury. A water budget study of Puget Sound and its subregions. Limnology and Oceanography, 17(2):237–247, March 1972. CODEN LIOCAH. ISSN 0024-

Fedorenko:1975:FCP

[Fed75a] Alice Y. Fedorenko. Feeding characteristics and predation impact of Chaoborus (Diptera, Chaoboridae) larvae in a small lake. Limnology and Oceanography, 19(4):250–258, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Fedorenko:1975:ISS

[Fed75b] Alice Y. Fedorenko. Instar and species-specific diets in two species of chaoborus. Limnology and Oceanography, 19(4):238–249, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Fee:1976:VSD

[Fee76] Everett J. Fee. The vertical and seasonal distribution of chlorophyll in lakes of the Experimental Lakes Area, northwestern Ontario: Implications for primary production estimates. Limnology and Oceanography, 21(6):767–783, November 1976. CODEN LI-OCAH. ISSN 0024-3590.

Fee:1979:RBL

[Fee79] Everett J. Fee. A relation between lake morphometry and primary productivity and its use in interpreting whole-lake eutrophication experiments. Limnology and Oceanography, 24(3):401–416, May 1979. CODEN LIOCAH. ISSN 0024-3590. See comment [Pat80].

Fenchel:1970:SDO

[Fen70] Tom Fenchel. Studies on the decomposition of organic detritus derived from the turtle grass Thalassia testudinum. Limnology and Oceanography, 15(1):14–20, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Fernald:1970:BRN

[Fer70] Robert L. Fernald. Book review: National estuary study. 1970. Bureau of Sport Fisheries and Wildlife, Washington, D.C. v. 1, $1.75; v. 2, $3.00; v. 3, $4.75; v. 4, $2.25; v. 5, $2.25; v. 6, $2.00; v. 7, $1.50. Limnology and Oceanography, 15(6):974, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Furnass:1975:ISD

[FF75] T. I. Furnass and W. C. Findley. An improved sorting device for zooplankton. Limnology and Oceanography, 19(4):295–297, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Feuillade:1979:CDA

[FF79] Jacques Feuillade and Mauricette Feuillade. A chemostat device adapted to planktonic Oscillatoria cultivation. Limnology and Oceanography, 24(3):562–564, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Flint:1975:EBG

[FG75] R. Warren Flint and Charles R. Goldman. The effects of a benthic grazer on the primary productivity of the littoral zone of Lake Tahoe. Limnology and Oceanography, 20(6):935–944, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Francko:1979:FDC

[FH79] David A. Francko and Robert T. Heath. Functionally distinct classes of complex phosphorus compounds in lake water. Limnology and Oceanography, 24(3):463–473, May 1979. CODEN LI-OCAH. ISSN 0024-3590.

Fisher:1979:CCA

[FHV79] Thomas R. Fisher, Jr., Evelyn B. Haines, and Richard J. Volk. A comment on the calculation of atom percent enrichment for stable isotopes. Limnology and Oceanography, 24(3):593–595, May 1979. CODEN LIOCAH. ISSN 0024-3590. See reply [SCA79].

Field:1972:SOR

[Fie72] J. G. Field. Some observations on the release of dissolved organic carbon by the sea urchin, Strongylocentrotus droebachiensis. Limnology and Oceanography, 17(5):759–761, September 1972. CO-DEN LIOCAH. ISSN 0024-3590.

Fitzgerald:1970:ALM

[Fit70] George P. Fitzgerald. Aerobic lake muds for the removal of phosphorus from lake waters. Limnology and Oceanography, 15(4): 550–555, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Fitzgerald:1975:MCO

[FL75] William F. Fitzgerald and W. Berry Lyons. Mercury concentrations in open-ocean waters: Sampling procedure. Limnology and Oceanography, 20(6):468–471, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Foulds:1978:CDM

[FM78] J. B. Foulds and K. H. Mann. Cellulose digestion in Mysis stenolepis and its ecological implications. Limnology and Oceanography, 23(4):760–766, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Folger:1972:SWS

[FMJC72] David W. Folger, Robert H. Meade, Blair F. Jones, and Robert L. Cory. Sediments and waters of Somes Sound, a fjordlike estuary in Maine. Limnology and Oceanography, 17(3):394–402, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Fliegel:1970:TBC

[FN70] Myron Fliegel and Ali A. Nowrooxi. Tides and bottom currents off the coast of northern California. Limnology and Oceanography, 15(4):615–624, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Fogg:1977:EOM

[Fog77] G. E. Fogg. Excretion of organic matter by phytoplankton. Limnology and Oceanography, 22(3):576–577, May 1977. CODEN LI-OCAH. ISSN 0024-3590.

Folger:1971:NTS

[Fol71] D. W. Folger. Nearshore tracking of seabed drifters. Limnology and Oceanography, 16(3):588–589, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Fott:1970:BRK

[Fot70] Bohuslav Fott. Book review: Kol, E. 1968. Kryobiologie. Biologie und Limnologie des Schnees und Eises I. Kryovegetation. Die Binnengewasser, v. 24. E. Schweizerbart’sche (Nägele & Obermiller), Stuttgart. viii + 216 p., 16 Plates. DM 66. Limnology and Oceanography, 15(4):660–661, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Fotheringham:1974:TCL

[Fot74] Nick Fotheringham. Trophic complexity in a littoral boulderfield. Limnology and Oceanography, 19(1):84–91, January 1974. CO-DEN LIOCAH. ISSN 0024-3590.

Fournier:1970:SPM

[Fou70] Robert O. Fournier. Studies on pigmented microorganisms from aphotic marine environments. Limnology and Oceanography, 15 (5):675–682, September 1970. CODEN LIOCAH. ISSN 0024-

Fournier:1971:SPM

[Fou71] Robert O. Fournier. Studies on pigmented microorganisms from aphotic marine environments. II. North Atlantic distribution. Limnology and Oceanography, 16(6):952–961, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Fournier:1973:SPM

[Fou73] Robert O. Fournier. Studies on pigmented microorganisms from aphotic marine environments. III. Evidence of apparent utilization by benthic and pelagic tunicata. Limnology and Oceanography, 18 (1):38–43, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Ferrante:1977:TDF

[FP77] J. G. Ferrante and J. I. Parker. Transport of diatom frustules by copepod fecal pellets to the sediments of Lake Michigan. Limnology and Oceanography, 22(1):92–98, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Ferguson:1979:DSB

[FP79a] Randolph L. Ferguson and Anthony V. Palumbo. Distribution of suspended bacteria in neritic waters south of Long Island during stratified conditions. Limnology and Oceanography, 24(4):697– 705, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Fisher:1979:ALW

[FP79b] Thomas R. Fisher, Jr. and Patricia E. Parsley. Amazon lakes: Water storage and nutrient stripping by algae. Limnology and Oceanography, 24(3):547–553, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Fisher:1979:ODP

[FPO79] John S. Fisher, James Pickral, and William E. Odum. Organic detritus particles: Initiation of motion criteria. Limnology and Oceanography, 24(3):529–532, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Ferguson:1976:CBS

[FR76] R. L. Ferguson and P. Rublee. Contribution of bacteria to standing crop of coastal plankton. Limnology and Oceanography, 21(1):141– 145, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Feigenbaum:1977:PDC

[FR77a] David Feigenbaum and M. R. Reeve. Prey detection in the Chaetognatha: Response to a vibrating probe and experimental determination of attack distance in large aquaria. Limnology and Oceanography, 22(6):1052–1058, November 1977. CODEN LI-OCAH. ISSN 0024-3590.

Fry:1977:CAE

[FR77b] John C. Fry and Angela J. Ramsay. Changes in the activity of epiphytic bacteria of Elodea canadensis and Chara vulgaris following treatment with the herbicide, paraquat. Limnology and Oceanography, 22(3):556–562, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Frank:1972:BRL

[Fra72] Peter W. Frank. Book review: Luther, H., and J. Rzóska [compilers]. 1971. Project aqua. A source book of inland waters proposed for conservation. IBP Handbook No. 21. F. A. Davis, Philadelphia (Blackwell Sci. Publ., Oxford and Edinburgh). xiv + 239 p. $7.00. Limnology and Oceanography, 17(5):803–804, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Frey:1971:WDE

[Fre71] David G. Frey. Worldwide distribution and ecology of eurycercus and saycia (cladocera). Limnology and Oceanography, 16(2):254– 308, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Frost:1972:ESC

[Fro72] B. W. Frost. Effects of size and concentration of food particles on the feeding behavior of the marine planktonic copepod Calanus pacificus. Limnology and Oceanography, 17(6):805–815, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Frost:1975:TFB

[Fro75] B. W. Frost. A threshold feeding behavior in Calanus pacificus. Limnology and Oceanography, 19(4):263–266, March 1975. CO-DEN LIOCAH. ISSN 0024-3590.

Frost:1977:FBC

[Fro77] Bruce W. Frost. Feeding behavior of Calanus pacificus in mixtures of food particles. Limnology and Oceanography, 22(3):472–491, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Frost:1978:SMC

[Fro78] Thomas M. Frost. In situ measurements of clearance rates for the freshwater sponge Spongilla lacustris. Limnology and Oceanography, 23(5):1034–1039, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Fry:1971:BRS

[Fry71] F. E. J. Fry. Book review: Spotte, S. H. 1970. Fish and invertebrate culture. Water management in closed systems. Wiley-Interscience, New York, xiv + 145 p. $8.95. Limnology and Oceanography, 16(6):1005, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Fedorenko:1972:CBC

[FS72a] Alice Y. Fedorenko and Michael C. Swift. Comparative biology of Chaoborus americanus and Chaoborus trivittatus in Eunice Lake, British Columbia. Limnology and Oceanography, 17(5):721–730, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Fowler:1972:SRE

[FS72b] Scott W. Fowler and Lawrence F. Small. Sinking rates of euphausiid fecal pellets. Limnology and Oceanography, 17(2):293– 296, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Fung:1973:ONS

[FT73] Y. C. Fung and L. B. Trott. The occurrence of a Noctiluca scintillans (Macartney) induced red tide in Hong Kong. Limnology and Oceanography, 18(3):472–476, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Fuhs:1973:IDC

[Fuh73] G. Wolfgang Fuhs. Improved device for the collection of sedimenting matter. Limnology and Oceanography, 18(6):989–993, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Filatova:1974:BRB

[FW74] Z. A. Filatova and B. Ya. Wilenkin. Book review: Brauer, R. W. [Ed.]. 1972. Barobiology and the experimental biology of the deep sea. Proceedings of a symposium. North Carolina Sea Grant Program, University of North Carolina, Chapel Hill, xiv + 428 p. $10. Limnology and Oceanography, 19(3):547–549, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Gunnison:1975:RSA

[GA75] Douglas Gunnison and Martin Alexander. Resistance and susceptibility of algae to decomposition by natural microbial communities. Limnology and Oceanography, 20(1):64–70, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Gagosian:1976:DVP

[Gag76] Robert B. Gagosian. A detailed vertical profile of sterols in the Sargasso Sea. Limnology and Oceanography, 21(5):702–710, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Gallagher:1975:SSF

[Gal75] John L. Gallagher. The significance of the surface film in salt marsh plankton metabolism. Limnology and Oceanography, 20 (1):120–123, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Garrett:1974:CLC

[Gar74] William D. Garrett. Comments on “Laboratory comparisons of four surface microlayer samplers” (R. F. Hatcher and B. C. Parker). Limnology and Oceanography, 19(1):166–167, January 1974. CODEN LIOCAH. ISSN 0024-3590. See [HP74].

Gardner:1975:RIM

[Gar75] Leonard Robert Gardner. Runoff from an intertidal marsh during tidal exposure — recession curves and chemical characteristics. Limnology and Oceanography, 20(1):81–89, January 1975. CO-DEN LIOCAH. ISSN 0024-3590.

Gardner:1977:IER

[Gar77] Wilford D. Gardner. Incomplete extraction of rapidly settling particles from water samplers. Limnology and Oceanography, 22 (4):764–768, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Gardner:1978:MMM

[Gar78] Wayne S. Gardner. Microfluorometric method to measure ammonium in natural waters. Limnology and Oceanography, 23(5): 1069–1072, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Gordon:1972:TML

[GB72a] Howard R. Gordon and Otis B. Brown. A theoretical model of light scattering by Sargasso Sea particulates. Limnology and Oceanography, 17(6):826–832, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Grenney:1972:FSM

[GB72b] William J. Grenney and David A. Bella. Field study and mathematical model of the slack-water buildup of a pollutant in a tidal river. Limnology and Oceanography, 17(2):229–236, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Gross:1972:RDN

[GB72c] M. Grant Gross and D. F. Bumpus. Residual drift of near-bottom waters in Long Island Sound, 1969. Limnology and Oceanography, 17(4):636–638, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Goulden:1974:APA

[GB74] P. D. Goulden and Peter Brooksbank. Automated phosphate analysis in the presence of arsenic. Limnology and Oceanography, 20 (2):705–707, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Gillespie:1979:MCC

[GB79] D. M. Gillespie and A. C. Benke. Methods of calculating cohort production from field data-some relationships. Limnology and Oceanography, 24(1):171–176, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Goodyear:1972:RBP

[GBB72] C. Phillip Goodyear, Claude E. Boyd, and Robert J. Beyers. Relationships between primary productivity and mosquitofish (Gambusia affinis) production in large microcosms. Limnology and Oceanography, 17(3):445–450, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Graham:1976:MNB

[GBK76] William F. Graham, Michael L. Bender, and Gary P. Klinkhammer. Manganese in Narragansett Bay. Limnology and Oceanography, 21(5):665–673, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Goering:1970:NDS

[GC70] John J. Goering and Joel D. Cline. A note on denitrification in seawater. Limnology and Oceanography, 15(2):306–309, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Geiling:1972:ETD

[GC72] W. T. Geiling and Robert S. Campbell. The effect of temperature on the development rate of the major life stages of Diaptomus pallidus Herrick. Limnology and Oceanography, 17(2):304–306, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Grancini:1973:DPF

[GC73] Gianfranco Grancini and Bruno Cescon. Dispersal processes of freshwaters in the Po River coastal area. Limnology and Oceanography, 18(5):705–710, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Goldman:1974:KAE

[GC74a] Joel C. Goldman and Edward J. Carpenter. A kinetic approach to the effect of temperature on algal growth. Limnology and Oceanography, 19(5):756–766, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Gonye:1974:PIB

[GC74b] Edward R. Gonye and Edward J. Carpenter. Production of iron-binding compounds by marine microorganisms. Limnology and Oceanography, 19(5):840–842, September 1974. CODEN LI-OCAH. ISSN 0024-3590.

Graham:1976:NSM

[GCB76] Linda B. Graham, Arthur D. Colburn, and John C. Burke. A new, simple method for gently collecting planktonic protozoa. Limnology and Oceanography, 21(2):336–341, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Gordon:1970:IEF

[GCL70] C. M. Gordon, R. A. Carr, and R. E. Larson. The influence of environmental factors on the sodium and manganese content of barnacle shells. Limnology and Oceanography, 15(3):461–466, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Galloway:1979:APM

[GCL79] James N. Galloway, Bernard J. Cosby, Jr., and Gene E. Likens. Acid precipitation: Measurement of pH and acidity. Limnology and Oceanography, 24(6):1161–1165, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Gallagher:1974:PPE

[GD74] John L. Gallagher and Franklin C. Daiber. Primary production of edaphic algal communities in a Delaware salt marsh. Limnology and Oceanography, 19(3):390–395, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Gagosian:1979:VIC

[GDHZ79] Robert B. Gagosian, Jerome P. Dean, Jr., Robert Hamblin, and Oliver C. Zafiriou. A versatile, interchangeable chamber seawater sampler. Limnology and Oceanography, 24(3):583–588, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Griffiths:1975:BOS

[GE75] Mary Griffiths and W. T. Edmondson. Burial of oscillaxanthin in the sediment of Lake Washington. Limnology and Oceanography, 20(6):945–952, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Gehrs:1974:HDA

[Geh74] Carl W. Gehrs. Horizontal distribution and abundance of Diaptomus clavipes schacht in relation to Potamogeton foliosus in a pond and under experimental conditions. Limnology and Oceanography, 19(1):100–104, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Gannon:1970:SSC

[GF70a] John E. Gannon and Everett J. Fee. Surface seiches and currents in Douglas Lake, Michigan. Limnology and Oceanography, 15(2): 281–288, March 1970. CODEN LIOCAH. ISSN 0024-3590.

George:1970:DMT

[GF70b] M. G. George and C. M. Fernando. Diurnal migration in three species of rotifers in Sunfish Lake, Ontario. Limnology and Oceanography, 15(2):218–223, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Gavis:1975:KCD

[GF75] Jerome Gavis and John F. Ferguson. Kinetics of carbon dioxide uptake by phytoplankton at high pH. Limnology and Oceanography, 19(4):211–221, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Gravitz:1975:PSR

[GG75a] Norman Gravitz and Linda Gleye. A photochemical side reaction that interferes with the phenolhypochlorite assay for ammonia. Limnology and Oceanography, 20(3):1015–1017, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Griffin:1975:FEC

[GG75b] John J. Griffin and Edward D. Goldberg. The fluxes of elemental carbon in coastal marine sediments. Limnology and Oceanography, 20(6):456–463, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Gat:1978:IHI

[GG78] J. R. Gat and J. R. Gat. Isotope hydrology of inland sabkhas in the Bardawil area, Sinai. Limnology and Oceanography, 23(5): 841–850, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Gough:1981:CPP

[GG81] Stephen B. Gough and Larry P. Gough. Comment on “Primary production of algae growing on natural and artificial aquatic plants: a study of interactions between epiphytes and their substrate” (Cattaneo and Kalff). Limnology and Oceanography, 27(4): 987–988, September 1981. CODEN LIOCAH. ISSN 0024-3590. See [CK79].

Gnaiger:1978:PFI

[GGW78] Erich Gnaiger, Gerhard Gluth, and Wolfgang Wieser. pH fluctuation in an intertidal beach in Bermuda. Limnology and Oceanography, 23(5):851–857, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Gerchakov:1972:ITA

[GH72] Sol M. Gerchakov and Patrick G. Hatcher. Improved technique for analysis of carbohydrates in sediments. Limnology and Oceanography, 17(6):938–943, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Gasith:1976:ALS

[GH76] A. Gasith and A. D. Hosier. Airborne litterfall as a source of organic matter in lakes. Limnology and Oceanography, 21(2):253– 258, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Green:1979:ROS

[GH79] T. Green and D. F. Houk. The removal of organic surface films by rain. Limnology and Oceanography, 24(5):966–970, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Garside:1978:DSC

[GHM78] Christopher Garside, Gretchen Hull, and Susan Murray. Determination of submicromolar concentrations of ammonia in natural waters by a standard addition method using a gas-sensing electrode. Limnology and Oceanography, 23(5):1073–1076, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Giddings:1977:CCP

[Gid77] Jeffrey M. Giddings. Chemical composition and productivity of Scenedesmus abundans in nitrogen limited chemostat cultures. Limnology and Oceanography, 22(5):911–918, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Gieskes:1970:ETP

[Gie70] Joris M. Gieskes. Effect of temperature on the pH of seawater. Limnology and Oceanography, 15(2):329, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Gallegos:1977:MRB

[GIK77] Charles L. Gallegos, George M. Ilornberger, and Mahlon G. Kelly. A model of river benthic algal photosynthesis in response to rapid changes in light. Limnology and Oceanography, 22(2):226–233, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:1971:SNC

[Gil71] John J. Gilbert. Some notes on the control of sexuality in the rotifer asplanchna sieboldi. Limnology and Oceanography, 16(2): 309–319, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Gosselink:1974:DSM

[GK74] James G. Gosselink and Conrad J. Kirby. Decomposition of salt marsh grass, Spartina alterniflora Loisel. Limnology and Oceanography, 19(5):825–832, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Graetz:1973:ESL

[GKA73] Donald A. Graetz, Dennis R. Keeney, and Ruben B. Aspiras. Eh status of lake sediment–water systems in relation to nitrogen transformations. Limnology and Oceanography, 18(6):908– 917, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Granhall:1971:NFL

[GL71] Ulf Granhall and Alf Lundgren. Nitrogen fixation in Lake Erken. Limnology and Oceanography, 16(5):711–719, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Gardner:1975:RAA

[GL75a] Wayne S. Gardner and G. Fred Lee. The role of amino acids in the nitrogen cycle of Lake Mendota. Limnology and Oceanography, 20 (4):379–388, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Gerhart:1975:EED

[GL75b] David Z. Gerhart and Gene E. Likens. Enrichment experiments for determining nutrient limitation: Four methods compared. Limnology and Oceanography, 20(6):649–653, July 1975. CODEN LI-OCAH. ISSN 0024-3590.

Galloway:1979:AEM

[GL79] James N. Galloway and Gene E. Likens. Atmospheric enhancement of metal deposition in Adirondack lake sediments. Limnology and Oceanography, 24(3):427–433, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Glover:1978:IMC

[Glo78] Hilary E. Glover. Iron in Maine coastal waters; seasonal variation and its apparent correlation with a dinoflagellate bloom. Limnology and Oceanography, 23(3):534–537, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Gorham:1974:SRB

[GLSD74] Eville Gorham, John W. G. Lund, Jon E. Sanger, and Walter E. Dean, Jr. Some relationships between algal standing crop, water chemistry, and sediment chemistry in the English Lakes. Limnology and Oceanography, 19(4):601–617, July 1974. CODEN LI-OCAH. ISSN 0024-3590.

Gerber:1974:IDL

[GM74] Ray P. Gerber and Nelson Marshall. Ingestion of detritus by the lagoon pelagic community at Eniwetok Atoll. Limnology and Oceanography, 19(5):815–824, September 1974. CODEN LI-OCAH. ISSN 0024-3590.

Grundmanis:1977:NDM

[GM77] Varis Grundmanis and James W. Murray. Nitrification and denitrification in marine sediments from Puget Sound. Limnology and Oceanography, 22(5):804–813, September 1977. CODEN LI-OCAH. ISSN 0024-3590.

Goldman:1978:SSG

[GM78] Joel C. Goldman and James J. McCarthy. Steady state growth and ammonium uptake of a fast-growing marine diatom. Limnology and Oceanography, 23(4):695–703, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Glover:1979:PCE

[GM79] Hilary E. Glover and Ian Morris. Photosynthetic carboxylating enzymes in marine phytoplankton. Limnology and Oceanography, 24(3):510–519, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Geesey:1978:SBI

[GMCG78] G. G. Geesey, R. Mutch, J. W. Costerton, and R. B. Green. Sessile bacteria: an important component of the microbial population in small mountain streams. Limnology and Oceanography, 23(6): 1214–1223, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Gundersen:1972:SCM

[GMT^{+}72] K. Gundersen, C. W. Mountain, Diane Taylor, R. Ohye, and J. Shen. Some chemical and microbiological observations in the Pacific Ocean off the Hawaiian Islands. Limnology and Oceanography, 17(4):524–532, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Goldman:1979:PDA

[GMTA79] Charles R. Goldman, Mark D. Morgan, Stephen T. Threlkeld, and Nadine Angeli. A population dynamics analysis of the cladoceran disappearance from Lake Tahoe, California-nevada. Limnology and Oceanography, 24(2):289–297, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Glooschenko:1972:SCP

[GMV72] Walter A. Glooschenko, James E. Moore, and Richard A. Vollenweider. The seasonal cycle of pheo-pigments in Lake Ontario with particular emphasis on the role of zooplankton grazing. Limnology and Oceanography, 17(4):597–605, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Gagosian:1979:TBS

[GN79] Robert B. Gagosian and Gale E. Nigrelli. The transport and budget of sterols in the western North Atlantic Ocean. Limnology and Oceanography, 24(5):838–849, September 1979. CODEN LI-OCAH. ISSN 0024-3590.

George:1978:NTS

[GO78] D. G. George and G. H. Owen. A new tube sampler for crustacean zooplankton. Limnology and Oceanography, 23(3):563–566, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Goldberg:1974:BRD

[Gol74a] Edward D. Goldberg. Book review: Data volume 1: Environmental of Port Valdez. Inst. Mar. Sci. Occas. Publ. 3A 800 p. $10.00. Univ. Alaska, Fairbanks. Limnology and Oceanography, 19(6):876, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Goldberg:1974:BRH

[Gol74b] Edward D. Goldberg. Book review: Hood, D. W., W. E. Shiels, and E. J. Kelley [Eds.]. 1973. Environmental studies of Port Valdez. Inst. Mar. Sci. Occas. Publ. 3. 495 p. $15.00. Limnology and Oceanography, 19(6):876–877, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Goldman:1977:BRK

[Gol77a] Joel C. Goldman. Book review: Kinne, O. [Ed.] 1976. Marine ecology. V. 3: Cultivation. Part 1. John Wiley and Sons, Inc., New York, xiii + 577 p. $49.00. Limnology and Oceanography, 22 (5):971, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Goldman:1977:TEP

[Gol77b] Joel C. Goldman. Temperature effects on phytoplankton growth in continuous culture. Limnology and Oceanography, 22(5):932– 936, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Golberg:1978:BRH

[Gol78] Edward D. Golberg. Book review: Hood, D. W., and D. C. Burrell [Eds.] 1976. Assessment of the Arctic marine environment. Inst. Mar. Sci. (Alaska) Occas. Publ. 4. xi + 486 p. Limnology and Oceanography, 23(3):582, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Goldberg:1979:BRR

[Gol79] Edward D. Goldberg. Book review: Riley, J. P., and R. Chester [Eds.]. 1978. Chemical oceanography, v. 7, 2nd Edition. Academic Press, London, xvi + 508 p. £19.60; US$40.50. Limnology and Oceanography, 24(6):1178, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Gorsline:1978:BRW

[Gor78] D. S. Gorsline. Book review: Wiley, M. [Ed.] 1977. Estuarine processes, v. 2: Circulation, sediments and transfer of material in the estuary. Academic Press, New York and London. xviii + 428 p. $19.50. Limnology and Oceanography, 23(3):577–578, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Gordon:1979:BRMa

[Gor79a] D. C. Gordon, Jr. Book review: Malins, D. C. [Ed.] 1977. Effects of petroleum on Arctic and subarctic marine environments and organisms. V. 1: Nature and fate of petroleum. V. 2: Biological effects. Academic Press, New York, San Francisco, and London. V. 1: xvii + 321 p. $14.50. V. 2: xx + 500 p. $14.50. Typewriter composition. Limnology and Oceanography, 24(1):199, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Gordon:1979:BRMb

[Gor79b] D. C. Gordon, Jr. Book review: Malins, D. C. [Ed.] 1977. Effects of petroleum on Arctic and subarctic marine environments and organisms. V. 1: Nature and fate of petroleum. V. 2: Biological effects. Academic Press, New York, San Francisco, and London. V. 1: xvii + 321 p. $14.50. V. 2: xx + 500 p. $14.50. Typewriter composition. Limnology and Oceanography, 24(4):785–787, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Goulden:1971:ECA

[Gou71] Clyde E. Goulden. Environmental control of the abundance and distribution of the chydorid cladocera. Limnology and Oceanography, 16(2):320–331, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Gaines:1972:AWP

[GP72a] Arthur G. Gaines, Jr. and Michael E. Q. Pilson. Anoxic water in the Pettaquamscutt River. Limnology and Oceanography, 17(1): 42–50, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Goering:1972:NFE

[GP72b] J. J. Goering and P. L. Parker. Nitrogen fixation by epiphytes on sea grasses. Limnology and Oceanography, 17(2):320–323, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Gallagher:1977:AMC

[GP77] John L. Gallagher and William J. Pfeiffer. Aquatic metabolism of the communities associated with attached dead shoots of salt marsh plants. Limnology and Oceanography, 22(3):562–565, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Greenlaw:1979:AEZ

[Gre79] Charles F. Greenlaw. Acoustical estimation of zooplankton populations. Limnology and Oceanography, 24(2):226–242, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Grigg:1972:OGF

[Gri72] Richard W. Grigg. Orientation and growth form of sea fans. Limnology and Oceanography, 17(2):185–192, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Griffiths:1978:SBG

[Gri78] Mary Griffiths. Specific blue-green algal carotenoids in sediments of esthwaite water. Limnology and Oceanography, 23(4):777–784, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Gross:1971:BRL

[Gro71] M. Grant Gross. Book review: Logan, B. W., G. R. Davies, J. F. Read, and D. E. Cebulski 1970. Carbonate sedimentation and environments, Shark Bay, Western Australia. Mem. 13, Amer. Ass. Petrol. Geol., Tulsa, Okla. viii + 223 p. $16.00. Limnology and Oceanography, 16(1):152, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Gross:1974:BRD

[Gro74a] M. Grant Gross. Book review: Degens, E. T., and D. A. Ross [eds.]. 1974. The Black Sea — Geology, chemistry, and biology. American Association of Petroleum Geologists, Tulsa, Oklahoma, ix + 633 p. $33.00. Limnology and Oceanography, 19(6):1016– 1017, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Gross:1974:BRL

[Gro74b] M. Grant Gross. Book review: Laking, P. N. 1974. The Black Sea: Its geology, chemistry, biology. A bibliography. Woods Hole Oceanographic Institution, Woods Hole, Massachusetts, xiv + 368 p. $10.00. Limnology and Oceanography, 19(6):1016, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Gross:1975:BRH

[Gro75a] M. Grant Gross. Book review: Hood, D. W., and E. J. Kelley [Eds.] 1974. Oceanography of the Bering Sea. Inst. Mar. Sci. Univ. Alaska, Fairbanks. Occas. Publ. 2. xxi + 623 p. $20.00. Limnology and Oceanography, 19(4):301–302, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Gross:1975:BRS

[Gro75b] M. Grant Gross. Book review: Sullivan, W. 1974. Continents in motion: The new earth debate. McGraw-Hill, New York, xiv + 399 p. $17.50. Limnology and Oceanography, 20(3):682–683, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Gross:1976:BRG

[Gro76a] M. Grant Gross. Book review: Goldberg, E. D. 1976. The health of the oceans. UNESCO, Paris. 172 p. $9.25. Limnology and Oceanography, 21(6):932, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Gross:1976:BRJ

[Gro76b] M. Grant Gross. Book review: J. P. Riley, and G. Skirrow [Eds.]. 1975. Chemical oceanography, v. 3. 2nd ed. Academic Press, New York and London, xviii + 564 p. $47.00. Limnology and Oceanography, 21(4):635, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Gross:1977:BRS

[Gro77] M. Grant Gross. Book review: Stanley, D. J., and D. J. Swift [Eds.]. 1976. Marine sediment transport and environmental management. John Wiley & Sons, New York, xv + 602 p. $35.00. Limnology and Oceanography, 22(1):174–175, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Gross:1978:BRB

[Gro78] M. Grant Gross. Book review: Burton, J. D., and P. S. Liss [Eds.] 1976. Estuarine chemistry. Academic Press, New York and London. xii + 229 p. $17.50. Limnology and Oceanography, 23(3): 575, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Gross:1979:BRS

[Gro79a] M. Grant Gross. Book review: Skreslet, S., R. Leinebø, J. B. L. Matthews, and E. Sakshaug [Eds.] 1976. Fresh water on the sea. Assoc. Norwegian Oceanographers, Oslo. 246 p. NKr. 50. Limnology and Oceanography, 24(4):794–795, July 1979. CODEN LI-OCAH. ISSN 0024-3590.

Grossnickle:1979:NFP

[Gro79b] Nevin E. Grossnickle. Nocturnal feeding patterns of Mysis relicta in Lake Michigan, based on gut content fluorescence. Limnology and Oceanography, 24(4):777–780, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Gibson:1972:CWC

[GS72a] C. E. Gibson and D. A. Stewart. Changes in the water chemistry of Lough Neagh over a 10-year period. Limnology and Oceanography, 17(4):633–635, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Gorham:1972:FPS

[GS72b] Eville Gorham and Jon E. Sanger. Fossil pigments in the surface sediments of a meromictic lake. Limnology and Oceanography, 17 (4):618–622, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Gibson:1973:ATCa

[GS73a] C. E. Gibson and Duncan A. Stewart. The annual temperature cycle of Lough Neagh. Limnology and Oceanography, 18(5):791– 792, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Gibson:1973:ATCb

[GS73b] C. E. Gibson and Duncan A. Stewart. The annual temperature cycle of Lough Neagh. Limnology and Oceanography, 18(5):795– 797, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Glynn:1973:DCR

[GS73c] Peter W. Glynn and Robert H. Stewart. Distribution of coral reefs in the Pearl Islands (Gulf of Panama) in relation to thermal conditions. Limnology and Oceanography, 18(3):367–379, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Gordon:1974:FSU

[GS74] Donald D. Gordon, Jr. and William H. Sutcliffe, Jr. Filtration of seawater using silver filters for particulate nitrogen and carbon analysis. Limnology and Oceanography, 19(6):989–993, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Gale:1974:ABS

[GT74] William F. Gale and J. Douglas Thompson. Aids to benthic sampling by SCUBA divers in rivers. Limnology and Oceanography, 19(6):1004–1007, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Griffiths:1978:MCU

[GTF78] Dilwyn J. Griffiths, Luong-Van Thinh, and Zoltan Florian. A modified chamber for use with an oxygen electrode to allow measurement of incident illumination within the reaction suspension.

Limnology and Oceanography, 23(2):368–372, March 1978. CO-DEN LIOCAH. ISSN 0024-3590.

Gust:1979:BRN

[Gus79] Giselher Gust. Book review: Nihoul, J. C. J. [Ed.] 1977. Bottom turbulence. Proc. 8th Int. Liege Colloq. Ocean Hydrodynamics, 1976. Elsevier Sci. Publ. Co., Amsterdam and New York, xiv + 306 p. $39.95. Typewriter composition. Limnology and Oceanography, 24(3):599–600, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Gieskes:1973:UME

[GV73] W. W. C. Gieskes and A. J. Van Bennekom. Unreliability of the ^{14}C method for estimating primary productivity in eutrophic Dutch coastal waters. Limnology and Oceanography, 18(3):494– 495, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Gillespie:1978:BBM

[GV78] Paul A. Gillespie and Ralph F. Vaccaro. A bacterial bioassay for measuring the copper-chelation capacity of seawater. Limnology and Oceanography, 23(3):543–548, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Goering:1970:NUP

[GWN70] J. J. Goering, D. D. Wallen, and R. M. Nauman. Nitrogen uptake by phytoplankton in the discontinuity layer of the eastern subtropical Pacific Ocean. Limnology and Oceanography, 15(5): 789–796, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Horner:1972:APA

[HA72] Rita Horner and V. Alexander. Algal populations in Arctic sea ice: an investigation of heterotrophy. Limnology and Oceanography, 17 (3):454–458, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Hager:1972:CSM

[HAG^{+}72] S. W. Hager, E. L. Atlas, L. I. Gordon, A. W. Mantyla, and P. K. Park. A comparison at sea of manual and autoanalyzer analyses of phosphate, nitrate, and silicate. Limnology and Oceanography, 17 (6):931–937, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Haines:1976:RBS

[Hai76] Evelyn B. Haines. Relation between the stable carbon isotope composition of fiddler crabs, plants, and soils in a salt marsh. Limnology and Oceanography, 21(6):880–883, November 1976. CO-DEN LIOCAH. ISSN 0024-3590.

Hairston:1979:ASC

[Hai79a] Nelson G. Hairston, Jr. The adaptive significance of color polymorphism in two species of Diaptomus (Copepoda). Limnology and Oceanography, 24(1):15–37, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Hairston:1979:RBP

[Hai79b] Nelson G. Hairston, Jr. The relationship between pigmentation and reproduction in two species of Diaptomus (Copepoda). Limnology and Oceanography, 24(1):38–44, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Hakanson:1976:BST

[Hak76] Lars Hakanson. A bottom sediment trap for recent sedimentary deposits. Limnology and Oceanography, 21(1):170–174, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Hale:1971:FUC

[Hal71] Alan M. Hale. The feasibility of using continuous dye injection for underwater flow visualization. Limnology and Oceanography, 16 (1):124–129, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Hammer:1975:UOB

[Ham75] William M. Hammer. Underwater observations of blue-water plankton: Logistics, techniques, and safety procedures for divers at sea. Limnology and Oceanography, 20(3):1045–1051, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Hamner:1979:BRA

[Ham79] W. M. Hamner. Book review: Aleyev, Yu. G. 1977. Nekton. Dr. W. Junk, The Hague, vi + 435 p. f.120. Limnology and Oceanography, 24(6):1173–1175, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Haney:1971:SMM

[Han71] James F. Haney. An in situ method for the measurement of zooplankton grazing rates. Limnology and Oceanography, 16(6):970– 977, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Hargrave:1970:EDF

[Har70] Barry T. Hargrave. The effect of a deposit-feeding amphipod on the metabolism of benthic microflora. Limnology and Oceanography, 15(1):21–30, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Hargrave:1971:EBD

[Har71] Barry T. Hargrave. An energy budget for a deposit-feeding amphipod. Limnology and Oceanography, 16(1):99–103, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Hargrave:1972:ADS

[Har72] Barry T. Hargrave. Aerobic decomposition of sediment and detritus as a function of particle surface area and organic content. Limnology and Oceanography, 17(4):583–586, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Harding:1973:DMC

[Har73a] Gareth C. H. Harding. Decomposition of marine copepods. Limnology and Oceanography, 18(4):670–673, July 1973. CODEN LI-OCAH. ISSN 0024-3590.

Hardy:1973:PET

[Har73b] John T. Hardy. Phytoneuston ecology of a temperate marine lagoon. Limnology and Oceanography, 18(4):525–533, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Harrison:1973:NRA

[Har73c] William G. Harrison. Nitrate reductase activity during a dinoflagellate bloom. Limnology and Oceanography, 18(3):457–465, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Hargis:1977:CTM

[Har77] Jack R. Hargis. Comparison of techniques for the measurement of zooplankton filtration rates. Limnology and Oceanography, 22(5): 942–945, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Hargrave:1978:BRC

[Har78a] B. T. Hargrave. Book review: Coull, B. C. [Ed.] 1977. Ecology of marine benthos. Belle W. Baruch Library Mar. Sci. No. 6. Univ. South Carolina Press, Columbia. xvii + 467 p. (Typewriter composition). Limnology and Oceanography, 23(3):580–581, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Harrison:1978:EMN

[Har78b] William G. Harrison. Experimental measurements of nitrogen remineralization in coastal waters. Limnology and Oceanography, 23(4):684–694, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Harrison:1972:RSC

[HB72a] Stanley M. Harrison and Allan W. H. Bé. A rotary system for calibrating flowmeters used for plankton sampling. Limnology and Oceanography, 17(1):152–156, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Harvey:1972:SMM

[HB72b] George W. Harvey and Linden A. Burzell. A simple microlayer method for small samples. Limnology and Oceanography, 17(1): 156–157, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Holland:1972:SES

[HB72c] Ruth E. Holland and Alfred M. Beeton. Significance to eutrophication of spatial differences in nutrients and diatoms in Lake Michigan. Limnology and Oceanography, 17(1):88–96, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Hargrave:1979:AST

[HB79] Barry T. Hargrave and Noel M. Burns. Assessment of sediment trap collection efficiency. Limnology and Oceanography, 24(6): 1124–1136, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Hurlbert:1976:LVL

[HBLP76] Stuart H. Hurlbert, Richard W. Berry, Matilde Lopez, and Silvia Pezzani. Lago Verde and Lago Flaco: Gypsum-bound lakes of the Chilean altiplano. Limnology and Oceanography, 21(5):637–645, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Hopkins:1973:MOC

[HBM73] Thomas L. Hopkins, Ronald C. Baird, and Dean M. Milliken. A messenger-operated closing trawl. Limnology and Oceanography, 18(3):488–490, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Head:1975:NFA

[HC75a] William D. Head and Edward J. Carpenter. Nitrogen fixation associated with the marine macroalga Codium fragile. Limnology and Oceanography, 20(4):815–823, September 1975. CODEN LI-OCAH. ISSN 0024-3590.

Holland:1975:HDP

[HC75b] Ruth E. Holland and Larry W. Claflin. Horizontal distribution of planktonic diatoms in Green Bay, mid-July 1970. Limnology and Oceanography, 20(4):365–378, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Hargrave:1978:DCS

[HC78] B. T. Hargrave and G. F. Connolly. A device to collect supernatant water for measurement of the flux of dissolved compounds across sediment surfaces. Limnology and Oceanography, 23(5):1005–1010, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Hamilton:1979:TSA

[HC79a] Eric I. Hamilton and Robert J. Clifton. Techniques for sampling the air–sea interface for estuarine and coastal waters. Limnology and Oceanography, 24(1):188–193, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Hamner:1979:CSA

[HC79b] W. M. Hamner and J. H. Carleton. Copepod swarms: Attributes and role in coral reef ecosystems. Limnology and Oceanography, 24(1):1–14, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Harbison:1979:ETS

[HC79c] G. R. Harbison and R. B. Campenot. Effects of temperature on the swimming of salps (Tunicata, Thaliacea): Implications for vertical migration. Limnology and Oceanography, 24(6):1081–1091, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Hall:1970:EAP

[HCW70] Donald J. Hall, William E. Cooper, and Earl E. Werner. An experimental approach to the production dynamics and structure of freshwater animal communities. Limnology and Oceanography, 15(6):839–928, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Hannan:1970:SMC

[HD70] Herbert H. Hannan and Troy C. Dorris. Succession of a macrophyte community in a constant temperature river. Limnology and Oceanography, 15(3):442–453, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Heaney:1977:DFI

[HD77] S. I. Heaney and W. Davison. The determination of ferrous iron in natural waters with 2,2’ bipyridyl. Limnology and Oceanography, 22(4):753–760, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Horne:1972:NFCb

[HDFG72] Alexander J. Horne, J. E. Dillard, D. K. Fujita, and C. R. Goldman. Nitrogen fixation in Clear Lake, California. II. Synoptic studies on the autumn anabaena bloom. Limnology and Oceanography, 17(5):693–703, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Hinga:1979:ECC

[HDS79] Kenneth R. Hinga, Paul G. Davis, and John McN. Sieburth. Enclosed chambers for the convenient reverse flow concentration and selective filtration of particles. Limnology and Oceanography, 24 (3):536–540, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Heath:1973:FCE

[Hea73] R. A. Heath. Flushing of coastal embayments by changes in atmospheric conditions. Limnology and Oceanography, 18(6):849–862, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Hedgpeth:1970:BRH

[Hed70] Joel W. Hedgpeth. Book review: Holdgate, M. W. [Ed.]. 1970. Antarctic ecology, v. 1. Academic Press, London and New York. xx + 604 p. £6. $18.50. Limnology and Oceanography, 15(6):973– 974, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Hassan:1975:CSD

[HES75] E. M. Hassan and M. I. El-Sabh. Circulation and salinity distribution in the southern part of the Suez Canal. Limnology and Oceanography, 20(5):667–671, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Hesslein:1976:SSC

[Hes76] Raymond H. Hesslein. An in situ sampler for close interval pore water studies. Limnology and Oceanography, 21(6):912–914, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Harman:1970:FYC

[HF70] Willard N. Harman and John L. Forney. Fifty years of change in the molluscan fauna of Oneida Lake, New York. Limnology and Oceanography, 15(3):454–460, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Horne:1972:NFCa

[HG72] Alexander J. Horne and Charles R. Goldman. Nitrogen fixation in Clear Lake, California. I. Seasonal variation and the role of heterocysts. Limnology and Oceanography, 17(5):678–692, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Harbison:1976:FRP

[HG76] G. R. Harbison and R. W. Gilmer. The feeding rates of the pelagic tunicate Pegea confederata and two other salps. Limnology and Oceanography, 21(4):517–528, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Haney:1973:SCD

[HH73] James F. Haney and Donald J. Hall. Sugar-coated Daphnia: a preservation technique for cladocera. Limnology and Oceanography, 18(2):331–333, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Helz:1978:VCB

[HH78] George R. Helz and Rong Y. Hsu. Volatile chloro- and bromocarbons in coastal waters. Limnology and Oceanography, 23(5): 858–869, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Holm-Hamen:1976:CBC

[HHGRW76] O. Holm-Hamen, C. R. Goldman, R. Richards, and P. M. Williams. Chemical and biological characteristics of a water column in Lake Tahoe. Limnology and Oceanography, 21(4):548–562, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Hobbie:1972:SDA

[HHHP^{+}72] John E. Hobbie, Osmund Holm-Hansen, Theodore T. Packard, Lawrence R. Pomeroy, Raymond W. Sheldon, James P. Thomas, and William J. Wiebe. A study of the distribution and activity of microorganisms in ocean water. Limnology and Oceanography, 17 (4):544–555, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Holm-Hansen:1970:ERF

[HHPP70] O. Holm-Hansen, T. T. Packard, and L. R. Pomeroy. Efficiency of the reverse-flow filter technique for concentration of particulate matter. Limnology and Oceanography, 15(5):832–835, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Hansen:1978:MHS

[HIJ78] Marianne Hartmann Hansen, Kjeld Ingvorsen, and Bo Barker Jøgensen. Mechanisms of hydrogen sulfide release from coastal marine sediments to the atmosphere. Limnology and Oceanography, 23(1):68–76, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Hildebrand:1974:RDB

[Hil74] Stephen G. Hildebrand. The relation of drift to benthos density and food level in an artificial stream. Limnology and Oceanography, 19(6):951–957, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Hitchcock:1978:SRL

[Hit78] Dian R. Hitchcock. Sulfide releases from Lake Ontario sediments. Limnology and Oceanography, 23(2):375–377, March 1978. CO-DEN LIOCAH. ISSN 0024-3590. See reply [NC78].

Horne:1971:FRT

[HJG71] Alexander J. Horne, Pavel Javornicky, and Charles R. Goldman. A freshwater “red tide” on Clear Lake, California. Limnology and Oceanography, 16(4):684–689, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Hecky:1973:DAS

[HK73] Robert E. Hecky and Peter Kilham. Diatoms in alkaline, saline lakes: ecology and geochemical implications. Limnology and Oceanography, 18(1):53–71, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Hecky:1974:ECP

[HK74] Robert E. Hecky and Peter Kilham. Environmental control of phytoplankton cell size. Limnology and Oceanography, 19(2):361– 366, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Huizenga:1979:PEM

[HK79] Douglas L. Huizenga and Dana R. Kester. Protonation equilibria of marine dissolved organic matter. Limnology and Oceanography, 24(1):145–150, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Harris:1973:OLC

[HL73a] G. P. Harris and J. N. A. Lott. Observations of Langmuir circulations in Lake Ontario. Limnology and Oceanography, 18(4): 584–589, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Hobbie:1973:OPD

[HL73b] John E. Hobbie and Gene E. Likens. Output of phosphorus, dissolved organic carbon, and fine particulate carbon from Hubbard Brook watersheds. Limnology and Oceanography, 18(5):734–742, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Hansmann:1971:DMM

[HLH71] Eugene W. Hansmann, Charles B. Lane, and James D. Hall. A direct method of measuring benthic primary production in streams. Limnology and Oceanography, 18(5):822–826, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Harrison:1975:DFE

[HM75] P. G. Harrison and K. H. Mann. Detritus formation from eelgrass (Zostera marina L.): The relative effects of fragmentation, leaching, and decay. Limnology and Oceanography, 20(6):924–934, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Harbison:1979:FFR

[HM79] G. R. Harbison and V. L. McAlister. The filter-feeding rates and particle retention efficiencies of three species of Cyclosalpa (Tunicata, Thaliacea). Limnology and Oceanography, 24(5):875–892, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Hamner:1975:UOG

[HMA^{+}75] W. M. Hamner, L. P. Madin, A. L. Alldredge, R. W. Gilmer, and P. P. Hamner. Underwater observations of gelatinous zooplankton: Sampling problems, feeding biology, and behavior. Limnology and Oceanography, 20(6):907–917, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Harbison:1977:RM

[HMS^{+}77] G. Richard Harbison, Laurence P. Madin, Neil R. Swanberg, Douglas C. Biggs, Edward M. Hulburt, Vicki L. McAlister, and James J. McCarthy. Rhizosolenia mats. Limnology and Oceanography, 22(4):739–741, July 1977. CODEN LIOCAH. ISSN 0024-

Harris:1972:PMF

[HMWW72] John E. Harris, Thomas R. Mckee, Raymond C. Wilson, Jr., and U. Grant Whitehouse. Preparation of membrane filter samples for direct examination with an electron microscope. Limnology and Oceanography, 17(5):784–787, September 1972. CODEN LI-OCAH. ISSN 0024-3590.

Horne:1975:PMM

[HNT75] Alexander J. Horne, J. Denis Newbold, and Max M. Tilzer. The productivity, mixing modes, and management of the world’s lakes. Limnology and Oceanography, 20(5):663–666, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Hobbie:1973:BRS

[Hob73] John E. Hobbie. Book review: Sorokin, Y. I., and H. Kadota [Eds.]. 1972. Techniques for the assessment of microbial production and decomposition in fresh waters. IBP Handbook 23. Blackwell Scientific Publ., Oxford and London. xvi + 112 p. $7.00. Limnology and Oceanography, 18(6):1011–1012, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Hobbie:1974:BRM

[Hob74] J. E. Hobbie. Book review: Melchiorri-Santolini, U., and J. W. Hopton [Eds.] 1972. Detritus and its role in aquatic ecosystems. Proc. IBP-UNESCO Symp. Mem. 1st. Ital. Idrobiol. vol. 29(suppl.). 548 p. $22.50. Limnology and Oceanography, 19(1): 169, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Hobbie:1979:BRR

[Hob79] John E. Hobbie. Book review: Reinheimer, G. [Ed.] 1977. Microbial ecology of a brackish water environment. Ecol. Stud. 25. Springer-Verlag, New York, Heidelberg, Berlin. xii + 288 p. $39.60. Limnology and Oceanography, 24(4):795–796, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Hoioarth:1978:RPM

[Hoi78] Robert W. Hoioarth. A rapid and precise method for determining sulfate seawater, estuarine waters, and sediment pore waters. Limnology and Oceanography, 23(5):1066–1069, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Hojerslev:1975:SLA

[Hoj75] Niels Hojerslev. A spectral light absorption meter for measurements in the sea. Limnology and Oceanography, 20(3):1024–1034, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Hojerslev:1978:SMU

[Høj78] Niels K. Højerslev. Solar middle ultraviolet (UV-B) measurement coastal waters rich in yellow substance. Limnology and Oceanography, 23(5):1076–1079, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Holmes:1970:SDT

[Hol70] Robert W. Holmes. The secchi disk in turbid coastal waters. Limnology and Oceanography, 15(5):688–694, September 1970. CO-DEN LIOCAH. ISSN 0024-3590.

Hoopes:1974:FRH

[Hoo74] Rickalon L. Hoopes. Flooding, as the result of Hurricane Agnes, and its effect on a macrobenthic community in an infertile headwater stream in central Pennsylvania. Limnology and Oceanography, 19(5):853–857, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Horne:1975:BRS

[Hor75] Alexander J. Horne. Book review: Salanki, J., and J. E. Ponyi [Eds.]. 1975. Limnology of shallow waters. Symp. Biol. Hung. v.

15. Akademiai Kiado, Budapest. 303 p. $19.00. Limnology and Oceanography, 20(3):1060–1061, November 1975. CODEN LI-OCAH. ISSN 0024-3590.

Horne:1976:BRB

[Hor76] Alexander J. Horne. Book review: Boney, A. D. 1975. Phytoplankton. Studies in Biology Ser., 52. Crane, Russak & Co., Inc., New York, vii + 116 p. $5.25. Limnology and Oceanography, 21(2):347, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Horne:1979:NFC

[Hor79] A. J. Horne. Nitrogen fixation in Clear Lake, California. 4. Diel studies on Aphanizomenon and Anabaena blooms. Limnology and Oceanography, 24(2):329–341, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Hough:1974:PPS

[Hou74] R. Anton Hough. Photorespiration and productivity in submersed aquatic vascular plants. Limnology and Oceanography, 19(6):912– 927, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Howells:1972:BRP

[How72] G. Howells. Book review: Pruter, A. T., and D. L. Alversos [Eds.]. 1972. The Columbia River estuary and adjacent ocean waters. Bioenvironmental studies. Univ. Wash. Press, Seattle. xiii + 896 p. $22.00. Limnology and Oceanography, 17(6):969–970, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Howells:1978:BRW

[How78] G. Howells. Book review: Wiley, M. [Ed.] 1976. Estuarine processes, v. 1: Uses, stresses, and adaption to the estuary. Academic Press, New York and London. xviii + 541 p. $22.00. Limnology and Oceanography, 23(3):576–577, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Hamilton:1970:OHA

[HP70] R. D. Hamilton and Janet E. Preslan. Observations on heterotrophic activity in the eastern tropical Pacific. Limnology and Oceanography, 15(3):395–401, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Hatcher:1974:LCF

[HP74] Roger F. Hatcher and Bruce C. Parker. Laboratory comparisons of four surface microlayer samplers. Limnology and Oceanography, 19(1):162–165, January 1974. CODEN LIOCAH. ISSN 0024-3590. See comment [Gar74].

Halpern:1976:ISW

[HP76] David Halpern and R. Dale Pillsbury. Influence of surface waves on subsurface current measurements in shallow water. Limnology and Oceanography, 21(4):611–616, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Heisey:1977:EAO

[HP77] Dennis Heisey and Karen G. Porter. The effect of ambient oxygen concentration on filtering and respiration rates of Daphnia galeata mendotae and Daphnia magna. Limnology and Oceanography, 22 (5):839–845, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Harrison:1972:SIP

[HPM72a] Michael J. Harrison, Robert E. Pacha, and Richard Y. Morita. Solubilization of inorganic phosphates by bacteria isolated from Upper Klamath Lake sediment. Limnology and Oceanography, 17 (1):50–57, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Hughes:1972:UMA

[HPM72b] Roger N. Hughes, D. L. Peer, and K. H. Mann. Use of multivariate analysis to identify functional components of the benthos in St. Margaret’s Bay, Nova Scotia. Limnology and Oceanography, 17 (1):111–121, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Huyer:1975:SVA

[HPS75] A. Huyer, R. D. Pillsbury, and R. L. Smith. Seasonal variation of the alongshore velocity field ovei the continental shelf off Oregon. Limnology and Oceanography, 20(1):90–95, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Hodgkin:1971:ICE

[HR71] E. P. Hodgkin and R. J. Rippingale. Interspecies conflict in estuarine copepods. Limnology and Oceanography, 16(3):573–576, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Halmann:1974:TUD

[HS74] M. Halmann and M. Stiller. Turnover and uptake of dissolved phosphate in freshwater. A study in Lake Kinneret. Limnology and Oceanography, 19(5):774–783, September 1974. CODEN LI-OCAH. ISSN 0024-3590.

Hirota:1975:STP

[HS75] Jed Hirota and James P. Szyper. Separation of total particulate carbon into inorganic and organic components. Limnology and Oceanography, 20(4):896–900, September 1975. CODEN LI-OCAH. ISSN 0024-3590.

Harris:1977:OSS

[HS77a] G. P. Harris and R. E. H. Smith. Observations of small-scale spatial patterns in phytoplankton populations. Limnology and Oceanography, 22(5):887–899, September 1977. CODEN LI-OCAH. ISSN 0024-3590.

Hitchcock:1977:BLN

[HS77b] Gary L. Hitchcock and Theodore J. Smayda. Bioassay of lower Narragansett Bay waters during the 1972–1973 winter-spring bloom using the diatom Skeletonema costatum. Limnology and Oceanography, 22(1):132–139, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Hitchcock:1977:ILI

[HS77c] Gary L. Hitchcock and Theodore J. Smayda. The importance of light in the initiation of the 1972–1973 winter-spring diatom bloom in Narragansett Bay. Limnology and Oceanography, 22(1): 126–131, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Home:1979:NFC

[HSC79] Alexander J. Home, James C. Sandusky, and Coen J. W. Carmiggelt. Nitrogen fixation in Clear Lake, California. 3. Repetitive synoptic sampling of the spring Aphanizomenon blooms. Limnology and Oceanography, 24(2):316–328, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Herrgesell:1976:SOD

[HSK76] Perry L. Herrgesell, Thomas H. Sibley, and Allen W. Knight. Some observations on dinoflagellate population density during a bloom in a California reservoir. Limnology and Oceanography, 21(4):619– 624, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Hsu:1975:WSN

[Hsu75] S. A. Hsu. Wind stress on nearshore and lagoonal waters of a tropical island. Limnology and Oceanography, 20(1):113–115, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Hendy:1972:DLC

[HSW72] C. H. Hendy, M. J. Selby, and A. T. Wilson. Deep Lake, Cape Barne — Antarctica. Limnology and Oceanography, 17(3):356– 362, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Hawks:1970:COF

[HT70] Robert C. Hawks and Arthur G. Troup. Chemistry and origin of freshwater ferromanganese concretions. Limnology and Oceanography, 15(5):702–712, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Howarth:1979:SRN

[HT79] Robert W. Howarth and John M. Teal. Sulfate reduction in a New England salt marsh. Limnology and Oceanography, 24(6): 999–1013, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Hu:1978:RBV

[Hu78] Vernon J. H. Hu. Relationships between vertical migration and diet in four species of euphausiids. Limnology and Oceanography, 23(2):296–306, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Hulburt:1976:LPS

[Hul76] Edward M. Hulburt. Limitation of phytoplankton species in the ocean off western Africa. Limnology and Oceanography, 21(2): 193–211, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Hunt:1975:BRK

[Hun75] John M. Hunt. Book review: Kaplan, I. R. [Ed.]. 1974. Natural gases in marine sediments. Plenum Press, New York and London. 324 p. $25.00. Limnology and Oceanography, 20(3):1058, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Huntley:1977:BRM

[Hun77] D. A. Huntley. Book review: Le Méhauté, B. 1976. An introduction to hydrodynamics and water waves. Springer-Verlag, New York, viii + 323 p. $24.80. Limnology and Oceanography, 22(5): 974–975, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Hurley:1973:FAB

[Hur73] Ann C. Hurley. Fecundity of the acorn barnacle Balanus pacificus Pilsbry: a fugitive species. Limnology and Oceanography, 18(3): 386–393, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Hutchinson:1970:CET

[Hut70] G. E. Hutchinson. The chemical ecology of three species of Myriophyllum (Angiospermae, Haloragaceae). Limnology and Oceanography, 15(1):1–5, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Hutchinson:1972:BRL

[Hut72] G. E. Hutchinson. Book review: Likens, G. E. [Ed.]. 1972. Nutrients and eutrophication. Amer. Soc. Limnol. Oceanogr. Spec.

Symp. 1. 328 p. $6.00, members; $8.00, nonmembers. Limnology and Oceanography, 17(6):965–968, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Hutchinson:1974:RP

[Hut74] G. Evelyn Hutchinson. De rebus planctonicis. Limnology and Oceanography, 19(2):360–361, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Huyer:1977:SVT

[Huy77] Adriana Huyer. Seasonal variation in temperature, salinity, and density over the continental shelf off Oregon. Limnology and Oceanography, 22(3):442–453, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Haury:1976:EES

[HW76] Loren Haury and Daniel Weihs. Energetically efficient swimming behavior of negatively buoyant zooplankton. Limnology and Oceanography, 21(6):797–803, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Hough:1978:PCC

[HW78] R. Anton Hough and Robert G. Wetzel. Photorespiration and CO_{2} compensation point in Najas flexilis. Limnology and Oceanography, 23(4):719–724, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Hall:1970:AAE

[HWL70] K. J. Hall, W. C. Weimer, and G. Fred Lee. Amino acids in an estuarine environment. Limnology and Oceanography, 15(1):162– 164, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Hynes:1971:BRW

[Hyn71] H. B. N. Hynes. Book review: Warren, C. E. 1971. Biology and water pollution control. W. B. Saunders Co., Philadelphia. xvi + 434 p. $11.00. Limnology and Oceanography, 16(3):593, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Hynes:1973:BRC

[Hyn73] H. B. N. Hynes. Book review: Clark, R. B., and R. J. Wootton [Eds.]. 1972. Essays in hydrobiology (presented to Leslie Harvey). University of Exeter. xii + 136 p. £4.00. Limnology and Oceanography, 18(4):694, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Hynes:1974:FSD

[Hyn74] H. B. N. Hynes. Further studies on the distribution of stream animals within the substratum. Limnology and Oceanography, 19 (1):92–99, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Hynes:1976:BRN

[Hyn76] H. B. N. Hynes. Book review: National Academy of Sciences. 1975. Productivity of world ecosystems. Proc. Symp. 5th General Assembly of the Special Committee for the International Biological Program, 31 August–1 September 1972, Seattle, Washington. NAS-NRC, Washington, D.C. vii + 166 p. $11.50. Limnology and Oceanography, 21(1):184–185, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Hynes:1978:BRA

[Hyn78] H. B. N. Hynes. Book review: Anderson, J. M., and A. Macfadyen [Eds.] 1976. The role of terrestrial and aquatic organisms in decomposition processes. British Ecol. Soc. Symp. 17, 1975. Blackwell Sci. Publ., Oxford. (Distributed in U.S.A. by Halsted Press, New York, and in Canada by J. B. Lippincott Co., Toronto.) xi + 474 p. £14. Limnology and Oceanography, 23(3):579–580, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Iverson:1976:LRD

[IBM76] Richard L. Iverson, Henry F. Bittaker, and Vernon B. Myers. Loss of radiocarbon in direct use of Aquasol for liquid scintillation counting of solutions containing ^{14}C-NaHCO_{3}. Limnology and Oceanography, 21(5):756–758, September 1976. CODEN LI-OCAH. ISSN 0024-3590.

Ichiye:1971:BRM

[Ich71] Takashi Ichiye. Book review: Marr, J. C. [Ed.]. 1970. The Kuroshio: a symposium on the Japan Current. East-West Center Press, Honolulu, x + 614 p. $17.50. Limnology and Oceanography, 16(6):1006, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Iverson:1974:SPB

[ICO^{+}74] Richard L. Iverson, Herbert C. Curl, Jr., Harold B. O’Connors, Jr., Deborah Kirk, and Karen Zakar. Summer phytoplankton blooms in Auke Bay, Alaska, driven by wind mixing of the water column. Limnology and Oceanography, 19(2):271–278, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Idso:1972:MSN

[Ids72] Sherwood B. Idso. Modifications of shielded net radiometers to measure solar radiation profiles in water. Limnology and Oceanography, 17(3):462–466, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Idso:1973:CLS

[Ids73] Sherwood B. Idso. On the concept of lake stability. Limnology and Oceanography, 18(4):681–683, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Imboden:1978:NRP

[IE78] D. M. Imboden and S. Emerson. Natural radon and phosphorus as limnologic tracers: Horizontal and vertical eddy diffusion in Greifensee. Limnology and Oceanography, 23(1):77–90, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Ingvorsen:1979:CMO

[IJ79] Kjeld Ingvorsen and Bo Barker Jorgensen. Combined measurement of oxygen and sulfide in water samples. Limnology and Oceanography, 24(2):390–393, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Kanari:1975:LPI

[iK75] Sei ichi Kanari. The long-period internal waves in Lake Biwa. Limnology and Oceanography, 20(4):544–553, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Illg:1975:BRF

[Ill75] P. L. Illg. Book review: Faulkner, D. 1974. This living reef. Quadrangle/The New York Times Book Co., New York. 179 p. $27.50. Limnology and Oceanography, 19(4):301, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Imboden:1974:PML

[Imb74] Dieter M. Imboden. Phosphorus model of lake eutrophication. Limnology and Oceanography, 19(2):297–304, March 1974. CO-DEN LIOCAH. ISSN 0024-3590.

Icanberry:1973:QSL

[IR73] John W. Icanberry and Roland W. Richardson. Quantitative sampling of live zooplankton with a filter-pump system. Limnology and Oceanography, 18(2):333–335, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Imboden:1977:LTG

[IWC^{+}77] Dieter M. Imboden, Ray F. Weiss, Harmon Craig, Robert L. Michel, and Charles R. Goldman. Lake Tahoe geochemical study.

1. Lake chemistry and tritium mixing study. Limnology and Oceanography, 22(6):1039–1051, November 1977. CODEN LI-OCAH. ISSN 0024-3590.

Jacobsen:1970:STC

[Jac70] VØGg H. Jacobsen. A simple tool for collecting burrowing animals in submerged areas. Limnology and Oceanography, 15(4):646–648, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Jackson:1977:NPG

[Jac77] George A. Jackson. Nutrients and production of giant kelp, Macrocystis pyrifera, off southern California. Limnology and Oceanography, 22(6):979–995, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Jannasch:1973:BCP

[Jan73] Holger W. Jannasch. Bacterial content of particulate matter in offshore surface waters. Limnology and Oceanography, 18(2):340– 342, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Jannasch:1974:SSC

[Jan74] Holger W. Jannasch. Steady state and the chemostat in ecology. Limnology and Oceanography, 20(2):716–720, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Jannasch:1975:MOL

[Jan75] Holger W. Jannasch. Methane oxidation in Lake Kivu (central Africa). Limnology and Oceanography, 20(4):860–864, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Jassby:1975:EAE

[Jas75] Alan D. Jassby. An evaluation of ATP estimations of bacterial biomass in the presence of phytoplankton. Limnology and Oceanography, 20(6):646–648, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Jones:1971:CWL

[JB71] Everett N. Jones and David G. Browning. Cold water layer in the southern Red Sea. Limnology and Oceanography, 16(3):503–509, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Jackson:1975:FTB

[JC75] John E. Jackson, Jr. and Richard W. Castenholz. Fidelity of thermophilic blue-green algae to hot spring habitats. Limnology and Oceanography, 20(3):305–322, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Jorgensen:1977:SLS

[JC77] Bo Barker Jørgensen and Yehuda Cohen. Solar Lake (Sinai). 5. The sulfur cycle of the bcnthic cyanobacterial mats. Limnology and Oceanography, 22(4):657–666, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Johannes:1970:RZN

[JCK70] R. E. Johannes, Stephen Lee Coles, and Nancy T. Kuenzel. The role of zooplankton in the nutrition of some scleractinian corals. Limnology and Oceanography, 15(4):579–586, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Jordan:1978:IFP

[JDL78] Marilyn J. Jordan, Ralph J. Daley, and Kenneth Lee. Improved filtration procedures for freshwater [^{35}S]SO_{4} uptake studies. Limnology and Oceanography, 23(1):154–157, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Jeffries:1970:SCT

[Jef70] H. Perry Jeffries. Seasonal composition of temperate plankton communities: fatty acids. Limnology and Oceanography, 15(3): 419–426, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Jeffries:1972:FAE

[Jef72] H. Perry Jeffries. Fatty-acid ecology of a tidal marsh. Limnology and Oceanography, 17(3):433–440, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Jernelov:1970:RMM

[Jer70] Arne Jernelöv. Release of methyl mercury from sediments with layers containing inorganic mercury at different depths. Limnology and Oceanography, 15(6):958–960, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Josephson:1972:CMS

[JF72] Robert K. Josephson and Karl W. Flessa. Cryolite: a medium for the study of burrowing aquatic organisms. Limnology and Oceanography, 17(1):134–135, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Jeffrey:1973:IDS

[JFH73] L. M. Jeffrey, A. D. Fredericks, and E. Hillier. An inexpensive deep-sea pumping system. Limnology and Oceanography, 18(2): 336–340, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Jeter:1972:EB

[JFKG72] Hewitt W. Jeter, Ernst Føyn, Michael King, and Louis I. Gordon. The expendable bathyoxymeter. Limnology and Oceanography, 17 (2):288–292, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Jassby:1974:LRL

[JG74] Alan D. Jassby and Charles R. Goldman. Loss rates from a lake phytoplankton community. Limnology and Oceanography, 20(2): 618–627, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Jorgensen:1979:MTS

[JKC79] Bo Barker Jorgensen, J. Gijs Kuenen, and Yehuda Cohen. Microbial transformations of sulfur compounds in a stratified lake (Solar Lake, Sinai). Limnology and Oceanography, 24(5):799–822, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1970:DLC

[JM70] Murray G. Johnson and Michael F. P. Michalski. Determination of low concentrations of inorganic carbon in lake water. Limnology and Oceanography, 15(3):481–482, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Jackson:1978:TMC

[JM78] George A. Jackson and James J. Morgan. Trace metal-chelator interactions and phytoplankton growth in seawater media: Theoretical analysis and comparison with reported observations. Limnology and Oceanography, 23(2):268–282, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1976:NOP

[Joh76] Bruce D. Johnson. Nonliving organic particle formation from bubble dissolution. Limnology and Oceanography, 21(3):444–446, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Jones:1974:SOD

[Jon74] J. Gwynfryn Jones. Some observations on direct counts of freshwater bacteria obtained with a fluorescence microscope. Limnology and Oceanography, 19(3):540–543, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Jones:1978:AFI

[Jon78] Roger I. Jones. Adaptations to fluctuating irradiance by natural phytoplankton communities. Limnology and Oceanography, 23(5): 920–926, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Jorgensen:1977:SCC

[Jør77] Bo Barker Jørgensen. The sulfur cycle of a coastal marine sediment (Limfjorden, Denmark). Limnology and Oceanography, 22(5):814– 832, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Jassby:1975:VPE

[JP75] Alan Jassby and Thomas Powell. Vertical patterns of eddy diffusion during stratification in Castle Lake, California. Limnology and Oceanography, 20(4):530–543, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Jassby:1976:MFR

[JP76] Alan D. Jassby and Trevor Platt. Mathematical formulation of the relationship between photosynthesis and light for phytoplankton. Limnology and Oceanography, 21(4):540–547, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Jordan:1978:SUM

[JP78] Marilyn J. Jordan and Bruce J. Peterson. Sulfate uptake as a measure of bacterial production. Limnology and Oceanography, 23 (1):146–150, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1979:BPE

[JPW79] Kenneth S. Johnson, Ricardo M. Pytkowicz, and C. S. Wong. Biological production and the exchange of oxygen and carbon dioxide across the sea surface in Stuart Channel, British Columbia. Limnology and Oceanography, 24(3):474–482, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Jones:1973:SCA

[JS73] Roger Jones and C. J. S. Stuart. A sampler for the chemical analysis of freshwaters using evacuated tubes. Limnology and Oceanography, 18(5):805–809, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Jones:1975:SOF

[JS75] J. G. Jones and B. M. Simon. Some observations on the fluorometric determination of glucose in freshwater. Limnology and Oceanography, 20(4):882–887, September 1975. CODEN LI-OCAH. ISSN 0024-3590.

Johnson:1979:CCS

[JS79] Paul W. Johnson and John McN. Sieburth. Chroococcoid cyanobacteria in the sea: a ubiquitous and diverse phototrophic biomass. Limnology and Oceanography, 24(5):928–935, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1971:LMI

[JSS^{+}71] Paul W. Johnson, John McN. Sieburth, Akella Sastry, C. R. Arnold, and Maxwell S. Doty. Leucothrix mucor infestation of benthic crustacea, fish eggs, and tropical algae. Limnology and Oceanography, 16(6):962–969, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Johannes:1970:MDC

[JW70] R. E. Johannes and William J. Wiebe. Method for determination of coral tissue biomass and composition. Limnology and Oceanography, 15(5):822–824, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Kimerle:1971:PBR

[KA71] Richard A. Kimerle and N. H. Anderson. Production and bioenergetic role of the midge glyptotendipes barbipes (staeger) in a waste stabilization lagoon. Limnology and Oceanography, 16(4): 646–659, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Kiefer:1974:EVP

[KA74] Dale A. Kiefer and Roswell W. Austin. The effect of varying phytoplankton concentration on submarine light transmission in the Gulf of California. Limnology and Oceanography, 19(1):55–64, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Knauer:1977:CCN

[KA77] George A. Knauer and Allyson V. Ayers. Changes in carbon, nitrogen, adenosine triphosphate, and chlorophyll-a in decomposing Thalassia testudinum leaves. Limnology and Oceanography, 22(3): 408–414, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Kaczynski:1971:RFS

[Kac71] V. W. Kaczynski. Rearing fairy shrimp in the laboratory. Limnology and Oceanography, 16(3):586–588, May 1971. CODEN LI-OCAH. ISSN 0024-3590.

Kimmel:1977:CRS

[KAG77] Bruce L. Kimmel, Richard P. Axler, and Charles R. Goldman. A closing, replicate-sample, sediment trap. Limnology and Oceanography, 22(4):768–772, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Kalff:1972:NPN

[Kal72] J. Kalff. Net plankton and nanoplankton production and biomass in a north temperate zone lake. Limnology and Oceanography, 17 (5):712–720, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Kaplan:1978:BRM

[Kap78] I. R. Kaplan. Book review: McCave, I. N. [Ed.] 1976. The benthic boundary layer. Plenum Publ. Corp., New York. x + 323 p. $27.50. Limnology and Oceanography, 23(3):581–582, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Karl:1978:DAM

[Kar78] David M. Karl. Distribution, abundance, and metabolic states of microorganisms in the water column and sediments of the Black Sea. Limnology and Oceanography, 23(5):936–949, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Katona:1971:ASE

[Kat71] Steven K. Katona. Aquatic science and the environmental crisis. Limnology and Oceanography, 16(1):148–151, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Katona:1973:ESP

[Kat73] Steven K. Katona. Evidence for sex pheromones in planktonic copepods. Limnology and Oceanography, 18(4):574–583, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Keirn:1971:NFR

[KB71] Michael A. Keirn and Patrick L. Brezonik. Nitrogen fixation ry bacteria in Lake Mize, Florida, and in some lacustrine sediments. Limnology and Oceanography, 18(5):720–731, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Kupferman:1976:CCM

[KB76] Stuart L. Kupferman and Vaughan T. Bowen. Comparison of 137Cs concentrations measured after in situ absorption with those determined by bulk water analyses or calculated from ^{90}Sr analyses. Limnology and Oceanography, 21(3):467–473, May 1976. CO-DEN LIOCAH. ISSN 0024-3590.

Killworth:1979:FBM

[KC79] Peter D. Killworth and Eddy C. Carmack. A filling-box model of river-dominated lakes. Limnology and Oceanography, 24(2):201– 217, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Krumbein:1977:SLS

[KCS77] Wolfgang E. Krumbein, Yehuda Cohen, and Moshe Shilo. Solar Lake (Sinai). 4. Stromatolitic cyanobacterial mats. Limnology and Oceanography, 22(4):635–656, July 1977. CODEN LIOCAH. ISSN 0024-3590.

King:1979:EVE

[KD79a] Frederick D. King and Allan H. Devol. Estimates of vertical eddy diffusion through the thermocline from phytoplankton nitrate uptake rates in the mixed layer of the eastern tropical Pacific. Limnology and Oceanography, 24(4):645–651, July 1979. CODEN LI-OCAH. ISSN 0024-3590.

Kinsey:1979:EEN

[KD79b] Donald W. Kinsey and Peter J. Davies. Effects of elevated nitrogen and phosphorus on coral reef growth. Limnology and Oceanography, 24(5):935–940, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Kelley:1970:CDS

[Kel70] John J. Kelley, Jr. Carbon dioxide in the surface waters of the north Atlantic Ocean and the Barents and Kara Seas. Limnology and Oceanography, 15(1):80–87, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Kellogg:1975:BRH

[Kel75] Thomas B. Kellogg. Book review: Herman, Y. [Ed.]. 1974. Marine geology and oceanography of the Arctic Seas. Springer-Verlag, New York, ix + 397 p. $34.80. Limnology and Oceanography, 20(3): 1057, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Kersting:1972:NCC

[Ker72] Kees Kersting. A nitrogen correction for caloric values. Limnology and Oceanography, 17(4):643–644, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Kerfoot:1977:ICP

[Ker77] W. Charles Kerfoot. Implications of copepod predation. Limnology and Oceanography, 22(2):316–325, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Kerfoot:1978:CBP

[Ker78] W. Charles Kerfoot. Combat between predatory copepods and their prey: Cyclops, Epischura, and Bosmina. Limnology and Oceanography, 23(6):1089–1102, November 1978. CODEN LI-OCAH. ISSN 0024-3590.

Kester:1972:EIP

[Kes72] Dana R. Kester. Effect of ion pairing on the pH of seawater. Limnology and Oceanography, 17(6):959–960, November 1972. CO-DEN LIOCAH. ISSN 0024-3590. See reply [Wan72b].

Kimmel:1978:RCM

[KGP^{+}78] Bruce L. Kimmel, Richard M. Gersberg, Larry J. Paulson, Richard P. Axler, and Charles R. Goldman. Recent changes in the meromictic status of Big Soda Lake, Nevada. Limnology and Oceanography, 23(5):1021–1025, September 1978. CODEN LI-OCAH. ISSN 0024-3590.

Kilham:1973:FGE

[KH73] Peter Kilham and Robert E. Hecky. Fluoride: geochemical and ecological significance in East African waters and sediments. Limnology and Oceanography, 18(6):932–945, November 1973. CO-DEN LIOCAH. ISSN 0024-3590.

Koenings:1976:ICO

[KH76] Jeffery P. Koenings and Frank F. Hooper. The influence of colloidal organic matter on iron and iron-phosphorus cycling in an acid bog lake. Limnology and Oceanography, 21(5):684–696, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Kelly:1974:CAM

[KHC74] Mahlon G. Kelly, George M. Hornberger, and B. J. Cosby. Continuous automated measurement of rates of photosynthesis and respiration in an undisturbed river community. Limnology and Oceanography, 19(2):305–312, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Kiefer:1972:PLT

[KHHG^{+}72] Dale A. Kiefer, Osmund Holm-Hansen, Charles R. Goldman, Robert Richards, and Thomas Berman. Phytoplankton in Lake Tahoe: deep-living populations. Limnology and Oceanography, 17 (3):418–422, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Kibby:1971:ETF

[Kib71] H. V. Kibby. Effect of temperature on the feeding behavior of Daphnia rosea. Limnology and Oceanography, 16(3):580–581, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Kilham:1971:HCS

[Kil71] Peter Kilham. A hypothesis concerning silica and the freshwater planktonic diatoms. Limnology and Oceanography, 16(1):10–18, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Kilham:1976:BRC

[Kil76] Peter Kilham. Book review: Cole, G. A. 1975. Textbook of limnology. Mosby, Saint Louis. 283 p. $12.00. Limnology and Oceanography, 21(1):186–187, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Kimball:1973:SFIa

[Kim73a] Kenneth D. Kimball. Seasonal fluctuations of ionic copper in Knights Pond, Massachusetts. Limnology and Oceanography, 18 (1):169–172, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Kimball:1973:SFIb

[Kim73b] Kenneth D. Kimball. Seasonal fluctuations of ionic copper in Knights Pond, Massachusetts. Limnology and Oceanography, 18 (1):172–175, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Kimor:1979:PNM

[Kim79] Baruch Kimor. Predation by Noctiluca miliaris Souriray on Acartia tonsa Dana eggs in the inshore waters of southern California. Limnology and Oceanography, 24(3):568–572, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Kinney:1970:BRH

[Kin70] Patrick Kinney. Book review: Hoult, D. P. [Ed.]. 1969. Oil on the sea. Plenum Press, New York and London. vii + 114 p. $8.75. Limnology and Oceanography, 15(3):489–490, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Kinsey:1978:ACC

[Kin78] Donald W. Kinsey. Alkalinity changes and coral reef calcification. Limnology and Oceanography, 23(5):989–991, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Kirchner:1975:EST

[Kir75a] W. B. Kirchner. An evaluation of sediment trap methodology. Limnology and Oceanography, 20(5):657–660, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Kirchner:1975:ERB

[Kir75b] W. B. Kirchner. An examination of the relationship between drainage basin morphology and the export of phosphorus. Limnology and Oceanography, 19(4):267–270, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Kjerfve:1978:BIN

[Kje78] Björn Kjerfve. Bathymetry as an indicator of net circulation in well mixed estuaries. Limnology and Oceanography, 23(4):816– 821, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Knoechel:1976:AGD

[KK76a] R. Knoechel and J. Kalff. The applicability of grain density autoradiography to the quantitative determination of algal species production: a critique. Limnology and Oceanography, 21(4):583– 590, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Knoechel:1976:TAM

[KK76b] R. Knoechel and J. Kalff. Track autoradiography: a method for the determination of phytoplankton species productivity. Limnology and Oceanography, 21(4):590–596, July 1976. CODEN LI-OCAH. ISSN 0024-3590.

Knoechel:1978:SSP

[KK78] R. Knoechel and J. Kalff. An in situ study of the productivity and population dynamics of five freshwater planktonic diatom species. Limnology and Oceanography, 23(2):195–218, March 1978. CO-DEN LIOCAH. ISSN 0024-3590.

Knoechel:1979:ADG

[KK79] R. Knoechel and J. Kalff. The advantages and disadvantages of grain density and track autoradiography. Limnology and Oceanography, 24(6):1170–1171, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Kamoisher:1975:EFS

[KL75] John Kamoisher and Kenneth Lawson. Electromagnetic flow sensors. Limnology and Oceanography, 20(2):174–182, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Krishnaswami:1973:GSS

[KLAS73] S. Krishnaswami, D. Lal, B. S. Amin, and A. Soutar. Geochronological studies in Santa Barbara Basin: ^{55}Fe as a unique tracer for particulate settling. Limnology and Oceanography, 18(5):763–770, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Kinney:1971:PDO

[KLG71] Patrick J. Kinney, Theodore C. Loder, and Joanne Groves. Particulate and dissolved organic matter in the Amerasian Basin of the Arctic Oceans. Limnology and Oceanography, 16(1):132–137, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Kanwisher:1974:ISC

[KLM74] John W. Kanwisher, K. D. Lawson, and L. R. McCloskey. An improved, self-contained polarographic dissolved oxygen probe. Limnology and Oceanography, 20(2):700–704, July 1974. CODEN LI-OCAH. ISSN 0024-3590.

Kemp:1972:DFN

[KM72a] A. L. W. Kemp and A. Mudrochova. Distribution and forms of nitrogen in a Lake Ontario sediment core. Limnology and Oceanography, 17(6):855–867, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Knauer:1972:MMP

[KM72b] George A. Knauer and John H. Martin. Mercury in a marine pelagic food chain. Limnology and Oceanography, 17(6):868–876, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Knauer:1973:SVC

[KM73] George A. Knauer and John H. Martin. Seasonal variations of cadmium, copper, manganese, lead, and zinc in water and phytoplankton in Monterey Bay, California. Limnology and Oceanography, 18(4):597–604, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Kitchen:1975:PSD

[KMP^{+}75] James C. Kitchen, David Menzies, Hasong Pak, J. Ronald, and V. Zaneveld. Particle size distributions in a region of coastal upwelling analyzed by characteristic vectors. Limnology and Oceanography, 20(4):775–783, September 1975. CODEN LI-OCAH. ISSN 0024-3590.

Krezoski:1978:IBM

[KMR78] John R. Krezoski, Samuel C. Mozley, and John A. Rohhins. Influence of benthic macroinvertebrates on mixing of profundal sediments in southeastern Lake Huron. Limnology and Oceanography, 23(5):1011–1016, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Knoechel:1977:ASG

[Kno77] R. Knoechel. Analyzing the significance of grazing in Lake Erken. Limnology and Oceanography, 22(5):967–969, September 1977. CODEN LIOCAH. ISSN 0024-3590. See reply [Cus77].

Kring:1976:ADP

[KO76] R. Lynn Kring and W. John O’Brien. Accommodation of Daphnia pulex to altered pH conditions as measured by feeding rate. Limnology and Oceanography, 21(2):313–315, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Koenings:1976:SED

[Koe76] Jeffery P. Koenings. In situ experiments on the dissolved and colloidal state of iron in an acid bog lake. Limnology and Oceanography, 21(5):674–683, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Kohn:1971:BRC

[Koh71a] Alan J. Kohn. Book review: Cheng, T. C. [Ed.]. 1971. Aspects of the biology of symbiosis. Univ. Park Press-Butterworths, Baltimore and London, x + 327 p. $14.50. Limnology and Oceanography, 18(5):842, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Kohn:1971:DUR

[Koh71b] Alan J. Kohn. Diversity, utilization of resources, and adaptive radiation in shallow-water marine invertebrates of tropical oceanic islands. Limnology and Oceanography, 16(2):332–348, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Kohn:1971:PBE

[Koh71c] Alan J. Kohn. Phylogeny and biogeography of Hutchinsonia: G. E. Hutchinson’s influence through his doctoral students. Limnology and Oceanography, 16(2):173–176, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Kohn:1972:BRE

[Koh72] Alan J. Kohn. Book review: Eltringham, S. K. 1971. Life in mud and sand. English Universities Press, London. vi + 218 p. Softcover £1.50; hardcover £2.95. Limnology and Oceanography, 17(3):503–504, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Kohn:1973:BRR

[Koh73] Alan J. Kohn. Book review: Rudloe, J. 1971. The erotic ocean. World Publ., New York. 448 p. $15.00. Limnology and Oceanography, 18(1):181–182, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Kohn:1974:BRZ

[Koh74] Alan J. Kohn. Book review: Zeitzschel, B. [Ed.]. 1973. The biology of the Indian Ocean. Ecological Studies 3. Springer-Verlag, New York-Heidelberg-Berlin. xiii + 549 p. $49.20. Limnology and Oceanography, 19(3):549–550, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Kohn:1977:BRP

[Koh77] Alan J. Kohn. Book review: Pielou, E. C. 1975. Ecological diversity. John Wiley & Sons, New York, viii + 165 p. $14.95. Limnology and Oceanography, 22(1):174, January 1977. CODEN LI-OCAH. ISSN 0024-3590.

Kohn:1979:BRS

[Koh79] Alan J. Kohn. Book review: Stoddart, D. R., and R. E. Johannes [Eds.] 1978. Coral reefs: Research methods. Monogr. Oceanogr. Methodol. No. 5. UNESCO, Paris, xv + 581 p. Limnology and Oceanography, 24(4):796–797, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Koshinsky:1970:MSL

[Kos70] Gordon D. Koshinsky. The morphometry of shield lakes in Saskatchewan. Limnology and Oceanography, 15(5):695–701, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Koslow:1979:VMS

[Kos79] J. Anthony Koslow. Vertical migrators see the light? Limnology and Oceanography, 24(4):783–784, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Kiefer:1979:RSM

[KOW79] D. A. Kiefer, R. J. Olson, and W. H. Wilson. Reflectance spectroscopy of marine phytoplankton. Part 1. Optical properties as related to age and growth rate. Limnology and Oceanography, 24 (4):664–672, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Kozloff:1970:BRK

[Koz70] Eugene N. Kozloff. Book review: Kinne, O., and H.-P. Bulnheim [Eds.] 1970. Cultivation of marine organisms and its importance for marine biology. (Int. Helgoland Symp.) Helgoländer Wiss. Meeresuntersuch. 20: 721 p. Biol. Anstalt Helgoland, Hamburg. $20.00. Limnology and Oceanography, 15(6):972, November 1970. CODEN LIOCAH. ISSN 0024-3590.

King:1975:RAR

[KP75] Frederick D. King and Theodore T. Packard. Respiration and the activity of the respiratory electron transport system in marine zooplankton. Limnology and Oceanography, 20(4):849–854, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Kraus:1979:BRH

[Kra79] Eric B. Kraus. Book review: Harvey, J. G. 1978. Atmosphere and ocean: Our fluid environments. Crane, Russak & Co., Inc., New York. (Artemis, 1976.) 143 p. $13.50; $8.95 paperback. Limnology and Oceanography, 24(3):598, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Kroes:1971:GIB

[Kro71] H. W. Kroes. Growth interactions between Chlamydomonas globosa Snow and Chlorococcum ellipsoideum Deason and Bold under different experimental conditions, with special attention to the role of pH. Limnology and Oceanography, 16(6):869–879, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Kroes:1972:GIB

[Kro72a] H. W. Kroes. Growth interactions between Chlamydomonas globosa Snow and Chlorococcum ellipsoideum Deason and Bold: the role of extracellular products. Limnology and Oceanography, 17 (3):423–432, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Kroger:1972:USC

[Kro72b] Richard L. Kroger. Underestimation of standing crop by the Surber sampler. Limnology and Oceanography, 17(3):475–478, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Kroopnick:1975:RPO

[Kro75] Peter M. Kroopnick. Respiration, photosynthesis, and oxygen isotope fractionation in oceanic surface water. Limnology and Oceanography, 20(4):988–992, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Kiefer:1970:CSP

[KS70] D. Kiefer and J. D. H. Strickland. A comparative study of photosynthesis in seawater samples incubated under two types of light attenuator. Limnology and Oceanography, 15(3):408–412, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Kahn:1978:PBT

[KS78] Norman Kahn and Elijah Swift. Positive buoyancy through ionic control in the nonmotile marine dinoflagellate Pyrocystis noctiluca Murray ex schuett. Limnology and Oceanography, 23(4):649–658, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Kemp:1971:SCM

[KSGM71] A. L. W. Kemp, H. A. Savile, C. B. Gray, and Alena Mudrochova. A simple corer and a method for sampling the mud-water interface. Limnology and Oceanography, 16(4):689–694, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Kharkar:1976:UTD

[KTTF76] D. P. Kharkar, J. Thomson, K. K. Turekian, and W. O. Forster. Uranium and thorium decay series nuclides in plankton from the Caribbean. Limnology and Oceanography, 21(2):294–299, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Kerfoot:1973:AEA

[KV73] William B. Kerfoot and Ralph F. Vaccaro. Adsorptive extraction for analysis of copper in seawater. Limnology and Oceanography, 18(4):689–693, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Kaplan:1979:DSM

[KVT79] Warren Kaplan, Ivan Valiela, and John M. Teal. Denitrification in a salt marsh ecosystem. Limnology and Oceanography, 24(4): 726–734, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Karanas:1979:MUB

[KVW79] James J. Karanas, Henry Van Dyke, and Robert C. Worrest. Midultraviolet (UV-B) sensitivity of Acartia clausii Giesbrecht (Copepoda). Limnology and Oceanography, 24(6):1104–1116, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Koehl:1977:MAG

[KW77] M. A. R. Koehl and S. A. Wainwright. Mechanical adaptations of a giant kelp. Limnology and Oceanography, 22(6):1067–1071, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Kelley:1975:RSS

[KWD75] J. C. Kelley, T. E. Whitledge, and R. C. Dugdale. Results of sea surface mapping in the Peru upwelling system. Limnology and Oceanography, 20(4):785–795, September 1975. CODEN LI-OCAH. ISSN 0024-3590.

Kaplan:1974:SWS

[KWK74] Eugene H. Kaplan, J. R. Welker, and M. Gayle Krause. A shallowwater system for sampling macrobenthic infauna. Limnology and Oceanography, 19(2):346–350, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Kamykowski:1977:DVM

[KZ77] Daniel Kamykowski and Sara-Joan Zentara. The diurnal vertical migration of motile phytoplankton through temperature gradients. Limnology and Oceanography, 22(1):148–151, January 1977. CODEN LIOCAH. ISSN 0024-3590.

LaFond:1975:BRJ

[LaF75] Eugene C. LaFond. Book review: Jerlov, N. G., and E. Steemaxn Nielsen [Eds.] 1974. Optical aspects of oceanography. Academic Press, London and New York, xv + 494 p. £10.50. Limnology and Oceanography, 19(4):302, March 1975. CODEN LIOCAH. ISSN 0024-3590.

LaFond:1977:BRC

[LaF77] Eugene C. LaFond. Book review: Coachman, L. K., K. Aagaard, and R. B. Tripp. 1976. Bering Strait: The regional physical oceanography. Univ. Washington Press, Seattle and London. 172 p. $20.00. Limnology and Oceanography, 22(5):975, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Lamon:1972:TTP

[Lam72] Douglas W. Lamon. Temperature, transparency, and phytoplankton productivity in Crater Lake, Oregon. Limnology and Oceanography, 17(3):410–417, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Lampert:1974:MDF

[Lam74] Winfried Lampert. A method for determining food selection by zooplankton. Limnology and Oceanography, 19(6):995–998, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Lampert:1978:RDO

[Lam78a] Winfried Lampert. Release of dissolved organic carbon by grazing zooplankton. Limnology and Oceanography, 23(4):831–834, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Lampitt:1978:CFS

[Lam78b] Richard S. Lampitt. Carnivorous feeding by a small marine copepod. Limnology and Oceanography, 23(6):1228–1231, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Lane:1970:GLT

[Lan70] Robert K. Lane. Great lakes thermal studies using infrared imagery. Limnology and Oceanography, 15(2):296–300, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Landry:1975:RBT

[Lan75a] Michael R. Landry. The relationship between temperature and the development of life stages of the marine copepod Acartia clausi Giesbr. Limnology and Oceanography, 20(4):854–857, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Landry:1975:STE

[Lan75b] Michael R. Landry. Seasonal temperature effects and predicting development rates of marine copepod eggs. Limnology and Oceanography, 20(6):434–440, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Landry:1978:PFB

[Lan78] M. R. Landry. Predatory feeding behavior of a marine copepod, Labidocera trispinosa. Limnology and Oceanography, 23(6):1103– 1113, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Landers:1979:DRE

[Lan79] Dixon Hamilton Landers. A durable, reusable enclosure system that compensates for changing water levels. Limnology and Oceanography, 24(5):991–994, September 1979. CODEN LI-OCAH. ISSN 0024-3590.

LaRow:1970:EOT

[LaR70] Edward J. LaRow. The effect of oxygen tension on the vertical migration of Chaoborus larvae. Limnology and Oceanography, 15 (3):357–362, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Lasenby:1975:DOD

[Las75] David C. Lasenby. Development of oxygen deficits in 14 southern Ontario lakes. Limnology and Oceanography, 20(4):993–999, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Lauff:1970:PAS

[Lau70] George H. Lauff. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 15(2): 330–332, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Laubier:1974:BRM

[Lau74] Lucien Laubier. Book review: Menzies, R. J., R. Y. George, and G. T. Rowe [Eds.]. 1973. Abyssal environment and ecology of the world oceans. Wiley-Inter-science, New York and London, xxiii + 488 p. Limnology and Oceanography, 19(6):877–878, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Lerman:1971:MMC

[LB71] Abraham Lerman and Gregg J. Brunskill. Migration of major constituents from lake sediments into lake water and its bearing on lake water composition. Limnology and Oceanography, 16(6): 880–890, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Lee:1977:DAA

[LB77] Cindy Lee and Jeffrey L. Bada. Dissolved amino acids in the equatorial Pacific, the Sargasso Sea, and Biscayne Bay. Limnology and Oceanography, 22(3):502–510, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Lean:1979:EEM

[LB79] D. R. S. Lean and B. K. Burnison. An evaluation of errors in the ^{14}C method of primary production measurement. Limnology and Oceanography, 24(5):917–928, September 1979. CODEN LI-OCAH. ISSN 0024-3590.

Lehman:1975:ARC

[LBL75] John T. Lehman, Daniel B. Botkin, and Gene E. Likens. The assumptions and rationales of a computer model of phytoplankton population dynamics. Limnology and Oceanography, 20(3):343– 364, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Lewin:1971:AIL

[LC71a] Joyce Lewin and Ching-Hong Chen. Available iron: a limiting factor for marine phytoplankton. Limnology and Oceanography, 16(4):670–675, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Long:1971:QCP

[LC71b] Edward B. Long and G. Dennis Cooke. A quantitative comparison of pigment extraction by membrane and glass-fiber filters. Limnology and Oceanography, 16(6):990–992, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Lee:1974:CSL

[LC74] Richard F. Lee and Lanna Cheng. A comparative study of the lipids of water-striders from marine, estuarine, and freshwater environments: Halobates, Rheumatobates, Gerris (Heteroptera: Gerridae). Limnology and Oceanography, 19(6):958–965, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Leach:1970:EAP

[Lea70] J. H. Leach. Epibenthic algal production in an intertidal mudflat. Limnology and Oceanography, 15(4):514–521, July 1970. CODEN LIOCAH. ISSN 0024-3590.

LeCren:1977:BRW

[LeC77] E. D. LeCren. Book review: Whitton, B. A. [Ed.]. 1975. River ecology. Studies in ecology, v. 2. University of California, Berkeley, x + 725 p. $40.00. Limnology and Oceanography, 22(1):173–174, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Lee:1977:DMS

[Lee77a] David Robert Lee. A device for measuring seepage flux in lakes and estuaries. Limnology and Oceanography, 22(1):140–147, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Leetmaa:1977:BRO

[Lee77b] A. Leetmaa. Book review: Officer, C. B. 1976. Physical oceanography of estuaries (and associated coastal waters). John Wiley and Sons, Inc., New York, xii + 465 p. $27.50. Limnology and Oceanography, 22(5):975–976, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1975:BRM

[Leh75] John T. Lehman. Book review: Middlebrooks, E. J., D. H. Falkenborg, and T. E. Maloney [Eds.]. 1973. Modeling the eutrophication process. Ann Arbor Science. 228 p. Limnology and Oceanography, 20(3):685–686, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1976:ENS

[Leh76a] John T. Lehman. Ecological and nutritional studies on Dinobryon Ehrenb.: Seasonal periodicity and the phosphate toxicity problem. Limnology and Oceanography, 21(5):646–658, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1976:FFO

[Leh76b] John T. Lehman. The filter-feeder as an optimal forager, and the predicted shapes of feeding curves. Limnology and Oceanography, 21(4):501–516, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1977:CDC

[Leh77] John T. Lehman. On calculating drag characteristics for decelerating zooplankton. Limnology and Oceanography, 22(1):170–172, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Levinton:1971:CTM

[Lev71a] Jeffrey S. Levinton. Control of Tellinacean (Mollusca: Bivalvia) feeding behavior by predation. Limnology and Oceanography, 16 (4):660–662, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Levy:1971:ACB

[Lev71b] Yitzhak Levy. Anomalies of Ca^{2+} and SO_{4}^{2} in the Bardawil Lagoon, Northern Sinai. Limnology and Oceanography, 16(6):983– 987, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1973:TRL

[Lew73] William M. Lewis, Jr. The thermal regime of Lake Lanao (Philippines) and its theoretical implications for tropical lakes. Limnology and Oceanography, 18(2):200–217, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1977:CAG

[Lew77] William M. Lewis, Jr. Comments on the analysis of grazing in Lake Erken. Limnology and Oceanography, 22(5):966–967, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Leyendekkers:1973:IAF

[Ley73] J. V. Leyendekkers. The ionic activity function of water and the activity coefficient of the hydrogen ion in seawater. Limnology and Oceanography, 18(5):784–787, September 1973. CODEN LI-OCAH. ISSN 0024-3590.

Lam:1976:MCF

[LF76] Ronald K. Lam and Bruce W. Frost. Model of copepod filtering response to changes in size and concentration of food. Limnology and Oceanography, 21(4):490–500, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Likens:1970:NQS

[LG70] Gene E. Likens and John J. Gilbert. Notes on quantitative sampling of natural populations of planktonic rotifers. Limnology and Oceanography, 15(5):816–820, September 1970. CODEN LI-OCAH. ISSN 0024-3590.

Loeb:1979:WNT

[LG79] Stanford L. Loeb and Charles R. Goldman. Water and nutrient transport via groundwater from Ward Valley into Lake Tahoe. Limnology and Oceanography, 24(6):1146–1154, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Loder:1972:DOC

[LH72] Theodore C. Loder and Donald W. Hood. Distribution of organic carbon in a glacial estuary in Alaska. Limnology and Oceanography, 17(3):349–355, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Lee:1973:WET

[LH73a] Richard F. Lee and Jed Hirota. Wax esters in tropical zooplankton and nekton and the geographical distribution of wax esters in marine copepods. Limnology and Oceanography, 18(2):227–239, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Lindberg:1973:MCP

[LH73b] Steven E. Lindberg and Robert C. Harriss. Mechanisms controlling pore water salinities in a salt marsh. Limnology and Oceanography, 18(5):788–791, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Lush:1973:FPF

[LH73c] Donald L. Lush and H. B. N. Hynes. The formation of particles in freshwater leachates of dead leaves. Limnology and Oceanography, 18(6):968–977, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Lien:1978:WDS

[LH78] S. L. Lien and J. A. Hoopes. Wind-driven, steady flows in Lake Superior. Limnology and Oceanography, 23(1):91–103, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Lewin:1973:CCS

[LhC73] Joyce Lewin and Ching hong Chen. Changes in the concentration of soluble and particulate iron in seawater enclosed in containers. Limnology and Oceanography, 18(4):590–596, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Lock:1976:FDO

[LI76] Maurice A. Lock and H. B. N. Ilynes. The fate of “dissolved” organic carbon derived from autumn-shed maple leaves (Acer saccharum) in a temperate hard-water stream. Limnology and Oceanography, 21(3):436–443, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Littler:1973:PHF

[Lit73] Mark M. Littler. The productivity of Hawaiian fringing-reef crustose Corallinaceae and an experimental evaluation of production methodology. Limnology and Oceanography, 18(6):946–952, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Liu:1970:SFW

[Liu70] Paul C. Liu. Some features of wind waves in Lake Michigan. Limnology and Oceanography, 15(2):257–272, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:1971:YPR

[Liv71] D. A. Livingstone. A 22,000-year pollen record from the plateau of Zambia. Limnology and Oceanography, 16(2):349–356, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:1972:BRF

[Liv72a] D. A. Livingstone. Book review: Flint, R. F. 1971. Glacial and Quaternary geology. Wiley. xiv + 892 p. $24.95. Limnology and Oceanography, 17(1):167, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:1972:BRG

[Liv72b] D. A. Livingstone. Book review: Garrels, R. M., and F. T. Mackenzie. 1971. Evolution of sedimentary rocks. W. W. Norton & Co., New York. xvi + 397 p. $11.50. Limnology and Oceanography, 17(1):165–166, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:1976:BRW

[Liv76] D. A. Livingstone. Book review: Wetzel, R. G. 1975. Limnology. W. B. Saunders Co., Philadelphia, London, and Toronto, xii + 743 p. $16.50. Limnology and Oceanography, 21(6):930–931, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:1977:BRB

[Liv77a] D. A. Livingstone. Book review: Beadle, L. C. 1974. The inland waters of tropical Africa. An introduction to tropical limnology. Longman Group, London, viii + 365 p. £7.00. Limnology and Oceanography, 22(5):972–973, September 1977. CODEN LI-OCAH. ISSN 0024-3590.

Livingstone:1977:BRG

[Liv77b] D. A. Livingstone. Book review: Golterman, H. L. 1975. Physiological limnology. Elsevier Sci. Publ. Co., Amsterdam, Oxford, and New York, xiii + 489 p. $51.95; Dfl 125. Limnology and Oceanography, 22(5):973–974, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Livingstone:1978:BRL

[Liv78] D. A. Livingstone. Book review: Likens, G. E., F. H. Bormann, R. S. Pierce, J. S. Eaton, and N. M. Johnson 1977. Biogeochemistry of a forested ecosystem. Springer-Verlag, New York, Heidelberg and Berlin. xii + 146 p. $9.80. Limnology and Oceanography, 23 (3):578, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Lerman:1973:TSS

[LJ73] Abraham Lerman and Blair F. Jones. Transient and steady-state salt transport between sediments and brine in closed lakes. Limnology and Oceanography, 18(1):72–85, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Lock:1979:EFP

[LJ79] Maurice A. Lock and Peter H. John. The effect of flow patterns on uptake of phosphorus by river periphyton. Limnology and Oceanography, 24(2):376–383, March 1979. CODEN LIOCAH. ISSN 0024-3590.

LaBarhera:1974:CEC

[LK74] Michael C. LaBarhera and Veter Kilham. The chemical ecology of copepod distribution in the lakes of East and central Africa. Limnology and Oceanography, 19(3):459–465, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Lasenby:1973:FAM

[LL73] David C. Lasenby and Raymond R. Langford. Feeding and assimilation of Mysis relicta. Limnology and Oceanography, 18(2): 280–285, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Lerman:1975:UMT

[LL75] A. Lerman and T. A. Lietzke. Uptake and migration of tracers in lake sediments. Limnology and Oceanography, 20(4):497–510, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Lam:1977:DAS

[LL77] Patricia Lam and Richard Levins. The dynamics of aquatic systems. 2. The effects of nutrient enrichment on model plankton communities. Limnology and Oceanography, 22(3):454–471, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Li:1979:SSF

[LL79] Judith L. Li and Hiram W. Li. Species-specific factors affecting predator-prey interactions of the copepod Acanthocyclops vernalis with its natural prey. Limnology and Oceanography, 24(4):613– 626, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Land:1975:POC

[LLS75] Lynton S. Land, Judith C. Lang, and Bruce N. Smith. Preliminary observations on the carbon isotopic composition of some reef coral tissues and symbiotic zooxanthellae. Limnology and Oceanography, 19(4):283–287, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Lock:1970:EVC

[LM70] Anthony R. Lock and Ian A. Mclaren. The effect of varying and constant temperatures on the size of a marine copepod. Limnology and Oceanography, 15(4):638–640, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Lai:1976:SSA

[LNS^{+}76] D. Lai, V. N. Nijampurkar, B. L. K. Somayajulu, M. Koide, and E. D. Goldberg. Silicon-32 specific activities in coastal waters of the world oceans. Limnology and Oceanography, 21(2):285–293, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Loden:1974:PCD

[Lod74] Michael S. Loden. Predation by chironomid (Diptera) larvae on oligochaetes. Limnology and Oceanography, 19(1):156–159, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Lohban:1974:SRM

[Loh74] Christopher S. Lohban. A simple, rapid method of solubilizing algal tissue for scintillation counting. Limnology and Oceanography, 19(2):356–359, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Lorenzen:1970:SCI

[Lor70] Carl J. Lorenzen. Surface chlorophyll as an index of the depth, chlorophyll content, and primary productivity of the euphotic layer. Limnology and Oceanography, 15(3):479–480, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Lorenzen:1979:URP

[Lor79] Carl J. Lorenzen. Ultraviolet radiation and phytoplankton photosynthesis. Limnology and Oceanography, 24(6):1117–1120, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Litchfield:1970:ADA

[LP70] Carol D. Litchfield and J. M. Prescott. Analysis by dansylation of amino acids dissolved in marine and freshwaters. Limnology and Oceanography, 15(2):250–256, March 1970. CODEN LIOCAH. ISSN 0024-3590.

LeBlond:1977:SEC

[LP77] P. H. LeBlond and T. R. Parsons. A simplified expression for calculating cohort production. Limnology and Oceanography, 22 (1):156–157, January 1977. CODEN LIOCAH. ISSN 0024-3590.

LeBlond:1978:RCW

[LP78] P. H. LeBlond and T. R. Parsons. Reply to comment by W. E. Ricker. Limnology and Oceanography, 23(2):380–382, March 1978. CODEN LIOCAH. ISSN 0024-3590. See [Ric78].

Lee:1972:TFS

[LR72] A. H. Lee and G. K. Rodgers. Temperature fine structure in Lake Ontario. Limnology and Oceanography, 17(5):672–677, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Lean:1974:THA

[LR74a] D. R. S. Lean and F. H. Rigler. A test of the hypothesis that abiotic phosphate complexing influences phosphorus kinetics in epilimnetic lake water. Limnology and Oceanography, 19(5):784– 788, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Lee:1974:CSS

[LR74b] A. H. Lee and G. K. Rodgers. Comments on “Static stability in freshwater lakes” (T. R. Osborn and P. H. LeBlond). Limnology and Oceanography, 19(3):546, May 1974. CODEN LIOCAH. ISSN 0024-3590. See [OL74].

Lovric:1972:RIT

[LS72] G. Lovrić and P. Strohal. Radioecological investigations in a thermal spring region. Limnology and Oceanography, 17(3):441–444, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Lai:1977:PTR

[LS77] D. Lai and B. L. K. Somayajulu. Particulate transport of radionuclides^{14}C and^{55}Fe to deep waters in the Pacific Ocean. Limnology and Oceanography, 22(1):55–59, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Livingston:1971:TEC

[LT71] H. D. Livingston and Geoffrey Thompson. Trace element concentrations in some modern corals. Limnology and Oceanography, 18 (5):786–796, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Liddicoat:1975:DAS

[LTB75] M. I. Liddicoat, Susan Tibhitts, and E. I. Butler. The determination of ammonia in seawater. Limnology and Oceanography, 20 (1):131–132, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Ludlam:1974:FGL

[Lud74] Stuart D. Ludlam. Fayetteville Green Lake, New York. 6. The role of turbidity currents in lake sedimentation. Limnology and Oceanography, 20(2):656–664, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Ludlam:1976:LSH

[Lud76] Stuart D. Ludlam. Laminated sediments in holomictic Berkshire lakes. Limnology and Oceanography, 21(5):743–746, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Lund:1979:BRW

[Lun79] J. W. G. Lund. Book review: Werner, D. [Ed.] 1977. The biology of diatoms. Bot. Monogr., V. 13. University of California Press, Berkeley and New York. vii + 498 p. $32.50. Limnology and Oceanography, 24(1):200, January 1979. CODEN LIOCAH. ISSN 0024-3590.

LaBarbera:1976:ITF

[LV76] Michael LaBarbera and Steven Vogel. An inexpensive thermistor flowmeter for aquatic biology. Limnology and Oceanography, 21 (5):750–756, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1973:RCT

[LWR73] A. G. Lewis, P. Whitfield, and A. Ramnarine. The reduction of copper toxicity in a marine copepod by sediment extract. Limnology and Oceanography, 18(2):324–326, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Lyman:1973:BRD

[Lym73] John Lyman. Book review: Dyrssen, D., and D. Jagner [EDs.] 1972. The changing chemistry of the oceans (20th Nobel Symposium). Wiley-Interscience, New York. 365 p. $21.50. Limnology and Oceanography, 18(2):350, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Lynch:1977:ZCP

[Lyn77] Michael Lynch. Zooplankton competition and plankton community structure. Limnology and Oceanography, 22(4):775–777, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Lynch:1979:PCZ

[Lyn79] Michael Lynch. Predation, competition, and zooplankton community structure: an experimental study. Limnology and Oceanography, 24(2):253–272, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Morel:1976:SRT

[MA76] Francois M. M. Morel and Donald M. Anderson. On the subject of red tide predictions from temperature patterns. Limnology and Oceanography, 21(4):625–627, July 1976. CODEN LIOCAH. ISSN 0024-3590. See reply [Bal76].

MacNamara:1970:SLO

[Mac70] E. E. MacNamara. Some limnological observations from Enderby Land, Antarctica. Limnology and Oceanography, 15(5):768–775, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Maclean:1977:OPB

[Mac77] J. L. Maclean. Observations on Pyrodinium bahamense Plate, a toxic dinoflagellate, in Papua New Guinea. Limnology and Oceanography, 22(2):234–254, March 1977. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:1978:APB

[Mac78a] Ferren MacIntyre. Additional problems in bubble and jet drop research. Limnology and Oceanography, 23(3):571–573, May 1978. CODEN LIOCAH. ISSN 0024-3590. See reply [BS78a].

MacIsaac:1978:DCI

[Mac78b] J. J. MacIsaac. Diel cycles of inorganic nitrogen uptake in a natural phytoplankton population dominated by Gonyaulax polyedra. Limnology and Oceanography, 23(1):1–9, January 1978. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:1979:BRA

[Mac79] Ferren MacIntyre. Book review: Andersen, N. R., and A. Malahoff [Eds.] 1977. The fate of fossil fuel CO_{2} in the oceans. Plenum Publ.

Corp., New York. xii + 749 p. $59.50. Limnology and Oceanography, 24(1):198, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Makemson:1972:IWS

[Mak72] John C. Makemson. An interstitial water sampler for sandy beaches. Limnology and Oceanography, 17(4):626–628, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Maly:1970:IPA

[Mal70] Edward J. Maly. The influence of predation on the adult sex ratios of two copepod species. Limnology and Oceanography, 15 (4):566–573, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Malone:1971:RIN

[Mal71] Thomas C. Malone. The relative importance of nannoplankton and netplankton as primary producers in tropical oceanic and neritic phytoplankton communities. Limnology and Oceanography, 16 (4):633–639, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Maly:1973:DSC

[Mal73] Edward J. Maly. Density, size, and clutch of two high altitude diaptomid copepods. Limnology and Oceanography, 18(6):840– 848, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Malone:1975:ECP

[Mal75] Thomas C. Malone. Environmental control of phytoplankton cell size. Limnology and Oceanography, 20(6):490, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Maly:1978:SFI

[Mal78] Edward J. Maly. Some factors influencing size of Diaptomus shoshone. Limnology and Oceanography, 23(4):835–837, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Mann:1970:BRS

[Man70] K. H. Mann. Book review: Steele, J. H. [Ed.] 1970. Marine food chains. Univ. Calif. Press, Berkeley and Los Angeles. viii + 552 p. $13.50. Oliver and Boyd, Edinburgh. £5. Limnology and Oceanography, 15(6):972–973, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Manheim:1972:BRW

[Man72a] Frank Manheim. Book review: Whitfield, M. 1971. Ion selective electrodes for the analysis of natural waters. AMSA Handbook, No. 2. Aust. Mar. Sci. Ass., Sydney. 130 p. A$4.00. Limnology and Oceanography, 17(1):165, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Mann:1972:CEI

[Man72b] K. H. Mann. On communicating ecological information to the general public. Limnology and Oceanography, 17(5):795, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Mann:1973:BRB

[Man73a] K. H. Mann. Book review: Barnes, J. S. K., and J. Green [Eds.] 1972. The estuarine environment. Applied Science Publishers Ltd. London. xii + 133 p. £3.50. Limnology and Oceanography, 18(2): 348–349, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Mantuani:1973:POA

[Man73b] Mark A. Mantuani. Postdepositional osmotic adjustments in sediments from Soap Lake, Washington. Limnology and Oceanography, 18(3):403–414, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Manheim:1975:BRG

[Man75a] F. T. Manheim. Book review: Gibbs, R. J. [Ed.]. 1974, Suspended solids in water. Marine Science Ser. 4. Plenum Press, New York and London, viii + 320 p. $29.50. Typewriter composition. Limnology and Oceanography, 20(3):1056–1057, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Mann:1975:BRP

[Man75b] K. H. Mann. Book review: Perkins, E. J. 1974. The biology of estuaries and coastal waters. Academic Press, London and New York, ix + 678 p. $14.60. Limnology and Oceanography, 20(3): 684–685, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Mann:1976:BRL

[Man76] K. H. Mann. Book review: Lieth, H., and R. H. Whittaker [Eds.]. 1975. Primary productivity of the biosphere. Ecological Studies, v.

14. Springer-Verlag, New York, vi + 339 p. $29.80. Limnology and Oceanography, 21(4):633–634, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Mann:1979:BRR

[Man79] K. H. Mann. Book review: Rzoska, J. 1978. On the nature of rivers, with case stories of the Nile, Zaire and Amazon. Dr. W.

Junk, The Hague. 67 p. f.20; US$10.25. (Softcover). Limnology and Oceanography, 24(6):1177–1178, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Martin:1970:PZR

[Mar70a] John H. Martin. Phytoplankton-zooplankton relationships in Narragansett Bay. IV. The seasonal importance of grazing. Limnology and Oceanography, 15(3):413–418, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Martin:1970:PTT

[Mar70b] John H. Martin. The possible transport of trace metals via moulted copepod exoskeletons. Limnology and Oceanography, 15 (5):756–761, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Martens:1974:MMD

[Mar74] Christopher S. Martens. A method for measuring dissolved gases in pore waters. Limnology and Oceanography, 19(3):525–530, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Marker:1977:SPA

[Mar77] A. F. H. Marker. Some problems arising from the estimation of chlorophyll-a and pheophytin-a in methanol. Limnology and Oceanography, 22(3):578–579, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Margalef:1978:BRD

[Mar78a] R. Margalef. Book review: Daget, J. 1976. Les modèles mathématiques en écologie. Masson, Paris. 172 p. 90F. Limnology and Oceanography, 23(3):575–576, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Marmorino:1978:WBW

[Mar78b] George Marmorino. Warm bottom water in Lake Michigan. Limnology and Oceanography, 23(5):1017–1020, September 1978. CO-DEN LIOCAH. ISSN 0024-3590.

Mason:1970:BRH

[Mas70] David T. Mason. Book review: Holdgate, M. W. [Ed.]. 1970. Antarctic ecology, v. 2. Academic Press, London. x + p. 607–998. £5. Limnology and Oceanography, 15(6):974–975, November 1970. CODEN LIOCAH. ISSN 0024-3590.

May:1973:EOD

[May73] Edwin B. May. Extensive oxygen depletion in Mobile Bay, Alabama. Limnology and Oceanography, 18(3):353–366, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Mayer:1976:CWS

[May76] Lawrence M. Mayer. Chemical water sampling in lakes and sediments with dialysis bags. Limnology and Oceanography, 21(6): 909–912, November 1976. CODEN LIOCAH. ISSN 0024-3590.

McRoy:1970:PAE

[MB70a] C. Peter McRoy and Robert J. Barsdate. Phosphate absorption in eelgrass. Limnology and Oceanography, 15(1):6–13, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Mullin:1970:ECF

[MB70b] Michael M. Mullin and Elaine R. Brooks. The effect of concentration of food on body weight, cumulative ingestion, and rate of growth of the marine copepod Calanus helgolandicus. Limnology and Oceanography, 15(5):748–755, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Maljkovic:1971:PSI

[MB71a] Dubravka Maljkovic and Marko Branica. Polarography of seawater II. Complex formation of cadmium with EDTA. Limnology and Oceanography, 18(5):779–785, September 1971. CODEN LI-OCAH. ISSN 0024-3590.

Mosser:1971:EWT

[MB71b] Jerry L. Mosser and Thomas D. Brock. Effect of wide temperature fluctuation on the blue-green algae of Bead Geyser, Yellowstone National Park. Limnology and Oceanography, 16(4):640–645, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Madigan:1976:QEB

[MB76a] Michael T. Madigan and Thomas D. Brock. Quantitative estimation of bacteriochlorophyll-c in the presence of chlorophyll-a in aquatic environments. Limnology and Oceanography, 21(3):462– 467, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Mullin:1976:SCD

[MB76b] Michael M. Mullin and Elaine R. Brooks. Some consequences of distributional heterogeneity of phytoplankton and zooplankton. Limnology and Oceanography, 21(6):784–796, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Martens:1977:IWC

[MB77] Christopher S. Martens and Robert A. Berner. Interstitial water chemistry of anoxic Long Island Sound sediments. 1. Dissolved gases. Limnology and Oceanography, 22(1):10–25, January 1977. CODEN LIOCAH. ISSN 0024-3590.

McCauley:1979:ZGP

[MB79] Edward McCauley and Frédéric Briand. Zooplankton grazing and phytoplankton species richness: Field tests of the predation hypothesis. Limnology and Oceanography, 24(2):243–252, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Madsen:1973:OUM

[MBB73] Bent Lauge Madsen, Johs. Bengtson, and Ilse Butz. Observations on upstream migration by imagines of some Plecoptera and Ephemeroptera. Limnology and Oceanography, 18(4):678–681, July 1973. CODEN LIOCAH. ISSN 0024-3590.

McRoy:1972:PCE

[MBN72] C. Peter McRoy, Robert J. Barsdate, and Mary Nebert. Phosphorus cycling in an eelgrass (Zostera marina L.) ecosystem. Limnology and Oceanography, 17(1):58–67, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Martens:1978:IWC

[MBR78] Christopher S. Martens, Robert A. Berner, and Jeffrey K. Rosenfeld. Interstitial water chemistry of anoxic Long Island Sound sediments. 2. Nutrient regeneration and phosphate removal. Limnology and Oceanography, 23(4):605–617, July 1978. CODEN LI-OCAH. ISSN 0024-3590.

Morse:1978:DFP

[MC78] John W. Morse and Neil Cook. The distribution and form of phosphorus in North Atlantic Ocean deep-sea and continental slope sediment. Limnology and Oceanography, 23(4):825–830, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Malone:1979:PFP

[MC79] Thomas C. Malone and Mira B. Chervin. The production and fate of phytoplankton size fractions in the plume of the Hudson River, New York Bight. Limnology and Oceanography, 24(4):683–696, July 1979. CODEN LIOCAH. ISSN 0024-3590.

McAlice:1971:PSS

[McA71] Bernard J. McAlice. Phytoplankton sampling with the Sedgwick– Rafter cell. Limnology and Oceanography, 16(1):19–28, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Means:1977:AGF

[MCA77] Jeffrey L. Means, David A. Crerar, and James L. Amster. Application of gel filtration chromatography to evaluation of organometallic interactions in natural waters. Limnology and Oceanography, 22(5):957–965, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Malone:1979:EMS

[MCB79] Thomas C. Malone, Mira B. Chervin, and David C. Boardman. Effects of 22-μ m screens on size-frequency distributions of suspended particles and biomass estimates of phytoplankton size fractions. Limnology and Oceanography, 24(5):956–960, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Mccarthy:1970:UMU

[Mcc70] James J. Mccarthy. A urease method for urea in seawater. Limnology and Oceanography, 15(2):309–313, March 1970. CODEN LIOCAH. ISSN 0024-3590.

McCarthy:1972:UUN

[McC72] James J. McCarthy. The uptake of urea by natural populations of marine phytoplankton. Limnology and Oceanography, 17(5): 738–748, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Mehrbach:1973:MAD

[MCHP73] C. Mehrbach, C. H. Culberson, J. E. Hawley, and R. M. Pytkowicx. Measurement of the apparent dissociation constants of carbonic acid in seawater at atmospheric pressure. Limnology and Oceanography, 18(6):897–907, November 1973. CODEN LIOCAH. ISSN 0024-3590.

McLachlan:1974:RMS

[McL74] A. J. McLachlan. Recovery of the mud substrate and its associated fauna following a dry phase in a tropical lake. Limnology and Oceanography, 19(1):74–83, January 1974. CODEN LIOCAH. ISSN 0024-3590.

McLachlan:1975:RAM

[McL75] A. J. McLachlan. The role of aquatic macrophytes in the recovery of the benthic fauna of a tropical lake after a dry phase. Limnology and Oceanography, 20(1):54–63, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Mcmahon:1973:MFR

[Mcm73] J. W. Mcmahon. Membrane filter retention — a source of error in the ^{14}C method of measuring primary production. Limnology and Oceanography, 18(2):319–324, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Megard:1979:ALD

[MCSK79] R. O. Megard, W. S. Combs, Jr., P. D. Smith, and A. S. Knoll. Attenuation of light and daily integral rates of photosynthesis attained by planktonic algae. Limnology and Oceanography, 24(6): 1038–1050, November 1979. CODEN LIOCAH. ISSN 0024-3590.

McGahee:1971:PRM

[MD71] Cheryl F. McGahee and Graham J. Davis. Photosynthesis and respiration in Myriophyllum spicatum L. as related to salinity. Limnology and Oceanography, 18(5):826–829, September 1971. CO-DEN LIOCAH. ISSN 0024-3590.

Muscatine:1978:URR

[MD78] L. Muscatine and C. F. D’Elia. The uptake, retention, and release of ammonium by reef corals. Limnology and Oceanography, 23(4): 725–734, July 1978. CODEN LIOCAH. ISSN 0024-3590.

McCarthy:1972:CCI

[ME72] James J. McCarthy and Richard W. Eppley. A comparison of chemical, isotopic, and enzymatic methods for measuring nitrogen assimilation of marine phytoplankton. Limnology and Oceanography, 17(3):371–382, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Mullin:1974:UDT

[ME74] Michael M. Mullin and Pamela M. Evans. The use of a deep tank in plankton ecology. 2. Efficiency of a planktonic food chain. Limnology and Oceanography, 19(6):902–911, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Megard:1972:PPP

[Meg72] Robert O. Megard. Phytoplankton, photosynthesis, and phosphorus in Lake Minnetonka, Minnesota. Limnology and Oceanography, 17(1):68–87, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Megard:1975:RCJ

[Meg75] Robert O. Megard. Reply to comment by J. Verduin. Limnology and Oceanography, 20(3):1053–1055, November 1975. CODEN LIOCAH. ISSN 0024-3590. See [Ver75].

Melack:1979:PGS

[Mel79] John M. Melack. Photosynthesis and growth of Spirulina platensis (Cyanophyta) in an equatorial lake (Lake Simbi, Kenya). Limnology and Oceanography, 24(4):753–760, July 1979. CODEN LI-OCAH. ISSN 0024-3590.

McElroy:1978:PRP

[MEW^{+}78] M. B. McElroy, J. W. Elkins, S. C. Wofsy, C. E. Kolb, A. P. Durán, and W. A. Kaplan. Production and release of N_{2}O from the Potomac Estuary. Limnology and Oceanography, 23(6):1168– 1182, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Meyers:1976:DFA

[Mey76] Philip A. Meyers. Dissolved fatty acids in seawater from a fringing reef and a barrier reef at Grand Cayman. Limnology and Oceanography, 21(2):315–319, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Meyer:1979:RSB

[Mey79] Judy L. Meyer. The role of sediments and bryophytes in phosphorus dynamics in a headwater stream ecosystem. Limnology and Oceanography, 24(2):365–375, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Morris:1971:SVD

[MF71] A. W. Morris and P. Foster. The seasonal variation of dissolved organic carbon in the inshore waters of the Menai Strait in relation to primary production. Limnology and Oceanography, 16(6):987– 989, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Maykut:1975:SDL

[MG75] Gary A. Maykut and Thomas C. Grenfell. The spectral distribution of light beneath first-year sea ice in the Arctic Ocean. Limnology and Oceanography, 20(4):554–563, July 1975. CODEN LI-OCAH. ISSN 0024-3590.

Martens:1978:EDT

[MG78] Christopher S. Martens and Martin B. Goldhaber. Early diagenesis in transitional sedimentary environments of the White Oak River Estuary, North Carolina. Limnology and Oceanography, 23 (3):428–441, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Mcroy:1973:NFA

[MGC73] C. Peter Mcroy, John J. Goering, and Barbara Chaney. Nitrogen fixation associated with seagrasses. Limnology and Oceanography, 18(6):998–1002, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Malone:1975:NUG

[MGHR75] T. C. Malone, C. Garside, K. C. Haines, and O. A. Roels. Nitrate uptake and growth of chaetoceros sp. in large outdoor continuous cultures. Limnology and Oceanography, 20(1):9–19, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Motten:1972:EFO

[MH72] Alexander F. Motten and Charles A. S. Hall. Edaphic factors override a possible gradient of ecological maturity indices in a small stream. Limnology and Oceanography, 17(6):922–926, November 1972. CODEN LIOCAH. ISSN 0024-3590.

McLerran:1974:DZC

[MH74] C. J. McLerran and Charles W. Holmes. Deposition of zinc and cadmium by marine bacteria in estuarine sediments. Limnology and Oceanography, 19(6):998–1001, November 1974. CODEN LI-OCAH. ISSN 0024-3590.

Monahan:1975:CVD

[MHK75] Edward C. Monahan, Bruce J. Higgins, and G. Thomas Kaye. A comparison of vertical drift-envelopes to conventional driftbottles. Limnology and Oceanography, 20(1):141–147, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Manheim:1972:SMS

[MHU72] Frank T. Manheim, John C. Hathaway, and Elazar Uchupi. Suspended matter in surface waters of the northern Gulf of Mexico. Limnology and Oceanography, 17(1):17–27, January 1972. CO-DEN LIOCAH. ISSN 0024-3590.

Miller:1970:SEC

[Mil70] Charles B. Miller. Some environmental consequences of vertical migration in marine zooplankton. Limnology and Oceanography, 15(5):727–741, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Mills:1971:DSA

[Mil71] Eric L. Mills. Deep-sea Amphipoda from the western north Atlantic Ocean. The family Ampeliscidae. Limnology and Oceanography, 16(2):357–386, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Miller:1973:BRAa

[Mil73a] Charles B. Miller. Book review: Anikouchine, W. A., and R. W. Sternberg. 1973. The world ocean. An introduction to oceanography. Prentice-Hall, Englewood Cliffs, New Jersey. xi + 338 p, $11.95. Limnology and Oceanography, 18(3):507–508, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Miller:1973:BRAb

[Mil73b] Charles B. Miller. Book review: Anikouchine, W. A., and R. W. Sternberg. 1973. The world ocean. An introduction to oceanography. Prentice-Hall, Englewood Cliffs, New Jersey. xi + 338 p, $11.95. Limnology and Oceanography, 18(3):507, May 1973. CO-DEN LIOCAH. ISSN 0024-3590.

Miller:1973:QPN

[Mil73c] John M. Miller. A quantitative push-net system for transect studies of larval fish and macrozooplankton. Limnology and Oceanography, 18(1):175–178, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Milliman:1977:BRR

[Mil77] John D. Milliman. Book review: Riley, J. P., and R. Chester [Eds.] 1976. Chemical oceanography, v. 5 and 6. 2nd ed. Academic Press, London. V. 5: xiv + 401 p. £,13.50. V. 6: xiv + 414 p. £14.00. Limnology and Oceanography, 22(6):1102–1103, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Miller:1978:BRS

[Mil78] Charles B. Miller. Book review: Steedman, H. F. [Ed.] 1976. Zooplankton fixation and preservation. Monogr. Oceanogr. Methodol.

4. UNESCO Press, Paris. 350 p. $16.50. Limnology and Oceanography, 23(3):583, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Miller:1977:GRM

[MJH77] Charles B. Miller, John Kenneth Johnson, and Donald R. Heinle. Growth rules in the marine copepod genus Acartia. Limnology and Oceanography, 22(2):326–335, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Melack:1974:PRP

[MK74] John M. Melack and Peter Kilham. Photosynthetic rates of phytoplankton in East African alkaline, saline lakes. Limnology and Oceanography, 19(5):743–755, September 1974. CODEN LI-OCAH. ISSN 0024-3590.

Mulholland:1979:OCE

[MK79] Patrick J. Mulholland and Edward J. Kuenzler. Organic carbon export from upland and forested wetland watersheds. Limnology and Oceanography, 24(5):960–966, September 1979. CODEN LIOCAH. ISSN 0024-3590.

McEnery:1970:TSC

[ML70] Marie McEnery and John J. Lee. Tracer studies on calcium and strontium mineralization and mineral cycling in two species of Foraminifera, Rosalina leei and Spiroloculina hyalina. Limnology and Oceanography, 15(2):173–182, March 1970. CODEN LI-OCAH. ISSN 0024-3590.

Millero:1979:DSS

[MLCG79] Frank J. Millero, Antonio Lo Surdo, Peter Chetirkin, and Norman L. Guinasso, Jr. The density and speed of sound of Orca basin waters. Limnology and Oceanography, 24(2):218–225, March 1979. CODEN LIOCAH. ISSN 0024-3590.

McLeod:1975:EVV

[MLKT75] Guy C. McLeod, Kaye V. Ladd, Kenneth Kustin, and David L. Toppen. Extraction of vanadium(V) from seawater by tunicates: a revision of concepts. Limnology and Oceanography, 20(6):491–493, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Monahan:1973:TDD

[MM73a] Edward C. Monahan and Elizabeth A. Monahan. Trends in drogue design. Limnology and Oceanography, 18(6):981–985, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Morcos:1973:CSD

[MM73b] Selim A. Morcos and S. N. Messieh. Circulation and salinity distribution in the southern part of the Suez Canal. Limnology and Oceanography, 18(1):121–130, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Marlowe:1975:PVD

[MM75] Christopher J. Marlowe and Charles B. Miller. Patterns of vertical distribution and migration of zooplankton at Ocean Station “P”. Limnology and Oceanography, 20(4):824–844, September 1975. CODEN LIOCAH. ISSN 0024-3590.

McKnight:1979:RWS

[MM79] Diane M. McKnight and Francois M. M. Morel. Release of weak and strong copper-complexing agents by algae. Limnology and Oceanography, 24(5):823–837, September 1979. CODEN LI-OCAH. ISSN 0024-3590.

McCullough:1979:BSC

[MMC79] Dale A. McCullough, G. Wayne Minshall, and Colbert E. Cushing. Bioenergetics of a stream “collector” organism, Tricorythodes minutus (Insecta: Ephemeroptera). Limnology and Oceanography, 24(1):45–58, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Mickelson:1979:NDD

[MMD79] Michael J. Mickelson, Helmut Maske, and Richard C. Dugdale. Nutrient-determined dominance in multispecies chemostat cultures of diatoms. Limnology and Oceanography, 24(2):298–315, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Manny:1971:UCB

[MMW71] B. A. Manny, M. C. Miller, and R. G. Wetzel. Ultraviolet com.bustion of dissolved organic nitrogen compounds in lake waters. Limnology and Oceanography, 16(1):71–85, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Mynn:1971:CP

[MNM71] Kay Mynn, I. Naught, and N. P. Mackz. Conservation piece. Limnology and Oceanography, 16(6):1003–1004, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Machan:1972:PMC

[MO72] Rudolf Machan and Jörg Ott. Problems and methods of continuous in situ measurements of redox potentials in marine sediments.

Limnology and Oceanography, 17(4):622–626, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Moed:1971:AOA

[Moe71] J. R. Moed. Aluminum oxide as adsorbent for natural watersoluble yellow material. Limnology and Oceanography, 16(1):140– 142, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Mofjeld:1973:BRD

[Mof73] Harold Mofjeld. Book review: Dyer, K. R. 1973. Estuaries: a physical introduction. Wiley-Interscience, New York and London. xv + I40 p. £3.00. Limnology and Oceanography, 18(6):1012, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Monheimer:1975:SUM

[Mon75] Richard H. Monheimer. Sulfate uptake by microplankton communities in western Lake St. Clair. Limnology and Oceanography, 20 (2):183–190, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Monheimer:1978:DIM

[Mon78] Richard H. Monheimer. Difficulties in interpretation of microbial heterotrophy from sulfate uptake data: Laboratory studies. Limnology and Oceanography, 23(1):150–154, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Mortimer:1971:CEB

[Mor71] C. H. Mortimer. Chemical exchanges between sediments and water in the Great Lakes — speculations on probable regulatory mechanisms. Limnology and Oceanography, 16(2):387–404, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Morales:1972:M

[Mor72] Rosario Morales. Models: 1. Limnology and Oceanography, 17(3): 499, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Mortimer:1973:LNM

[Mor73] C. H. Mortimer. The Loch Ness monster — limnology or paralimnology? Limnology and Oceanography, 18(2):343, March 1973. CODEN LIOCAH. ISSN 0024-3590. See reply [SK73a].

Morcos:1975:TSC

[Mor75] Selim A. Morcos. A transitional stage in the current regime in the Suez Canal. Limnology and Oceanography, 20(5):672–679, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Moss:1970:SCT

[Mos70] Brian Moss. Seston composition in two freshwater pools. Limnology and Oceanography, 15(4):504–513, July 1970. CODEN LI-OCAH. ISSN 0024-3590.

Moskowitz:1976:ASV

[Mos76] Paul D. Moskowitz. An analysis of salinity variations within Great South Bay, New York. Limnology and Oceanography, 21(5):740– 742, September 1976. CODEN LIOCAH. ISSN 0024-3590.

MacCaull:1977:DVP

[MP77a] Wendy A. MacCaull and Trevor Platt. Diel variations in the photosynthetic parameters of coastal marine phytoplankton. Limnology and Oceanography, 22(4):723–731, July 1977. CODEN LI-OCAH. ISSN 0024-3590.

Morel:1977:AVO

[MP77b] Anclré Morel and Louis Prieur. Analysis of variations in ocean color. Limnology and Oceanography, 22(4):709–722, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Mayzaud:1978:ITF

[MP78] Patrick Mayzaud and Serge A. Poulet. The importance of the time factor in the response of zooplankton to varying concentrations of naturally occurring particulate matter. Limnology and Oceanography, 23(6):1144–1154, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Marvin:1970:SEF

[MPN70] K. T. Marvin, R. R. Proctor, Jr., and R. A. Neal. Some effects of filtration on the determination of copper in freshwater and saltwater. Limnology and Oceanography, 15(2):320–325, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Marvin:1972:SEF

[MPN72] K. T. Marvin, R. R. Proctor, Jr., and R. A. Neal. Some effects of filtration on the determination of nutrients in fresh and salt water. Limnology and Oceanography, 17(5):777–784, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Meyers:1971:IBF

[MQ71] Philip A. Meyers and James G. Quinn. Interaction between fatty acids and calcite in seawater. Limnology and Oceanography, 16(6): 992–997, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Meyers:1974:MAF

[MQM74] Philip A. Meyers, James G. Quinn, and Nelson Marshall. A method for analysis of fatty acids in coral. Limnology and Oceanography, 19(5):846–848, September 1974. CODEN LI-OCAH. ISSN 0024-3590.

Magnuson:1970:SWS

[MS70] John J. Magnuson and Warren E. Stuntz. A siphon water sampler for use through the ice. Limnology and Oceanography, 15(1):156– 158, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Megard:1974:MRG

[MS74a] Robert O. Megard and Paul D. Smith. Mechanisms that regulate growth rates of phytoplankton in Shagawa Lake, Minnesota. Limnology and Oceanography, 19(2):279–296, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Morel:1974:RBT

[MS74b] André Morel and Raymond C. Smith. Relation between total quanta and total energy for aquatic photosynthesis. Limnology and Oceanography, 19(4):591–600, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Mullin:1975:IPG

[MSF75] Michael M. Mullin, Elizabeth Fuglister Stewart, and Frederick J. Fuglister. Ingestion by planktonic grazers as a function of concentration of food. Limnology and Oceanography, 19(4):259–262, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Millero:1974:TTA

[MSH74] Frank J. Millero, Sheldon R. Schrager, and Lee D. Hansen. Thermometric titration analysis of seawater for chlorinity, sulfate, and alkalinity. Limnology and Oceanography, 20(2):711–715, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Mattson:1974:CMD

[MSJG74] James S. Mattson, Carroll A. Smith, Thomas T. Jones, and Sol M. Gerchakov. Continuous monitoring of dissolved organic matter by UV-visible photometry. Limnology and Oceanography, 19(3):530– 535, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Murphy:1975:EFE

[MSR^{+}75] E. B. Murphy, K. A. Steidinger, B. S. Roberts, J. Williams, and J. W. Jolley, Jr. An explanation for the Florida east coast Gymnodinium breve red tide of November 1972. Limnology and Oceanography, 20(6):481–486, May 1975. CODEN LIOCAH. ISSN 0024-3590.

McCarthy:1977:NNP

[MTT77] James J. McCarthy, W. Rowland Taylor, and Jay L. Taft. Nitrogenous nutrition of the plankton in the Chesapeake Bay. 1. Nutrient availability and phytoplankton preferences. Limnology and Oceanography, 22(6):996–1011, November 1977. CODEN LI-OCAH. ISSN 0024-3590.

Muller:1974:SPP

[Mul74] Pamela Hallock Muller. Sediment production and population biology of the benthic foraminifer Amphistegina madagascariensis. Limnology and Oceanography, 19(5):802–809, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Mullin:1979:DPC

[Mul79] Michael M. Mullin. Differential predation by the carnivorous marine copepod, Tortanus discaudatus. Limnology and Oceanography, 24(4):774–777, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Murphy:1971:CPM

[Mur71] Garth I. Murphy. Clarifying a production model. Limnology and Oceanography, 16(6):981–983, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Murray:1972:TDO

[Mur72] Stephen P. Murray. Turbulent diffusion of oil in the ocean. Limnology and Oceanography, 17(5):651–660, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Murray:1976:LWC

[Mur76] Declan A. Murray. A light weight corer for sampling soft subaqueous deposits. Limnology and Oceanography, 21(2):341–344, March 1976. CODEN LIOCAH. ISSN 0024-3590.

McKinley:1977:TOU

[MW77] Kelton R. McKinley and Robert G. Wetzel. Tritium oxide uptake by algae: an independent measure of phytoplankton photosynthesis. Limnology and Oceanography, 22(2):377–380, March 1977. CODEN LIOCAH. ISSN 0024-3590.

McKinley:1977:MOM

[MWW77] Kelton R. McKinley, Amelia K. Ward, and Robert G. Wetzel. A method for obtaining more precise measures of excreted organic carbon. Limnology and Oceanography, 22(3):570–573, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Moreth:1970:SMD

[MY70] Clarice M. Moreth and Charles S. Yentsch. A sensitive method for the determination of open ocean phytoplankton phycoerythrin pigments by fluorescence. Limnology and Oceanography, 15(2): 313–317, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Morris:1972:NMC

[MY72] Ian Morris and Charles S. Yentsch. A new method for concentrating phytoplankton by filtration with continuous stirring. Limnology and Oceanography, 17(3):490–493, May 1972. CODEN LI-OCAH. ISSN 0024-3590.

Morris:1971:PSR

[MYY71a] Ian Morris, Clarice M. Yentsch, and Charles S. Yentsch. The physiological state with respect to nitrogen of phytoplankton from lownutrient subtropical water as measured by the effect of ammonium ion on dark carbon dioxide fixation. Limnology and Oceanography, 16(6):859–868, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Morris:1971:RBL

[MYY71b] Ian Morris, Clarice M. Yentsch, and Charles S. Yentsch. Relationship between light carbon dioxide fixation and dark carbon dioxide fixation by marine algae. Limnology and Oceanography, 16(6): 854–858, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Naiman:1976:PPS

[Nai76] Robert J. Naiman. Primary production, standing stock, and export of organic matter in a Mohave Desert thermal stream. Limnology and Oceanography, 21(1):60–73, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Natarajan:1970:DSV

[Nat70] K. V. Natarajan. Distribution and significance of vitamin B_{12} and thiamine in the subarctic Pacific Ocean. Limnology and Oceanography, 15(4):655–659, July 1970. CODEN LIOCAH. ISSN 0024-

Nriagu:1976:ESL

[NC76] J. O. Nriagu and R. D. Coker. Emission of sulfur from Lake Ontario sediments. Limnology and Oceanography, 21(4):485–489, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Nriagu:1978:RCD

[NC78] J. O. Nriagu and R. D. Coker. Reply to comments by D. R. Hitchcock. Limnology and Oceanography, 23(2):377–379, March 1978. CODEN LIOCAH. ISSN 0024-3590. See [Hit78].

Nriagu:1979:TPR

[NCK79] J. O. Nriagu, R. D. Coker, and A. L. W. Kemp. Thiosulfate, polythionates, and rhodanese activity in lakes erie and Ontario sediments. Limnology and Oceanography, 24(2):383–389, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Nelson:1978:ASA

[NG78] David M. Nelson and John J. Goering. Assimilation of silicic acid by phytoplankton in the Baja California and northwest Africa upwelling systems. Limnology and Oceanography, 23(3):508–517, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Nichols:1974:STD

[Nic74] Frederic H. Nichols. Sediment turnover by a deposit-feeding polychaete. Limnology and Oceanography, 19(6):945–950, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Nissenbaum:1972:CIE

[NK72] Arie Nissenbaum and I. R. Kaplan. Chemical and isotopic evidence for the In Situ origin of marine humic substances. Limnology and Oceanography, 17(4):570–582, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Neihof:1972:SCP

[NL72] Rex A. Neihof and George I. Loeb. The surface charge of particulate matter in seawater. Limnology and Oceanography, 17(1): 7–16, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Nival:1971:SDS

[NN71] Suzanne Nival and Paul Nival. A simple device for the sorting of living planktonic copepods. Limnology and Oceanography, 16(6): 977–980, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Nival:1976:PRE

[NN76] P. Nival and S. Nival. Particle retention efficiencies of an herbivorous copepod, Acartia clausi (adult and copepodite stages): Effects on grazing. Limnology and Oceanography, 21(1):24–38, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Nival:1979:CPR

[NN79] Paul Nival and Suzanne Nival. Calculations of particle retention efficiency. Limnology and Oceanography, 24(5):995–998, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Neal:1971:TMC

[NND71] Victor T. Neal, Stephen J. Neshyba, and Warren W. Denner. Temperature microstructure in Crater Lake, Oregon. Limnology and Oceanography, 16(4):695–700, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Neal:1972:VTS

[NND72] Victor T. Neal, Stephen J. Neshyba, and Warren W. Denner. Vertical temperature structure in Crater Lake, Oregon. Limnology and Oceanography, 17(3):451–454, May 1972. CODEN LIOCAH. ISSN 0024-3590.

North:1975:PAC

[Nor75] Barbara B. North. Primary amines in California coastal waters: Utilization by phytoplankton. Limnology and Oceanography, 20 (1):20–27, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Northby:1976:CRM

[Nor76] J. A. Northby. A comment on rate measurements in open systems. Limnology and Oceanography, 21(1):180–182, January 1976. CO-DEN LIOCAH. ISSN 0024-3590.

Nriagu:1978:DSP

[Nri78] Jerome O. Nriagu. Dissolved silica in pore waters of Lakes Ontario, Erie, and Superior sediments. Limnology and Oceanography, 23 (1):53–67, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Naiman:1978:TNC

[NS78] Robert J. Naiman and John R. Sibert. Transport of nutrients and carbon from the Nanaimo River to its estuary. Limnology and Oceanography, 23(6):1183–1193, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Noble:1970:ATS

[NW70] Vincent E. Noble and John C. Wilkerson. Airborne temperature surveys of Lake Michigan, October 1966 and 1967. Limnology and Oceanography, 15(2):289–295, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Noshkin:1975:TOR

[NWEG75] V. E. Noshkin, K. M. Wong, R. J. Eagle, and C. Gatrousis. Transuranics and other radionuclides in Bikini Lagoon: Concentration data retrieved from aged coral sections. Limnology and Oceanography, 20(4):729–742, September 1975. CODEN LI-OCAH. ISSN 0024-3590.

OConnor:1971:BRS

[O’C71] John T. O’Connor. Book review: Stumm, W., and J. J. Morgan. 1970. Aquatic chemistry. Wiley-Interscience, New York, xv + 583 p. $24.95. Limnology and Oceanography, 16(3):595–596, May 1971. CODEN LIOCAH. ISSN 0024-3590.

OConnor:1974:ASG

[O’C74] J. F. O’Connor. An apparatus for sampling gravel substrate in streams. Limnology and Oceanography, 19(6):1007–1011, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Ogura:1970:PDO

[Ogu70] Norio Ogura. On the presence of 0.1–0.5-μ dissolved organic matter in seawater. Limnology and Oceanography, 15(3):476–479, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Otsuki:1972:PDOa

[OH72a] Akira Otsuki and Takahisa Hanya. Production of dissolved organic matter from dead green algal cells. I. Aerobic microbial decomposition. Limnology and Oceanography, 17(2):248–257, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Otsuki:1972:PDOb

[OH72b] Akira Otsuki and Takahisa Hanya. Production of dissolved organic matter from dead green algal cells. II. Anaerobic microbial decomposition. Limnology and Oceanography, 17(2):258–264, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Osborn:1974:SSF

[OL74] Thomas R. Osborn and Paul H. LeBlond. Static stability in freshwater lakes. Limnology and Oceanography, 19(3):544–545, May 1974. CODEN LIOCAH. ISSN 0024-3590. See comments [LR74b].

Omori:1979:GFM

[Omo79] M. Omori. Growth, feeding, and mortality of larval and early postlarval stages of the oceanic shrimp Sergestes similis Hansen. Limnology and Oceanography, 24(2):273–288, March 1979. CO-DEN LIOCAH. ISSN 0024-3590.

Oremland:1979:MAP

[Ore79] Ronald S. Oremland. Methanogenic activity in plankton samples and fish intestines a mechanism for in situ methanogenesis in oceanic surface waters. Limnology and Oceanography, 24(6): 1136–1141, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Orians:1973:BRD

[Ori73] Gordon H. Orians. Book review: Deevey, E. S. [Ed.]. 1972. Growth by intussusception. Ecological essays in honor of G. Evelyn Hutchinson. Trans. Conn. Acad. Arts Sci. 44. Archon Books, The Shoestring Press, Hamden, Conn. 443 p. $25.00. Limnology and Oceanography, 18(2):347–348, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Orians:1977:BRM

[Ori77] Gordon Orians. Book review: Margalef, R. 1975. Ecología. Ediciones Omega, S.A., Barcelona. $41.00. Limnology and Oceanography, 22(6):1100–1102, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Oglesby:1978:PLL

[OS78] R. T. Oglesby and W. R. Schaffner. Phosphorus loadings to lakes and some of their responses. Part 2. Regression models of summer phytoplankton standing crops, winter total P, and transparency of New York lakes with known phosphorus loadings. Limnology and Oceanography, 23(1):135–145, January 1978. CODEN LIOCAH. ISSN 0024-3590.

OBrien:1979:ABM

[OS79] W. John O’Brien and David Schmidt. Arctic Bosmina morphology and copepod predation. Limnology and Oceanography, 24(3):564– 568, May 1979. CODEN LIOCAH. ISSN 0024-3590.

OConnors:1976:PSM

[OSD76] Harold B. O’Connors, Lawrence F. Small, and Percy L. Donaghay. Particle-size modification by two size classes of the estuarine copepod Acartia clausi. Limnology and Oceanography, 21(2):300–308, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Ohwada:1972:VBT

[OT72] Kouichi Ohwada and Nobuo Taga. Vitamin B_{12}, thiamine, and biotin in Lake Sagami. Limnology and Oceanography, 17(2):315– 320, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Oremland:1977:DFC

[OT77] Ronald S. Oremland and Barrie F. Taylor. Diurnal fluctuations of O_{2}, N_{2}, and CH_{4} in the rhizosphere of Thalassia testudinum. Limnology and Oceanography, 22(3):566–570, May 1977. CODEN LIOCAH. ISSN 0024-3590.

OBrien:1978:PPE

[OV78] W. John O’Brien and Gary L. Vinyard. Polymorphism and predation: The effect of invertebrate predation on the distribution of two varieties of Daphnia carinata in South India ponds. Limnology and Oceanography, 23(3):452–460, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Otsuki:1972:CPC

[OW72] Akira Otsuki and Robert G. Wetzel. Coprecipitation of phosphate with carbonates in a marl lake. Limnology and Oceanography, 17 (5):763–767, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Otsuki:1973:IYO

[OW73] Akira Otsuki and Robert G. Wetzel. Interaction of yellow organic acids with calcium carbonate in freshwater. Limnology and Oceanography, 18(3):490–493, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Otsuki:1974:RDO

[OW74] Akira Otsuki and Robert G. Wetzel. Release of dissolved organic matter by autolysis of a submersed macrophyte, Scirpus subterminalis. Limnology and Oceanography, 19(5):842–845, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Owen:1974:OEA

[Owe74] Robert W. Owen, Jr. Optically effective area of particle ensembles in the sea. Limnology and Oceanography, 19(4):584–590, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Palmer:1970:SEM

[PA70] Douglas S. Palmer and Lawrence J. Albright. Salinity effects on the maximum hydrostatic pressure for growth of the marine psychrophilic bacterium, Vibrio marinus. Limnology and Oceanography, 15(3):343–347, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Pytkowicz:1975:BIS

[PA75] R. M. Pytkowicz and E. Atlas. Buffer intensity of seawater. Limnology and Oceanography, 19(4):222–229, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Packard:1973:LDN

[Pac73] Theodore T. Packard. The light dependence of nitrate reductase in marine phytoplankton. Limnology and Oceanography, 18(3): 466–469, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Paeri:1974:BUD

[Pae74] Hans W. Paeri. Bacterial uptake of dissolved organic matter in relation to detrital aggregation in marine and freshwater systems. Limnology and Oceanography, 19(6):966–972, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1978:MOC

[Pae78] Hans W. Paerl. Microbial organic carbon recovery in aquatic ecosystems. Limnology and Oceanography, 23(5):927–935, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Paffenhofer:1976:FGF

[Paf76] Gustav-Adolf Paffenhofer. Feeding, growth, and food conversion of the marine planktonic copepod Calanus helgolandicus. Limnology and Oceanography, 21(1):39–50, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Paine:1970:BRS

[Pai70] R. T. Paine. Book review: Spinner, G. P. 1969. The wildlife wetlands and shellfish areas of the Atlantic Coastal Zone. Serial Atlas of the Marine Environment. Folio 18. American Geographical Society, New York. 4 p. + 12 Plates. $12.00. Limnology and Oceanography, 15(1):170, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Paine:1971:EFN

[Pai71] Robert T. Paine. Energy flow in a natural population of the herbivorous gastropod Tegula funerralis. Limnology and Oceanography, 16(1):86–98, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Paine:1975:BRM

[Pai75a] R. T. Paine. Book review: Mariscal, R. N. [Ed.] 1974. Experimental marine biology. Academic Press, New York, x + 373 p. $18.50. Limnology and Oceanography, 19(4):303, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Paine:1975:BRT

[Pai75b] R. T. Paine. Book review: Tasch, P. 1973. Paleobiology of the invertebrates. Data retrieval from the fossil record. Wiley and Sons, New York. 946 p. Limnology and Oceanography, 19(4):303, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Paloheimo:1974:CIB

[Pal74] Jyri E. Paloheimo. Calculation of instantaneous birth rate. Limnology and Oceanography, 20(2):692–694, July 1974. CODEN LI-OCAH. ISSN 0024-3590.

Pamatmat:1971:OCS

[Pam71] Mario M. Pamatmat. Oxygen consumption by the seabed IV. Shipboard and laboratory experiments. Limnology and Oceanography, 16(3):536–550, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Parsons:1970:BRV

[Par70] T. R. Parsons. Book review: Vollenweider, R. A. [Ed.]. 1969. A manual on methods for measuring primary production in aquatic environments. IBP Handbook No. 12. F. A. Davis Co., Philadelphia, Penn. 213 p. $8.90. (Also available from Blackwell Sci. Publ., Oxford. 45s.). Limnology and Oceanography, 15(1):168, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Parsons:1972:BRC

[Par72] T. R. Parsons. Book review: Costlow, J. D. [ed.]. 1971. Fertility of the sea. Gordon and Breach, New York, Vol. 1, xi + p. 1–308.

$19.50. Vol. 2, xi + p. 309–622. $14.50. Limnology and Oceanography, 17(5):799–800, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Parsons:1973:BRM

[Par73] T. R. Parsons. Book review: Martin, D. F. 1972. Marine chemistry, v. 1. Analytical methods. Marcel Dekker, New York. xi + 389 p. $9.50. Limnology and Oceanography, 18(1):181, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Parsons:1975:BRH

[Par75] T. R. Parsons. Book review: Hanson, J. A. [Ed.]. 1974. Open sea mariculture. Perspectives, problems, and prospects. Dowden, Hutchinson & Ross, Inc. Strouds-berg, Penn, xi + 410 p. $24.00. Limnology and Oceanography, 20(3):683–684, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Parker:1977:VBL

[Par77] Michael Parker. Vitamin B_{12} in Lake Washington, USA: Concentration and rate of uptake. Limnology and Oceanography, 22(3): 527–538, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Parsons:1978:BRB

[Par78] T. R. Parsons. Book review: Bougis, P. [Ed.] 1976 Océanographie biologique appliquee: L’Exploitation de la vie marine. Masson, Paris, New York, Barcelona and Milan, xii + 320 p. Limnology and Oceanography, 23(5):1086, September 1978. CODEN LI-OCAH. ISSN 0024-3590.

Patrick:1971:EIL

[Pat71] Ruth Patrick. The effects of increasing light and temperature on the structure of diatom communities. Limnology and Oceanography, 16(2):405–421, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Patalas:1980:CRB

[Pat80] K. Patalas. Comment on “A relation between lake morphometry and primary productivity and its use in interpreting whole-lake eutrophication experiments” (E. J. Fee). Limnology and Oceanography, 25(6):1147–1149, November 1980. CODEN LIOCAH. ISSN 0024-3590. See [Fee79].

Pamatmat:1973:AML

[PB73] Mario M. Pamatmat and Ashok M. Bhagwat. Anaerobic metabolism in Lake Washington sediments. Limnology and Oceanography, 18(4):611–627, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Pickett:1977:ORC

[PB77] R. L. Pickett and S. Bermick. Observed resultant circulation of Lake Ontario. Limnology and Oceanography, 22(6):1071–1076, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Pak:1970:LSV

[PBHC70] Hasong Pak, George F. Beardsley, Jr., G. Ross Heath, and Herbert Curl. Light-scattering vectors of some marine particles. Limnology and Oceanography, 15(5):683–687, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Prochazkova:1970:CCI

[PBK70] L. Procházková, P. Blažka, and M. Králová. Chemical changes involving nitrogen metabolism in water and particulate matter during primary production experiments. Limnology and Oceanography, 15(5):797–807, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Peterson:1974:PSR

[PBS74] Bruce J. Peterson, John P. Barlow, and Alice E. Savage. The physiological state with respect to phosphorus of Cayuga Lake phytoplankton. Limnology and Oceanography, 19(3):396–408, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Poisson:1976:PMV

[PC76] Alain Poisson and Jacques Chanu. Partial molal volumes of some major ions in seawater. Limnology and Oceanography, 21(6):853– 861, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Park:1970:NCD

[PCWR70] P. Kilho Park, Magdalena Catalfomo, George R. Webster, and Barry H. Reid. Nutrients and carbon dioxide in the Columbia River. Limnology and Oceanography, 15(1):70–79, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Patterson:1978:NMC

[PDF^{+}78] R. J. Patterson, L. Dijkes, S. Frape, R. A. McLeod, and J. Wickens. A new method for collecting water from beneath the ice.

Limnology and Oceanography, 23(5):1029, September 1978. CO-DEN LIOCAH. ISSN 0024-3590.

Parker:1976:CDF

[PE76] Jon I. Parker and David N. Edgington. Concentration of diatom frustules in Lake Michigan sediment cores. Limnology and Oceanography, 21(6):887–893, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Pearre:1979:ASV

[Pea79] Sifford Pearre, Jr. On the adaptive significance of vertical migration. Limnology and Oceanography, 24(4):781–782, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Pechlaner:1970:PSO

[Pec70] Roland Pechlaner. The phytoplankton spring outburst and its conditions in Lake Erken (Sweden). Limnology and Oceanography, 15(1):113–130, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Peckarsky:1979:BID

[Pec79] Barbara L. Peckarsky. Biological interactions as determinants of distributions of benthic invertebrates within the substrate of stony streams. Limnology and Oceanography, 24(1):59–68, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Pedrick:1974:NES

[Ped74] Robert A. Pedrick. Nonmetallic elutriation and sieving device for benthic macrofauna. Limnology and Oceanography, 19(3):535–538, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Pennak:1973:WCR

[Pen73] Robert W. Pennak. Water chemistry and the reliability of electronic field instruments. Limnology and Oceanography, 16(5):811– 812, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Perry:1976:PUO

[Per76] M. J. Perry. Phosphate utilization by an oceanic diatom in phosphorus-limited chemostat culture and in the oligotrophic waters of the central North Pacific. Limnology and Oceanography, 21 (1):88–107, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Peters:1971:EWV

[Pet71] Robert Peters. Ecology and the world view. Limnology and Oceanography, 16(1):143–148, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Peters:1975:PEM

[Pet75] Robert Henry Peters. Phosphorus excretion and the measurement of feeding and assimilation by zooplankton. Limnology and Oceanography, 20(4):858–859, September 1975. CODEN LI-OCAH. ISSN 0024-3590.

Peterka:1978:RRL

[Pet78a] John J. Peterka. Relationship of rainfall and lake groundwater seepage. Limnology and Oceanography, 23(4):821–825, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Peterson:1978:RUR

[Pet78b] Bruce J. Peterson. Radiocarbon uptake: Its relation to net particulate carbon production. Limnology and Oceanography, 23(1): 179–184, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Platt:1973:SVP

[PF73] Trevor Platt and Christiane Filion. Spatial variability of the productivity: biomass ratio for phytoplankton in a small marine basin. Limnology and Oceanography, 18(5):743–749, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1972:HAD

[PG72] Hans W. Paerl and Charles R. Goldman. Heterotrophic assays in the detection of water masses at Lake Tahoe, California. Limnology and Oceanography, 17(1):145–148, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Pasciak:1974:TLN

[PG74] Walter J. Pasciak and Jerome Gavis. Transport limitation of nutrient uptake in phytoplankton. Limnology and Oceanography, 19 (6):881–888, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Pasciak:1975:TLN

[PG75] Walter J. Pasciak and Jerome Gavis. Transport limited nutrient uptake rates in Ditylum brightwellii. Limnology and Oceanography, 20(6):604–617, July 1975. CODEN LIOCAH. ISSN 0024-

Pollock:1971:CCI

[PH71] Leland W. Pollock and William D. Hummon. Cyclic changes in interstitial water content, atmospheric exposure, and temperature in a marine beach. Limnology and Oceanography, 16(3):522–535, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Pytkowicz:1974:BCI

[PH74] R. M. Pytkowicz and J. E. Hawley. Bicarbonate and carbonate ion-pairs and a model of seawater at 25^{o}C. Limnology and Oceanography, 19(2):223–234, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Peterson:1978:DGN

[PHH78] Bruce J. Peterson, John E. Hobbie, and James F. Haney. Daphnia grazing on natural bacteria. Limnology and Oceanography, 23(5): 1039–1044, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Phleger:1970:FPM

[Phl70] Fred B. Phleger. Foraminiferal populations and marine marsh processes. Limnology and Oceanography, 15(4):522–534, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Phleger:1973:BRM

[Phl73] Fred B. Phleger. Book review: Murray, J. W. 1973. Distribution and ecology of living benthic foraminiferids. Crane Russak & Co., New York. xiii + 274 p. $24.75. Limnology and Oceanography, 18 (6):1011, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Packard:1971:VDA

[PHR71] T. T. Packard, M. L. Healy, and F. A. Richards. Vertical distribution of the activity of the respiratory electron transport system in marine plankton. Limnology and Oceanography, 16(1):60–70, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Palmer:1972:LMP

[PI72] M. D. Palmer and J. Bryan Izatt. Lake movements with partial ice cover. Limnology and Oceanography, 17(3):403–409, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Platt:1973:CCP

[PI73] Trevor Platt and Brian Irwin. Caloric content of phytoplankton. Limnology and Oceanography, 18(2):306–310, March 1973. CO-DEN LIOCAH. ISSN 0024-3590.

Pickett:1976:LOC

[Pic76] R. L. Pickett. Lake Ontario circulation in November. Limnology and Oceanography, 21(4):608–611, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Pilson:1974:AUR

[Pil74] Michael E. Q. Pilson. Arsenate uptake and reduction by Pocillopora verrucosa. Limnology and Oceanography, 19(2):339–341, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Pytkowicz:1974:IAE

[PIM74] R. M. Pytkowicz, S. E. Ingle, and C. Mehrbach. Invariance of apparent equilibrium constants with pH. Limnology and Oceanography, 20(2):665–669, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Pingree:1973:CLS

[Pin73] R. D. Pingree. A component of Labrador sea water in the Bay of Biscay. Limnology and Oceanography, 18(5):711–718, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Pearcy:1974:RSW

[PK74] W. G. Pearcy and Donald F. Keene. Remote sensing of water color and sea surface temperatures off the Oregon coast. Limnology and Oceanography, 19(4):573–583, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Pearcy:1977:BTZ

[PKC77] W. G. Pearcy, E. E. Krygier, and N. H. Cutshall. Biological transport of zinc-65 into the deep sea. Limnology and Oceanography, 22(5):846–855, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Peters:1973:CSP

[PL73] Robert Peters and David Lean. The characterization of soluble phosphorus released by limnetic zooplankton. Limnology and Oceanography, 18(2):270–279, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Platford:1978:BRL

[Pla78] R. F. Platford. Book review: Leyendekkers, J. V. 1976. Thermodynamics of seawater. Part 1: Entropy, volume, expansibility, compressibility. Mar. Sci. Ser. 3. Marcel Dekker, Inc., New York. xi + 496 p. $24.50. Limnology and Oceanography, 23(3):577, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1977:CSD

[PM77] Hans W. Paerl and Lincoln A. Mackenzie. A comparative study of the diurnal carbon fixation patterns of nannoplankton and net plankton. Limnology and Oceanography, 22(4):732–738, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Prager:1973:BMC

[PME73] Earl W. Davey Prager, Myra J. Morgan, and Stanton J. Erickson. A biological measurement of the copper complexation capacity of seawater. Limnology and Oceanography, 18(6):993–997, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Page:1979:DHI

[PN79] T. L. Page and D. A. Neitzel. A device to hold and identify rock-filled baskets for benthic sampling in large rivers. Limnology and Oceanography, 24(5):988–990, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Pomeroy:1974:BRU

[Pom74] Lawrence R. Pomeroy. Book review: UNESCO. 1973. A guide to the measurement of marine primary production under some special conditions. Monogr. Oceanogr. Method. 3. UNESCO, Paris. 73 p. $3.00. Limnology and Oceanography, 19(1):170, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Pomeroy:1979:BRG

[Pom79] Lawrence R. Pomeroy. Book review: Good, R. E., D. F. Whigham, and R. L. Simpson [Eds.] 1978. Freshwater wetlands (Ecological Processes and management potential). Academic Press, New York, xvii + 378 p. $17.50. Limnology and Oceanography, 24(4):796, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Porter:1972:MLS

[Por72] Karen Glaus Porter. A method for the In Situ study of zooplankton grazing effects on algal species composition and standing crop. Limnology and Oceanography, 17(6):913–917, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Poulet:1973:GPM

[Pou73] S. A. Poulet. Grazing of Pseudocalanus minutus on naturally occurring particulate matter. Limnology and Oceanography, 18 (4):564–573, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Poulet:1978:CBF

[Pou78] S. A. Poulet. Comparison between five coexisting species of marine copepods feeding on naturally occurring particulate matter. Limnology and Oceanography, 23(6):1126–1143, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Palmer:1976:HHP

[PP76] Mew D. Palmer and Donald J. Poulton. Hamilton harbor: Periodicities of the physicochemical process. Limnology and Oceanography, 21(1):118–127, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Porter:1977:QSD

[PP77] James W. Porter and Karen Glaus Porter. Quantitative sampling of demersal plankton migrating from different coral reef substrates. Limnology and Oceanography, 22(3):553–556, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Peters:1973:PRD

[PR73] R. H. Peters and F. H. Rigler. Phosphorus release by daphnia. Limnology and Oceanography, 18(6):821–839, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Prepas:1978:EDD

[PR78] E. Prepas and F. H. Rigler. The enigma of Daphnia death rates. Limnology and Oceanography, 23(5):970–988, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Pravdic:1970:SCC

[Pra70] V. Pravdić. Surface charge characterization of sea sediments. Limnology and Oceanography, 15(2):230–233, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Prepas:1978:SFD

[Pre78] E. Prepas. Sugar-frosted Daphnia: an improved fixation technique for cladocera. Limnology and Oceanography, 23(3):557–559, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Price:1978:CDO

[PRG78] C. A. Price, E. M. Reardon, and R. R. L. Guillard. Collection of dinoflagellates and other marine microalgae by centrifugation in density gradients of a modified silica sol. Limnology and Oceanography, 23(3):548–553, May 1978. CODEN LIOCAH. ISSN 0024-

Prakash:1973:IHS

[PRJR73] A. Prakash, M. A. Rashid, Arne Jensen, and D. V. Subha Rao. Influence of humic substances on the growth of marine phytoplankton: diatoms. Limnology and Oceanography, 18(4):516–524, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1975:SNC

[PRLG75] Hans W. Paerl, Robert C. Richards, Robert L. Leonard, and Charles R. Goldman. Seasonal nitrate cycling as evidence for complete vertical mixing in Lake Tahoe, California-nevada. Limnology and Oceanography, 20(1):1–8, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Proctor:1971:CGT

[Pro71] Vernon W. Proctor. Chara globularis Thuillier (= C. vragilis Desvaux): breeding patterns within a cosmopolitan complex. Limnology and Oceanography, 16(2):422–436, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Pearl:1973:PFP

[PS73] Hans W. Pearl and Susan L. Shimp. Preparation of filtered plankton and detritus for study with scanning electron microscopy. Limnology and Oceanography, 18(5):802–805, September 1973. CO-DEN LIOCAH. ISSN 0024-3590.

Penhale:1977:EDO

[PS77] Polly A. Penhale and Walker O. Smith, Jr. Excretion of dissolved organic carbon by eelgrass (Zostera marina) and its epiphytes. Limnology and Oceanography, 22(3):400–407, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1979:DGD

[PS79] Hans W. Paerl and Elisabeth Ann Stull. In defense of grain density autoradiography. Limnology and Oceanography, 24(6):1166–1169, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Peters:1978:CCN

[PSK78] K. E. Peters, R. E. Sweeney, and I. R. Kaplan. Correlation of carbon and nitrogen stable isotope ratios in sedimentary organic matter. Limnology and Oceanography, 23(4):598–604, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Pyle:1970:SVZ

[PT70] Thomas E. Pyle and Thomas T. Tieh. Strontium, vanadium, and zinc in the shells of pteropods. Limnology and Oceanography, 15 (1):153–154, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Parsons:1973:ECP

[PT73] T. R. Parsons and M. Takahashi. Environmental control of phytoplankton cell size. Limnology and Oceanography, 18(4):511–515, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Parsons:1974:RCH

[PT74] T. R. Parsons and M. Takahashi. A rebuttal to the comment by hecky and kilham. Limnology and Oceanography, 19(2):366–368, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Parsons:1975:RCM

[PT75] T. R. Parsons and M. Takahashi. A response to the comment by malone. Limnology and Oceanography, 20(6):490, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Pugh:1970:LSC

[Pug70] P. R. Pugh. Liquid scintillation counting of ^{14}C-diatom material on filter papers for use in productivity studies. Limnology and Oceanography, 15(4):652–655, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Pugh:1973:ELS

[Pug73] P. R. Pugh. An evaluation of liquid scintillation counting techniques for use in aquatic primary production studies. Limnology and Oceanography, 18(2):310–319, March 1973. CODEN LI-OCAH. ISSN 0024-3590.

Pugh:1977:GGB

[Pug77] D. T. Pugh. Geothermal gradients in British lake sediments. Limnology and Oceanography, 22(4):581–596, July 1977. CODEN LI-OCAH. ISSN 0024-3590.

Pytkowicx:1971:AOU

[Pyt71] Ricardo M. Pytkowicx. On the apparent oxygen utilization and the preformed phosphate in the oceans. Limnology and Oceanography, 16(1):39–42, January 1971. CODEN LIOCAH. ISSN 0024-

Pytkowicz:1972:CCS

[Pyt72] Ricardo M. Pytkowicz. Comments on “The control of seawater pH by ion pairing” (P. J. Wangersky). Limnology and Oceanography, 17(6):958–959, November 1972. CODEN LIOCAH. ISSN 0024-3590. See [Wan72a].

Pytkowicz:1975:ACB

[Pyt75] R. M. Pytkowicz. Activity coefficients of bicarbonates and carbonates in seawater. Limnology and Oceanography, 20(4):971–975, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Pytkowicz:1976:BRH

[Pyt76a] R. M. Pytkowicz. Book review: Hissel, J. 1975. La chimie des eaux. La practique du calcul des équilibres. Centre Beige d’etude et de Documentation des Eaux (CEBADEAU), Liége. 209 p. Limnology and Oceanography, 21(1):185, January 1976. CODEN LI-OCAH. ISSN 0024-3590.

Pytkowicz:1976:BRR

[Pyt76b] Ricardo M. Pytkowicz. Book review: Riley, J. P., and G. Skirrow [Eds.]. 1975. Chemical oceanography, v. 1. 2nd ed. Academic Press, New York and London, xx + 606 p. $49.00. £18.50. Limnology and Oceanography, 21(2):345–347, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Quinn:1971:RDO

[QM71] James G. Quinn and Philip A. Meyers. Retention of dissolved organic acids in seawater by various filters. Limnology and Oceanography, 16(1):129–131, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Quinn:1970:MMU

[QPB70] William H. Quinn, Daniel A. Panshin, and Wayne V. Burt. Method for measuring upper-level winds at sea from small ships. Limnology and Oceanography, 15(5):836–838, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Qasim:1970:PPO

[QS70] S. Z. Qasim and V. N. Sankaranarayanan. Production of particulate organic matter by the reef on Kavaratti Atoll (Laccadives). Limnology and Oceanography, 15(4):574–578, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Richards:1971:CPO

[RAC71] Francis A. Richards, James J. Andwso, and Joel D. Cline. Chemical and physical observations in Golfo, Dulce, an anoxic basin on the Pacific coast of Costa Rica. Limnology and Oceanography, 16 (1):43–50, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Rodel:1977:SHA

[RAH77] M. G. Rodel, D. E. Armstrong, and R. F. Harris. Sorption and hydrolysis of added organic phosphorus compounds in lake sediments. Limnology and Oceanography, 22(3):415–422, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Ramsay:1976:EAM

[Ram76] Angela J. Ramsay. The effect of acidification on measurements of microbial uptake of radioactive glucose in freshwater. Limnology and Oceanography, 21(6):922–926, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Ratzlaff:1974:SMA

[Rat74] Willis Ratzlaff. Swarming in Moina affinis. Limnology and Oceanography, 19(6):993–995, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Rau:1978:CNP

[Rau78] Greg H. Rau. Conifer needle processing in a subalpine lake. Limnology and Oceanography, 23(2):356–358, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Rusness:1970:ARN

[RB70] Donald Rusness and R. N. Burris. Acetylene reduction (nitrogen fixation) in Wisconsin lakes. Limnology and Oceanography, 15(5): 808–813, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Richards:1971:CCI

[RB71] F. A. Richards and W. W. Broenkow. Chemical changes, including nitrate reduction, in Darwin Bay, Galapagos Archipelago, over a 2-month period, 1969. Limnology and Oceanography, 18(5):758– 765, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Ramm:1974:SPA

[RB74] Alan E. Ramm and David A. Bella. Sulfide production in anaerobic microcosms. Limnology and Oceanography, 19(1):110–118, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Rhoads:1971:SPC

[RC71] Donald C. Rhoads and Steven Cande. Sediment profile camera for In Situ study of organism–sediment relations. Limnology and Oceanography, 16(1):110–114, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Roff:1972:LCS

[RC72] John C. Roff and John C. H. Cartel. Life cycle and seasonal abundance of the copepod Limnocalanus macrurus Sars in a High Arctic lake. Limnology and Oceanography, 17(3):363–370, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Rigler:1974:UFD

[RC74] F. H. Rigler and J. M. Cooley. The use of field data to derive population statistics of multivoltine copepods. Limnology and Oceanography, 20(2):636–655, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Raymond:1976:BMS

[RD76] James A. Raymond and Arthur L. DeVries. Bioluminescence in McMurdo Sound, Antarctica. Limnology and Oceanography, 21 (4):599–602, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Reed:1973:FIC

[RE73] R. K. Reed and William P. Elliott. Freshwater input to coastal waters off the Pacific northwest. Limnology and Oceanography, 18 (4):683–686, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Redfield:1975:PCP

[Red75] Alfred C. Redfield. A piston corer for peat. Limnology and Oceanography, 20(3):1042–1045, November 1975. CODEN LI-OCAH. ISSN 0024-3590.

Rees:1974:DDS

[Ree74] Colin P. Rees. A device for determining substratum selection under flowing-water conditions. Limnology and Oceanography, 19(6): 1012–1014, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Reimold:1972:MPT

[Rei72a] Robert J. Reimold. The movement of phosphorus through the salt marsh cord grass, Spartina alterniflora Loisel. Limnology and Oceanography, 17(4):606–611, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Reiswig:1972:SPO

[Rei72b] Henry M. Reiswig. The spectrum of particulate organic matter of shallow-bottom boundary waters of Jamaica. Limnology and Oceanography, 17(3):341–348, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Remsen:1971:DUC

[Rem71] Charles C. Remsen. The distribution of urea in coastal and oceanic waters. Limnology and Oceanography, 18(5):732–740, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Reynolds:1978:PIC

[Rey78] Robert C. Reynolds, Jr. Polyphenol inhibition of calcite precipitation in Lake Powell. Limnology and Oceanography, 23(4):585–597, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Rudd:1976:FCM

[RFFH76] John W. M. Rudd, Akira Furutani, R. J. Flett, and R. D. Hamilton. Factors controlling methane oxidation in shield lakes: The role of nitrogen fixation and oxygen concentration. Limnology and Oceanography, 21(3):357–364, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Robbins:1976:SEE

[RG76a] John A. Robbins and John Gustinis. A squeezer for efficient extraction of pore water from small volumes of anoxic sediment. Limnology and Oceanography, 21(6):905–909, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Roger:1976:PFW

[RG76b] Claude Roger and Rene Grandperrin. Pelagic food webs in the tropical Pacific. Limnology and Oceanography, 21(5):731–735, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Rudd:1975:TSM

[RH75] John W. M. Rudd and R. D. Hamilton. Two samplers for monitoring dissolved gases in lake water and sediments. Limnology and Oceanography, 20(4):902–906, September 1975. CODEN LI-OCAH. ISSN 0024-3590.

Reed:1976:OCU

[RH76] R. K. Reed and D. Halpern. Observations of the California undercurrent off Washington and Vancouver Island. Limnology and Oceanography, 21(3):389–398, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Reeburgh:1977:MMC

[RH77] William S. Reeburgh and David T. Heggie. Microbial methane consumption reactions and their effect on methane distributions in freshwater and marine environments. Limnology and Oceanography, 22(1):1–9, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Rudd:1978:MCE

[RH78] John W. M. Rudd and R. D. Hamilton. Methane cycling in a eutrophic shield lake and its effects on whole lake metabolism. Limnology and Oceanography, 23(2):337–348, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Rudd:1974:MMO

[RHC74] John W. M. Rudd, R. D. Hamilton, and N. E. R. Campbell. Measurement of microbial oxidation of methane in lake water. Limnology and Oceanography, 19(3):519–524, May 1974. CODEN LI-OCAH. ISSN 0024-3590.

Rhee:1972:CBA

[Rhe72] G-Yull Rhee. Competition between an alga and an aquatic bacterium for phosphate. Limnology and Oceanography, 17(4):505– 514, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Rhee:1978:EPA

[Rhe78] G-Yull Rhee. Effects of N : P atomic ratios and nitrate limitation on algal growth, cell composition, and nitrate uptake. Limnology and Oceanography, 23(1):10–25, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Robinson:1973:RBH

[RHG73] Gordon G. C. Robinson, Leonard L. Hendzel, and Douglas C. Gillespie. A relationship between heterotrophic utilization of organic acids and bacterial populations in West Blue Lake, Manitoba. Limnology and Oceanography, 18(2):264–269, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Radford:1971:SSS

[RHR71] D. S. Radford and R. Hartland-Rowe. Subsurface and surface sampling of benthic invertebrates in two stream. Limnology and Oceanography, 16(1):114–120, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Richards:1979:APA

[RHW79] Martin C. Richards and Christine M. Happey-Wood. The application of pattern analysis to freshwater phytoplankton communities. Limnology and Oceanography, 24(5):950–956, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Richardson:1975:MLP

[Ric75] Jonathan L. Richardson. Morphometry and lacustrine productivity. Limnology and Oceanography, 20(5):661–663, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Ricker:1978:CP

[Ric78] W. E. Ricker. On computing production. Limnology and Oceanography, 23(2):379–380, March 1978. CODEN LIOCAH. ISSN 0024-3590. See reply [LP78].

Richey:1979:PPS

[Ric79] Jeffrey E. Richey. Patterns of phosphorus supply and utilization in Lake Washington and Findley Lake. Limnology and Oceanography, 24(5):906–916, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Riemann:1978:CIS

[Rie78] B. Riemann. Carotenoid interference in the spectrophotometry determination of chlorophyll degradation products from natural populations of phytoplankton. Limnology and Oceanography, 23 (5):1059–1066, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Rigler:1972:BRW

[Rig72] Frank Rigler. Book review: Winberg, G. G. [Ed.] 1971. Methods for the estimation of production of aquatic animals. Academic Press. xii + 175 p. £3.00, $9.00. Limnology and Oceanography, 17 (2):325, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Rigler:1976:BRP

[Rig76] F. H. Rigler. Book review: Patten, B. C. [Ed.].1975. Systems analysis and simulation in ecology, volume 3. Academic Press, Inc., New York, xv + 601 p. $39.50. Limnology and Oceanography, 21(3):481–483, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Riley:1971:BRH

[Ril71a] Gordon A. Riley. Book review: Hood, D. W. [Ed.]. 1970. Organic matter in natural waters. Inst. Mar. Sci. Occas. Publ. 1. Univ. Alaska, College. xiv + 625 p. Hardcover, $10.00; softcover, $8.00. Limnology and Oceanography, 16(1):152, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Riley:1971:I

[Ril71b] Gordon A. Riley. Introduction. Limnology and Oceanography, 16 (2):177–179, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Riley:1973:BRS

[Ril73] G. A. Riley. Book review: Schlieper, C. [Ed.]. 1972. Research methods in marine biology. University of Washington Press, Seattle, xiii + 356 p. $12.50. Limnology and Oceanography, 18(4): 694–695, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Riley:1974:BRP

[Ril74] G. A. Riley. Book review: Parsons, T. R., and M. Takahashi. 1973. Biological oceanographic processes. Pergamon Press, Oxford, New York, Toronto, Sydney, Braunschweig, x + 186 p. $8.75. Limnology and Oceanography, 19(6):875, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Riley:1975:TCR

[Ril75] Gordon A. Riley. Transparency-chlorophyll relations. Limnology and Oceanography, 20(1):150–152, January 1975. CODEN LIOCAH. ISSN 0024-3590. See reply [Ban75].

Riley:1976:MPP

[Ril76] Gordon A. Riley. A model of plankton patchiness. Limnology and Oceanography, 21(6):873–880, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Riley:1977:BRG

[Ril77] J. P. Riley. Book review: Grasshoff, K. [Ed.] 1976. Methods of seawater analysis. Verlag Chemie, Weinheim and New York, xv + 317 p. $43.60. Limnology and Oceanography, 22(6):1103, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Riley:1978:BRG

[Ril78a] Gordon A. Riley. Book review: Goldberg, E. D., I. N. McCave, J. J. O’Brien, and J. H. Steele [Eds.] 1977. The sea (Ideas and observations on progress in the study of the seas). V. 6: Marine modeling. John Wiley and Sons, New York. xiv + 1048 p. $49.50. Limnology and Oceanography, 23(3):574–575, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Riley:1978:BRK

[Ril78b] Gordon A. Riley. Book review: Koppelman, L. E., P. K. Weyl, M. G. Gross, and DeW. S. Davies. 1976. The urban sea: Long Island Sound. Praeger Publishers, New York, viii + 223 p. Limnology and Oceanography, 23(5):1085, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Risk:1973:SPC

[Ris73] Michael J. Risk. Settling plates of cold-cure acrylic plastic replicated from natural surfaces. Limnology and Oceanography, 18(5): 801–802, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Reshkin:1979:LSP

[RK79] Stephan J. Reshkin and George A. Knauer. Light stimulation of phosphate uptake in natural assemblages of phytoplankton. Limnology and Oceanography, 24(6):1121–1124, November 1979. CO-DEN LIOCAH. ISSN 0024-3590.

Richman:1975:RMP

[RLS75] Sumner Richman, Y. Loya, and Lawrence B. Sloboclkin. The rate of mucus production by corals and its assimilation by the coral reef copepod Acartia negligens. Limnology and Oceanography, 20 (6):918–923, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Rueter:1979:TEC

[RMC79] John G. Rueter, James J. McCarthy, and Edward J. Carpenter. The toxic effect of copper on Oscillatoria (Trichodesmium) theibautii. Limnology and Oceanography, 24(3):558–562, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Richerson:1970:CEA

[RMG70] Peter J. Richerson, Gerald A. Moshiri, and Gordon L. Godshalk. Certain ecological aspects of pollen dispersion in Lake Tahoe (California–Nevada). Limnology and Oceanography, 15(1):149– 153, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Ritchie:1973:DRR

[RMG73] Jerry C. Ritchie, J. Roger McHenry, and Angela C. Gill. Dating recent reservoir sediments. Limnology and Oceanography, 18(2): 254–263, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Roberts:1977:ESC

[RMS77] H. H. Roberts, S. P. Murray, and J. N. Suhayda. Evidence for strong currents and turbulence in a deep coral reef groove. Limnology and Oceanography, 22(1):152–156, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Ruby:1978:SCS

[RN78] E. G. Ruby and K. H. Nealson. Seasonal changes in the species composition of luminous bacteria in nearshore seawater. Limnology and Oceanography, 23(3):530–533, May 1978. CODEN LI-OCAH. ISSN 0024-3590.

Roelofs:1970:EOP

[RO70] T. D. Roelofs and R. T. Oglesby. Ecological observations on the planktonic cyanophyte Gleotrichia echinulata. Limnology and Oceanography, 15(2):224–229, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Rodhe:1974:PPP

[Rod74] Wilhelm Rodhe. Plankton, planktic, planktonic. Limnology and Oceanography, 19(2):360, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Rogers:1979:PSC

[Rog79] Caroline S. Rogers. The productivity of San Cristobal Reef, Puerto Rico. Limnology and Oceanography, 24(2):342–349, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Roman:1978:IBG

[Rom78] Michael R. Roman. Ingestion of the blue-green alga Trichodesmium by the harpactacoid copepod, Macrosetella gracilis. Limnology and Oceanography, 23(6):1245–1248, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Rosenfeld:1979:AAD

[Ros79a] Jeffrey K. Rosenfeld. Amino acid diagenesis and adsorption in nearshore anoxic sediments. Limnology and Oceanography, 24(6): 1014–1021, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Rosenfeld:1979:AAN

[Ros79b] Jeffrey K. Rosenfeld. Ammonium adsorption in nearshore anoxic sediments. Limnology and Oceanography, 24(2):356–364, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Rivkin:1979:DVP

[RS79] Richard B. Rivkin and Elijah Swift. Diel and vertical patterns of alkaline phosphatase activity in the oceanic dinoflagellate Pyrocystis noctiluca. Limnology and Oceanography, 24(1):107–116, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Reid:1978:SDP

[RSEG78] F. M. H. Reid, E. Stewart, R. W. Eppley, and D. Goodman. Spatial distribution of phytoplankton species in chlorophyll maximum layers off southern California. Limnology and Oceanography, 23 (2):219–226, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Rakusa-Suszczewski:1976:RAC

[RSMC76] S. Rakusa-Suszczewski, M. A. McWhinnie, and M. O. Cahoon. Respiration of the Antarctic copepod, Rhincalanus gigas. Limnology and Oceanography, 21(5):763–765, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Rent:1978:SLM

[RSS78] Jonathan D. Rent, Alan L. Shanks, and Mary Wilcox Silver. In situ and laboratory measurements on macroscopic aggregates in Monterey Bay, California. Limnology and Oceanography, 23(4): 626–635, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Ruppert:1972:EQM

[Rup72] Edward E. Ruppert. An efficient, quantitative method for sampling the meiobenthos. Limnology and Oceanography, 17(4):629– 631, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Roger:1971:PFC

[RW71] T. Roger and S. Wilson. A portable flow-cell membrane salinometer. Limnology and Oceanography, 16(3):581–586, May 1971. CO-DEN LIOCAH. ISSN 0024-3590.

Reeve:1978:LSI

[RWI78] M. R. Reeve, M. A. Walter, and T. Ikeda. Laboratory studies of ingestion and food utilization in lobate and tentaculate ctenophores.

Limnology and Oceanography, 23(4):740–751, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Rzoska:1979:BRM

[Rzo79] Julian Rzoska. Book review: Mordukhai-Boltovskoi, Ph. D. 1979. The River Volga and its life. Monogr. Biol. 33. Dr. W. Junk, The Hague, xiii + 473 p. f.140; US$70. Limnology and Oceanography, 24(6):1176–1177, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Schell:1970:IIG

[SA70] Donald M. Schell and Vera Alexander. Improved incubation and gas sampling techniques for nitrogen fixation studies. Limnology and Oceanography, 15(6):961–962, November 1970. CODEN LI-OCAH. ISSN 0024-3590.

Stansson:1971:NDR

[SA71] Unnsteinn Stánsson and Larry P. Atkinson. Nutrient-density relationships in the Western North Atlantic between Cape Lookout and Bermuda. Limnology and Oceanography, 16(1):51–59, January 1971. CODEN LIOCAH. ISSN 0024-3590.

Speir:1974:UED

[SA74] Jon A. Speir and N. H. Anderson. Use of emergence data for estimating annual production of aquatic insects. Limnology and Oceanography, 19(1):154–156, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Swift:1976:UNS

[SA76] James H. Swift and Knut Aagaard. Upwelling near samalga pass. Limnology and Oceanography, 21(3):399–408, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Schindler:1972:PIO

[SAH72] J. E. Schindler, J. J. Alberts, and K. R. Honick. A preliminary investigation of organic-inorganic associations in a stagnating system. Limnology and Oceanography, 17(6):952–957, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Salonen:1979:VMR

[Sal79] Kalevi Salonen. A versatile method for the rapid and accurate determination of carbon by high temperature combustion. Limnology and Oceanography, 24(1):177–183, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Sanford:1977:MGI

[San77] Thomas B. Sanford. Measurements by geomagnetic induction of volume transport in a salt marsh drainage channel. Limnology and Oceanography, 22(6):1082–1089, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Saunders:1972:PHN

[Sau72a] George W. Saunders. Potential heterotrophy in a natural population of Oscillatoria agardhii Var. Isothrix Skuja. Limnology and Oceanography, 17(5):704–711, September 1972. CODEN LI-OCAH. ISSN 0024-3590.

Saunders:1972:PAS

[Sau72b] George W. Saunders, Jr. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 17(1):161–163, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Shannon:1972:LCN

[SB72] Earl E. Shannon and Patrick L. Brezonik. Limnological characteristics of north and central Florida lakes. Limnology and Oceanography, 17(1):97–110, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Simpson:1973:NMD

[SB73] H. James Simpson and Wallace S. Broecker. A new method for determining the total carbonate ion concentration in saline waters. Limnology and Oceanography, 18(3):426–440, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Scott:1975:ASS

[SB75] T. G. Scott and A. J. S. Ball. An apparatus for studying syntrophic interactions between microorganisms. Limnology and Oceanography, 19(4):291–295, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Scranton:1978:CDM

[SB78a] Mary I. Scranton and Peter G. Brewer. Consumption of dissolved methane in the deep ocean. Limnology and Oceanography, 23(6): 1207–1213, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Smith:1978:BOS

[SB78b] Raymond C. Smith and Karen S. Baker. The bio-optical state of ocean waters and remote sensing. Limnology and Oceanography, 23(2):247–259, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Smith:1978:OCN

[SB78c] Raymond C. Smith and Karen S. Baker. Optical classification of natural waters. Limnology and Oceanography, 23(2):260–267, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Sick:1979:CIM

[SB79] Lowell V. Sick and Garry J. Baptist. Cadmium incorporation by the marine copepod Pseudodiaptomus coronatus. Limnology and Oceanography, 24(3):453–462, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Shonting:1970:SDT

[SC70] David H. Shonting and Gerald S. Cook. On the seasonal distribution of temperature and salinity in Rhode Island Sound. Limnology and Oceanography, 15(1):100–112, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Smith:1976:URM

[SC76a] Raymond C. Smith and John Calkins. The use of the Robertson meter to measure the penetration of solar middle-ultraviolet radiation (UV-B) into natural waters. Limnology and Oceanography, 21(5):746–749, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Sozanski:1976:EDM

[SC76b] A. G. Sozanski and D. S. Cronan. Environmental differentiation of morphology of ferromanganese oxide concretion in Shebandowan Lakes, Ontario. Limnology and Oceanography, 21(6): 894–898, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Slawyk:1977:UIS

[SCA77] Gerd Slawyk, Yves Collos, and Jean-Christian Auclair. The use of the^{13}C and^{15}N isotopes for the simultaneous measurement of carbon and nitrogen turnover rates in marine phytoplankton. Limnology and Oceanography, 22(5):925–932, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Slawyk:1979:RCF

[SCA79] G. Slawyk, Y. Collos, and C. Auclair. Reply to comment by Fisher et al. Limnology and Oceanography, 24(3):595–597, May 1979. CODEN LIOCAH. ISSN 0024-3590. See [FHV79].

Smith:1976:FVM

[SCE^{+}76] K. L. Smith, Jr., C. H. Clifford, A. H. Eliason, B. Walden, G. T. Rowe, and J. M. Teal. A free vehicle for measuring benthic community metabolism. Limnology and Oceanography, 21(1):164–170, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Schegg:1970:NBS

[Sch70] Ernst Schegg. A new bacteriological sampling bottle. Limnology and Oceanography, 15(5):820–822, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Schafer:1971:SSD

[Sch71] Charles T. Schafer. Sampling and spatial distribution of benthonic Foraminifera. Limnology and Oceanography, 16(6):944– 951, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Schoener:1972:FPM

[Sch72] Amy Schoener. Fecundity and possible mode of development of some deep-sea ophiuroids. Limnology and Oceanography, 17(2): 193–199, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Schindler:1973:BRH

[Sch73] D. W. Schindler. Book review: Hrbáček, j., and M. Straškraba [Eds.]. 1973. Hydrobiological studies 2 and 3. Academia, Prague. Vol. 2, 348 p., Kčs 90.; vol. 3, 310 p., Kčs 85. Limnology and Oceanography, 16(5):819–820, September 1973. CODEN LI-OCAH. ISSN 0024-3590.

Schell:1974:URF

[Sch74] Donald M. Schell. Uptake and regeneration of free amino acids in marine waters of Southeast Alaska. Limnology and Oceanography, 19(2):260–270, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Schindler:1975:FAG

[Sch75a] D. W. Schindler. Factors affecting gas exchange in natural waters. Limnology and Oceanography, 20(3):1053, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Schindler:1975:BRB

[Sch75b] David W. Schindler. Book review: Broecker, W. S. 1974. Chemical oceanography. Harcourt, Brace, Jovanovich, Inc., New York. 214 p. $7.95. Limnology and Oceanography, 19(4):299–300, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Schnitzer:1976:BRG

[Sch76] M. Schnitzer. Book review: Gjessing, E. T. 1976. Physical and chemical characteristics of aquatic humus. Ann Arbor Science, Ann Arbor, Michigan, vii + 120 p. $14.50. Limnology and Oceanography, 21(6):932–933, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Scheffer:1977:BRC

[Sch77a] Victor B. Scheffer. Book review: Coffey, D. J. 1977. Dolphins, whales and porpoises: an encyclopedia of sea mammals. Macmillan Publ. Co., New York (designed and produced by George Rainbird, Ltd., London). 223 p. $17.95. Limnology and Oceanography, 22(6): 1105, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Schwab:1977:IFO

[Sch77b] David J. Schwab. Internal free oscillations in Lake Ontario. Limnology and Oceanography, 22(4):700–708, July 1977. CODEN LI-OCAH. ISSN 0024-3590.

Schindler:1978:FRP

[Sch78a] D. W. Schindler. Factors regulating phytoplankton production and standing crop in the world’s freshwaters. Limnology and Oceanography, 23(3):478–486, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Schindler:1978:PEM

[Sch78b] D. W. Schindler. Predictive eutrophication models. Limnology and Oceanography, 23(5):1080–1081, September 1978. CODEN LIOCAH. ISSN 0024-3590. See reply [Tho78].

Seliger:1971:BPS

[SCL^{+}71] H. H. Seliger, J. H. Carpenter, M. Loftus, W. H. Biggley, and W. D. McElroy. Bioluminescence and phytoplankton successions in Bahía Fosforescente, Puerto Rico. Limnology and Oceanography, 16(4):608–622, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Seliger:1970:MAH

[SCLM70] H. H. Seliger, J. H. Carpenter, M. Loftus, and W. D. McElroy. Mechanisms for the accumulation of high concentrations of dinoflagellates in a bioluminescent bay. Limnology and Oceanography, 15(2):234–245, March 1970. CODEN LIOCAH. ISSN 0024-

Silver:1978:LAA

[SD78] Mary Wilcox Silver and Peter J. Davoll. Loss of ^{14}C activity after chemical fixation of phytoplankton: Error source for autoradiography and other productivity measurements. Limnology and Oceanography, 23(2):362–368, March 1978. CODEN LIOCAH. ISSN 0024-3590. See comment [Bro78a].

Strong:1978:SOC

[SE78] Alan E. Strong and Brian J. Eadie. Satellite observations of calcium carbonate precipitations in the Great Lakes. Limnology and Oceanography, 23(5):877–887, September 1978. CODEN LI-OCAH. ISSN 0024-3590.

Shapiro:1971:CCC

[SEA71] Joseph Shapiro, W. T. Edmondson, and David E. Allison. Changes in the chemical composition of sediments of Lake Washington, 1958–1970. Limnology and Oceanography, 16(2):437–452, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Seckbach:1971:PAU

[Sec71] Joseph Seckbach. Pigmentation of algae under pressure. Limnology and Oceanography, 16(3):567–572, May 1971. CODEN LI-OCAH. ISSN 0024-3590.

Senft:1978:DLS

[Sen78] W. Herbert Senft. Dependence of light-saturated rates of algal photosynthesis on intracellular concentrations of phosphorus. Limnology and Oceanography, 23(4):709–718, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Serruya:1974:LKS

[SEPS74] Colette Serruya, Miriam Edelstein, Utsa Pollingher, and Salvador Serruya. Lake Kinneret sediments: Nutrient composition of the pore water and mud water exchanges. Limnology and Oceanography, 19(3):489–508, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Serruya:1971:LKN

[Ser71] C. Serruya. Lake Kinneret: the nutrient chemistry of the sediments. Limnology and Oceanography, 16(3):510–521, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Serruya:1974:MPJ

[Ser74] Salvador Serruya. The mixing patterns of the Jordan River in Lake Kinneret. Limnology and Oceanography, 19(2):175–181, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Swift:1975:SAP

[SF75] Michael C. Swift and Alice Y. Fedorenko. Some aspects of prey capture by Chaoborus larvae. Limnology and Oceanography, 20 (4):418–425, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Sanger:1970:DPL

[SG70] Jon E. Sanger and Eville Gorham. The diversity of pigments in lake sediments and its ecological significance. Limnology and Oceanography, 15(1):59–69, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Sholkovitz:1971:PCS

[SG71] Edward R. Sholkovitz and Joris M. Gieskes. A physical-chemical study of the flushing of the Santa Barbara Basin. Limnology and Oceanography, 16(3):479–489, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Sanger:1972:SFP

[SG72] Jon E. Sanger and Eville Gorham. Stratigraphy of fossil pigments as a guide to the postglacial history of Kirchner Marsh, Minnesota. Limnology and Oceanography, 17(6):840–854, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Stephens:1976:PPG

[SG76] D. W. Stephens and D. M. Gillespie. Phytoplankton production in the Great Salt Lake, Utah, and a laboratory study of algal response to enrichment. Limnology and Oceanography, 21(1):74– 87, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Schindler:1971:ORD

[SH71] J. E. Schindler and K. R. Honick. Oxidation-reduction determinations at the mud-water interface. Limnology and Oceanography, 18(5):837–840, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Smith:1972:CGS

[SH72] K. L. Smith, Jr. and James D. Howard. Comparison of a grab sampler and large volume corer. Limnology and Oceanography, 17 (1):142–145, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Sperling:1973:PRU

[SH73] Jon A. Sperling and Geraldine M. Hale. Patterns of radiocarbon uptake by a thermophilic blue-green alga under varying conditions of incubation. Limnology and Oceanography, 18(4):658–662, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Slovacek:1977:VFD

[SH77] Rudolf E. Slovacek and Patrick J. Hannan. In vivo fluorescence determinations of phytoplankton chlorophyll-a. Limnology and Oceanography, 22(5):919–925, September 1977. CODEN LI-OCAH. ISSN 0024-3590.

Sandusky:1978:PAC

[SH78] James C. Sandusky and Alexander J. Horne. A pattern analysis of Clear Lake phytoplankton. Limnology and Oceanography, 23 (4):636–648, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Shapiro:1973:BRA

[Sha73a] Joseph Shapiro. Book review: Allen, H. E., and J. R. Kramer [Eds.]. 1972. Nutrients in natural waters. Wiley-Interscience, New York. x + 457 p. $21.50. Limnology and Oceanography, 18(3): 509–510, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Shapiro:1973:FFT

[Sha73b] Joseph Shapiro. A field fixation technique for dissolved phosphate in lake water. Limnology and Oceanography, 18(1):143–145, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Sharp:1973:SCO

[Sha73c] Jonathan H. Sharp. Size classes of organic carbon in seawater. Limnology and Oceanography, 18(3):441–447, May 1973. CODEN LIOCAH. ISSN 0024-3590.

SharafElDin:1974:LST

[Sha74a] S. H. Sharaf El Din. Longshore sand transport in the surf zone along the Mediterranean Egyptian coast. Limnology and Oceanography, 19(2):182–189, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Sharp:1974:IAP

[Sha74b] Jonathan H. Sharp. Improved analysis for “particulate” organic carbon and nitrogen from seawater. Limnology and Oceanography, 19(6):984–989, November 1974. CODEN LIOCAH. ISSN 0024-3590.

SharafElDin:1977:EAH

[Sha77a] S. H. Sharaf El Din. Effect of the Aswan High Dam on the Nile flood and on the estuarine and coastal circulation pattern along the Mediterranean Egyptian coast. Limnology and Oceanography, 22(2):194–207, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Sharp:1977:EOM

[Sha77b] Jonathan H. Sharp. Excretion of organic matter by marine phytoplankton: Do healthy cells do it? Limnology and Oceanography, 22(3):381–399, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Sharp:1978:RCA

[Sha78] J. H. Sharp. Reply to comment by S. Aaronson. Limnology and Oceanography, 23(4):839–840, July 1978. CODEN LIOCAH. ISSN 0024-3590. See [Aar78].

Sheldon:1972:SSM

[She72] R. W. Sheldon. Size separation of marine seston by membrane and glass-fiber filters. Limnology and Oceanography, 17(3):494– 498, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Sheldon:1979:MPG

[She79] R. W. Sheldon. Measurement of phytoplankton growth by particle counting. Limnology and Oceanography, 24(4):760–767, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Sholkovitz:1970:FVB

[Sho70] Edward R. Sholkovitz. A free vehicle bottom-water sampler. Limnology and Oceanography, 15(4):641–643, July 1970. CODEN LI-OCAH. ISSN 0024-3590.

Shouting:1974:COW

[Sho74] David H. Shouting. Current observations from the western Mediterranean during COBLAMED 69. Limnology and Oceanography, 19(5):866–874, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Shuter:1978:SDP

[Shu78] B. J. Shuter. Size dependence of phosphorus and nitrogen subsistence quotas in unicellular microorganisms. Limnology and Oceanography, 23(6):1248–1255, November 1978. CODEN LI-OCAH. ISSN 0024-3590.

Sommers:1970:DTO

[SHW^{+}70] L. E. Sommers, R. F. Harris, J. D. H. Williams, D. E. Armstrong, and J. K. Syers. Determination of total organic phosphorus in lake sediments. Limnology and Oceanography, 15(2):301–303, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Silver:1975:HSF

[Sil75] Mary Wilcox Silver. The habitat of Salpa fusiformis in the California Current as defined by indicator assemblages. Limnology and Oceanography, 19(4):230–237, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Silvert:1979:PCF

[Sil79] William Silvert. Practical curve fitting. Limnology and Oceanography, 24(4):767–773, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Sioli:1977:BRR

[Sio77] Harald Sioli. Book review: Rzóska, J. [Ed.] 1976. The Nile. Biology of an ancient river. Monogr. Biol. 29. Junk, The Hague, xx + 418 p. Dfl. 120. Limnology and Oceanography, 22(6):1100, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Sheldon:1972:PDM

[SK72] R. W. Sheldon and S. R. Kerr. The population density of monsters in Loch Ness. Limnology and Oceanography, 17(5):796–798, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Sheldon:1973:LNM

[SK73a] R. W. Sheldon and S. R. Kerr. The Loch Ness monster: Reply to comments of C. H. Mortimer. Limnology and Oceanography, 18(2):345–346, March 1973. CODEN LIOCAH. ISSN 0024-3590. See [Mor73].

Sorokin:1973:PDO

[SK73b] Yu. I. Sorokin and I. W. Konovalova. Production and decomposition of organic matter in a bay of the Japan Sea during the winter diatom bloom. Limnology and Oceanography, 18(6):962–967, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Smith:1975:CDM

[SK75] S. V. Smith and G. S. Key. Carbon dioxide and metabolism in marine environments. Limnology and Oceanography, 20(6):493– 495, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Small:1974:EDF

[SKF74] Lawrence F. Small, Stjepan KečkeŠ, and Scott W. Fowler. Excretion of different forms of zinc by the prawn Palaemon serratus (Pennant). Limnology and Oceanography, 19(5):789–793, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Stockner:1970:LAP

[SL70] John G. Stockner and J. W. G. Lund. Live algae in postglacial lake deposits. Limnology and Oceanography, 15(1):41–58, January 1970. CODEN LIOCAH. ISSN 0024-3590.

Shuman:1975:QDC

[SL75] F. Randolph Shuman and Carl J. Lorenzen. Quantitative degradation of chlorophyll by a marine herbivore. Limnology and Oceanography, 20(4):580–586, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Shoaf:1976:IEC

[SL76] W. Thomas Shoaf and Bruce W. Lium. Improved extraction of chlorophyll-a and b from algae using dimethyl sulfoxide. Limnology and Oceanography, 21(6):926–928, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Sunda:1978:ECN

[SL78] William G. Sunda and Jo Ann M. Lewis. Effect of complexation by natural organic ligands on the toxicity of copper to a unicellular alga, Monochrysis lutheri. Limnology and Oceanography, 23(5): 870–876, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Smith:1973:OCP

[SM73] Stephen V. Smith and James A. Marsh, Jr. Organic carbon production on the windward reef flat of Eniwetok Atoll. Limnology and Oceanography, 18(6):953–961, November 1973. CODEN LI-OCAH. ISSN 0024-3590.

Stiller:1974:CEC

[SM74] Mariana Stiller and Mordechai Magaritz. Carbon-13 enriched carbonate in interstitial waters of Lake Kinneret sediments. Limnology and Oceanography, 19(5):849–853, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Sayles:1976:ASD

[SM76] Frederick L. Sayles and Paul C. Mangelsdorf, Jr. The analysis of SO_{4}^{2−} in seawater by difference chromatography. Limnology and Oceanography, 21(6):899–905, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Sansone:1978:MOC

[SM78] Francis J. Sansone and Christopher S. Martens. Methane oxidation in Cape Lookout Bight, North Carolina. Limnology and Oceanography, 23(2):349–355, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Sapienxa:1979:DPA

[SM79] Carmen Sapienxa and Timothy H. Mague. DNA polymerase activity and growth rates in Artemia salina nauplii. Limnology and Oceanography, 24(3):572–576, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Smayda:1974:BGP

[Sma74a] Theodore J. Smayda. Bioassay of the growth potential of the surface water of lower Narragansett Bay over an annual cycle using the diatom Thalassiosira pseudonana (oceanic clone, 13-1). Limnology and Oceanography, 19(6):889–901, November 1974. CO-DEN LIOCAH. ISSN 0024-3590.

Smayda:1974:SES

[Sma74b] Theodore J. Smayda. Some experiments on the sinking characteristics of two freshwater diatoms. Limnology and Oceanography, 20 (2):628–635, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Slawyk:1976:INU

[SMD76] G. Slawyk, J. J. MacIsaac, and R. C. Dugdale. Inorganic nitrogen uptake by marine phytoplankton under in situ and simulated in situ incubation conditions: Results from the northwest African upwelling region. Limnology and Oceanography, 21(1):149–152, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Smith:1971:DSI

[Smi71] K. L. Smith, Jr. A device for sampling immediately above the sediment-water interface. Limnology and Oceanography, 16(4): 675–677, July 1971. CODEN LIOCAH. ISSN 0024-3590.

Smith:1972:PCC

[Smi72] S. V. Smith. Production of calcium carbonate on the mainland shelf of southern California. Limnology and Oceanography, 17(1): 28–41, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Smith:1973:FOW

[Smi73a] Ned P. Smith. The fall overturn and winter cooling in Lake Superior. Limnology and Oceanography, 18(3):483–487, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Smith:1973:CDD

[Smi73b] S. V. Smith. Carbon dioxide dynamics: a record of organic carbon production, respiration, and calcification in the Eniwetok Reef Flat community. Limnology and Oceanography, 18(1):106–120, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Smith:1974:BRM

[Smi74a] Stephen V. Smith. Book review: Milliman, J. D. 1974. Marine carbonates. Recent sedimentary carbonates part 1. Springer-Verlag, New York, Heidelberg, & Berlin, xv + 375 p. $25.50. Limnology and Oceanography, 19(6):876, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Smith:1974:ODS

[Smi74b] K. L. Smith, Jr. Oxygen demands of San Diego Trough sediments: an in situ study. Limnology and Oceanography, 19(6):939–944, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Smith:1975:BRH

[Smi75] Stephen V. Smith. Book review: Hickling, C. F., and P. Lancaster Brown. 1974. The sea and oceans in color. Macmillan Co., New York, x + 249 p. $5.95. Limnology and Oceanography, 19(4):302– 303, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Smith:1976:BRW

[Smi76a] S. V. Smith. Book review: Wood, E. J. F., and R. E. Johannes [Eds.]. 1975. Tropical marine pollution. Elsevier Oceanography Series, v. 12. Elsevier Sci. Publ. Co., Amsterdam and New York. ix + 192 p. $26.95. Limnology and Oceanography, 21(4):634–635, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Smith:1976:BRK

[Smi76b] Stephen V. Smith. Book review: King, C. A. M. 1975. Introduction to physical and biological oceanography. Crane, Russak & Co., Inc., New York, ix + 372 p. 29.00 (paperback, $13.50). Limnology and Oceanography, 21(1):185–186, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Smith:1976:BRR

[Smi76c] Stephen V. Smith. Book review: Riley, J. P., and G. Skirrow [Eds.]. 1975. Chemical oceanography, v. 2. 2nd ed. Academic Press, New York and London, xx + 647 p. $51.50. £19.50. Limnology and Oceanography, 21(2):347, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Smith:1977:NWT

[Smi77] Ned P. Smith. A note on winter temperature variations in a shallow seagrass flat. Limnology and Oceanography, 22(6):1079–1082, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Smith:1978:BRFb

[Smi78a] Stephen V. Smith. Book review: Flemming, N. C. [Ed.] 1977. The undersea. Macmillan Publ. Co., Inc., New York. 319 p. $27.50. Limnology and Oceanography, 23(5):1088, September 1978. CO-DEN LIOCAH. ISSN 0024-3590.

Smith:1978:BRFa

[Smi78b] Stephen V. Smith. Book review: Flügel, E. [Ed.] 1977. Fossil algae: Recent results and developments. Springer-Verlag, New York, xiii + 375 p. $36.10. Limnology and Oceanography, 23(5):1087– 1088, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Smith:1979:NDP

[Smi79] Val H. Smith. Nutrient dependence of primary productivity in lakes. Limnology and Oceanography, 24(6):1051–1064, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Stephens:1971:ECA

[SN71] Grover C. Stephens and Barbara B. North. Extrusion of carbon accompanying uptake of amino acids by marine phytoplankters. Limnology and Oceanography, 18(5):752–757, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Sasseville:1975:PHG

[SN75] Dennis R. Sasseville and Stephen A. Norton. Present and historic geochemical relationships in four Maine lakes. Limnology and Oceanography, 20(4):699–714, September 1975. CODEN LI-OCAH. ISSN 0024-3590.

Snow:1972:ESA

[Sno72] Norman B. Snow. The effect of season and animal size on the caloric content of Daphnia pulicaria Forbes. Limnology and Oceanography, 17(6):909–912, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Schaffner:1978:PLL

[SO78] W. R. Schaffner and R. T. Oglesby. Phosphorus loadings to lakes and some of their responses. Part 1. A new calculation of phosphorus loading and its application to 13 New York lakes. Limnology and Oceanography, 23(1):120–134, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Sutcliffe:1976:DAM

[SOHH76] W. H. Sutcliffe, Jr., E. A. Orr, and O. Holm-Hansen. Difficulties with ATP measurements in inshore waters. Limnology and Oceanography, 21(1):145–149, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Sandholm:1973:USA

[SOP73] M. Sandholm, H. E. Oksanen, and L. Pesonen. Uptake of selenium by aquatic organisms. Limnology and Oceanography, 18(3):496– 499, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Sorokin:1973:FSS

[Sor73] Yu. I. Sorokin. On the feeding of some scleractinian corals with bacteria and dissolved organic matter. Limnology and Oceanography, 18(3):380–386, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Strohal:1973:TWA

[SP73] P. Strohal and T. Pinter. Thorium in water and algae from the Adriatic Sea. Limnology and Oceanography, 18(2):250–253, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Silvert:1978:EFP

[SP78] William Silvert and Trevor Platt. Energy flux in the pelagic ecosystem: a time-dependent equation. Limnology and Oceanography, 23(4):813–816, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Sheldon:1972:SDP

[SPS72] R. W. Sheldon, A. Prakash, and W. H. Sutcliffe, Jr. The size distribution of particles in the ocean. Limnology and Oceanography, 17(3):327–340, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Schurr:1977:DCE

[SR77] J. M. Schurr and J. Ruchti. Dynamics of O_{2} and CO_{2} exchange, photosynthesis, and respiration in rivers from time-delayed correlations with ideal sunlight. Limnology and Oceanography, 22(2): 208–225, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Schroeder:1979:ETF

[SR79] Lauren Alfred Schroeder and Patricia S. Reeder. Effects of temperature and fasting on the rate of sinking of Hydra oligactis. Limnology and Oceanography, 24(6):1092–1103, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Seyfried:1979:DSS

[SRR79] William E. Seyfried, Robert J. Rosenbauer, and James L. Rischoff. Determination of sulfate in seawater and natural brines by ^{133}Ba and membrane dialysis. Limnology and Oceanography, 24(2):393– 396, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Schelske:1974:RPL

[SRSS74] Claire L. Schelske, Edward D. Rothman, Eugene F. Stoermer, and Mila A. Santiago. Responses of phosphorus limited Lake Michigan phytoplankton to factorial enrichments with nitrogen and phosphorus. Limnology and Oceanography, 19(3):409–419, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Staresinic:1978:MVF

[SRSW78] Nick Staresinic, Gilbert T. Rowe, Dan Shaughnessey, and Albert J. Williams III. Measurement of the vertical flux of particulate organic matter with a free-drifting sediment trap. Limnology and Oceanography, 23(3):559–563, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Schiemer:1970:NBS

[SS70] E. W. Schiemer and J. R. Schubel. A near-bottom suspended sediment sampling system for studies of resuspension. Limnology and Oceanography, 15(4):644–646, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Satake:1974:CDC

[SS74] Ken’ichi Satake and Yatsuka Saijo. Carbon dioxide content and metabolic activity of microorganisms in some acid lakes in Japan. Limnology and Oceanography, 19(2):331–338, March 1974. CO-DEN LIOCAH. ISSN 0024-3590.

Stanley:1977:AUA

[SS77] Patricia M. Stanley and James T. Staley. Acetate uptake by aquatic bacterial communities measured by autoradiography and filterable radioactivity. Limnology and Oceanography, 22(1):26–37, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Sheldon:1978:GTH

[SS78a] R. W. Sheldon and W. H. Sutcliffe, Jr. Generation times of 3 h for Sargasso Sea microplankton determined by ATP analys. Limnology and Oceanography, 23(5):1051–1055, September 1978. CO-DEN LIOCAH. ISSN 0024-3590.

Storch:1978:PER

[SS78b] Thomas A. Storch and George W. Saunders. Phytoplankton extracellular release and its relation to the seasonal cycle of dissolved organic carbon in a eutrophic lake. Limnology and Oceanography, 23(1):112–119, January 1978. CODEN LIOCAH. ISSN 0024-3590.

Szekielda:1972:RVS

[SSA72] K.-H. Szekielda, V. Salomonson, and L. J. Allison. Rapid variations of sea surface temperature in the Persian Gulf as recorded by Nimbus 2 HRIR. Limnology and Oceanography, 17(2):307–309, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Salonen:1976:REO

[SSHV76] Kalevi Salonen, Jouko Sarvala, Ilpo Hakala, and Maija-Liisa Viljanen. The relation of energy and organic carbon in aquatic invertebrates. Limnology and Oceanography, 21(5):724–730, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Sieburth:1978:PES

[SSL78] John McN. Sieburth, Victor Smetacek, and Jürgen Lenz. Pelagic ecosystem structure: Heterotrophic compartments of the plankton and their relationship to plankton size fractions. Limnology and Oceanography, 23(6):1256–1263, November 1978. CODEN LI-OCAH. ISSN 0024-3590.

Swift:1976:SGR

[SSM76] Elijah Swift, Marc Stuart, and Valerie Meunier. The in situ growth rates of some deep-living oceanic dinoflagellates: Pyrocystis fusiformis and Pyrocystis noctiluca. Limnology and Oceanography, 21(3):418–426, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Sheldon:1973:PPS

[SSP73] R. W. Sheldon, W. H. Sutcliffe, Jr., and A. Prakash. The production of particles in the surface waters of the ocean with particular reference to the Sargasso Sea. Limnology and Oceanography, 18(5): 719–733, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Strayer:1978:SMP

[ST78] Richard F. Strayer and James M. Tiedje. In situ methane production in a small, hypereutrophic, hard-water lake: Loss of methane from sediments by vertical diffusion and ebullition. Limnology and Oceanography, 23(6):1201–1206, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Shanks:1979:MSM

[ST79] Alan L. Shanks and Jonathan D. Trent. Marine snow: Microscale nutrient patches. Limnology and Oceanography, 24(5):850–854, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Stavn:1971:HVD

[Sta71] Robert Hans Stavn. The horizontal-vertical distribution hypothesis: Langmuir circulations and Daphnia distributions. Limnology and Oceanography, 16(2):453–466, March 1971. CODEN LI-OCAH. ISSN 0024-3590.

Staley:1972:BRG

[Sta72a] James T. Staley. Book review: Gordon, R. W. 1972. Field and laboratory microbial ecology. Wm. Brown, Dubuque. vii + 166 p. $3.95. Limnology and Oceanography, 17(1):167–168, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Staley:1972:BRR

[Sta72b] James T. Staley. Book review: Reinheimer, G. 1971. Mikrobiologie der Gewässer. Gustav Fischer, Jena. 184 p. 22.40M. Limnology and Oceanography, 17(2):325–326, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Stainton:1974:AMD

[Sta74a] M. P. Stainton. An automated method for determination of chloride and sulfate in freshwater using cation exchange and measurement of electrical conductance. Limnology and Oceanography, 20 (2):707–711, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Stanier:1974:BRF

[Sta74b] R. Y. Stanier. Book review: Fogg, G. E., W. D. P. Stewart, P. Fay, and A. E. Walsby, 1973. The blue-green algae. Academic Press, London and New York. vi + 459 p. $24.00. Limnology and Oceanography, 19(3):550–552, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Stanier:1975:BRS

[Sta75] R. Y. Stanier. Book review: Sieburth, J. M. 1975. Microbial seascapes. A pictorial essay on marine microorganisms and their environments. University Park Press, Baltimore. Unpaged. $9.50. Limnology and Oceanography, 20(3):1059–1060, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Starkweather:1976:ILR

[Sta76] Peter L. Starkweather. Influences of light regime on postembryonic development in two strains of Daphnia pulex. Limnology and Oceanography, 21(6):830–837, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Starkweather:1978:DVF

[Sta78] Peter L. Starkweather. Diel variation in feeding behavior of Daphnia pulex. Influences of food density and nutritional history on mandibular activity. Limnology and Oceanography, 23(2):307–317, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Sternberg:1970:BRW

[Ste70] Richard W. Sternberg. Book review: Weyl, P. K. 1970. Oceanography. An introduction to the marine environment. John Wiley, New York. xvii + 535 p. $12.50. Limnology and Oceanography, 15 (4):661–662, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Stearns:1971:M

[Ste71] Stephen C. Stearns. Murphy. Limnology and Oceanography, 16(6): 1000–1002, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Stearns:1972:M

[Ste72] Stephen C. Stearns. Models: 2. Limnology and Oceanography, 17 (3):500–501, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Stenson:1973:PHG

[Ste73a] Jan A. E. Stenson. On predation and Holopedium gibberum (Zaddach) distribution. Limnology and Oceanography, 18(6):1005– 1010, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Stewart:1973:DTV

[Ste73b] Kenton M. Stewart. Detailed time variations in mean temperature and heat content of some Madison lakes. Limnology and Oceanography, 18(2):218–226, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Stearns:1975:LRD

[Ste75] Stephen C. Stearns. Light responses of Daphnia pulex. Limnology and Oceanography, 20(4):564–570, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Stenson:1976:SPI

[Ste76] Jan A. E. Stenson. Significance of predator influence on composition of bosmina spp. populations. Limnology and Oceanography, 21(6):814–822, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Steele:1979:BRK

[Ste79a] John H. Steele. Book review: Kremer, J. N., and S. W. Nixon. 1977. A coastal marine ecosystem. Simulation and analysis. Springer-Verlag, New York, Heidelberg, and Berlin, xi + 217 p. $32.00. Limnology and Oceanography, 24(3):598, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Stewart:1979:BRK

[Ste79b] R. W. Stewart. Book review: Kamenkovich, V. M. 1977. Fundamentals of ocean dynamics [transl. R. Radok]. Elsevier Ocean-ogr. Ser., 16. Elsevier Sci. Publ. Co., Amsterdam and New York, x + 249 p. $49.00. Limnology and Oceanography, 24(2):400, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Smith:1973:OPC

[STG73] Raymond C. Smith, John E. Tyler, and Charles R. Goldman. Optical properties and color of Lake Tahoe and Crater Lake. Limnology and Oceanography, 18(2):189–199, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Sasseville:1974:LVI

[STND74] D. R. Sasseville, A. P. Takacs, S. A. Norton, and R. B. Davis. A large-volume interstitial water sediment squeezer for lake sediments. Limnology and Oceanography, 19(6):1001–1004, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Stromgren:1975:LMG

[Str75a] Tor Stromgren. Linear measurements of growth of shells using laser diffraction. Limnology and Oceanography, 20(4):845–848, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Stross:1975:BRP

[Str75b] Raymond G. Stross. Book review: Palmer, J. D. 1974. Biological clocks in marine organisms. Wiley-Interscience, New York, x + 173 p. $12.95. Limnology and Oceanography, 20(3):1061–1062, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Stromgren:1976:SGH

[Str76] Tor Stromgren. Skeleton growth of the hydrocoral Millepora complanata Lamarck in relation to light. Limnology and Oceanography, 21(1):156–160, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Strickler:1977:OSP

[Str77] J. R. Strickler. Observation of swimming performances of planktonic copepods. Limnology and Oceanography, 22(1):165–170, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Stumm:1971:BRR

[Stu71] Werner Stumm. Book review: Reley, J. P., and R. Chester. 1971. Introduction to marine chemistry. Academic Press, London and New York, xiv + 465 p. £6.00. $17.50. Limnology and Oceanography, 16(6):1006–1007, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Stocker:1972:FCM

[SW72] Z. S. J. Stocker and D. Dudley Williams. A freezing core method for describing the vertical distribution of sediments in a streambed. Limnology and Oceanography, 17(1):136–138, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Scheider:1973:AMC

[SW73] W. Scheider and P. Wallis. An alternate method of calculating the population density of monsters in Loch Ness. Limnology and Oceanography, 18(2):343–345, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Sweers:1970:ICC

[Swe70a] H. E. Sweers. An improved code to classify the location of marine and terrestrial data. Limnology and Oceanography, 15(5):827–832, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Sweers:1970:VDC

[Swe70b] H. E. Sweers. Vertical diffusivity coefficient in a thermocline. Limnology and Oceanography, 15(2):273–280, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Sweeney:1976:DFT

[Swe76] Bernard W. Sweeney. A diurnally fluctuating thermal system for studying the effect of temperature on aquatic organisms. Limnology and Oceanography, 21(5):758, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Sweeney:1978:BDR

[Swe78] Bernard W. Sweeney. Bioenergetic and developmental response of a mayfly to thermal variation. Limnology and Oceanography, 23 (3):461–477, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Smith:1978:NEO

[SWLH78] K. L. Smith, Jr., G. A. White, M. B. Laver, and J. A. Haugsness. Nutrient exchange and oxygen consumption by deep-sea benthic communities: Preliminary in situ measurements. Limnology and Oceanography, 23(5):997–1005, September 1978. CODEN LI-OCAH. ISSN 0024-3590.

Swallow:1978:PDC

[SWM^{+}78] K. C. Swallow, J. C. Westall, D. M. McKnight, N. M. L. Morel, and F. M. M. Morel. Potentiometric determination of copper complexation by phytoplankton exudates. Limnology and Oceanography, 23(3):538–542, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Tobin:1978:TTI

[TA78] Richard S. Tobin and D. H. J. Anthony. Tritiated thymidine incorporation as a measure of microbial activity in lake sediments.

Limnology and Oceanography, 23(1):161–165, January 1978. CO-DEN LIOCAH. ISSN 0024-3590.

Tabata:1979:VWP

[Tab79] S. Tabata. On the variability of water properties in the surface mixed layer at Station P. Limnology and Oceanography, 24(1): 158–163, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Taft:1976:RCH

[Taf76] J. L. Taft. Recovery of ^{14}C–CO_{2} at high flow rates from a CHN analyzer. Limnology and Oceanography, 21(1):161–164, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Taylor:1978:MP

[Tay78] D. Taylor. Multiple publication. Limnology and Oceanography, 23 (2):382, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Theodorsson:1975:ABM

[TB75] Pall Theodorsson and Jon Orn Bjarnason. The acid-bubbling method for primary productivity measurements modified and tested. Limnology and Oceanography, 20(3):1018–1019, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Twilley:1977:PAT

[TBD77] Robert R. Twilley, Mark M. Brinson, and Graham J. Davis. Phosphorus absorption, translocation, and secretion in Nuphar luteum. Limnology and Oceanography, 22(6):1022–1032, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Thompson:1975:MSL

[TBOO75] R. Thompson, R. W. Battarbee, P. E. O’Sullivan, and F. Oldfield. Magnetic susceptibility of lake sediments. Limnology and Oceanography, 20(5):687–698, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Thomas:1974:RNM

[TC74] N. D. Thomas and D. J. Clarke. Resonant and nonresonant motion in a spindle-shaped basin with an entrance. Limnology and Oceanography, 19(2):199–208, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Tett:1975:PQC

[TCTW75] Paul Tett, Janet C. Cottrell, D. O. Trew, and B. J. B. Wood. Phosphorus quota and the chlorophyll: carbon ratio in marine phytoplankton. Limnology and Oceanography, 20(4):587–603, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Thomas:1972:NDT

[TD72] William H. Thomas and Anne N. Dodson. On nitrogen deficiency in tropical Pacific Oceanic phytoplankton. II. Photosynthetic and cellular characteristics of a chemostat-grown diatom. Limnology and Oceanography, 17(4):515–523, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Teal:1978:BRC

[Tea78] John M. Teal. Book review: Chapman, V. J. [Ed.] 1977. Wet coastal ecosystems. Elsevier, Amsterdam and New York. xii + 428 p. $49.95. Limnology and Oceanography, 23(5):1086, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Takahashi:1970:PPG

[TeI70] Masayuki Takahashi and Shun ei Ichimura. Photosynthetic properties and growth of photosynthetic sulfur bacteria in lakes. Limnology and Oceanography, 15(6):929–944, November 1970. CO-DEN LIOCAH. ISSN 0024-3590.

Tempel:1973:IRT

[Tem73] Neal R. Tempel. An inexpensive, recording tide gauge. Limnology and Oceanography, 18(1):178–180, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Tenore:1977:GCC

[Ten77] Kenneth R. Tenore. Growth of Capitella capitata cultured on various levels of detritus derived from different sources. Limnology and Oceanography, 22(5):936–941, September 1977. CODEN LI-OCAH. ISSN 0024-3590.

Terrell:1972:ISV

[TG72] Robert E. Terrell and Theodore Green III. Investigations of the surface velocity structure of lake currents. Limnology and Oceanography, 17(1):158–160, January 1972. CODEN LIOCAH. ISSN 0024-3590.

Tucker:1977:TDM

[TG77] W. A. Tucker and A. W. Green. A time-dependent model of the lake-averaged, vertical temperature distribution of lakes. Limnology and Oceanography, 22(4):687–699, July 1977. CODEN LI-OCAH. ISSN 0024-3590.

Tett:1978:RAS

[TGK^{+}78] Paul Tett, Charles Gallegos, Mahlon G. Kelly, George M. Hornberger, and B. J. Cosby. Relationships among substrate, flow, and benthic microalgal pigment density in the Mechums River, Virginia. Limnology and Oceanography, 23(4):785–797, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Thorpe:1974:EKH

[TH74] S. A. Thorpe and A. J. Hall. Evidence of Kelvin–Helmholtz billows in Loch Ness. Limnology and Oceanography, 19(6):973–976, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Tenore:1979:EON

[THDW79] Kenneth R. Tenore, Roger B. Hanson, Bruce E. Dornseif, and Craig N. Wiederhold. The effect of organic nitrogen supplement on the utilization of different sources of detritus. Limnology and Oceanography, 24(2):350–355, March 1979. CODEN LIOCAH. ISSN 0024-3590.

Theodorsson:1975:SPF

[The75] Pall Theodorsson. The study of ^{14}C penetration into filters in primary productivity measurements using double side counting. Limnology and Oceanography, 19(4):288–291, March 1975. CO-DEN LIOCAH. ISSN 0024-3590.

Thomas:1970:EAN

[Tho70a] William H. Thomas. Effect of ammonium and nitrate concentration on chlorophyll increases in natural tropical Pacific phytoplankton populations. Limnology and Oceanography, 15(3):386– 394, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Thomas:1970:NDT

[Tho70b] William H. Thomas. On nitrogen deficiency in tropical Pacific Oceanic phytoplankton: photosynthetic parameters in poor and rich water. Limnology and Oceanography, 15(3):380–385, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Thomann:1977:CLP

[Tho77] Robert V. Thomann. Comparison of lake phytoplankton models and loading plots. Limnology and Oceanography, 22(2):370–373, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Thomann:1978:RCD

[Tho78] R. V. Thomann. Reply to comment by D. W. Schindler. Limnology and Oceanography, 23(5):1082–1083, September 1978. CODEN LIOCAH. ISSN 0024-3590. See [Sch78b].

Threlkeld:1979:ECB

[Thr79] Stephen T. Threlkeld. Estimating cladoceran birth rates: The importance of egg mortality and the egg age distribution. Limnology and Oceanography, 24(4):601–612, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Tilzer:1973:DPP

[Til73] Max M. Tilzer. Diurnal periodicity in the phytoplankton assemblage of a high mountain lake. Limnology and Oceanography, 18 (1):15–30, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Timms:1972:MLA

[Tim72] B. V. Timms. A meromictic lake in Australia. Limnology and Oceanography, 17(6):918–922, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Timms:1975:MCV

[Tim75] B. V. Timms. Morphometric control of variation in annual heat budgets. Limnology and Oceanography, 20(1):110–112, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Titman:1975:FTM

[Tit75] David Titman. A fluorometric technique for measuring sinking rates of freshwater phytoplankton. Limnology and Oceanography, 20(4):869–875, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Tuttle:1972:OTT

[TJ72] Jon H. Tuttle and Holger W. Jannasch. Occurrence and types of thiobacillus-like bacteria in the sea. Limnology and Oceanography, 17(4):532–543, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Tuttle:1976:MUT

[TJ76] Jon H. Tuttle and Holger W. Jannasch. Microbial utilization of thiosulfate in the deep sea. Limnology and Oceanography, 21(5): 697–701, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Tuttle:1979:MDA

[TJ79] Jon H. Tuttle and Holger W. Jannasch. Microbial dark assimilation of CO_{2} in the Cariaco Trench. Limnology and Oceanography, 24(4):746–753, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Titman:1976:SFP

[TK76] David Titman and Peter Kilham. Sinking in freshwater phytoplankton: Some ecological implications of cell nutrient status and physical mixing processes. Limnology and Oceanography, 21(3): 409–417, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Turekian:1973:TET

[TKC73] Karl K. Turekian, Amitai Katz, and Lui Chan. Trace element trapping in pteropod tests. Limnology and Oceanography, 18(2): 240–249, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Tett:1975:MSM

[TKH75] Paul Tett, Mahlon G. Kelly, and George M. Hornberger. A method for the spectrophotometric measurement of chlorophyll-a and pheophytin a in benthic microalgae. Limnology and Oceanography, 20(4):887–896, September 1975. CODEN LIOCAH. ISSN 0024-3590.

Tett:1977:ECP

[TKH77] Paul Tett, Mahlon G. Kelly, and George M. Hornberger. Estimation of chlorophyll-a and pheophytin-a in methanol. Limnology and Oceanography, 22(3):579–580, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Tilman:1976:MCA

[TKK76] David Tilman, Susan Soltau Kilham, and Peter Kilham. Morphometric changes in Asterionella formosa colonies under phosphate and silicate limitation. Limnology and Oceanography, 21(6):883– 886, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Toth:1975:SLO

[TL75] David J. Toth and Abraham Lerman. Stratified lake and oceanic brines: Salt movement and time limits of existence. Limnology and Oceanography, 20(4):715–728, September 1975. CODEN LI-OCAH. ISSN 0024-3590.

Torrey:1976:NFL

[TL76] Marguerite Sherman Torrey and G. F. Lee. Nitrogen fixation in Lake Mendota, Madison, Wisconsin. Limnology and Oceanography, 21(3):365–378, May 1976. CODEN LIOCAH. ISSN 0024-3590.

Therriault:1978:SVP

[TLP78] Jean-Claude Therriault, Donald J. Lawrence, and Trevor Piatt. Spatial variability of phytoplankton turnover in relation to physical processes in a coastal environment. Limnology and Oceanography, 23(5):900–911, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Tietjen:1970:GCP

[TLR^{+}70] J. H. Tietjen, J. J. Lee, J. Rullman, A. Greengart, and J. Trompeter. Gnotobiotic culture and physiological ecology of the marine nematode Rhabditis marina Bastian. Limnology and Oceanography, 15(4):535–543, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Taft:1977:PUP

[TLT77] Jay L. Taft, Michael E. Loftus, and W. Rowland Taylor. Phosphate uptake from phosphomonoesters by phytoplankton in the Chesapeake Bay. Limnology and Oceanography, 22(6):1012–1021, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Toetz:1971:DUN

[Toe71] Dale W. Toetz. Diurnal uptake of N0_{3} and NH_{4} by a ceratophyllum-periphyton community. Limnology and Oceanography, 18(5):819–822, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Tonolli:1971:ARA

[Ton71] Livia Tonolli. Allomorphotic relationship of antenna to body length in eudiaptomus vulgaris schmeil. Limnology and Oceanography, 16(2):467–472, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Tonolli:1972:BRE

[Ton72] Livia Tonolli. Book review: Edmondson, W. T., and G. G. Winberg [Eds.]. 1971. A manual on methods for the assessment of secondary productivity in fresh waters. IBP Handbook No. 17. F.

A. Davis, Philadelphia (Blackwell Sci. Publ., Oxford and Edinburgh). xxiv + 358 p. $13.50. Limnology and Oceanography, 17 (5):800–801, September 1972. CODEN LIOCAH. ISSN 0024-3590.

Towle:1973:PGK

[TP73] David W. Towle and John S. Pearse. Production of the giant kelp, Macrocystis, estimated by in situ incorporation of ^{14}C in polyethylene bags. Limnology and Oceanography, 18(1):155–159, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Tunzi:1974:CAC

[TP74] Milton G. Tunzi and Donald B. Porcella. Carbon-14 assimilation, chlorophyll, and particulate organic matter in steady state systems at Lake Tahoe. Limnology and Oceanography, 19(3):420–428, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Therriault:1978:SHP

[TP78] Jean-Claucle Therriault and Trevor Piatt. Spatial heterogeneity of phytoplankton biomass and related factors in the near-surface waters of an exposed coastal embayment. Limnology and Oceanography, 23(5):888–899, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Tilzer:1977:SVA

[TPG77] Max M. Tilzer, Hans W. Paerl, and Charles R. Goldman. Sustained viability of aphotic phytoplankton in Lake Tahoe (california-nevada). Limnology and Oceanography, 22(1):84–91, January 1977. CODEN LIOCAH. ISSN 0024-3590.

Topinka:1976:ENA

[TR76] Jerry A. Topinka and Jon V. Robbins. Effects of nitrate and ammonium enrichment on growth and nitrogen physiology in Fucus spiralis. Limnology and Oceanography, 21(5):659–664, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Tyler:1978:AST

[TS78] Mary Altalo Tyler and H. H. Seliger. Annual subsurface transport of a red tide dinoflagellate to its bloom area: Water circulation patterns and organism distributions in the Chesapeake Bay. Limnology and Oceanography, 23(2):227–246, March 1978. CODEN LIOCAH. ISSN 0024-3590.

Taghon:1978:PPS

[TSJ78] Gary L. Taghon, Robert F. L. Self, and Peter A. Jumars. Predicting particle selection by deposit feeders: a model and its implications. Limnology and Oceanography, 23(4):752–759, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Torgersen:1977:NMP

[TTC^{+}77] T. Torgersen, Z. Top, W. B. Clarke, W. J. Jenkins, and W. S. Broecker. A new method for physical limnology — tritium-helium-3 ages — results for Lakes Erie, Huron, and Ontario. Limnology and Oceanography, 22(2):181–193, March 1977. CODEN LI-OCAH. ISSN 0024-3590.

Thiel:1975:UTS

[TTW75] Hjalmar Thiel, David Thistle, and George D. Wilson. Ultrasonic treatment of sediment samples for more efficient sorting of meiofauna. Limnology and Oceanography, 20(6):472–473, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Turner:1978:CPR

[Tur78] R. Eugene Turner. Community plankton respiration in a salt marsh estuary and the importance of macrophytic leachates. Limnology and Oceanography, 23(3):442–451, May 1978. CODEN LI-OCAH. ISSN 0024-3590.

Teal:1979:NFR

[TVB79] John M. Teal, Ivan Valiela, and Diane Berlo. Nitrogen fixation by rhizosphere and free-living bacteria in salt marsh sediments. Limnology and Oceanography, 24(1):126–132, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Thayer:1975:LCQ

[TWPC75] Gordon W. Thayer, Richard B. Williams, Thomas J. Price, and David R. Colby. A large corer for quantitatively sampling benthos in shallow water. Limnology and Oceanography, 20(6):474–481, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Tyler:1971:BRK

[Tyl71] John E. Tyler. Book review: Kinne, O. [Ed.] 1970. Marine ecology, v. 1. Environmental factors, part 1. A comprehensive, integrated treatise on life in oceans and coastal waters. Wiley-Interscience, New York, xv + 681 p. $34.50. Limnology and Oceanography, 18 (5):841–842, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Tyler:1973:LVQ

[Tyl73] John E. Tyler. Lux vs. quanta. Limnology and Oceanography, 16 (5):810, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Tyler:1975:SQE

[Tyl75] John E. Tyler. The in situ quantum efficiency of natural phytoplankton populations. Limnology and Oceanography, 20(4):976– 980, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Vallentyne:1971:BRB

[Val71a] J. R. Vallentyne. Book review: Brinkhurst, R. O., and D. A. Chant. 1971. This good, good earth: Our fight for survival. The Macmillan Company of Canada, Toronto, x + 174 p. $6.95. Limnology and Oceanography, 16(6):1005, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Vallentyne:1971:DHL

[Val71b] J. R. Vallentyne. A description of Hutchinson Lake, Ontario. Limnology and Oceanography, 16(2):473–478, March 1971. CODEN LIOCAH. ISSN 0024-3590.

Vallentyne:1974:ATL

[Val74a] J. R. Vallentyne. Authors and titles in Limnology and Oceanography volume 19, 1974. Limnology and Oceanography, 19(5):1018– 1023, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Vallentyne:1974:BRW

[Val74b] J. R. Vallentyne. Book review: Who’s who in ecology 1973. 1973. Special Reports, Inc., 8 West 40th St., New York. 291 p. $35.00. Limnology and Oceanography, 19(5):875–876, September 1974. CODEN LIOCAH. ISSN 0024-3590.

VanSickle:1977:MES

[Van77] John Van Sickle. Mortality estimates from size distributions: a critique of Smith’s model. Limnology and Oceanography, 22(4): 774–775, July 1977. CODEN LIOCAH. ISSN 0024-3590.

Vanderborght:1975:VDN

[VB75] Jean-Pierre Vanderborght and Gilles Billen. Vertical distribution of nitrate concentration in interstitial water of marine sediments with nitrification and denitrification. Limnology and Oceanography, 20(6):953–961, November 1975. CODEN LIOCAH. ISSN 0024-3590.

VonDamm:1979:CCM

[VBT79] Karen L. Von Damm, Lurry K. Benninger, and Karl K. Turekian. The^{210}Pb chronology of a core from Mirror Lake, New Hampshire. Limnology and Oceanography, 24(3):434–439, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Venrick:1971:SS

[Ven71] E. L. Venrick. The statistics of subsampling. Limnology and Oceanography, 18(5):811–818, September 1971. CODEN LI-OCAH. ISSN 0024-3590.

Venrick:1974:DSR

[Ven74] E. L. Venrick. The distribution and significance of Richelia intracellularis Schmidt in the North Pacific central gyre. Limnology and Oceanography, 19(3):437–445, May 1974. CODEN LIOCAH. ISSN 0024-3590.

Verduin:1975:RCD

[Ver75] Jacob Verduin. Rate of carbon dioxide transport across air-water boundaries in lakes. Limnology and Oceanography, 20(3):1052– 1053, November 1975. CODEN LIOCAH. ISSN 0024-3590. See reply [Meg75].

Vanderhoef:1974:NFA

[VHMW74] Larry N. Vanderhoef, Chi-Ying Huang, Robert Musil, and Jeffery Williams. Nitrogen fixation (acetylene reduction) by phytoplankton in Green Bay, Lake Michigan, in relation to nutrient concentrations. Limnology and Oceanography, 19(1):119–125, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Vlymen:1970:EES

[Vly70] William J. Vlymen. Energy expenditure of swimming copepods. Limnology and Oceanography, 15(3):348–356, May 1970. CODEN LIOCAH. ISSN 0024-3590.

Vlymen:1977:RCJ

[Vly77] William J. Vlymen. Reply to comment by j. t. enright. Limnology and Oceanography, 22(1):162–165, January 1977. CODEN LIOCAH. ISSN 0024-3590. See [Enr77c].

Vuceta:1977:HCI

[VM77] Jasenka Vuceta and James J. Morgan. Hydrolysis of Cu(II). Limnology and Oceanography, 22(4):742–746, July 1977. CODEN LI-OCAH. ISSN 0024-3590.

Valiela:1976:PDE

[VTP76] Ivan Valiela, John M. Teal, and Norma Y. Persson. Production and dynamics of experimentally enriched salt marsh vegetation: Belowground biomass. Limnology and Oceanography, 21(2):245– 252, March 1976. CODEN LIOCAH. ISSN 0024-3590.

Valiela:1978:NPF

[VTV^{+}78] Ivan Valiela, John M. Teal, Suzanne Volkmann, Deborah Shafer, and Edward J. Carpenter. Nutrient and particulate fluxes in a salt marsh ecosystem: Tidal exchanges and inputs by precipitation and groundwater. Limnology and Oceanography, 23(4):798–812, July 1978. CODEN LIOCAH. ISSN 0024-3590.

VanRaalte:1976:PES

[VVT76] Charlene D. Van Raalte, Ivan Valiela, and John M. Teal. Production of epibenthic salt marsh algae: Light and nutrient limitation. Limnology and Oceanography, 21(6):862–872, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Vanderborght:1977:KMDa

[VWB77a] Jean-Pierre Vanderborght, Roland Wollas, and Gilles Bitten. Kinetic models of diagenesis in disturbed sediments. Part 1. Mass transfer properties and silica diagenesis. Limnology and Oceanography, 22(5):787–793, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Vanderborght:1977:KMDb

[VWB77b] Jean-Pierre Vanderborght, Roland Wollas, and Gilles Bitten. Kinetic models of diagenesis in disturbed sediments. Part 2. Nitrogen diagenesis. Limnology and Oceanography, 22(5):794–803, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Wium-Andersek:1972:CDC

[WAA72a] S. Wium-Andersek and Jens M. Andersen. Carbon dioxide content of the interstitial water in the sediment of Grane Langsø, a Danish Lobelia lake. Limnology and Oceanography, 17(6):943–947, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Wium-Andersen:1972:IVR

[WAA72b] S. Wium-Andersen and Jens M. Andersen. The influence of vegetation on the redox profile of the sediment of Grane Langsø, a Danish Lobelia lake. Limnology and Oceanography, 17(6):948–952, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Wakeham:1977:HBL

[Wak77] Stuart G. Wakeham. Hydrocarbon budgets for Lake Washington. Limnology and Oceanography, 22(5):952–957, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Walsby:1973:PAM

[Wal73] A. E. Walsby. A portable apparatus for measuring relative gas vacuolation, the strength of gas vacuoles, and turgor pressure in planktonic blue-green algae and bacteria. Limnology and Oceanography, 18(4):653–658, July 1973. CODEN LIOCAH. ISSN 0024-3590.

Walker:1974:SML

[Wal74a] K. F. Walker. The stability of meromictic lakes in central Washington. Limnology and Oceanography, 19(2):209–222, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1974:BRS

[Wal74b] John J. Walsh. Book review: Steele, J. H. 1974. The structure of marine ecosystems. Harvard University Press, Cambridge, Massachusetts, x + 128 p. $7.95. Limnology and Oceanography, 19(6): 1015–1016, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Walker:1975:SPC

[Wal75] K. F. Walker. The seasonal phytoplankton cycles of two saline lakes in central Washington. Limnology and Oceanography, 20(1): 40–53, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1976:BRS

[Wal76a] John J. Walsh. Book review: Steemann Nielsen, E. 1975. Marine photosynthesis with special emphasis on the ecological aspects. Elsevier Sci. Publ. Co., Amsterdam and New York, x + 142 p. $21.75. Limnology and Oceanography, 21(4):633, July 1976. CO-DEN LIOCAH. ISSN 0024-3590.

Walsh:1976:HFP

[Wal76b] John J. Walsh. Herbivory as a factor in patterns of nutrient utilization in the sea. Limnology and Oceanography, 21(1):1–13, January 1976. CODEN LIOCAH. ISSN 0024-3590.

Wangersky:1972:CSP

[Wan72a] Peter J. Wangersky. The control of seawater pH by ion pairing. Limnology and Oceanography, 17(1):1–6, January 1972. CODEN LIOCAH. ISSN 0024-3590. See comments [Pyt72].

Wangersky:1972:IPP

[Wan72b] Peter J. Wangersky. Ion pairing and pH: a reply. Limnology and Oceanography, 17(6):960–962, November 1972. CODEN LIOCAH. ISSN 0024-3590. See [Kes72].

Wangersky:1974:POC

[Wan74] Peter J. Wangersky. Particulate organic carbon: Sampling variability. Limnology and Oceanography, 19(6):980–984, November 1974. CODEN LIOCAH. ISSN 0024-3590.

Warshaw:1972:EAA

[War72] Stephen J. Warshaw. Effects of alewives (Alosa pseudoharengus) on the zooplankton of Lake Wononskopomuc, Connecticut. Limnology and Oceanography, 17(6):816–825, November 1972. CO-DEN LIOCAH. ISSN 0024-3590.

Wartenberg:1978:SEI

[War78] Daniel E. Wartenberg. Spectrophotometric equations: an intercalibration technique. Limnology and Oceanography, 23(3):566–570, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Ward:1979:STW

[War79] Peter R. B. Ward. Seiches, tides, and wind set-up on Lake Kariba. Limnology and Oceanography, 24(1):151–157, January 1979. CO-DEN LIOCAH. ISSN 0024-3590.

Worswick:1974:EAM

[WB74] Joseph M. Worswick, Jr. and Michael T. Barbour. An elntriation apparatus for macroinvertebrates. Limnology and Oceanography, 19(3):538–540, May 1974. CODEN LIOCAH. ISSN 0024-3590.

White:1977:DMA

[WBM^{+}77] David C. White, Ronald J. Bobbie, Susan J. Morrison, Darlene K. Oosterhof, Cecile W. Taylor, and Duane A. Meeter. Determination of microbial activity of estuarine detritus by relative rates of lipid biosynthesis. Limnology and Oceanography, 22(6):1089–1098, November 1977. CODEN LIOCAH. ISSN 0024-3590.

Waters:1973:APS

[WC73] Thomas F. Waters and Gary W. Crawford. Annual production of a stream mayfly population: a comparison of methods. Limnology and Oceanography, 18(2):286–296, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Wakeham:1976:AHS

[WC76] Stuart G. Wakeham and Roy Carpenter. Aliphatic hydrocarbons in sediments of Lake Washington. Limnology and Oceanography, 21(5):711–723, September 1976. CODEN LIOCAH. ISSN 0024-3590.

Wolf:1971:ASM

[WCR71] E. G. Wolf, C. E. Cushing, and F. W. Rabe. An automated system for multiple recording of diurnal pH. Limnology and Oceanography, 16(3):577–580, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Whitledge:1972:CS

[WD72] T. E. Whitledge and R. C. Dugdale. Creatine in seawater. Limnology and Oceanography, 17(2):309–314, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Wallace:1978:TPO

[WD78] Gordon T. Wallace, Jr. and Robert A. Duce. Transport of particulate organic matter by bubbles in marine waters. Limnology and Oceanography, 23(6):1155–1167, November 1978. CODEN LI-OCAH. ISSN 0024-3590.

Whitney:1979:MDC

[WD79] David E. Whitney and W. Marshall Darley. A method for the determination of chlorophyll-a in samples containing degradation products. Limnology and Oceanography, 24(1):183–186, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Webb:1975:EEA

[WDW^{+}75] K. L. Webb, W. D. DuPaul, W. Wiebe, W. Sottile, and R. E. Johannes. Enewetak (Eniwetok) Atoll: Aspects of the nitrogen cycle on a coral reef. Limnology and Oceanography, 20(2):198– 210, March 1975. CODEN LIOCAH. ISSN 0024-3590.

Weaver:1979:CFI

[Wea79] Sylvia S. Weaver. Ceratium in Fire Island Inlet, Long Island, New York (1971–1977). Limnology and Oceanography, 24(3):553–558, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Webster:1974:ISS

[Web74] R. M. Webster. An inexpensive sampler for studying the microbiology of river mud. Limnology and Oceanography, 19(1):159–161, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Weiler:1973:IWC

[Wei73a] R. R. Weiler. The interstitial water composition in the sediments of the Great Lakes. I. Western Lake Ontario. Limnology and Oceanography, 18(6):918–931, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Weiss:1973:SBL

[Wei73b] Richard L. Weiss. Survival of bacteria at low pH and high temperature. Limnology and Oceanography, 18(6):877–883, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Wells:1970:EAP

[Wel70] LaRue Wells. Effects of alewife predation on zooplankton populations in Lake Michigan. Limnology and Oceanography, 15(4): 556–565, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Welch:1974:MRA

[Wel74] Harold E. Welch. Metabolic rates of Arctic lakes. Limnology and Oceanography, 19(1):65–73, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Welch:1972:RPP

[WEMD72] Eugene B. Welch, Richard M. Emery, Robert I. Matsuda, and William A. Dawson. The relation of periphytic and planktonic algal growth in an estuary to hydrographic factors. Limnology and Oceanography, 17(5):731–737, September 1972. CODEN LI-OCAH. ISSN 0024-3590.

Wetzel:1970:RPP

[Wet70] Robert G. Wetzel. Recent and postglacial production rates of a marl lake. Limnology and Oceanography, 15(4):491–503, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Wetzel:1975:BRR

[Wet75] Robert G. Wetzel. Book review: Rodewald-Rudescu, L. 1974. Das Schilfrohr. Phragmites communis Trinius. Die Binnengewasser Band 27. E. Schweizerbart’sche, Stuttgart, ix + 302 p. + 11 plates. DM97.50. Limnology and Oceanography, 20(3):1059, November 1975. CODEN LIOCAH. ISSN 0024-3590.

Wetzel:1976:BRH

[Wet76] Robert G. Wetzel. Book review: Hutchinson, G. E. 1975. A treatise on limnology, v. 3. Limnological botany. John Wiley & Sons, New York, London, Sydney, and Toronto, xii + 660 p. $30.00. Limnology and Oceanography, 21(6):929–930, November 1976. CO-DEN LIOCAH. ISSN 0024-3590.

Wiegert:1972:EYT

[WF72] Richard G. Wiegert and Peter C. Fraleigh. Ecology of Yellowstone thermal effluent systems: net primary production and species diversity of a successional blue-green algal mat. Limnology and Oceanography, 17(2):215–228, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Wada:1971:NME

[WH71] Eitaro Wada and Akihiko Hatton. Nitrite metabolism in the euphotic layer of the central north Pacific Ocean. Limnology and Oceanography, 18(5):766–772, September 1971. CODEN LI-OCAH. ISSN 0024-3590.

Wright:1978:CSN

[WH78] Richard F. Wright and Arne Henriksen. Chemistry of small Norwegian lakes, with special reference to acid precipitation. Limnology and Oceanography, 23(3):487–498, May 1978. CODEN LI-OCAH. ISSN 0024-3590.

Whetten:1973:BRP

[Whe73] John T. Whetten. Book review: Pettijohn, F. J., P. E. Potter, and R. Siever. 1972. Sand and sandstone. Springer-Verlag, New York. xvi + 618 p. $31.10. Limnology and Oceanography, 18(1): 182–183, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Wheeler:1975:FCS

[Whe75] John R. Wheeler. Formation and collapse of surface films. Limnology and Oceanography, 20(3):338–342, May 1975. CODEN LI-OCAH. ISSN 0024-3590.

Wheeler:1976:FMW

[Whe76] John R. Wheeler. Fractionation by molecular weight of organic substances in Georgia coastal water. Limnology and Oceanography, 21(6):846–852, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Wheeler:1977:DOC

[Whe77] John R. Wheeler. Dissolved organic carbon: Spectral relationships in coastal waters. Limnology and Oceanography, 22(3):573–575, May 1977. CODEN LIOCAH. ISSN 0024-3590.

Whitfield:1971:CPS

[Whi71] M. Whitfield. A compact potentiometric sensor of novel design. In Situ determination of pH, pS^{2−}, and Eh. Limnology and Oceanography, 18(5):829, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Whitfield:1972:ECN

[Whi72] M. Whitfield. The electrochemical characteristics of natural redox cells. Limnology and Oceanography, 17(3):383–393, May 1972. CODEN LIOCAH. ISSN 0024-3590.

Whitfield:1974:TLU

[Whi74a] M. Whitfield. Thermodynamic limitations on the use of the platinum electrode in Eh measurements. Limnology and Oceanography, 19(5):857–865, September 1974. CODEN LIOCAH. ISSN 0024-3590.

Whitfield:1974:IAM

[Whi74b] Michael Whitfield. The ion-association model and the buffer capacity of the carbon dioxide system in seawater at 25^{o}C and 1 atmosphere total pressure. Limnology and Oceanography, 19(2): 235–248, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Whitfield:1975:EIA

[Whi75] M. Whitfield. The effect of increases in the atmospheric carbon dioxide content on the carbonate ion concentration of surface ocean water at 25^{o}C. Limnology and Oceanography, 20(1): 103–108, January 1975. CODEN LIOCAH. ISSN 0024-3590.

Wickstrom:1971:CGE

[Wic71] Conrad E. Wickstrom. The credibility gap and the ecology reaction. Limnology and Oceanography, 16(6):998–1000, November 1971. CODEN LIOCAH. ISSN 0024-3590.

Wiebe:1970:SSS

[Wie70] Peter Howard Wiebe. Small-scale spatial distribution in oceanic zooplankton. Limnology and Oceanography, 15(2):205–217, March 1970. CODEN LIOCAH. ISSN 0024-3590.

Wiebe:1971:CMS

[Wie71] Peter H. Wiebe. A computer model study of zooplankton patchiness and its effects on sampling error. Limnology and Oceanography, 16(1):29–38, January 1971. CODEN LIOCAH. ISSN 0024-

Wilson:1972:PNA

[Wil72a] D. O. Wilson. Phosphate nutrition of the aquatic agiosperm, Myriophyllum exalbescens Fern. Limnology and Oceanography, 17 (4):612–616, July 1972. CODEN LIOCAH. ISSN 0024-3590.

Wilson:1972:CMA

[Wil72b] Mildred Stratton Wilson. Copepods of marine affinities from mountain lakes of western North America. Limnology and Oceanography, 17(5):762–763, September 1972. CODEN LI-OCAH. ISSN 0024-3590.

Williams:1973:VAS

[Wil73] P. J. Leb. Williams. The validity of the application of simple kinetic analysis to heterogeneous microbial populations. Limnology and Oceanography, 18(1):159–165, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Wilson:1978:EMD

[Wil78] Donald E. Wilson. An equilibrium model describing the influence of humic materials on the speciation of Cu^{2+}, Zn^{2+}, and Mn^{2+} in freshwaters. Limnology and Oceanography, 23(3):499–507, May 1978. CODEN LIOCAH. ISSN 0024-3590.

Windom:1971:FCC

[Win71] Herbert L. Windom. Fluoride concentration in coastal and estuarine waters of Georgia. Limnology and Oceanography, 18(5): 806–810, September 1971. CODEN LIOCAH. ISSN 0024-3590.

Wilson:1977:EPC

[WK77] Donald E. Wilson and Patrick Kinney. Effects of polymeric charge variations on the proton-metal ion equilibria of humic materials. Limnology and Oceanography, 22(2):281–289, March 1977. CO-DEN LIOCAH. ISSN 0024-3590.

Wilson:1979:RSM

[WK79] W. H. Wilson and D. A. Kiefer. Reflectance spectroscopy of marine phytoplankton. Part 2. A simple model of ocean color. Limnology and Oceanography, 24(4):673–682, July 1979. CODEN LI-OCAH. ISSN 0024-3590.

Walsh:1974:SBC

[WKW^{+}74] John J. Walsh, James C. Kelley, Terry E. Whitledge, Jane J. MacIsaac, and Susan A. Huntsman. Spin-up of the Baja California upwelling ecosystem. Limnology and Oceanography, 19(4):553– 572, July 1974. CODEN LIOCAH. ISSN 0024-3590.

Weimer:1973:SCC

[WL73] Walter C. Weimer and F. Fred Lee. Some considerations of the chemical limnology of meromictic Lake Mary. Limnology and Oceanography, 18(3):414–425, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Wefer:1978:CPB

[WL78] Gerold Wefer and Gerhard F. Lutze. Carbonate production by benthic Foraminifera and accumulation in the western Baltic. Limnology and Oceanography, 23(5):992–996, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Welton:1979:TST

[WL79] J. S. Welton and M. Ladle. Two sediment trap designs for use in small rivers and streams. Limnology and Oceanography, 24(3): 588–592, May 1979. CODEN LIOCAH. ISSN 0024-3590.

Wetzel:1972:DDO

[WM72a] Robert G. Wetzel and Bruce A. Manny. Decomposition of dissolved organic carbon and nitrogen compounds from leaves in an experimental hard-water stream. Limnology and Oceanography, 17 (6):927–931, November 1972. CODEN LIOCAH. ISSN 0024-3590.

Wulff:1972:LSA

[WM72b] Barry L. Wulff and C. David McIntire. Laboratory studies of assemblages of attached estuarine diatoms. Limnology and Oceanography, 17(2):200–214, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Williams:1978:SSS

[WM78] D. Dudley Williams and J. H. Mundie. Substrate size selection by stream invertebrates and the influence of sand. Limnology and Oceanography, 23(5):1030–1033, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Wilcox:1974:POS

[WMM74] J. Ross Wilcox, Robert P. Meek, and David Mook. A pneumatically operated slurp gun. Limnology and Oceanography, 19(2): 354–355, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Ward:1971:CGM

[WN71] F. J. Ward and Masami Nakanishi. A comparison of geiger-mueller and liquid scintillation counting methods in estimating primary productivity. Limnology and Oceanography, 16(3):560–563, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Wheeler:1977:UGN

[WNLS77] P. Wheeler, B. North, M. Littler, and G. Stephens. Uptake of glycine by natural phytoplankton communities. Limnology and Oceanography, 22(5):900–910, September 1977. CODEN LI-OCAH. ISSN 0024-3590.

Wheeler:1974:AAU

[WNS74] Patricia A. Wheeler, Barbara B. North, and Grover C. Stephens. Amino acid uptake by marine phytoplankters. Limnology and Oceanography, 19(2):249–259, March 1974. CODEN LIOCAH. ISSN 0024-3590.

Wolfe:1971:FCC

[Wol71] Douglas A. Wolfe. Fallout cesium-137 in clams (Rangia cuneata) from the Neuse River Estuary, North Carolina. Limnology and Oceanography, 18(5):797–805, September 1971. CODEN LI-OCAH. ISSN 0024-3590.

Wong:1979:APS

[Won79] George T. F. Wong. Alkalinity and pH in the southern Chesapeake Bay and the James River estuary. Limnology and Oceanography, 24(5):970–977, September 1979. CODEN LIOCAH. ISSN 0024-3590.

Woodcock:1971:NCP

[Woo71] Alfred H. Woodcock. Note concerning physalia behavior at sea. Limnology and Oceanography, 16(3):551–552, May 1971. CODEN LIOCAH. ISSN 0024-3590.

Woodin:1973:BRV

[Woo73] S. A. Woodin. Book review: Vagners, J. [Supervisor], and P. Mar [Coordinator]. 1972. Oil on Puget Sound. An interdisciplinary study in systems engineering. Univ. Washington Press, Seattle and London. xvii + 648 p. $10.00. Limnology and Oceanography, 18 (3):508–509, May 1973. CODEN LIOCAH. ISSN 0024-3590.

Woodwell:1977:BRN

[Woo77] G. M. Woodwell. Book review: Nriagu, J. O. [Ed.] 1976. Environmental biogeochemistry. V. 1: Carbon, nitrogen, phosphorus, sulfur and selenium cycles. V. 2: Metals transfer and ecological mass balances. Ann Arbor Sci. Publ., Inc., Ann Arbor, Mich. 797 p. $29.50. Limnology and Oceanography, 22(5):971, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Wooster:1979:BRD

[Woo79] Warren S. Wooster. Book review: Deacon, M. B. [Ed.] 1978. Oceanography: Concepts and history. Benchmark Papers Geol.

35. Dowden, Hutchinson, and Ross, Strouds-burg. xvii + 394 p. $25.00. Limnology and Oceanography, 24(4):792–793, July 1979. CODEN LIOCAH. ISSN 0024-3590.

Wormuth:1977:FTH

[Wor77] John H. Wormuth. A flow-through heavy fraction zooplankton sorting column. Limnology and Oceanography, 22(5):945–948, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Wright:1976:VTS

[WP76] W. R. Wright and C. E. Parker. A volumetric temperature/ salinity census for the Middle Atlantic Bight. Limnology and Oceanography, 21(4):563–571, July 1976. CODEN LIOCAH. ISSN 0024-3590.

Webster:1978:SSD

[WP78] Katherine Emerson Webster and Robert Henry Peters. Some size-dependent inhibitions of larger cladoceran filterers in filamentous suspensions. Limnology and Oceanography, 23(6):1238–1245, November 1978. CODEN LIOCAH. ISSN 0024-3590.

Wyatt:1973:AME

[WQ73] T. Wyatt and S. Z. Qasim. Application of a model to an estuarine ecosystem. Limnology and Oceanography, 18(2):301–306, March 1973. CODEN LIOCAH. ISSN 0024-3590.

Wright:1975:SGA

[Wri75] Richard T. Wright. Studies on glycolic acid metabolism by freshwater bacteria. Limnology and Oceanography, 20(6):626–633, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Warnke:1973:CNE

[WRO73] Detlef A. Warnke, Joseph Richter, and Carl H. Oppenheimer. Characteristics of the nearshore environment off the south coast of Anvers Island, Antarctic Peninsula. Limnology and Oceanography, 18(1):131–142, January 1973. CODEN LIOCAH. ISSN 0024-3590.

Wooster:1970:SVT

[WS70] Warren S. Wooster and Hellmuth A. Sievers. Seasonal variations of temperature, drift, and heat exchange in surface waters off the west coast of South America. Limnology and Oceanography, 15 (4):595–605, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Watt:1979:AOC

[WSR79] W. D. Watt, D. Scott, and S. Ray. Acidification and other chemical changes Halifax County lakes after 21 years. Limnology and Oceanography, 24(6):1154–1161, November 1979. CODEN LI-OCAH. ISSN 0024-3590.

Wu:1979:BRA

[Wu79] T. Y. Wu. Book review: Aleyev, Yu. G. 1977. Nekton. Dr. W. Junk, The Hague, vi + 435 p. f.120. Limnology and Oceanography, 24(6):1172–1173, November 1979. CODEN LIOCAH. ISSN 0024-3590.

Wunsch:1973:MDLa

[Wun73a] Carl Wunsch. On the mean drift in large lakes. Limnology and Oceanography, 18(5):793–794, September 1973. CODEN LI-OCAH. ISSN 0024-3590.

Wunsch:1973:MDLb

[Wun73b] Carl Wunsch. On the mean drift in large lakes. Limnology and Oceanography, 18(5):798–800, September 1973. CODEN LI-OCAH. ISSN 0024-3590.

White:1973:MST

[WW73] W. Sedgefield White and Robert G. Wetzel. A modified sedimentation trap. Limnology and Oceanography, 18(6):986–988, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Williams:1974:CTU

[WW74] D. Dudley Williams and N. E. Williams. A counterstaining technique for use in sorting benthic samples. Limnology and Oceanography, 19(1):152–154, January 1974. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1978:WEF

[WWB^{+}78] John J. Walsh, Terry E. Whitledge, Frank W. Barvenik, Creighton D. Wirick, Steven O. Howe, Wayne E. Esaias, and Jon T. Scott. Wind events and food chain dynamics within the New York Bight. Limnology and Oceanography, 23(4):659–683, July 1978. CODEN LIOCAH. ISSN 0024-3590.

Walker:1970:MMO

[WWH70] K. F. Walker, W. D. Williams, and U. T. Hammer. The Miller method for oxygen determination applied to saline lakes. Limnology and Oceanography, 15(5):814–815, September 1970. CODEN LIOCAH. ISSN 0024-3590.

Whitney:1975:NFF

[WWH75] D. E. Whitney, G. M. Woodwell, and R. W. Howarth. Nitrogen fixation in flax pond: a Long Island salt marsh. Limnology and Oceanography, 20(6):640–643, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Woodvoell:1977:FPE

[WWHH77] G. M. Woodvoell, D. E. Whitney, C. A. S. Hall, and R. A. Houghton. The flax pond ecosystem study: Exchanges of carbon in water between a salt marsh and Long Island Sound. Limnology and Oceanography, 22(5):833–838, September 1977. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1977:FTS

[WWK^{+}77] John J. Walsh, Terry E. Whitledge, James C. Kelley, Susan A. Huntsman, and R. Dale Pillsbury. Further transition states of the Baja California upwelling ecosystem. Limnology and Oceanography, 22(2):264–280, March 1977. CODEN LIOCAH. ISSN 0024-3590.

Ward:1970:LSP

[WWR70] F. J. Ward, Brian Wong, and G. G. C. Robinson. A liquid scintillation procedure for determining the effect of size on self-absorption of ^{14}C in Daphnia pulex. Limnology and Oceanography, 15(4): 648–652, July 1970. CODEN LIOCAH. ISSN 0024-3590.

Young:1973:RMQ

[YK73] Thomas C. Young and Darrell L. King. A rapid method of quantifying algal carbon uptake kinetics. Limnology and Oceanography, 18(6):978–981, November 1973. CODEN LIOCAH. ISSN 0024-3590.

Li:1975:PMS

[yLKR75] Chih yen Li, Kenneth M. Kiser, and Ralph R. Rumer. Physical model study of circulation patterns in Lake Ontario. Limnology and Oceanography, 20(3):323–337, May 1975. CODEN LIOCAH. ISSN 0024-3590.

Yoder:1979:CBC

[YY79] James A. Yoder and James A. Yoder. A comparison between the cell division rate of natural populations of the marine diatom Skeletonema costatum (Greville) Cleve grown in dialysis culture and that predicted from a mathematical model. Limnology and Oceanography, 24(1):97–106, January 1979. CODEN LIOCAH. ISSN 0024-3590.

Yentsch:1972:APA

[YYP72] Clarice M. Yentsch, Charles S. Yentsch, and James P. Perras. Alkaline phosphatase activity in the tropical marine bluegreen alga, Oscillatoria erythraea (“Trichodesmium”). Limnology and Oceanography, 17(5):772–774, September 1972. CODEN LI-OCAH. ISSN 0024-3590.

Zaret:1972:PIP

[Zar72] Thomas M. Zaret. Predators, invisible prey, and the nature of polymorphism in the cladocera (class crustacea). Limnology and Oceanography, 17(2):171–184, March 1972. CODEN LIOCAH. ISSN 0024-3590.

Zaret:1978:BRH

[Zar78] Thomas M. Zaret. Book review: Hurlbert, S. H. [Ed.] 1977. Biota acuatica de Sudamerica austral. Aquatic Biota — SDSU Found., c/o S. H. Hurlbert, Dept. Biol., San Diego State Univ., San Diego, Calif. 92182. xiv + 342 p. $13.00 (prepaid); $8.00 in Latin America (prepaid). Limnology and Oceanography, 23(5):1087, September 1978. CODEN LIOCAH. ISSN 0024-3590.

Zirino:1970:IZC

[ZH70] Alberto Zirino and Michael L. Healy. Inorganic zinc complexes in seawater. Limnology and Oceanography, 15(6):956–958, November 1970. CODEN LIOCAH. ISSN 0024-3590.

Zirino:1971:VMZ

[ZH71] Alberto Zirino and Michael L. Healy. Voltammetric measurement of zinc in the northeastern tropical Pacific Ocean. Limnology and Oceanography, 18(5):773–778, September 1971. CODEN LI-OCAH. ISSN 0024-3590.

Zirino:1975:MAP

[Zir75] Alberto Zirino. Measurement of the apparent pH of seawater with a combination microelectrode. Limnology and Oceanography, 20 (6):654–657, July 1975. CODEN LIOCAH. ISSN 0024-3590.

Zo:1973:BGC

[Zo73] Zinntae Zo. Breeding and growth of the Chaetognath Sagitta elegans in Bedford Basin. Limnology and Oceanography, 18(5):750– 756, September 1973. CODEN LIOCAH. ISSN 0024-3590.

Zaret:1976:VMZ

[ZS76] Thomas M. Zaret and J. Samuel Suffern. Vertical migration in zooplankton as a predator avoidance mechanism. Limnology and Oceanography, 21(6):804–813, November 1976. CODEN LIOCAH. ISSN 0024-3590.

Zirino:1972:PDM

[ZY72] Alberta Zirino and Sachio Yamamoto. A pH-dependent model for the chemical speciation of copper, zinc, cadmium, and lead in seawater. Limnology and Oceanography, 17(5):661–671, September 1972. CODEN LIOCAH. ISSN 0024-3590.